A Complete Bibliography of the Journal of Statistical Physics: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

01 March 2019 Version 1.05

Title word cross-reference

(1 + 1) [Cor87]. (1/2, 1/2) [Mar87]. (2 + 1) [Cor88]. (d + 1) [Sch86a]. 1 − μ||x||^{z} [de 88]. 1/2 [TS81, KLS88a]. 1/f [MS83b, NT81]. 1/r^{n} [KH82]. 1/|x − y|^{2} [ACCN88]. 100 ≤ Γ ≤ 160 [HMV84]. 2^{n} [Kai82]. 3 [BDS86, DC89, Zha89a]. 4 [Lin87]. 8 [Lin87]. ^{3} [HP85, deV85]. ^{4} [HP85, Tou89]. _{2} [BR80, BWdF88]. _{3} [BWdF88, CR86]. _{z} [BWdF88]. A [Rue86a, Vai89]. A_{g}RB_{f−g} [Spo83]. A_{n} [KY88b]. A_{n}^{(1)} [KY88b]. B_{6} [Kra82b]. B_{7} [Kra82b]. β [GGP86]. β^{2} < 8π [LSM89]. β ≥ 2/δ [New87]. d [BL85, O’C84a, OB84, dSVT87, YHM80]. d = 2 [GS85]. d = 2 + 1 [BP89]. d = 2 + ϵ [SH82]. d = 3 [ST81b]. D_{n} [KY88b]. D_{n}^{(1)} [KY88b]. δ [OP89]. dv = 2Δ − γ [ST81b]. g_{0}: ϕ^{4}:_{d} [BK81]. γ [MIH86, SD86]. Γ = 2 [CFS83, FJS83]. H [AC84, BS89b, Del84b, Mar81, Pia87]. i(π/2)kT [Wu86]. K [GGM86, MT84]. L^{1} [Cer88, Pol89b]. L^{∞} [Ark83]. λ(∇ϕ)^{4} [GKT84]. ≤ 2 [GM88b, GM96]. M [de 88, KST89]. Z^{k−1} [May85]. μ [EC81]. N [RG86a, AZ85, Bau83, Bur86, FTZ84, HR83a, Kal81, KD82a, KPS89, KST89, Kur83b, Lav82b, Sch87c, WS82]. n^{−2} [Sim81a]. O(n)

[ABCT86, Dup87, Lar86]. Ω [Bur86]. p [TVW89]. Φ^{4} [MHB86]. ϕ_{3}^{4} [Wat89]. ϕ_{3}^{6} [MS86a]. ϕ_{4}^{4} [GK82]. q [AB87b, Bin81, Haa89, LG87]. R [SD85]. R^{3} [Pol88a, Pol89b]. R^{n} [CEK81]. →_{t→∞} [DF80]. S [DS83b, Sha82b]. SO(N) [Ito82]. SU(3) [GGP86]. T [GS86, Hub83]. T < T_{c} [Ko88a]. T = 0 [Ruj82a]. T > T_{c} [Ko88b]. t^{−3/2} [Eva80]. T_{1} [BS87b]. T_{2} [BS87b]. T_{c} [Nel89b]. θ [AF81]. U [CR86]. U(1) [BP89]. φ^{4} [Har87, HT87a, Sor88]. x ≥ y ≥ z [WZ89]. XY [BD85, YHOF85]. y [BG80b]. Z [AV87, AV03]. Z(N) [Tru86]. Z(N^{2}) [Tru86]. Z_{3} [Dot84]. Z_{n} [RT86, Tra86].

-adic [TVW89]. -Bernoulli [Mar87]. -body [Bur86]. -city [RG86a]. -Dimensional [dSVT87, YHM80, BL85, Cor87, Cor88, O’C84a, OB84, Sch86a]. -expansion [Bur86]. -expansion-like [BG80b]. -flows [GGM86]. -function [OP89]. -furcations [de 88]. -invariant [Ito82]. -matrix [DS83b, Hub83]. -mode [FTZ84]. -particle [KST89, Sch87c, KST89]. -partitioning [LG87]. -perturber [HR83a]. -points [GS86]. -shifts [MT84]. -spin [Tru86]. -state [Bin81, Haa89]. -step [WS82]. -theorem [BS89b, Mar81, Pia87]. -trees [Lav82b]. -Trinomial [AB87b]. -vector [AZ85, Bau83, Kal81, Kur83b]. -vertex [Tru86].

1. [Nos86a]. 100 [Sug86]. 13th [Ano87a]. 17th [Ano88a]. 1979 [PD81]. 1984 [JKHV85]. 19th [Ano81o].

42nd [Ano80n]. 43rd [Ano80o]. 44th [Ano81o]. 45th [Ano88o, Ano89o, Ano82n]. 46th [Ano82o]. 47 [AV03]. 47th [Ano82p]. 48th [Ano83s]. 49th [Ano83t].

50th [Ano84u]. 51st [Ano85q]. 52nd [Ano85r]. 54th [Leb86b]. 55th [Leb86c]. 56th [Ano87p]. 57th [Ano87p]. 58th [Leb88d]. 59th [Leb89b].

60th [Leb89c]. 61st [Leb89d].

Abelian [BN87]. above [Abr85, PSMZ84]. Absence [AJ81, Bau80, BC89a, BS86, Gro81d, RT81, Sim81a, ST88, vEvH85, KLS81, Pic83, Kle89]. absolutely [CD89b]. absorbers [PPK84]. absorbing [BT81b, BT81a, Tit84, Wei86d, WT89a]. absorption [Sor89]. Abstracts [GBD^{+}86, LLE^{+}83, dAHK^{+}89]. Acceleration [LWS88, BH88a, DBK^{+}86, DK86b, ME82]. according [Fre87, Spe84]. acid [MV86]. acoustic [DP85, Hor85, Lin85, Sch89b, SSM85]. action [GGP86]. actions [May85]. Activated [FCB86, Cha86, KDMS89]. adaptive [dH84b]. Addendum [Han86a]. additional [Ano87q, VB89]. additive [Gro81b, hGH86]. Adiabatic [DF86b, DF87, SSD82, BNMT83, BOG87, GP86b, GG87, Bur86, Gro89b]. adic [TVW89]. adjoint [KP85]. adjustable [AKS89]. adsorbate [SBS88]. adsorbed [YS85]. Adsorption [SMK88, Cor89a, HFJ86, PAF84, PAF85, Pol89a, RLB86, WM83]. Advanced [Che86]. advection [KK88]. against [BP85a, LFtH88, Van88]. aggregates [HKW^{+}85]. Aggregation [EF87, AS89, BW84, Coh88, Vos84, WP82, vD89a]. Alder [EvB81, Oka86]. Alexander [PSZ87]. algebra [Dot84]. Algebraic [HCM85]. algorithm [AKvL88, BDS86, BKLP89, Gid85, Her86, MS88a, WK84, Zha89a]. algorithms [FRY87, MS87, PGS87]. alkali [HGR87]. Alley [EvB81]. alloy [Bur87, MM87]. alloys [HGR87, LPS85]. almost [CD89b, EH84, SS84b]. along [For86b]. also [Mar87]. alternating [For86b]. always [Lan84b]. American [Ano80o, SS84b]. among [Cuk83, MC81]. amplitudes [NL87]. Analog [SM89, FGMZ85, Var81]. Analogue [SLM89, SMM89]. Analogy [KWL87a, BM89, Kür88]. Analysis [Luc87, AKvL88, Ala87b, AN87, ADFJ89, BK88a, BLS82, CL89, CS83, Dic89, Enz81, FL85, FN83, GKN84, GV89a, HHB87, HD83, Kad86, LN84, Mag87, MHB86, NMM80, NGH82, SS82a, Uim84, VM87, VM88, VV89]. Analytic [GP89a, PS82c, ST81b, SMYO83, YMS83, GP85a, Mis85, NG87, Rue81]. Analytical [HMOV87, PL87, DS87]. Analyticity [CCO83, CFS84, CFGK87, Kes81, O’C84a, OB84, Zah87, BGN89, Bor87, Nau87, RS84, vEF89]. and/or [DGP84]. Anderson [BK88b, Bov89, DLS85, GOSS86, KLS89, Sim85, Sin87, Sor89]. angle [CG80]. angle-dependent [CG80]. angular [LA89]. anharmonic [BMPP80, Fri86]. anharmonicity [vHZ88]. animals [DS86c, FP87, LD87, LH88a]. anisotropic [BW89, CT84, Dun85, HMOV87, Hey82, MK85b, TS81, YHOF85, vBK87]. anisotropy [ZT85]. Annealed [VTH86]. annealing [AKvL88, Gid85, Kir84a, Mas89]. annealings [RG86a]. annihilation [Ken83, KHNA83]. ANNNI [BD82, BD88, FdF86b, MZ89, Uim84]. Announcement [Ano84a, Ano86a, Ano87b, Ano89a, Ano80m, Ano81n, Ano85n, Ano85o, Ano86n]. Announcements [Ano81a, Ano83a, Ano82m, Ano83b, Ano83p, Ano83q, Ano84q]. annuli [MPH87]. anomalies [SG87]. Anomalous [GN87, Mac86, MK89, WB87, WF88, Aha84, Hor85, Sta84]. anomaly [Suz88]. ansatz [BBP87, DdVG89]. antiferromagnet [BSF89, KLT88]. antiferromagnetic [Cal83b, RHG89, KU86]. antiferromagnets [KLS88a, KY89, Kna88, Shl80]. any [Dup89]. aperiodic [GM85, SHKM86]. Appearance [Del85]. Application [LG87, MO82, Muk82, Sta84, TC84, BGM86, EN85, KTP88, KB85, KVGZ87, PP83, Rai81, Ros84, WKUK88, vBD80a, vK87a]. Applications [MEvBD84, Nos86b, Bar85, Bau83, BGM84, BT85, Che86, Gre82a, RS87a, Rei84, RBHL86, Suz86a, Wes85a, Wol82, ZS88, DFGW89, GHKM89b, RS85b]. applied [BBF86, CF88a, Gee82]. Approach [Cos80, Dic85, GLR86, Pia83, AZ85, CCR81, CGOV84, CCR85, cC87, FFGM88, FFG^{+}88, Fai81, FL85, GKK87, GG87, Gra81a, HP81, HR85, HHO86, Jen81, KMST86b, KW83, Lam88, LN84, LSM89, Mac83c, Mas89, MR87, MR88a, NM84, Par80, PW86, PC86, PC87, Ram85, SS80, ST84a, Sto86, SYG83, TS81, Tej87, VA81, Wat89, vDG89]. approximant [AH89, OMS86]. approximants [AM84, DLT88, Gib81]. Approximate [WRL83, DP87]. approximating [HW82]. Approximation [Wei82, CPRV89, DPST88, FdF86b, FdF86a, GK82, GK84, GvD89, GPS86, GP89a, Gro81b, Gzy82a, Hub83, KEKS83, LRS89, Mac83a, Per80, PL87, Ros84, Sek83, Sen88, Yip89, Zha89b]. approximations [Fye86, GBP88a, KMST86a, Mal82, Sch84, Sch85a]. aqueous [RSK86]. Arbitrarily [GG88]. arbitrary [LN88, Per82]. area [KW88, MPH87]. areas [IKY88]. argon [dSvLH89]. argument [MM84, vE87a]. arguments [SS81, SS82b]. arising [MV86]. array [Fel86, Rub86, Sug86]. arrays [GV89b]. Ashkin [Igl89, Pfi82, Sal88, dOB89]. Aspects [DCW85, SKO87, HR83a, HR83b, Nos83, Per87]. associated [Bax89, BGM84, Dot84, Mar87, Sch86a, de 88]. associations [HH85]. asymmetric [BCFN89, DN87, DS83b, Sch86b]. asymmetry [Nie84]. Asymptotic [BP85a, BMV82, Kle80, MT82, MV82, Pol88a, PS89, Rav88, RS85a, RS85b, Spo85b, Sta88a, TM88, VMB82, BCT86, BL88, DF86b, DG83, DS86c, Fon81, Fri85, Hey82, HWW82, MR83, MS86a]. Asymptotically [Maj87, O’C84b]. Asymptotics [MZ89, ADG88, Jau85, LS87]. atom [HW86, HJ87, Koh88]. atomic [Shl85]. atoms [SM86]. attraction [Wer86a]. Attractive [AV03, AV87, Wer84a, Wer84b, Wer86a, Wer86b]. attractor [Aur87a, Aur87b, Bro89, Cos80, Cur81, IM86, Kad86, KWL87a]. Attractors [dO88, Fal87, Gra81b, GBP88b, HK85, MR83, OWY84, RP85, ST89c, Tél83, Vai89]. Au-type [LPS85]. Aubry [VT89]. Augmented [Ram85, RL86]. Australian [Ano87o, Ano87o]. autobiography [Wei86b]. autocatalytic [PP87]. autocorrelation [Kir89, VH87, VH89, vBS83]. automata [AF86b, BS89d, CM89, Di 84, GL89a, hGH86, Jen86a, Jen86b, PW85, Pan89, PTB^{+}89, Ruj87, Tak89, VTH86, Zab88, dS88a, dAHK^{+}89]. automaton [CH87, Fro89, Gra86, RK88b, Wol86]. Auxiliary [Koo86, MB89]. Auxiliary-field [Koo86]. Average [NZ86, Hub83, dH84a]. averages [SF89]. averaging [ABG^{+}89]. avoiding [AH89, BS85, BBAH83, CG85, Dom83, Gla88, Kle80, Law88, MS87, Mad88, MS88a, NCF88, SD86, Whi83, dFKP86, dFKP87]. Award [Fis84a]. axial [BD82]. Axiom [Rue86a, Vai89]. Axiom- [Vai89]. axisymmetric [SWYM87].

B [BvH89, Gro89c, Kei84]. backbone [LMJ87]. background [CJB88, For86a, GP89a, Tit88]. bacteria [TC84]. baker [EK85]. Balance [OS81, AFV85, DM81]. Banach [MR88a]. Band [YMS83, Alb89b, LGP85, Lor84, NL87, SMYO83]. band-splitting [SMYO83]. bands [lHyZ82]. barrier [FGS84, FCB86, GH82, WJS84]. basal [BWdF88]. based [CC86, HW88, HR85, Kür88]. Basic [EB88, Wol86, Tak89]. Basics [Rap87]. basis [Kra81, LS82b]. bath [FLO88, GLR86, NLSV89]. baths [Che87]. Baxter [Dav87, Gut86, RT86, Tra86]. BBGKY [GLM80]. bcc [Mou82, KMS87]. be [Rue81]. Behavior [DL89, FJS83, HM84, Lav87, NLSV89, AN84, AF86a, AB87a, ABCN81, ACT82, APSJ85, BNHF83, BCSY84, BF88, CGOV84, CCD86, CkM82, Daw83, DIP82, DFIP82, DF86b, DS86c, Dot84, FLNO88, FTZ84, FGN88, FN83, GS84a, Hey82, HWW82, IP86, Kad83, KBZ84, KMST86a, LPS88, tLC86a, tLC86b, Lib84, LPRV87a, MM87, Mei83, MR88b, NNK89, NT81, Nie84, NL85, NC85, ORD86, PPKP89, PAF84, PAF85, PP82, Pol88a, SK82, Tak89, Tra86, VA81, Wei85b, WO86, ZS80]. behaviors [YMS83]. behaviour [Fri85]. Belavin [RT86]. Belousov [RDRS87]. below [PSMZ84]. Benard [vBC88, EH89, VA81, ZL81]. Bernoulli [KLS88b, Mar87]. Bethe [GM96, BBP87, DdVG89, GW87b, GM88b, HS82b, PPKP89, SW87, dO89]. between [AGL88, BL88, BS87b, DH85b, HZ86, Kei83, KWL87a, Man84a, NR82, Pan84, RV81, Rod89, Sha82a, Van83b, VKK85, dFSZ87, ’t 88]. beyond [Sha86]. BGK [BS84, GP89b, PG83]. bias [ADG88]. biased [BR86, Dom88, KB85, LMN84]. bibliographies [HL82]. Bifurcation [GKN84, BC89b, Fed88a, FL85, GS86, LLS88, ME87, SH89a, ST89c, Tél88, VMB82]. Bifurcations [BS89a, Gra89, KW83, CEK81, EJT87, Fox83, Fra80, GT83, KY85, LN84, McL81, MV88c]. biggest [LW86]. billiard [Mac83b, MR88b]. billiards [Lév86]. Bimodality [MBG82, CPPV89]. binary [DE89b, GDx82, HGR87, MM87, Mei88a, NL85, NLC^{+}86, RL80, SZ88, VD89b, Woo89]. binary-collision [DE89b]. biochemistry [HS82a]. biological [Bar85, Nos83]. biologically [PGDT86]. biology [Dem83, Lec85, Wes83a, Wes84a, Wes85a]. biomolecules [Han83a]. biophysics [HS82a]. birth [KK84]. Bistability [LV88, LTL87]. Bistable [POD87, BM88, CCRG80, CCR81, GB84, Gou86, HNMK80, HZ82, IVMA83, KWL87a, PC87]. BKW [Tou83]. black [Ale85]. Bloch [WKC83]. Block [Wat89, DvE89, GvD89, Nea80, vDG89]. Blocking [LBF83]. blood [Row85, WP82]. Blume [BL87, DMMSR88, FE87, WW88b]. Bobylev [HP81]. bodies [Bav89, SWYM87]. body [AM83, Bur86, CF89, CWD82, CWD83, Gri88b, Gro81b, GOSS86, HV85, Hoo86, kM83, MIH86, Wag89]. boiling [DPT80]. Boltzmann [Mas89, Ark83, BG86a, BFG86, Bar88b, BL88, BBS83, Cer87, Cor80, CG80, CG81a, Cor86, Cor87, Cor88, Cor89b, DF80, Del82a, Del82b, DH86a, DH86b, Dud89, EP88, Ern84, Fis84a, FBG86, Gzy82a, Gzy82b, HP81, HN82, HD87, Hoo86, Hug83, IT88, LH85, kM83, McL89, MIH86, PG83, Pol88a, Pom85, Spo81, Str87, Tou85, Tsa88, Uch88, Uey80b]. Bond [RG86b, vVE87, CO84, EvV86, Ern87, RK88a]. bonding [LS84]. Book [AM87a, Agm88, Bak82, Bar85, Bar89, Ben85a, Ber83, BL84, Bru88, CS86a, Col86a, Col86b, Cuk82, Dar82, Dom87, Dor83, Dor85b, DZ87, Fam88, Fox83, Git88, Git89, Gli83, Gre82a, Gre82b, Gre85, Gre86, Gre89, Gri88a, Kes88, KY85, Leb86a, Leb89a, Lec85, Lee83, Lin84, LL86, Lin88, Lin89b, MBAWP87, Mas89, Nos84, Nos86b, Nos86a, Opp80, OP81, OI82, Opp82, Opp86a, Opp86b, Opp88a, Opp88b, Opp89, Pap83, Par87b, PM82, Per84, Rap87, Rap88, Rap89, Red89, Rei82, Rei84, Rei88a, Sen82, Sen86, Shl85, Shl88, Sil81, Sil82, Spo85a, Spo86a, Ste87, Wei82, Wei83, Wei84a, Wei85a, Wei86a, Wei86b, Wei86c, WW87, Wei87b, Wei89, WC89, Wes83a, Wes83b, Wes84a, Wes84b, Wes85a, Wes85b, Wes86a, Wes86b, WO86, Wes87, Wes88, Wes89, WP87, Wol82, Zwa81]. Book [bA88a, bA88b, dST86, dH89b, vK81a]. Boolean [GM85, Sta87b]. bootstrap [vE87a]. Borel [GKT84]. Bose [BP85b, FP80, MBG82, OP89, Par85, dS86a]. boson [FF87, LR88a, Loh86, Rag88, vdB83, vdBLdS84]. bottlenecks [Kog86]. bound [DG83, Ken82, McL82a, McL82b, Pic84, Pri81b, Sim80, Siu85, Sla83, TH84]. boundaries [BT81a, RW86]. Boundary [IT88, KMST86a, AKS89, ADM83, BT81b, BT81a, CP87, Cor89b, GKI85, JS83, Key83, Kur83a, Nau87, NKD84, Sch89b, ST89a, ST84a, VtLZ89, WT89a, WT89b]. bounded [KK88, Per80]. Bounds [BB80, GP85a, Mai87, vE85, ADS80, De 85b, FM87, Fri87, Gro83a, Gro83c, Guj82, KK88, Lav82a, Lav82b, Mon85, Pea81, RR82, SS89, Sok82, Tas85, vE87b]. Boyle [Shl85]. branched [DC89]. branching [Spo85b]. breakdown [KLS81, LN85, Pic83, PW84, RK88a]. breaking [BF81a, BC89a, KR85, LN84, LLS88, NB87, Sim81a, ST89c]. breath [FKRB88, FKRB89]. brick [YNS86]. brickwork [FdF86a]. Brillouin [SC87]. broadening [HR83a, HR83b]. Broadwell [Uch88]. Brownian [CDK89, CH88b, Cha82, CR81, Cuk82, Gee82, KM87, MK83a, MK84, MK85b, Maz81, Oka86, RV85, Rou87, RB88, Rue86b, SS80, SSD82, SB88, Skr84, Spo87, ST87, Wei86a, WT89b, vK86]. bubble [KSF89]. Bulk [XS89, MS82a, SKS82, Tor86b]. bump [BCD89]. Bunde [dH85a]. Bunimovich [Lév86]. Burgers [DFV89]. Burnett [HT88]. Búttiker [BL90, SH89c]. BZ [Fif85].

C [Kei84]. C. [Wag83]. Cahn [NC85]. calculate [DZVS84, Tou85]. calculating [KSF89, PS87]. Calculation [GV89b, kM81, DL82, NHH88, SD86, SF89, ZBS86, dSCKP89]. Calculations [Mut83, BBP87, EW81, EW84, EW85, HMV84, PPK84, Reb86, SM86, SIK86, VH89]. calculus [Fox87]. call [Wei87b]. can [Doi89, GN83]. Canberra [Ano87o]. Canonical [CT80, FO89, LH88b, Bon85]. Cantor [Süt89, VT89]. capable [HH85]. Capel [BL87, DMMSR88, FE87]. capillarity [Per84]. capture [Ken83, PPK84]. caricature [BDS83, DML82, DPST86, DPST88, KSS87]. Carleman [CDPP89]. Carlo [Rap87, Ber86, BS85, BK88a, BK80, BW89, dAB87, BF88, BSF89, BBAH83, CC86, CLWH82, CDM86, CG85, Cep86, CS86b, DL82, DC89, FOU86, Gla88, GWU86, Gut84, HW82, HW86, HGR87, HHB87, HR82, HMV84, JS83, JMS83, KW82, KTB86, KL83, Koo86, Kun86, Loh86, MS87, MS88a, Mei83, MGJ89, MS86b, Mut83, Neg86, PSZ87, PGS87, PA88, Rap89, RBHL86, SS89, SM86, SF89, Sug86, STA88b, Suz86a, Swe84, TWB86, WZK84, WK84, Zha89a]. Cascades [Uim84]. case [AKvL88, HHO86, Jan89b, Jan89c, KMS87, Maj89, VD89b, ZD88, KN87].

Casimir [HT87b, HT88, Tit88]. catalysts [PAF84, PAF85]. Causal [SG81]. cavity [ACS89, BG82, CGVG83]. Cayley [HKW^{+}85, HKLW87, HM84, ITH83, MTA85, MH81, Pri86, Tho82a]. CECAM [JKHV85, dAHK^{+}89]. CECAM-Workshop [JKHV85]. cell [BM84b, Han83b, HK85, KH86, KK85, OY87, Sta88a, Sti80, WP82, vBC88]. cells [BvH89, Row85]. Cellular [Ruj87, Wol86, BGWY85, BS89d, CM89, CH87, Di 84, Fro89, GL89a, Gra86, hGH86, Jen86a, Jen86b, PW85, Pan89, PTB^{+}89, RK88b, Tak89, VTH86, Zab88, dS88a, dAHK^{+}89]. cellular-automaton [RK88b]. center [AS89, KW83]. center-manifold [KW83]. centers [JW89]. Central [CG81b, FHT88, Kür88, Rez81]. certain [Bra88, BS86, GG83, KP85, Kna88, Req82, SG88, Spi85, vEvH85]. chain [AF86b, AP89, Cal83a, Fra86, Gut84, KPS89, KH82, Mei88a, Pan84, PT86, Red81, Ros86]. chain-preferred [Pan84]. chains [BTZ85, Bro82, BT89b, Con88, EW88, Gid85, Guj82, HCM85, HGPRS89, KS85, KW81, Kun86, NL85, NLC^{+}86, NL87, Sin82, VB86]. Chance [Red89, Wei86b]. change [HR82]. changes [FLS86, KFB89, OWY84]. channels [KN85]. Chaos [Git89, Nos86a, Okn88, Shl88, FKP88, Git88, Gri88a, GS84b, MKK89, RFH88, Sch87a, Sch87b, SYOM84, Ste89, Tél83, TG81, Wes85b, Wes89, YMS83, Nos86a, Wes86a]. Chaotic [LPRV87a, NNN89, Wei85b, ACT82, BM84a, CPV88, Doi89, GN83, GR89, GNKW89, lHyZ82, IM86, Lib84, McL81, OWY84, RP85, ST89c, WO86]. Chapman [qC81, qCeY81, LM86, dH89a, vK87a]. character [PS82a]. characteristic [Ala89, Sch87a, Sch87b, dO88]. characteristics [LGP85, Sch85c]. characterization [HN80, Tor86b]. characterizations [Fei87b]. charge [Ken82, MY80, Nie84]. Charged [Bor87, For86a, AJ84, EP88, For86b, For88, LM84, Pia84]. checkerboard [LW88, Lin89a]. Chemical [Hyn86, Ade86, AKN84, Blo81, BGM86, BM88, DM81, DBW81, Els84, Fox83, FN83, GHKM89a, GHKM89b, Jan89a, Jan89b, Jan89c, MPH87, OY87, POD87, RR82, SHKM86, Yeu89, ZR87, Wes86b]. chemically [Uey80a]. chemistry [RBHL86, Wei83, Wei84a, Wei86c, Wes85a]. chemotaxis [TC84]. Chippewa [dF86a]. chiral [Bax88, Bax89]. chromatic [LG87]. chromodynamics [GKT^{+}86, Kog84]. Circle [RL80, JJ86, Lan84c, May85]. city [RG86a]. Clarkson [Ano88o]. class [ABF87, BGM86, CP82, Del85, Gie89, HN80, HD83, iKOS85, iK87, LOS88, MT84, OB87, PW81, ST88, ST89b, SS83, Tra88]. Classical [DB89, Hee82, Jan82a, Jan82b, Pol88a, Rad84, Sch89a, AJ84, AM85, Ala87a, Ala87b, ACH88, AFV85, ABCT86, BP85a, Bau83, BCG^{+}80, CLWH82, Cal83b, Cha86, CC84, CFG80, CPR86, CPR87, CPRV89, DD85, FVV82a, Fon81, FM84, FP80, FS81, Gie89, Gri88b, GLM80, HW85, HMV84, Jan84, Jen81, Kal81, KIM86, Ken82, Kie89, KSF89, LO85, LM84, MNT85, MCDK85, MKDC85, MY80, MMSP84, NLSV89, O’C84b, Par82, PS88a, Per80, Per82, Per86, Per88, PT87, Rad86, Roe81, RRO88, Sch85a, Sch86a, Sin83, TG88, Tar87, Zwa83, Omn88c]. Classification [Bro81, Jau85]. Clausius [FFC83]. Clerk [Fis88]. clos [LBF83]. close [Kho87, SI87, Skr84]. closed [SV80, ZS80]. closure [Blu80]. closures [DK86b]. cloud [Pet87]. Cluster [FFC82a, KW81, AA88, ADS80, An88, BW84, BWdF88, BK80, BPK82, CF88c, DS84, FFC82b, HW82, HR83b, JKHV85, Lu87, MN84, Mor84, Ngu88, Oli88, Par82, Phi82, Red82, Sch84, Sch88b, Tas87, PS88b]. cluster-variation [Sch84]. Clustering [Ito82, BG84, LPW81]. clusters [Coh88, DS83a, GM84, HKLW87, Lam89, Man84b, NH88, NS81, PLM^{+}81, TWB86, Web84]. coagulating [vDE87, vD87]. Coagulation [HEZ83, Lu87, vDE86, vDE88]. Coarse [GKK87]. Coarse-graining [GKK87]. Coast [Ano81p, Ano82q, Ano85a, Ano86q, Ano87a, Ano88b, Ano89b, Ano84v]. codimension [GS86]. Coding [Kre80, BS89d, WZK84, WK84, Zha89a]. coefficient [Gib81, NL85, VH87, VH89]. coefficients [AB87b, Del84a, GH82, Kra82a, Kra85, Rai81, SSD82, Spe86]. coexist [DdGM89]. Coexistence [MV87, MV88b, BKL83, Nar82, Req83, Req88]. coherence [Row85]. coherent [Suz88, Gro81c]. coherent-anomaly [Suz88]. coil [PS88d]. Coincidence [Wer88]. Collapse [DS86c, LD87, Lam89, Sal86, Sti80]. Collective [Bis82, Per89, DZ87, YS85, dSC82]. collision [BGM86, CdS87a, CdS87b, DE89b, kM83, McL82b, MPH87, HR83a]. collisionless [VKK85]. collisions [KLPS83a, KLPS83b, Kiv88, Rie85]. Colored [KDMS89, AN87, DMM^{+}89, DM87, HDL89, SS80, Tou89, Val85, vK89, FFGM88]. colored-noise [SS80]. colors [dHK85]. column [SSM85]. combination [CD89c]. Combinatorial [VV89, LG87, Mas89]. combinatorics [BGH84]. combined [Iye86]. cometary [Pet87]. Commensurate [Sin82, UP84]. Commensurate-incommensurate [Sin82]. Comment [BCT86, EJT87, Eu84, KTP88, Man84a, SF89, Sza89, dH85a, Gut86]. Commentary [Yip89, Opp89]. Comments [MS83a, Ros84, Ros86, Kei84]. Committees [Ano84b]. Common [FGN88]. communication [Lan89]. communities [KN87]. commutativity [Gut86]. commutator [FO89]. compact [FP87]. comparative [IVMA83]. compared [PB83]. Comparison [FCB86, GG87, JS83, KL83, DH85b]. competing [FT88, GLM87, ITH83, MTA85, Mie89, PT86, Red81, Ruj82a, Ruj82b, Ruj84, Zha88]. Competitive [Wei86d]. complete [EW87, MG83, Uim84, dCD87]. Completely [DS87, BN88, BCFN89]. completeness [BP83]. Complex [GPS86, NS80, Jan89a, Jan89b, Jan89c, KP85, Lop88, MMSP84, WBKS83, dH84b]. Complexes [GvD89, vDG89]. component [ADM83, CLWH82, CFG80, CFS83, CJB88, Dic85, FJS83, For84, For88, For89a, HW85, HMV84, HS86, Jan84, Jan85, JLM85, JMM87, JLM95, LW83, Mac83a, PL87, Ros84, Sch83, SKA86, SS88b, Sin83, Sor88, VA81, Vie85, VB89, dAC87]. composite [GP83, SG81]. composites [JMF88]. composition [MS88b]. compression [Abd87]. comprising [Coh88]. Computation [KH86, Lan89, Lav82c]. Computational [Ker86, dST86]. computationally [Jer87]. computations [HR82]. Computer [BCFN89, CMM86, FKRB88, FKRB89, JE89, NNK89, Pan84, Pan89, WBL89, Ben80, CPRV89, Chr86, DST83, De 85a, FZ85, KM88, KL83, LCY83, PLM^{+}84, Rap88, TC84, WZK84, dAHK^{+}89]. computers [Ben80, HBB^{+}84, SV80]. Computing [Vai89, BKLP89, Cur81, ED81, HH85, Lax86, Mas89]. concatenation [BGWY85]. concavity [Syl83]. concentrated [Keh83, vBK87]. Concentration [MC81]. concepts [Ade86, Sta84]. concurrent [FOU86]. Condensation [vdBLdS84, dS86a, vdB83]. condensed [Ade86, Gro85, Sca86]. condition [Gut86, KSB85, Kur83a, MG83, Ngu87b, Oli88]. conditions [AKS89, ADM83, Cor89b, GKI85, HN82, JS83, Key83, PS82a, Sch89b, VtLZ89, Vie85]. conducting [CDM86, FGS84, MB81]. conduction [ABF^{+}85, FM87, KSS87, SS85]. conductivity [DZVS84, HK86, Hoo86, KH87, Muk82]. conductor [MV87, MV88b]. conductors [Die85, KLS84]. Conference [Ano81o, Ano83b, Ano88a, GBD^{+}86, Ano81p, Ano82q, Ano85a, Ano86q, Ano87a, Ano88b, Ano89b]. configurations [BCD89, CHM83, KD82a, KW88, WR80]. confined [Bal87]. confinement [AF81]. confining [Cor86]. confluence [DC89]. Conformal [BT89a, BPZ84, Dot84, dFSZ87]. conformation [HR82]. conformations [Bro82, BMS82]. conjecture [CC84, EvB81]. conjectures [EW87]. conjugacies [AH88]. connected [Row86]. connecting [LBF83]. connection [Red81]. connectivities [CCS87]. consequence [Büt89, Had81]. consequences [SBS88]. conservation [DB89, Lee87, Pan84]. conservative [ZWLV88]. conserving [MS87]. considerations [BR85, LW86, Sen88]. Consistent [Gri84, Mor84, HGR87, KK85, Mac83a, NM84, Yip89]. constant [Cia85, CF88b, CF89, Der83, FFC82a, FFC82b, KW88, MM87, Rod89]. constants [Fai81, PS87, WM83]. constitutive [Spe84]. constrained [ERH85, Kle80, PGDT86, WR80]. constraints [MTG87, Whi83]. constructing [Fed88a, FZ85]. Construction [AM84, Cor88, Fal87, Tar87, Yeu89, ZAT88, Lav82a]. constructions [Man84c]. Constructive [DS87, GBP88a]. consumption [Nak81]. contact [GMO89, GLR86, NLSV89, Sch85b, Sch86b]. containing [JMF88]. continuation [GP85a, PS82c]. continued [MR87, MR88a, VR85]. continued-fraction [MR87, MR88a]. Continuity [Meu84, BL86, Bru88, PG83]. continuous [BK81, CD89b, DDdG83, Del85, FP80, Gie89, KLS81, KLS89, MV88a, Sch85d, Sim81a, Süt89, TL82, WWH89, Zwa83]. continuous-spin [BK81]. continuous-time [MV88a, WWH89]. Continuum [MF83, Dom89, Kle86, WB87]. contour [BF81b, ST88]. contribution [Ano86b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano88c, Ano88d]. contributions [Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano82l, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l, Ano83m, Ano83n, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano85i, Ano85j, Ano85k, Ano85l, Ano85m, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano86m, Ano87h, Ano87i]. contributions [Ano87j, Ano87k, Ano87l, Ano87m, Ano87n, Ano88e, Ano88f, Ano88g, Ano88h, Ano88i, Ano88j, Ano88k, Ano88l, Ano88m, Ano88n, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k, Ano89l, Ano89m, Ano89n, Ket84]. control [Har87, HT87a, Rub82]. controlled [AP89, Cuk83, MY85, MC81, WF88]. Convection [Wes89, EH89, Enz81, Ste89]. convective [AMC89, Bau85, KC83, vBC88]. convectively [Dei89]. Convergence [BPT83, Bry86, Nea80, Gid85, KD84b, Oli88, ST89b, dlL82]. Convergent [O’C84b, O’C84a, OB84]. convex [Bav89, SG88]. Convexity [GG83]. cooperative [Daw83, PPK84, Suz88]. coordinate [OWY84]. copolymers [GvD89, vDG89]. core [BP85b, CW81, FH85, KY88a, KvB85, MPH87, Ros84, RL80, Shi83]. corner [KP89a, ZAT88]. corners [BPP84]. Correction [AV03, Hee82]. Corrections [mMH87, Ham87, HBB88, Har87, HT87a, Kra85, Row86, SPA88, SS80]. Correlated [dlSLW88, BLM87, HJM89, JM86, KLW86, KWL87b, LV88, Wei84b]. Correlation [BF81b, CF80, Dun85, DST89, FP80, Gee82, Gra82a, iKOS85, MTG87, MS81b, Ngu87a, Sal88, Shl80, Skr84, Sor88, APSJ85, Ber86, BOP87, Cia85, CK88b, DMMSR88, DD88, Ern87, EW81, Eva80, FJS83, GS85, Kal81, KY89, Ko88a, Ko88b, KI83, iK87, KSB85, LvB89a, LvB89b, LCY83, Lin89a, LS89b, MK85a, Nos86b, O’C84a, OB84, O’C84b, Req82, Req83, SC85b, Sch88c, SS88b, Sok82, VB80, Wer88, YAA82, Zha88, dSEC81, dSCKP89, vBLLS80, vE85, vE87b]. Correlations [Rad86, SYOM84, TG81, AGK84, AM85, BW89, BW82, FLM88, FGS85, For89a, FJ86, GDx82, GG88, HV85, Jan85, JLM85, JMM87, JLM95, KGA84, KST89, LA89, LLS88, LYAA83, Mac83b, MR86, MZ89, PL87, Pic84, RW86, Sch87c, Sim81b, Tas85, TCB89, YNS86, ZWLV88, Smi89]. Corrigendum [BCM88, Pen87]. cosmology [KLK^{+}85]. Coulomb [AF81, CPR86, CPR87, CPRV89, CJ87, FM84, FILS80, FS81, GN85, HV85, Hug85, Jan82a, Jan82b, Ken84, Lav87, LSM89, MP89, MP82, MY80, Nie84, RLB86, Smi88, Smi89, Spe86, Yan87, dFSZ87]. Coulombic [CH88c]. counterflow [SH89a]. Coupled [CM89, BMV82, BCT86, Blo81, GRS80, HW85, HD87, Kun86, RRO88, SHKM86, VMB82]. coupling [BCG^{+}80, BM84b, DR83, MM87, Meu85, Sch83, Yip89, dSvLH89, vB84, vB85]. couplings [CEGM84]. Courbage [dlL82]. Covering [Dom89]. creating [GT83]. Creutz [CR86]. crisis [RDRS87]. criteria [LS82a]. criterion [Jau86, RS87a]. Critical [AMR82, BNHF83, BCSY84, CCD86, DDdG83, Daw83, Dot84, Dup87, Nie84, OMS86, SK82, Tra86, Van83a, Zab88, dFKP87, Abr85, AN84, AF86a, AB87a, AFV85, ABCN81, Bar89, BPZ84, Bin81, BW89, BF88, BLP89, CLRT83, Coo83, DdVG89, Fam84, FB85a, FLNO88, GS81, HW82, HC88, KZ87, Kes88, tLC86a, tLC86b, Mei83, Mon85, Mou82, New86, New87, Ngu87a, Sal88, Sok81a, Sok81b, Sto86, TB88, Tas87, Tas89, YMS83, dFSZ87, BK80]. Critical-point [DDdG83]. criticality [HKLW87, SI87]. critique [Eu84, Kei84]. cross [CG80, Req82]. cross-correlation [Req82]. crossing [FCB86, Wu86]. Crossover [Pan84, RK88a, LO85]. Crystal [AG85, Ima86, Kir89, NR82, Ruj84, WKUK88]. crystalline [AJ81, ADM83, BGBS88, KS85]. crystallines [YHOF85]. Crystallinity [Wag83]. crystallization [BPU85]. Crystallographic [Che82]. Crystals [ZAT88, ATZ83, Cha84, Gat88, KFB89, KI83, LCG85, MK89, Zia86, LCG86]. Cu [BWdF88, LPS85]. cubic [AM84, BSF89, EvV86, Ern87, FCK83, Guj82, HMOV87, Huc83, YHOF85]. cumulant [HT80]. cumulants [HS88a]. curdling [Di 84]. Curie [AZ85, EE88]. Current [LMN84, Unn89]. current-driven [Unn89]. curved [WW86b]. curves [KSF89, Tél83]. cusp [BSF89, VMB82]. CVM [FdF86b, FdF86a, Ros87]. Cyber [BDS86]. cycle [BMV82, Kai82]. cycles [BvH89, Di 84]. cyclomatic [LD87]. cylinder [vBD80b, vBD82]. cylinders [Kle80].

D [BDS86, DC89, Zha89a]. Damage [dSH88, CP88]. damped [BT89b, SG87]. damping [SSD82]. data [Cer88, DK86a, kM81]. Davies [ABF^{+}85]. dealing [Gri82]. death [KK84]. Debye [CPRV89, Yan87]. Debye-Hückel [CPRV89]. Decay [AM85, BW82, LA89, MR86, VB86, CCS87, FGS85, GG88, HCM85, LO85, LN88, Mac83b, MZ89, MSW82, PL87, Pic84, RNRT83, Req82, Req83, RW86, ZWLV88, vE85, vE87b]. decaying [DS86b, Kle86]. decomposition [CP82, DV87a, EF87, KL83, NC85, SS88a]. deconfining [GKT^{+}86]. Decorated [Sha88, KPS89]. decoupling [Igl89]. decreasing [CG81a, KN85]. decrescence [Gat88]. dedicated [Leb81]. Dedication [Ano84c]. Defect [cCM89]. defects [BT89a, Cla83, Ko88a, Ko88b, TG88, YHOF85]. defined [Nea80]. deformation [SW88]. degeneracy [AL81, ARS81, PS89]. degenerate [IT88, O’C84b, ZD88]. degenerating [Fed88b]. degree [EF87, WRL83]. degrees [BDE89, ED81, PPP84]. Delay [ME87, GM85]. Delft [SIvL^{+}88]. delocalization [Bil88]. Delta [Rai81, KLW86]. delta-correlated [KLW86]. Delta-function [Rai81]. dense [CL88, LAA87, LvB89b, MK85a, NCF88, SS88b, VH87, dSvLH89]. densities [AB86, BA86, Ben84, Bon87, GBP88a, Hug85, WDG86]. Density [CCR85, JK89, Ala88, BP85a, CC84, Cia85, CFS84, CFGK87, CS83, DMM^{+}89, FF86, GMO89, Gro83c, HGR87, HS88b, KLS89, LS82c, Mai87, NL85, Ram89, RV81, Tit84, Ull81, WWH89, dSvLH89, vBLLS80]. Department [Ano81o, Ano85s]. dependence [BHR^{+}88, MCDK85, MC81, ORD86, dSvLH89]. dependent [Büt89, CG80, Ern87, FHW88, GP83, JLM85, JMM87, LOS88, SM89, vVE87, JLM95].

Depolarized [Kiv88]. depth [KN85]. Derivation [Cor89b, FM88, Hug83, Str88, CR81, CWD82, CWD83, DdGM89, Els84, GCF82, MGL87, PS87, PS82b, Ram86, ST81a, Vis85, dCD87]. derivatives [BB80]. Derrida [GMP89]. described [KK88]. Description [MT84, AZ85, BvH89, DS87, FK83, Koh88, RP85, RL86, RPSS85, Rub86, Van83a]. descriptions [BG80a, BG80b, Kra88a, RSRD83]. Detailed [AFV85, DM81]. detection [LCG85, LCG86]. determinant [GOSS86]. Determination [VH87, WM83]. determined [VH89]. Deterministic [AMC89, Nos86a, PTB^{+}89, RFH88, Sch86c, DGL83, EFN83, Els84, hGH86, GNKW89, Okn88, ’t 88]. developed [BMO^{+}84, GS84b]. Deviation [NT81, PSZ87]. Deviations [ZS80, DIP82, DFIP82, DV87b, Vai89]. devil [Uim84]. dew [KSF89]. dew/bubble [KSF89]. diagonal [Bov89]. diagonalization [FLO88]. diagram [BN87, Cho88a, LPS85, MTA85, Pfi87, Tél83]. Diagrammatic [Jan87]. diagrams [BC89b, GS88, SGL89, Tar87, Zah87]. diatomic [MS81b]. dichotomic [Van83b, FFG^{+}88]. dichotomous [BSTK87]. Dielectric [CF88b, Ram84b, CPR86, CPR87, CF89, Cuk82, FFC82a, FFC82b, FFC83, KC85, MB84, PS87, PW84, Shl84]. Difference [Lee87, PS88b, PS88c, Sin87]. different [IVMA83, JS83, MM83, Nar82, vEF89]. differentiability [Pra83]. Differential [DS80, YHM80, ABKS89, De 85b, ER81, Fed88a, Fox89, GSH88, HS88a, Ito84b, Nak86, VMA83]. Diffusion [Agm86, BS83, CCR81, Cuk83, Cuk86, EEEF84, GGL88, GB84, Gou86, Heu86, IS88, IVMA83, Keh83, MS81a, MHB87, PSMZ84, PC87, Aha84, Ben85a, BCD^{+}86, Bur86, Büt89, CCRG80, CH88c, DF85, DFL86, DFL87, DL89, Der83, Ebe86, FJ86, Fro89, GM83, GG86, HL89, HZ86, HZ82, Ker88b, KLPS83a, KLPS83b, KMST86a, KI83, KHM^{+}83, KvB85, LS82b, LS82c, MK85b, MY85, MC81, NOR89, NMM80, Pia81b, Rod89, SM89, SLW84, Szá80, TG87, WB87, WJS84, WSL83, WKL89, WMvBL88, WF88, WE82, WM83, Zwa82, dH89a, vBK87, vD88]. Diffusion-controlled [Cuk83, MY85, MC81]. Diffusion-influenced [Cuk86]. diffusion-limited [TG87, WKL89]. diffusion-reaction [BCD^{+}86, Ker88b]. diffusional [BM84b, Kos86]. diffusions [Man84b]. Diffusive [Gar87, KK85, KLSZ86, tLC86a, tLC86b, tL88, ORD86, WBL89, vB84, vB85]. dilute [CS86b, CK87b, Geo81, HP85, NCF88, RHG89, Wat80, Wu80]. diluted [BCSY84, BF88, BSF89, DN87]. dimension [Ala87b, BB82, BN88, CLP87, EC86, Fal87, Fam84, Gra81b, KGA84, KLPS83a, KLPS83b, Mis85, Nau85, OWY84, PF85, PPK84, RP85, Rad84, RW84, dS88a]. Dimensional [dSVT87, WB99, YHM80, Agm86, AF81, AF86a, ABL87, ACCN88, AJ81, AJ84, ADM83, ADG88, ADFJ89, AV81, ABG^{+}89, BM89, Ban87, BD82, BPP84, BD88, Bau85, Bax82, BZ87, BG82, BPP87, BT89a, BS83, BNHF83, BS87a, Bil88, Bin81, BDL81, Bis82, BDS83, BW88a, BPK82, BOP87, BL85, BK88b, Bov89, Boy88, Bra88, BLP81a, BH88b, BW88b, Bur87, Bur88, CLWH82, Cal83a, CCO83, CGVG83, CCD86, CCS87, CkM82, CFS83, Cho88a, CS86b, CK87b, Cia85, Cor87, Cor88, CJ87, CJB88, Cor89a, CMM86, CPV88, DP87, DF85, DL82, DLS85, DL89, Der83, DFP86, Deu83, DB80, Dic85, DPST86, DPST88, Dom89, EP88, ERH85, EF87, EN85, Fed88b, FKP88, FF86, Fel86, Fer81, FdF86b, FH87, FLNO88, FJS83, FM87, Fri85, Fri87]. dimensional [GN85, GMO89, Gar87, GD85, Gat88, GMAK84, GP86a, GS84b, HV85, Har87, HT87a, HCKS84, Hög88, IM86, Ima86, Ito82, IKY88, Jan84, Jen86b, Jer87, Jon86, KP89a, KC85, KD84b, KTB86, KLT88, KZ87, Kle86, KY88a, Ko88a, Ko88b, KU86, Kor86, KLS88b, Kur83b, KvB85, LH88a, Lav87, LS84, LPT87, LW83, LPRV87b, LMR88, LP86, kM83, MP89, MP82, Mar81, MK87, MLSK85, MM84, MOS89, MM83, MO82, MJ85, MSW82, Mür87, NNK89, Ngu88, Nie84, NR88, NHH88, O’C84a, OB84, OMS86, OP89, PW85, PV89, PMBB^{+}84, PSZ87, Par80, Per82, PP83, PL87, PTB^{+}89, Pol89a, PS85, PW88a, PW88b, Pri81a, Pri81b, Red81, RR82, Rez81, RT81, Rub84, SB84, Sal86, SD86, SI87, Sch88a, Sch85a, Sch86a, SKO87, Sel89, Ser86, Sha88, SS86]. dimensional [Sim81a, Sin82, Sin87, Spe86, Sta87a, Sta88a, Stö88, ST84b, SWYM87, TG88, Tej87, TCB89, UP84, VB86, VD89b, Wei81, Wu86, Yan87, Zol87, Zwa82, dFSZ87, vBS83, vEvH85, vE87b]. dimensionality [DV87a, Man84a, Man84b, PS88d]. dimensions [ACH88, AH88, BW84, BPZ84, BPTV88, dAB87, BR85, CG85, Dup87, Eis87, KD82a, KLS81, LMJ87, Law88, PV89, PSMZ84, PS88c, SH82, SL80, SPA88, Tas89, VM87, VM88, Wag83, dFKP87, vVE87]. Dimer [YNS86, BNHF83, EN85, Jer87]. dimers [Pri81a, Pri81b]. dipolar [LW85]. dipole [BGN89, Fon81, FS81, GP89a]. dipoles [PHB81, VB80]. Dirac [FP80]. Direct [Par89, BOP87, Gib81, Wer88]. Directed [FP87, Zha89b, IS88, LD87, PS88d, Whi89]. directional [LS84, Wer84a, Wer84b, Wer86a, Wer86b]. Discontinuity [ACCN88, MF83]. discontinuous [MV87, MV88b]. Discovery [SV80]. Discrete [BA87, MV88a, THFT87, BZ87, CIPS88, Cor87, Cor88, Far87, KV82, KH82, Kür88, MR83, Okn88, Vis81]. discrete-time [MR83]. discretized [Sen88]. discrimination [BSZ89]. disk [Ito88, KD82b, LYAA83, LCY83]. Disks [HK86, HR80, Kra81, KD82b, Rad81]. dislocations [WS83]. disorder [AP89, BK88b, Bov89, Con88, DB80, EvV86, FLNO88, HGPRS89, LGP85, LN88, NE85, PIP89, RHHG88, Ruj82a, Ruj82b, Ruj84, ZD88, vHZ88]. Disordered [Pas82, Pas83, bA88b, AAK84, BH88a, CF88c, CFGK87, CD89c, DS88b, Ern87, Fel86, FC89, FSW86, GM83, GP85b, HGPRS89, KLT88, KPS89, KFS83, LMN84, Mac83c, MGP^{+}83, Muk82, Muk83, NM84, Rub84, Ste81, Tor86b, Zwa82]. disparate [Joh89]. disparate-mass [Joh89]. dispersed [JMF88]. dispersion [KC85, Rei80b]. displacement [LHFW88]. dissimilar [FG88, FG89]. Dissipation [Lan89, Meu85, RL86, SD85, Str88]. dissipative [APSJ85, BDE89, KK85, MR83, SHKM86]. dissociating [McL89]. distances [KRS89, Lav87]. distinct [WS82]. Distinctive [MPH87]. distributed [KK85, VD89b, dH84a]. Distribution [MG88, NG87, ADS80, CLRT83, CWD82, CWD83, DM81, Fed88a, FF87, GP86a, Gol85, GM88b, GM96, HDL89, HKLW87, IKY88, LD87, Lav87, Lu87, Luc86a, MGL87, Pel88, Pel89, Per80, PP82, Req83, RV81, SC88, Spo83, Sta87a, Tél88, WW86a, WP82, dH89a]. Distributions [NH88, BC89b, HGR87, KB85, Kle80, KLS89, MB84, PP87, Pol88b, Sta89, VKK85]. disturbances [Fre87]. divergence [GRS80]. divisibility [GW87b]. do [Nel89a]. Dobrushin [Gro81d, Pra83, RS87a]. Does [Kir89, vL89]. Domain [CK87b, KT85, NNK89, Nau87, PAF84, PAF85, Sch88a]. domains [LHFW88]. dominant [BS89c]. Dormant [BvH89]. double [RV85, Ryt87, VR85]. double-well [RV85, Ryt87, VR85]. doubling [CEK81, mMH87, McL81]. doubly [CJB88]. down [AFV85]. drag [vBD80a, vBD80b]. drift [KP85, SKO87]. Driven [VtLZ89, BSTK87, BLS82, BT89b, CS89, DMM^{+}89, DM87, HJM89, HL89, KLSZ86, LKMJ85, tLC86a, tLC86b, tL88, LV88, Rei88b, Unn89, WBL89]. driving [Ala89, KvB85, LSSW80b]. droplet [MM86, MM88]. dropletlike [RS84]. Droplets [Sel89, WT89a]. drunkard [WZ89]. drying [SIvL^{+}88]. dual [HWW82]. dual-medium [HWW82]. due [KK88, Mac86]. Duffing [FGMZ85]. Dynamic [HHB87, JMS83, MGJ89, SBS88, Bin81, Bro81, MO82, Wei87a, YHOF85, Nos86b]. Dynamical [BD88, GDx82, LYAA83, Mac87, SD81, AM83, AN87, Bon85, CLP87, CL89, De 85a, DK86a, Fei87a, FSW86, GMC81, GP81, Gri88b, HL89, HN80, JE89, KY85, KN87, LLE^{+}83, Maj87, Mar85, MOS89, NHH88, OHI88, Oji89, OY87, Req82, Rue86a, Sch87a, Sch87b, SV80, Str88, ST87, Wic85]. dynamically [HGPRS89]. Dynamics [BC89b, KM88, Orb84, Pia84, SB84, SIvL^{+}88, Unn89, dH84b, AMR82, AB83, BTZ85, BCG^{+}80, BDL81, dAB87, Cal83a, CGOV84, Cha86, cC87, CT80, CPV88, Daw83, Dei89, DGL83, Ebe86, EK85, EW81, EW84, EW85, Fei88, FB85a, FCK83, Fri84, GN83, Git89, GP83, GN87, GG87, GNKW89, Gun84, HD87, HEH89, HMH84, HK86, Jen81, KTB86, KL82, KL83, KH87, LH85, LKMJ85, LPRV87b, Maj89, MNT85, MK87, MKB86, Muk83, MB89, NNN89, Ram89, Sch86a, SKA86, Spo87, Suz86b, THFT87, TVW89, Val85, Vil87, VH87, VH89, YAR89, ZR87, Zha88, ZWLV88, Zha89a, Zwa83, Wes84a]. dynamo [Gla85, Ito88]. dynamos [Wei85b]. Dyson [HS88a, Spo87].

earth [Bau85, Ito88]. Eastern [Ano81o]. easy [ZT85]. easy-plane [ZT85]. Eden [BMS87]. edge [Kur83b, NL87]. Editor [LP89, vBEL88, Leb85a, Leb85b, Leb88a, Leb88b]. Effect [DB80, OP87, SS85, SMM89, AN87, Cia85, Cor85, HW85, HMC87, KMD89, LM87, Oka86, Pia81a, ZL81, dS86a]. Effective [HP85, JMF88, Cho88a, GP85a, HM84, Suz88]. Effects [AP89, AMC89, BW89, CDM86, Fro89, Gra86, GOSS86, KC85, KD84a, Kle89, ME87, MV86, NKD84, Pes85, PF83, Req88, Tou89, Wei86d, dSvLH89, vBC88]. efficient [MS88a]. Eigen [Leu87]. eigenfunctions [Dör85a, RV85]. eigenvalue [Ull81]. Eigenvalues [RV85, Gol85]. eigenvectors [DdVG89, SF88]. Eight [ABF84, AO87, BBP87, DdVG89, DS83b, FB85b, Ham87]. Eight-parameter [AO87]. Eight-vertex [ABF84, BBP87, DdVG89, DS83b, FB85b, Ham87]. Eighth [Lee83, Ano82q]. Einstein [FP80, KF85, Omn88b, Par85, Rod89, SL80, dS86a]. elastic [BCF86, Bav86, HR88, JMF88, Kho87, Nos84, Tej89, WS84]. elasticity [Rou87, ZBS86]. electrical [Muk82]. electrode [Cor89a]. electrohydrodynamic [KFB89]. electrolytes [CW81]. electrolytic [KEKS83]. electromechanical [BMV85]. Electron [KD83, AJ81, Ala87a, Ala87b, ACH88, CDM86, Deu83, HGR87, KP86, MCDK85, MKDC85, MK85a, MKD87a, MKD87b, PFvdMZ89, SS85, SC88, SIK86, STA88b, TWB86, Wat80]. electron-electron [CDM86]. electronic [CFGK87]. electrophoretic [HM88]. electrothermal [MB81]. Elementary [Ram86, Fre87, Gra86, Kna88, Tak89, Zab88]. Eleventh [Ano85a]. eliminating [vK87a]. elimination [Bur86, Jan89a, Jan89b, Jan89c, SSD82, Shi81]. Elliott [Opp86b]. ellipsoid [Lan84a, WW82]. elliptic [AB86]. elongational [RB88]. Emden [MS82b]. Emery [WW88b]. emission [Nie88, Sek83]. Empirical [SK82]. enclose [KW88]. End [Mad88]. energies [NS80, Pel88, Pel89]. Energy [BZ83, CH87, FVV82a, Kna88, LPW81, ACS89, ADM83, BB80, Bax88, BS87a, Bre87, BL86, CCP87, Coo83, DS84, Don87, DGP84, FC89, GMP89, Gol85, Har84, NR88, Per88, PP82, PS89, Rag88, Sch88a, SS89, SS81, SS82b]. Energy-entropy [FVV82a]. Enhanced [Eva80]. Enhancement [Str82]. Enigmas [Wei86b]. ensemble [Bon85, CT80, DHB88]. ensembles [ERH85]. Enskog [qC81, qCeY81, BL88, Cer88, CK88a, LYAA83, LCY83, LM86, MK83a, Pia87, Pla86, Pol89b, Sza89, YAA82, dH89a, vK87a]. Entropic [BN88, LM87]. entropies [BPTV88, Les82]. Entropy [OHI88, Oji89, Ale85, BKLP89, Cur81, Don87, Eva89, FVV82a, FLS86, GP81, GS88, Guj82, Had81, Hoo86, KMD89, Lan84b, Len88, kM81, Mei83, MS83b, Nak81, Opp80, SS81, SS82b, Spo85a, SGL89, Tou83, Tou85]. entropy-energy [SS81, SS82b]. enumeration [Red82]. environment [cC87, IS88, KK84, KLS88b, ST84b, bARC89]. environments [DFGW89]. enzyme [Agm88]. Equal [LSSW80b]. Equal-time [LSSW80b]. Equation [WB99, Ala87a, Ark83, ABF^{+}85, Bar88a, BG86a, BFG86, Bar88b, BP83, Blu80, BBS83, BCD^{+}86, BT88, BS84, BK87, BW88b, BT81b, BT81a, CC86, CDPP89, Cer87, Cer88, CFG80, CR81, Con82, Cor80, CG80, CG81a, Cor86, Cor89b, CW81, DLT88, DFV89, Dei85, Del82a, Del82b, DF86b, DS86b, DH86a, DH86b, Dud89, EP88, Els84, Eng88, Ern84, EW84, EW85, FFGM88, FFG^{+}88, Fai81, Far87, Fed88b, FFdlM87, FK87, FL85, FM88, FBG86, Gra81a, GP89b, Gri82, GLM80, Gzy82a, HZ86, HP81, HN82, HR85, Hug83, IT88, Ken83, LH85, LRS88, LRS89, LN84, kM83, Maa84, MKDC85, MGL87, MK83a, McL89, Mei88a, MS82b, MS83a, MS81b, NMM80, NC85, PS82b, Pia87, Pla86, PG83, Pol88a, Pol89b, Pom85, RV85, RPSS85, SS80]. equation [ST84a, Spo81, TC84, TM88, Tit84, Uch88, Uey80b, Val83, Var81, Vis85, VR85, Wat80, WW86a, WKC83, WC81, YAA82, vDE86, vDE88, vK89].

Equations [AV03, AC84, ABKS89, AV87, BF81a, BMV82, BCT86, BL88, BJLS89, Blo81, CH88a, Che82, CK88a, De 85b, DFL86, DFL87, DGK81, Der87b, DG83, DR83, DB89, ER81, EHH82, Fed88a, FM84, Fox89, FT81, FTZ84, FGN88, GMLC87, GM85, GSH88, Gro84, Gzy81, Gzy82b, HL89, HS88a, HEZ83, HD87, HT88, Ito84b, KP85, Lan84c, Lee87, LS89c, MCDK85, McL82a, MK83b, Mis85, Nak86, OS81, PS88b, PS88c, Ram86, Rei80a, ST81a, Sch83, Spe84, TL82, Tho82b, Uey80a, VMA83, VMB82, Vig82, Wer84b, WRL83, ZS82, vK81c]. Equilibrium [ATZ83, Cha82, Coh88, FM84, GLM80, MJ83, Shi83, Spo86b, SW87, vBLLS80, BTZ85, Blo81, BPU85, CH88a, CHM83, DdGM89, DIP82, DFIP82, DPV86, Dic85, GW87a, GL89a, GLR86, HD83, HR86, KD82a, Kie89, KLS81, Lan84b, LS89a, LR88a, LPRV87a, LV80, Maa84, Maj89, Opp88a, Opp88b, PS82a, Saw84, Skr84, ST89b, Spo85a, SN84, WP82, Zia86, ZS80, dSCKP89, dlL82, Wer86b]. Equipartition [ACS89, TM89]. Equivalence [PW86, SG88, ’t 88, BL88, Gzy81, Man84a, OB87, Spo85b]. equivalent [MT84, Tru86]. erased [Law88]. Ergodic [Col86b, Eis87, SS86, Tak89, AB83, BOG87, CGVG83]. Ergodicity [Ito84a, NTW88]. Erratum [Ano84d, Del82a, EW85, GM96, Han86a, JLM95, LCG86, MPPT80, Pas83, PAF85, WB99]. error [Cep86]. Escape [BR86, DPV86, Han86b, HJM89, Cha82, Gar83, KB85, THFT87]. Essay [Fis88]. essentially [TM88]. Estimates [Lav82c, Lav82a, Lav82b, Ark83, BMPP80, Bre87, BF81b, Far87, Gro89a, Kes81, LT86, Par89]. estimating [Boy88]. Eu [Kei84]. Euclidean [CS83]. Euler [BPP87, BW88a, Sch85c]. eutectic [DML82]. evaluating [Pol88b]. Evaluation [YAA82, LS89b, Phi82, vVE87]. even [iKOS85, iK87]. event [Nie88]. events [LS87]. Evidence [DV87a, Pel88, Pel89, BSF89, ORD86, PSZ87, ST81b]. Evolution [Ste87, AC84, CC82, EEEF84, FN83, Fri85, Leu87, Lu87, MP82, MO83, PS85, Pul82, SS82a, Sch86c, Skr84, WRL83]. Exact [AKS89, ABCT86, ABG^{+}89, Cor87, CJB88, DH87, DD88, EFN83, Eng88, Ern84, EH84, FGS85, FLO88, FP87, FLNO88, For86b, For89a, GS85, GR87, hGH86, HR82, HZ82, HS86, Jan84, KH82, Luc86a, PPKP89, PPK84, PS88d, Red81, Sch84, Sta89, SS83, SPA88, TW85, Vis81, Wu80, Wu85, Zia86, Ala89, Cor88, DGK81, DB89, FdF86a, FB85b, Gzy81, Sha88, Sta87a, VD89b, ZS88, Ruj84]. Exactly [BD85, DFP86, AS86, Bax82, For84, KT85, KMP82, MB81, OS84, Pri81a, vD88]. Examination [RR82]. Example [GM83, GG82, RHM86]. examples [AF81, McK87]. Excess [NE85]. exchange [Gib81, MO82, Zha89a]. exchanges [ACS89, WBL89]. Excitation [KHNA83, BGWY85, Hub83, Iye86, KFS83]. excitations [Kna88, MSW82, Muk83, Web84]. exciton [Ken83, KL82]. excluded [FK83]. Exclusion [MM86, MM88, BCFN89, Bra88, Pol89a]. exclusions [Huc83]. exhibiting [SS83]. exhibits [DIP82, DFIP82, Wei87a]. Existence [CP87, Gie89, KLS88a, Par87a, CC82, Cal83b, CCP87, Cer87, Cer88, DS84, GP89b, Pla86, PG83, Pol89b, Dud89]. exit [HDL89, MOS89]. Expansion [MK83b, ANFdL^{+}83, BG80b, BCG^{+}80, Ben85b, BK88b, Bov89, Bur86, qC81, qCeY81, cCM89, CF88c, Del82a, Del82b, DS86c, FFC82a, FFC82b, Fon81, Gou86, Jan87, KW81, Kra82a, LM86, MST84, MIH86, Muk82, O’C84a, OB84, Oli88, Par82, Pol89a, Rai81, SGL89, Ull81, ZD88]. Expansions [HS88b, WB99, Bry86, BK87, BW88b, HR83b, MN84, O’C84b, Phi82, RS87b]. expectance [CS83]. expected [WS82]. experience [vL89]. Experiment [RDRS87, FCB86, Omn88b]. Experimental [Vid87]. Experiments [Lib84, WC89, Kop86, WCD85]. Explicit [Kra85, PHB81, Yeu89]. explosion [BNMT83, GP86b, GG87, MV86]. explosive [FN83, NB87, vK87b]. exponent [Boy88, BLP89, DC89, KH86, LH88a, MM87, New86, New87, Ngu87a, NHH88, Rav89, SD86, Tas87, Vai89]. Exponential [CCS87, ST89b, JM86]. exponents [BPTV88, CPV88, DF86b, DF87, Dup87, HC88, KZ87, LR88b, LPRV87b, Mar85, MR88b, Ngu87b, Sok81a, Sok81b, ZD88, dAC87, dFKP87, dO88]. expressions [GS85]. extended [AP89, Ban87, Eu84, GCF82, GCdHR^{+}84, Kei83, Kei84, Pul82, YHOF85]. Extension [KWL87b]. External [BDE89, Dei89, BSTK87, BS89a, Cor86, De 87, DF85, EP88, ENO81, Fra86, FM88, GP86a, Imr84, KFB89, LM87, MS82a, MMSP84, OP87, Per82, PWvH84, RPSS85, SW85, SS84a, SH89a, SH89b, SMM89, Uim84, ZL81]. Extrema [SL80]. Extremality [Zeg86]. extremely [RV85].

face [OB87]. faceting [AG85]. factor [AM84]. factorizable [DS83b, Sha82b]. factors [KI83, de 88]. falloff [Sim81b]. Falls [dF86a]. families [CEK81]. family [Lop88]. far [Blo81, GW87a, LV80, Saw84]. far-from-equilibrium [GW87a, LV80]. Fast [Her86, BDS86, Chr86, GG86, Jan89a, Jan89b, Jan89c, KLS84, KLSZ86, MV87, MV88b, Shi81, Zha89a, vK87a]. fast-relaxing [Shi81]. FCC [HMV84]. feed [BSZ89]. feed-forward [BSZ89]. Feigenbaum [Aur87a, Aur87b, EW87, Fra80, Nau87, TM88]. Fermi [AM87b, FP80, GP83, Hug83, Rei80a, Rei80b]. fermion [Bak82, Dav87, DL82, Kun86, OP87, ST89a]. Fermionic [Chu85b, Che87, cC87]. Fermions [WB99, JS86, Kog86, BW88b]. ferrimagnets [BF88]. ferro [BF88]. ferro-and [BF88]. ferrofluid [FH85]. ferromagnet [AZ85, Gra82a, Gra82b, HC88, IP86, Sto86, ZT85]. ferromagnetic [BB82, BLP81b, Mon88, PIP89, Pea81, Siu85]. ferromagnetism [ARS81]. ferromagnets [BL86, DdG84, De 87, Geo81, iKOS85, iK87, MMSP84, Mon80, Sim80, Syl83, VTB86, vHBW84]. Feynman [CC86]. fiber [Kod85]. Fibonacci [Süt89, Tra88]. Fick [LS82b]. Field [Ano80o, Di 84, Fer81, tLC86a, tLC86b, Par80, PYC85, RW84, AJ84, Ala87a, Ala87b, ACH88, ATZ83, BK81, BR80, BZ87, BBF86, BCD89, BLM87, BK88c, BHR^{+}88, Cam81, CH88c, DBK^{+}86, Daw83, De 87, DF85, DGK81, EP88, ENO81, FVV82b, FFS84, FT88, Fra86, FM88, GP86a, GPS86, Gla85, HMC87, Ito88, JMM87, Jan87, Ken84, Kir84b, Koo86, LM84, LM87, LRS89, Mag87, MKD87a, MKD87b, MTA85, MS82a, MMSP84, MY85, MHB86, MR88b, Nea80, OP87, Pea81, Per82, PWvH84, Pia83, SI87, SW85, SS88b, ST81b, SVBO88, Sim80, Sla83, Sok82, Str82, Suz86b, Suz88, TB88, TH84, TF86, Tho82b, Uim84, Van83a, Wag89, Wat89, Wu85, Wu86, ZS82]. Field-induced [RW84, Pia83]. Field-theoretic [Par80]. Fields [Mis87, BCSY84, BHR^{+}88, FHW88, HM84, Imr84, MN84, MB89, Pap83, KY85]. fifteenth [RS87b, Ano89b]. Figures [FKRB89, FKRB88]. filling [EH84, EN85]. films [KD84a]. final [Sek83]. fine [Req88]. Finite [BW89, Bra89, DHB88, Fro89, Ham87, MHB86, PF83, RGJ85, SPA88, AKS89, AMR82, BBP87, BP89, BT88, BS89c, BMS87, CPRV89, CF80, GM88a, HBB88, HW85, MTG87, MGTM88, MOS89, MM83, Oli88, RPSS85, RG86b, SR86, VM87, VM88, WHB85]. finite- [MOS89]. Finite-size [BW89, Bra89, DHB88, Fro89, Ham87, MHB86, PF83, RGJ85, SPA88, BT88, HBB88, Oli88, VM87, VM88]. finite-temperature [BP89]. finite-width [AKS89]. finiteness [Kra85]. First [FHW88, Wei81, Wei84b, WDG86, BDKW88, CCD86, DP87, GKK87, KT85, KMST86b, LW86, MW88, NR82, PLM^{+}81, PLM^{+}84, PS82c, PF83, Str88, THFT87, WSL83, WHM89, vBC88, Ano88o, LSSW80a]. first-order [DP87, GKK87, KT85, NR82, PLM^{+}81, PLM^{+}84, PS82c, PF83]. five [BF81a, cCM89, TL82]. five-mode [BF81a, TL82]. five-thirds [cCM89]. Fixed [FTZ84, Fei88, IP86, Vis81]. FKG [BR80, Zeg86]. flat [Doi89]. flexibility [Bro82, KS85]. flexible [HH85]. flights [CS83]. Floquet [Rei88b]. Flory [AH89]. flow [Bau85, BPP87, BS89a, Del82a, Del82b, HMH84, KMP82, Kle89, LFtH88, MD83, MIH86, RSRD83, RB88, Rub86, Wei86a]. flows [Ala89, Bab86, GGM86, KHM^{+}83, Kre80, Mis87, vBD80a, vBD80b, vBD82]. Fluctuating [Uey80a, Agm88, EJT87, GMLC87, HZ86, Kei83, KW83]. Fluctuation [DD85, PS87, Str88, Wei87b]. fluctuation-dissipation [Str88]. Fluctuations [BW88a, Bur86, FOT85, GMP89, LS89a, SC85a, SC85b, SC87, Spo81, vDE87, vD87, vD88, Abr85, AJ84, BSV87, BPZ84, BA87, CLRT83, Daw83, De 87, DH85b, EC81, Fox83, GMO89, Gee82, GVV89, GN87, Gra81a, Gri82, Gro81c, HM88, KC83, KLW86, KWL87b, KR85, LS82a, LM84, MD83, MK87, MY80, NZ86, NT81, PW88a, PW88b, Ram85, RSRD83, SS84a, SVBO88, SN84, SMM89, Suz87, TvdB88, Uey80b, Van83a, WC89, YKS81, ZS80, vD89a, vK87b]. Fluid [Par87b, ANFdL^{+}83, BP85a, BR85, CW81, DS81, Deu83, Jon86, Kle89, LHFW88, LvB89b, LYAA83, LCY83, Nel89a, Per80, Per82, Pia84, Ram89, RB88, SC85a, SC85b, SC87, Sch89b, SIvL^{+}88, Ste89, Str82, VH87, VH89, WW86a, WW86b, WC81, XS89, ZS88]. fluid-to-solid [ANFdL^{+}83]. Fluids [Wer84a, Wer84b, Wer86a, Wer86b, AM85, BG80a, BG80b, CF88b, ERH85, EvB81, FCD86, GBD^{+}86, KSF89, KVGZ87, LAA87, LA89, LS89a, LvB89a, MG88, MTG87, MGTM88, MK87, MF83, NKD84, Per86, Per88, Pia81b, Rap85, RW86, RV81, Ros86, RK88b, Tej89, Wer88, Wes89, Wol86, dSEC81, dSC82, dSCKP89]. Fluorescence [Gee82]. flux [Nel89b, SC85a, SC85b, SC87]. Fluxes [Shi81]. Fock [Zha89b]. Fokker [BP83, BT81b, BT81a, Eng88, FFGM88, FFG^{+}88, Fed88b, FFdlM87, Gra81a, GT84a, GT84b, HR85, Jau85, MGL87, MK83b, MS83a, RV85, RSRD83, TC84, Val83, VR85]. folding [Go83, MJ83]. Foldy [FC89]. following [WS83]. force [KvB85, LSSW80b, VH87, VH89]. forced [BBLT81, McL81]. forces [Ala89, AM85, BLS82, CFG80, Cor86, GLM80, Pia81b, VB89, Wer84a, Wer84b, Wer86a, Wer86b]. forcing [BS89a, Sta87b]. Ford [Row86]. Foreword [BN81, Doo85, Leb84]. forgetting [DN87]. formal [PGDT86, Ala87a]. Formalism [RHM86, Gie89, MS83b, Opp80, PV89, RV81, SG81, GHKM89a]. formation [DPT80, HV85, JKHV85, KMD89, RZ88, SB84, Spo86b, WKL89]. forming [BS88]. formula [CC86, FFC83, FC89, Val83, WJS84]. formulas [BW88b, FE87, PS87, WB99]. FORTRAN [Red82]. forward [BSZ89]. foundation [Opp82]. foundations [GCdHR^{+}84, Omn88a]. four [CF88a, HBB88, Har87, HT87a, PP83, Wu86]. four-dimensional [Har87, HT87a, PP83]. four-spin [Wu86]. four-state [HBB88]. four-vortex [CF88a]. Fourier [BH88a, DBK^{+}86, FCD86, PA88]. Fourteenth [Ano88b]. Fractal [AKN84, AH88, DDI83, Fam84, HMS82, HMS83, KBZ84, LHFW88, Nau87, Pro84, dS88a, EC86, Gra81b, Hol88, KGA84, LMJ87, Le 84, Man84b, Man84c, Orb84, Pen86, Pen87, PAF84, PAF85, Ram84a, Rap85, Rub84, Shl84, Sta84, Vos84, WS84]. Fractals [Man84b, AGK84, Aha84, AH89, Fal87, HBAM84, Kop86, MS83b]. fraction [MR87, MR88a]. fractions [VR85]. fracton [Man84a]. fracton/spectral [Man84a]. fractured [IKY88]. Franco [SS84b]. Free [Bax88, Don87, Per88, ADM83, BB80, BS87a, Coo83, Dav87, DS84, EN85, FGP80, Kur83a, NS80, PS89, Rag88, SS89, Smi89, vdB83]. free-boson [vdB83]. free-fermion [Dav87]. Fréedericksz [SMM89]. freedom [BDE89, ED81, MS86a, PPP84, WRL83]. freezing [Bau87]. French [Ano80o]. frequencies [NLC^{+}86, NL87]. frequency [Far87, LTL87, LSSW80b, Maz81, Tej89]. frequency-modulated [Maz81]. friction [FGMZ85, GB84, Gou86, Hyn86, Ryt87, VH87, VH89]. Fritz [PS85]. Frobenius [Dör85a]. Front [Bra88, Ker88b]. Frontiers [Wes88]. fronts [Ker86]. frustrated [GR87, KU86, Kor86]. Frustration [Mie89, Süt80]. fugacity [Fon81]. fully [BMO^{+}84, GS84b, SR86]. function [BPTV88, BOP87, Cep86, Cia85, CK88b, Eva80, Fei88, Gra82a, Gro83b, GGP86, Gut84, KC85, Ken82, Kir89, Ko88a, Ko88b, Kre80, Lav87, LCY83, Lev88, LV80, MGTM88, Mon83, Nau87, OP89, RP85, Rai81, Req83, SG81, Ull81, VH87, VH89, WW86a, Wer88, Zha88, dSEC81, vBS83]. Functional [Gro84, Jen81, Lan84c, LV88, Wol82, Ala88, Ala89, CCR81, CC84, CCR85, cC87, JK89, LSM89, RV81]. functions [Ala88, AB86, APSJ85, BSV87, Ber86, CF80, CWD82, CWD83, DD88, Ern87, EW81, Fed88a, Fei88, FJS83, GS85, GDx82, Kes81, KSB85, LvB89a, LvB89b, LS89b, MG88, MTG87, MK85a, MS81b, Muk82, NG87, Okn88, Req82, Rue81, Sal88, SS88b, Skr84, Sta87b, Ste81, VB80, YAA82, dSEC81, dSCKP89, vBLLS80]. functions-Application [Muk82]. Fundamental [Spo86a]. fundamentals [Ban87]. furcations [de 88]. Further [FB85b, HMV84]. fused [CW81]. fusion [Wes85b]. Future [Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano82l, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l, Ano83m, Ano83n, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano85i, Ano85j, Ano85k, Ano85l, Ano85m, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano86m, Ano86b, Ano87h]. Future [Ano87c, Ano87d, Ano87i, Ano87e, Ano87f, Ano87j, Ano87k, Ano87l, Ano87m, Ano87n, Ano87g, Ano88e, Ano88c, Ano88d, Ano88f, Ano88g, Ano88h, Ano88i, Ano88j, Ano88k, Ano88l, Ano88m, Ano88n, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k, Ano89l, Ano89m, Ano89n, Ket84]. fuzzy [HH85].

G. [HW86, dH85a]. galactic [SS82a]. galvanomagnetic [Pes85]. games [Ste87]. Gap [Ngu87b, LPW81, ST88]. gaps [Kna88]. gas [AM87b, Ala87a, Ala87b, ACH88, AB86, AB87b, Ban87, BET80, BA86, Ben85b, BGN89, BBLT81, BS89a, BBS83, BL85, BH88b, BP85b, Cer86, CL88, Cia85, CK88b, CG81a, CJ87, DS83a, Dic85, FH87, For89a, Fre87, FJ86, GN85, GM83, HV85, Hoo86, Huc83, Hug83, IT88, Joh89, JE89, KLS84, Ker86, KD83, KY88a, KSS87, KvB85, LH85, Lav87, LSM89, MR86, MNT85, MCDK85, MKDC85, MK85a, McL82a, McL82b, McL89, Mei88a, Mei83, MHB87, MIH86, Nie84, NGH82, OP89, PFvdMZ89, PB83, Pia81a, PP82, Pia83, SZ88, STA88b, ST87, dSVT87, dFSZ87, vBK87, vdB83, vdBLdS84]. gas-liquid [Mei88a]. gaseous [Bru88]. gases [CL81, qC81, qCeY81, DF85, EB88, EvV89, Fon81, FS81, FH85, Gie89, Keh83, KD83, MP89, MLSK85, Roe81, Rot89, RL80, Spe84, Uey80a, ZWLV88]. Gates [Kat85]. gauge [BP89, Bor87, BS86, CR86, Chr86, CFGK87, GGP86, Kes88, Kog84, Kog86]. gauge-invariant [CFGK87]. Gauss [MR87, MR88a]. Gaussian [BLM87, CS89, De 87, FFGM88, FFG^{+}88, GK84, Gro83b, Ito84b, KH82, Lav82c, NT81, Tas85, Van83b, WRL83, YHM80]. gelation [HEZ83, Pan84, vDE86]. Gel’fand [HR85]. general [ABF87, BKL83, DL89, FE87, Lav82c, Lav82a, Lav82b, Tas87, vHZ88, BG80a, BG80b, CS83, Gou86, Kor86, SC85a, vHGHK88a]. general-spin [FE87]. Generalization [FC89, AB86, AB87b, BA86, FFC83, Haa89, OP89, Tsa88, WJS84]. generalizations [FB85b, Wei84b]. Generalized [AMV85, BS88, BPTV88, CPV88, LYAA83, NGH82, PT87, Rou87, ABF84, AZ85, ABCT86, BMS87, CC86, CCP87, CR81, EE88, GMP89, Gri82, HS88a, HK85, KH86, KW81, LW88, Lin89a, PPKP89, SH89c, YAA82, vBD80a, Der87b, BL90]. generated [HKLW87, Lop88, TG81]. generating [Ala88, BBAH83]. Generation [LCG85, LCG86]. generator [PGS87]. genetic [KTP88]. Geometric [Tas87, Gra84]. Geometrical [MGL87, FLS86, Whi83]. Geometry [BGH84, CHM83, HFJ86, McK87, FJS83]. GF11 [BDW86]. ghost [MB89]. GHS [ENO81, TH84]. Gibbs [GM96, BLP81b, GM88b, Kle86, Kre80, MN84, Par87a, Pra83, Tsa88, Zeg86]. Gibbsian [CH88a, DvE89]. Ginzburg [BJLS89, Dei85, Fri87, FM88]. given [Ano87q, Lav82a]. glass [ARS81, BS88, BF86, CEGM84, GD85, Kam86, Wre83, vE87b]. glass-forming [BS88]. glasses [BM89, BM84a, DS88b, SYG83, Tho82b, You84, Zeg87, vE85]. Glauber [DB80]. Global [Cer87, Jen86a, Pol89b, GS84a, GP89b, Kad86, PS82a, Pla86, Zeg86]. Going [Sha86]. golden [Lan84c]. Goldstone [LPW81]. goodness [BOG87]. Gordon [BBLT81, BT89b, LSM89]. Gordon/ [LSM89]. governing [LS82a, TC84]. Grad [MS83a, Uch88]. Gradient [DS81, Mür87, dSH88]. graining [GKK87]. grand [Fam84, LH88b]. Graph [dO89, AN84, Kra82a, Kra82b, LG87, Lav82a, Wu88, vDG89]. graph-theoretical [vDG89]. graphical [PW86]. graphs [Lav82c, Lav82a, Lav82b, Row86, Spo85b, Whi89]. gravitating [Kie89]. gravitational [ATZ83, MR88b, Req88]. Green [BMO^{+}84, Cep86, KSB85, SG81, Woo89]. gridlock [Nag85]. Griffiths [Bau83, Gra82b, WS88, WW88b]. gross [GHT80]. Ground [ARS81, Bur88, FH87, Gro89a, Gro89b, Gro89c, KP86, AL81, BBF86, Bre87, BS89c, HR80, KIM86, KD84b, KLT88, KMS87, Nis81, Rad81, Rad84, Rad86, SS83, ZT85]. Ground-state [FH87, Bre87]. group [ADFJ89, AMV85, AO87, BK81, BB80, BS87a, CL89, DGK81, Fam88, FK83, GWU86, JMS83, Kad86, Lar86, LS84, Mac83c, MO82, Mis85, MGJ89, Mut83, RGJ85, Swe84, TS81, Vis85, dOB89]. groups [Che82]. growing [Cho88b, MK89]. grown [HKW^{+}85]. Growth [FF86, Gat88, PLM^{+}81, AMV85, BMS87, CK87b, DGP84, FKRB88, FKRB89, KT85, NNK89, Pan84, PYC85, Ruj84, Sch88a, WKUK88, WC89, WT89a, vD88].

H [CL81, dH85a]. H-theorem [CL81]. hadron [DDL87, Reb86]. Half [BP83, CDK89, CIPS88, Pfi87]. half-infinite [Pfi87]. Half-range [BP83]. half-space [CDK89, CIPS88]. Hamilton [BK87, BW88b]. Hamiltonian [WB99, Alb89a, BD82, Bax89, BG86b, Ben80, Ben82, BP89, Dup87, DD88, ED81, FLO88, LPRV87a, RL86, SZ88, Süt89, TM89]. Hamiltonians [DF87]. Handbook [Wei84a, Wei86c]. Hard [ANFdL^{+}83, BG80a, BG80b, BET80, DE89b, HK86, KH87, AB86, AB87b, BP85a, Bax81, BA86, Blu80, BDS83, BS84, BP85b, EW81, EW84, EW85, EC86, FH85, Gib81, Haa89, Huc83, Kir89, KY88a, KD82b, Kra88b, KvB85, LAA87, LA89, LvB89a, LvB89b, LYAA83, LCY83, LW85, Mar81, MM83, Mei83, PS88a, PP82, Pia83, Pia84, Pol84, Pom85, RS87a, Ros84, RL80, Shi83, TGN87, VB80, WW86a, WW86b, Wer88, XS89, ZS88, dSEC81, dSCKP89]. hard-core [FH85, KvB85, Ros84, RL80, Shi83]. hard-disk [KD82b, LYAA83, LCY83]. Hard-particle [BG80a, BG80b, Pol84, Wer88]. hard-point [Mar81]. Hard-sphere [ANFdL^{+}83, DE89b, KH87, EW84, Kir89, LAA87, LA89, LvB89a, PP82, XS89, ZS88, dSCKP89]. Hard-square [BET80, Haa89, Mei83]. harmonic [AP89, BPT83, DPST86, DPST88, LSSW80b, NL85, NLC^{+}86, NL87, Phi82]. Hartree [Zha89b]. Hausdorff [Fal87, Gra81b, PV89]. Havlin [dH85a]. Heat [HK86, KMP82, KSS87, ABF^{+}85, CT84, Che87, Fer81, FLO88, GLR86, HMH84, KH87, NLSV89, SC85a, SC85b, SC87, Wei87b]. Heavy [ORD86]. Heisenberg [Alb89a, Bau83, KIM86, KLS88a, KPS89, Nis81, Sto86, TS81, ZT85]. helium [SH89a]. Hemmen [BC89b]. Henon [Cur81, FR81]. Hermite [cCM89]. Herring [DdGM89]. heteroclinic [GS86]. heterogeneous [RHHG88]. heteropolymers [Gro85]. heuristic [BF86]. hexagon [AB86, AB87b, Bax81, BA86, TGN87]. hexagonal [GV89b, KIM86, dFKP86]. Hierarchical [BM84a, BGN89, BK88c, CD89c, DDI83, GK82, GKT84, HC88, Kam86, KCH88, LMR88, MP89, MS86a, Vil87, GK84]. hierarchies [HT80]. Hierarchy [lHyZ82, MK83b, dAC87]. Higgs [BN87]. High [Bar88a, Pra83, RS87b, AF86a, ABL87, ACH88, BW84, Che86, CPV88, Deu83, Far87, GB84, IM86, LPRV87b, MZ89, Nel89b, O’C84a, PF85, Sok81b, WS83, Yan87, Zeg87]. high- [Nel89b]. high-dimension [PF85]. high-dimensional [AF86a, ABL87, CPV88, IM86, LPRV87b]. High-pressure [Bar88a]. high-speed [Che86]. High-temperature [Pra83, RS87b, ACH88, O’C84a, Sok81b]. high-velocity [WS83]. Higher [MH81, Eis87, FJ86]. Higher-order [MH81]. highly [Chr86, MS88a, Wer84a, Wer84b, Wer86a, Wer86b]. Hilbert [qC81, qCeY81, Del82a, Del82b, MR87]. Hilbert-Enskog-Chapman [qC81, qCeY81]. Hilliard [NC85]. histories [Gri84]. hit [CS83]. HNC [KEKS83]. HNC-type [KEKS83]. hole [Ale85, ERH85]. Holes [DMM^{+}89]. Holsztynski [Mon88]. homeomorphic [AH88]. homeomorphisms [Lan84c]. Homoclinic [GW87a, GN83, GKN84, GS84a]. Homogeneous [HMH84, MB81, ABvT87, Ark83, Che82, Cor85, GM88a, JK89]. honeycomb [HS86, WW88b, Zia86]. Hopf [FL85, LLS88]. Hopfield [KP89b]. hopping [Der83, MGP^{+}83, NE85, SS85, VM86]. horseshoes [GT83]. Houches [Kes88]. Hove [LCY83, Oji89]. Hubbard [Hir86, Sha88]. Hückel [CPRV89]. human [Row85]. Hurst [Gra82b]. Hydrodynamic [AV87, LvB89a, MKD87a, MKD87b, Wic89, BS89d, Fri87, FM88, GMLC87, Kra88a, LPS88, Mür87, SS86, Uey80a, Uey80b, vBD80a, vBD80b, vBD82, AV03]. hydrodynamical [DIP82, DFIP82]. hydrodynamics [BDS83, DF84, DPST86, DPST88, Dud89, HLC88, IT88, Joh89, Kad85, KV82, Kiv88, Vis81]. hyperbolic [GBP88b, Rie85]. hypercubic [NCF88, Sor88]. hypernetted [Ros86]. hyperscaling [RK88a, ST81b]. hypersurface [SW88]. hypotheses [Sch86c]. hysteresis [LTL87].

I. [Ano85p, GP85b]. ice [MK89, OB87]. ice-type [OB87]. ideal [AM87b, MNT85, MBG82, dSVT87]. identities [ABF84, Bax81, FB85b]. identity [FB85a]. II [MV86, Wes84a, WB99, Ala87b, AB86, Ano89o, BMV82, BG80b, BMS82, BW88b, qCeY81, CPR87, DFIP82, Der87b, Dri82, Dud89, Eis87, FFG^{+}88, Fer81, FP80, FILS80, GHKM89b, GM85, Gro89b, HGR87, HR83b, HS89, Jan82b, Jan89b, Ko88b, Kor86, KD82b, Leb86a, LRS89, MEvBD84, Mat84, MR88a, MIH86, MGJ89, Omn88b, PPK84, Pri81b, PS88c, Rei80b, RS85b, Ruj82b, SC85b, Smi89, SGL89, TvdB88, Wer84b, WC81, Zeg86, dHK85, vBD80b, vD87, vHGHK88b]. III [ACH88, AB87b, CW81, Gro81c, Gro89c, Jan89c, Lav82b, Omn88c, Ruj84, SC87, Wer86a, vBD82]. illumination [BK88a]. illustration [ACT82]. Ilya [Kag85]. Immiscible [RK88b]. immune [BvH89, Kür88, Pan89]. immunity [Lan89]. impact [WS83]. imperfect [vdBLdS84]. implicit [Kra85]. Important [Ano80m, Ano81n, Ano85n, Ano85o, Ano86n]. Improved [GWU86, KK88, LT86, BKLP89, GGP86, NGH82, Pri81b]. improvement [Gra82b]. impurities [ST81c]. Impurity [Cal83a, GOSS86, LGP85, Ric83, SS85]. INCA [VTH86]. inclined [SPA88]. inclusions [FFC83]. incoherent [dSEC81]. incommensurate [Sin82]. incompressible [FT81, LFtH88]. increases [FTZ84]. independence [ADM83]. Independent [FLO88]. induced [BH88b, DH85b, Doi89, EH89, FCK83, HHB87, KFB89, MT83a, MT83b, Mat84, Pia83, RW84, Ryt87, SH89b, Wes85a]. inelastic [Koh88]. inequalities [AN84, Bau83, BF81b, DdG84, DMMSR88, Dun85, DT89b, FVV82a, FP80, Gra82a, HS87, Kal81, KZ87, KY89, iKOS85, iK87, New86, Sch88c, Shl80, Sok81a, Sok82, Sor88, Tas87, dHK85]. inequality [BR80, ENO81, Gra82b, New87, Sok81b, TH84]. Infinite [BPZ84, GW87b, NS81, dAC87, ABCN81, ABG^{+}89, BPT83, CGVG83, DV87b, EW88, FF87, Fri84, KD84b, MPPT78, MPPT80, MP82, MO83, Mar81, MOS89, Pfi87, PS88d, RHHG88, SKW82, Skr84, Wic85]. infinite-dimensional [CGVG83, MOS89]. infinite-dimensionality [PS88d]. infinite-particle [DV87b]. infinitely [CC82, DPV86, HBAM84, Pul82]. infinitesimal [Fre87]. Influence [KvB85, LTL87, Mei88a, PIP89, VM86, BA87, cC87, Fed88a, IP86, LS82a, MKH81, Sek83]. influenced [Cuk86]. Information [MHV89, MT87, Sch87a, Sch87b, vHGHK88b]. Infrared [MS86a]. Inhomogeneous [Cor86, FVV82b, ZP89, DFP86, HGR87, JLM85, JLM95, KdH83, LM86, MB81, RW86, RS85b, ST81c, Tej87, VTH86, dS88a, TvdB88]. Initial [FE87, MCDK85, Ala89, DH86a, HN82, KSB85]. Initiation [CP88]. inputs [HH85]. Instabilities [Fox83, Gri88a, Les82, DT89a, DBW81, Wie85]. instability [Enz81, ED81, KFB89, KC83, MB81, SSM85, Sti80, TvdB88, Van83a, vBC88]. insulator [KD84a]. integrability [Git89, Sha88, TW85]. integrable [BG86b]. integral [Ber86, CCR81, FFC82b, FCD86, Jen81, LSM89, LV88, SIK86, TWB86, WW86a, Wer84b, bA88a]. integrals [Gro83a, Gro83c, Gro84, Lav82c, Nak86, Phi82]. integrands [Woo89]. Integration [Gzy82a, De 85b, GMLC87, Gre82a, GSH88, Wol82]. Interacting [Pri86, BR86, Bau83, CPPV89, CR81, CWD82, CWD83, DPV86, DFL86, DFL87, FH87, FJ86, Heu86, MKDC85, NOR89, Skr84, Spo87, SS83, Wic89, WMvBL88]. Interaction [LLS88, AM87b, BBF86, BDE89, Bra89, BS84, CDM86, GV89b, Ker88a, OB87, Per82, Pic83, Pic84, Ric83, Ros87, Zeg87, Zol87]. interaction-round-a-face [OB87]. interactions [AB87a, BC89a, BMS82, CCO83, Cho88a, CG80, DS87, FT88, FILS80, HP85, ITH83, Ito82, KH82, Kle86, KP86, MTA85, Mie89, Par85, PPK84, Pen86, PT86, Red81, Ruj82a, Ruj82b, Ruj84, SBS88, Smi88, Smi89, Wag89, Wu86, dS86a, vEF89, Pen87]. interatomic [Bar88a, PA88]. Interface [Bil88, Ker88a, Rui89, BPU85, DFV89, Fer88, HGR87, HP87, KM87, LM87, NGH82, PIP89, Req88, RZ88, vL89]. interfaces [BR85, Gun84, HKZ88, Høy87, PS89, SIvL^{+}88, SPA88]. Interfacial [BS87a, tL88, Cho88b, Par87b, Wid88a]. Interferences [Omn88b]. intermediate [AGL88, GB84]. intermittency [Dei85, EH89, Meu84]. intermittent [SYOM84]. Internal [Sim87, WR80, BK88a, SS84a]. Internally [ERH85]. International [Ano88a, GBD^{+}86]. Interplay [Abr85]. interpretable [d’E89a]. Interpretation [For84, Gra84, Gri84, WS84]. Intersecting [Kra81, KD82b]. intersection [Par89]. intersections [BLP89]. interval [Boy87, Cos80]. intervals [Dom89]. interview [Fei85]. intractable [Jer87]. Intrinsic [NB87, PPP84, vK87b, GMC81, MGL87, SKS82]. Introduction [Ano83o, Ano86o, Lin84, Lin88, Rap88, Rap89, bA88a]. Invariance [Fer88, BT89a, DFGW89, Spe84, dFSZ87]. Invariant [FGS84, Jen86b, Tél83, BPP87, BK88b, CFGK87, GBP88a, GBP88b, Ito82, MS88b, OWY84]. invariants [BOG87, DF86b, Pet87]. Inverse [CK87a, ABKS89, Bax89, Bra89, CC84, Gou86, O’C84a, OB84, O’C84b, TCB89]. inversion [Bax82]. investigation [VH87]. investigations [CMM86]. Ion [CH88c, Hor85, DP85, GP89a, HGR87, Lav87, Lin85, SSM85]. ion-dipole [GP89a]. ion-ion [Lav87]. ionic [BH83, GDx82, KLS84, MV87, MV88b]. ions [Blu80, PHB81, VB80]. IRF [Ruj82b]. iron [MV86]. irrational [Sar84]. irreducible [Jan89c]. irregularity [PAF84, PAF85]. Irrelevant [SH82]. Irreversibility [KB86, GGM86, Len88, MT87, Omn88c, Spo85a, dlL82]. irreversible [Del84b, Eu84, EH84, GCF82, GCdHR^{+}84, Kei83, Kei84, LS82a, MK89, Pan84, Ram85]. Ising [AKS89, ABvT87, AF86a, ABL87, ABF87, ACCN88, AB87a, AF86b, AM84, AMR82, BK81, BB82, BD82, BPP84, BZ87, BT89a, Ber85, BDS86, BBF86, BS87a, Bil88, BOP87, BF88, BSF89, BLP81a, BL86, BK88c, Cam81, CCS87, CkM82, CS86b, CK87b, CF80, CHM83, CMM86, DD81, DV87a, Dav82, Dav87, DdG84, Fer81, FFS84, Fra86, FE87, GLM87, GS85, GR87, GW87b, Gra82a, Gra82b, Gra84, HMOV87, HCKS84, Her86, HS87, HM84, ITH83, IP86, JS83, Jan87, KP89a, KW82, Ker88a, Kir84b, KY89, Ko88a, Ko88b, KMS87, Kur83a, KM87, Kur83b, Lin87, LW88, LM88, Lin89a, Mal82, MTA85, MJ85, Mon80, Mon85, Mon88, MH81, Mou82, MS86b, O’C84a, OB84, PPKP89, PIP89, Per89, Pfi87, Red81, Ros87, SW85, Sch88a, ST81b, Sel89, Sim81b]. Ising [Siu85, SD85, Süt80, SPA88, Tej87, Tho82a, Tra88, VTB86, WL88, WZK84, WK84, Wu85, Wu86, Zha88, ZP89, ZS82, Zia86, Zol86]. Ising-like [Cam81]. Ising-type [ABF87]. isomerization [EFN83]. isospin [PPKP89]. isothermal [FE87, Kie89, MV86, MS82b]. isotherms [For88]. isotropic [GS85, KLT88, LFtH88]. isotropie [BGBS88]. issue [Leb81]. Itô [vK81b]. Iterated [BMV85, Ben85b]. iterates [MT82]. Iterations [Cos80]. Iterative [FBG86, BH88a]. IUPAP [Ano88a]. IV [Lav82c, Omn89, Wer86b].

J [AV03]. Jacobi [WB99, BK87, BW88b, CK87a]. James [Fis88]. jellium [MM84]. Joint [Szá80]. Jones [BDL81, Bis82, VH87, VH89]. Josephson [LKMJ85, Lom85, Sok81b, Unn89]. Journal [Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano82l, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l, Ano83m, Ano83n, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano85i, Ano85j, Ano85k, Ano85l, Ano85m, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano86m, Ano86b, Ano87h]. Journal [Ano87c, Ano87d, Ano87i, Ano87e, Ano87f, Ano87j, Ano87k, Ano87l, Ano87m, Ano87n, Ano87g, Ano88e, Ano88c, Ano88d, Ano88f, Ano88g, Ano88h, Ano88i, Ano88j, Ano88k, Ano88l, Ano88m, Ano88n, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k, Ano89l, Ano89m, Ano89n, Ket84]. Julia [BGH84, BGM84]. jump [KMST86a, KMST86b, vBK87]. junction [Unn89]. junctions [LKMJ85, Lom85].

Kac [ABCN81, CC86, Cor85, Dre87, Fei85, Kas87]. Kac-type [ABCN81]. Kadanoff [Lar86, TS81]. Kagomé [GR87]. KAM [CF88a, LT86, SK82]. Kampen [Van88]. Kauffman [CP88, KTP88, Lam88, Sta87b, Stö88]. Kawasaki [Zha89a]. KDV [McK87]. Keizer [Eu84]. Kelvin [EF87]. kicked [DGP84, Pel88, Pel89]. kicked-rotator [Pel88, Pel89]. kind [Str88, Wei87b]. Kinematics [BBLT81]. Kinetic [KC83, MKDC85, MK85a, McL82a, SS88b, Wat80, vBD80a, vBD80b, vBD82, Ala87a, Ala87b, ACH88, BS89b, BS84, BMS87, BT81b, BT81a, Cer86, CP87, CL88, CIPS88, qC81, qCeY81, CdS87a, CdS87b, DV87a, DR83, DB89, EHH82, GCF82, JKHV85, LS89c, LWS88, MCDK85, MK83a, MK83b, Pan84, PT86, PS82b, Rei80a, SS88a, Sch88a, ST84a, Spe84, SD81, Sza89, WT89a]. Kinetics [MM83, PB83, PLM^{+}84, WM83, Zif80, vDE84, Ade86, Agm86, AKN84, AMV85, BA87, Dor83, EFN83, EF87, EH84, EN85, KK85, Pol89a, SBS88]. kinks [BT89b]. Kirkwood [Zag82]. Klein [BT89b, ST84a, Tit84]. know [Nel89a]. Knudsen [Bab86, DH86a, DH86b]. Kolmogorov [EK85, Mis87]. Kondo [TW85]. Koopmanism [MR87]. Kosterlitz [BP89, Dun85, MP89]. Kramers [CCR82, MST84, ST84a, Tit84, WJS84]. Krook [Cor80]. krypton [Bar88a]. Kubo [FRY87, Woo89].

lacunary [HMS83]. Lagrangian [SB88]. Landau [BJLS89, Dei85, Fri87, FM88, Shl85]. Landauer [SH89c]. lane [RHM86, MS82b]. Langevin [BMV82, BCT86, CR81, DBK^{+}86, Der87b, Fed88a, FK87, Gri82, KP85, Maa84, Ram85, RL86, RSRD83, SKA86, VMA83, VMB82, vK81c, vK89]. Langevin-like [vK89]. Langevin-type [Fed88a, VMA83]. Laplace [AV81, Had81]. Laplacian [FOT85, vdB83]. Large [Bre87, DV87b, GMO89, KPS89, PP82, Reb86, ADG88, Che86, CPV88, DvE89, ENO81, ED81, FHT88, For89b, GM87, Jau85, KW82, KN87, MM87, MIH86, Sim81b, SS83, Vai89, Vig82]. Large- [KPS89]. large-block [DvE89]. large-coupling-constant [MM87]. Large-density [GMO89]. Large-order [Bre87]. Large-scale [Reb86, Che86, ED81]. larger [FGN88]. largest [KH86]. laser [YAR89]. lasing [HNMK80]. lateral [SBS88]. Lattice [AB86, AB87b, BA86, Chr86, Ker86, Kog84, LSSW80a, Sha86, AL81, AFV85, ADFJ89, AV81, ABG^{+}89, BET80, Bax82, BS89a, BP89, Bor87, BN87, BKL83, BH88b, BS86, CR86, CG85, CCR85, CF80, Cia85, CFS84, DBK^{+}86, DF85, DL82, DS83a, DS86c, Dic89, DPST86, DPST88, DD88, EW82, EB88, EvV89, FVV82a, FH87, Fon81, FP87, Fri87, FJ86, FH85, GR87, GW87b, GLO86, GM88b, GM96, GKT^{+}86, GG83, GS88, Guj82, GGP86, GV89b, HR82, HMV84, Heu86, Hey82, HWW82, HKZ88, Huc83, HS86, HS82b, KD84b, KLS84, Keh83, KTP88, KPS89, KY88a, Kog86, KU86, KMS87, KvB85, LD87, LH88a, LS84, LN85, Lin87, LW88, LM88, Lin89a, MPPT78, MPPT80, MLSK85, Mei83, MHB86, MS86a, O’C84b, Oli88, PMBB^{+}84, PPKP89, Par82, PB83]. lattice [PW81, Pol89a, Pra83, Pri81a, Pri81b, Roe81, Ros87, Rot89, RG86b, RL80, Sch85a, SZ88, SW87, SGL89, SWYM87, Tar87, Tas85, VM86, VM87, VM88, VtLZ89, WHB85, Wes82, WW88b, YNS86, Zah87, ZWLV88, Zwa82, dFKP86, dO89, vBK87, FILS80]. Lattice-gas [Ker86, SZ88]. lattices [AO87, AAK84, CM89, CD89c, EvV86, Ern87, EH84, EN85, Hal83, HC88, KdH83, KIM86, KI83, KHNA83, KFS83, Mac83c, NCF88, NR88, Pen86, Pen87, Per89, RS85b, TCB89, Zia86, dH84a, dHK85, vBK87]. law [AL81, BS88, Bra89, cCM89, DDdG83, Del84a, EK85, FHT88, Kes81, LS82b, Mac83b, MR88b, NNN89, Rav89, ZWLV88]. laws [DB89, GS81, GBP88b, Lee87, LR88b]. Lax [FC89]. layer [BT81b, BT81a, DH86a, Req83, ST84a, WT89a, WT89b]. layered [LM88]. layers [Req88]. leading [DC89]. learn [GN83]. Learning [DN87, PGDT86, Ste87]. Lebesgue [Süt89]. Lee [De 87, Kur83b, MS84, VTB86]. length [MS87, Mei88a, Ngu87a, O’C84a, OB84]. length-conserving [MS87]. lengths [DST89, O’C84b]. Lennard [BDL81, Bis82, VH87, VH89]. Lennard-Jones [BDL81, Bis82, VH87, VH89]. level [Jan89a, Jan89b, Jan89c, LOS88, Sek83]. levels [Har84]. Levitan [HR85]. Levy [Ben84, MB84, ZBKS89]. Lewis [HW86]. Liapunov [BPTV88, Boy88, DF87, EW88, LR88b, LPRV87b, MM87, ZD88]. lifetime [BMS87]. lifetimes [NS80]. Lifschitz [Mez87, Sim85, Sim87]. Lifshitz [Ano85p, GP85b, Kag85, KS86, LN88, NLC^{+}86, NL87]. ligand [Han83a]. Light [SSM85, BK88a, CLRT83, HM88, Kiv88, MK89, Nos84, Nos86b]. like [BG80a, BG80b, Cam81, Fye86, vK89]. likely [WZ89]. Limit [Gol85, ACH88, AZ85, BMV82, BW88a, Bon87, BCM87, CD89a, CG81b, DDdG83, DdVG89, Di 84, DvE89, FOT85, FHT88, For89b, FTZ84, Fri87, FP80, FGMZ85, GT84a, GT84b, Gro89b, Gro89c, HR83a, Igl89, KPS89, KLSZ86, MK83a, MK85b, Meu85, Mür87, Oji89, Omn88c, PS88d, Rez81, RHHG88, SS86, TF86, Uch88, VMB82, Wic89, vBLLS80, vEvH83, BCM88]. limited [KS85, TG87, WKL89, vD89a]. limits [GK84, MK84, PF85]. line [BLP81a, Fed88b, For86b, For88, HR83a, HR83b, ORD86, Szá80, WH84]. Linear [CS89, KS81, VD89b, Agm86, CP82, CdS87a, CdS87b, Coh88, Con88, DH87, DR83, EP88, FvdM89, KW81, Ko88a, Ko88b, Muk82, Sal86, Str88, Sza89, Tél83, Van88, Der87a]. linearity [WE82]. linearization [ER81, Suz87, VMA83]. Linearized [Gzy82b, Cer86, Dud89, BLS82, Rei80a]. lines [Loh86, MV87, MV88b, Nel89b, Ruj82a, Ruj82b, Ruj84, SC87]. Liouville [DLT88]. Liquid [BGBS88, Bru88, Cha84, HGR87, HP87, HEH89, KFB89, Key83, KS85, LCG85, LCG86, MK85b, Mei88a, NGH82, SH89a, Tou89]. liquid-crystalline [KS85]. liquid-vapor [HGR87, HP87]. liquid-wall [Key83]. liquid/gas [NGH82]. liquids [BS88, Cha86, DPT80, HW86, Kiv88, MK84]. List [Ano84p, Ano85p, Ano86p, Row86]. Local [BS89b, Fed88a, GS84a, GV89b, KY88b, Nar82, Pia87, Skr84, Tho82a, AB86, BA86, BLP81a, DIP82, DFIP82, Der87a, Der87b, GGL88, MS87, NM84, PG83, Rue86b, SS89, ST81c, d’E89a]. Localization [Far87, FSW86, HMC87, KC85, Spe88, Zol86, Abr85, DLS85, Sin87, Sor89, TWB86]. Localized [KFS83, SLW84, TG88, RLB86]. locally [GLM87, MNT85]. Locating [Rue86a]. locking [DH87]. locomotion [Nos83]. log [For89a]. log-gas [For89a]. Logarithmic [HBB88, HS87, Har87, HT87a]. Logical [Omn88a, Omn88b, Omn88c, Omn89]. logistic [MKH81, TG81]. Long [ADG88, EMDvB84, EvV89, Gra86, MEvBD84, MK83a, MK85b, NKD84, NKO^{+}89, TCB89, Wes84a, Woo89, AM85, Bau80, BZ83, CCO83, CFG80, Eva80, GLM80, HZ86, Ito82, Ker88a, LvB89b, LMN84, LN88, Mac86, MK85a, MK84, Mei88a, MBG82, Pen86, Pen87, Pia81b, Pic83, Pic84, RK88a, RT81, SG87, SC88, SF89, WWH89, Zeg87, vEvH83, vEvH85, vE85, vE87b]. Long-range [Gra86, NKD84, NKO^{+}89, TCB89, AM85, Bau80, BZ83, CCO83, CFG80, GLM80, Ito82, Ker88a, MBG82, Pen86, Pen87, Pia81b, Pic83, Pic84, RK88a, RT81, Zeg87, vEvH83, vEvH85, vE85, vE87b]. long-tail [SC88]. long-tailed [WWH89]. Long-time [ADG88, EvV89, MK83a, MK85b, Wes84a, Woo89, Eva80, LN88, SG87, SF89]. Longitudinal [HGR87]. Loop [Law88, vL89]. Loop-erased [Law88]. Lorentz [BBS83, BL85, CK88b, EvB81, EB88, EvV89, GM83, JE89, KD83, KSS87, LH85, MR86, MHB87, Pia81a, vBS83]. Lorenz [Fra80, Ito84a, KWL87a, Ste89, ZL81]. Lorenz-type [Ste89]. loss [CH88a, Sch87a, Sch87b]. Lovett [MG83]. Low [CkM82, DS86c, Fon81, PF85, SG87, Uim84, FGMZ85, HKZ88, Ito82, Jon86, KD83, Lav87, LS82c, OB84, Ryt87, Tar87, Zah87, vBLLS80, SGL89]. Low-and [PF85]. low-density [vBLLS80]. low-dimensional [Ito82]. low-friction [FGMZ85]. Low-fugacity [Fon81]. Low-temperature [CkM82, DS86c, SG87, Uim84, Jon86, OB84, Tar87, Zah87]. Lower [ADS80, Gro83a, Guj82, Ken82, Pri81b]. Luttinger [HSU80, OP87]. Lyapunov [CPV88, DF86b, Kai82, KH86, Mar85, MR88b, NL85, Rav89, Vai89].

M [Ano85p, GP85b]. machines [Ben80, Ben82, Mas89]. macrodynamics [KN87]. Macromolecular [Bro82, BMS82, HM88]. macromolecules [Fam88]. Macroscopic [PW88a, PW88b, EK85, Rub86, ST81a, Jan89a]. Magnetic [AGK84, Gla85, GLO86, FFS84, HMC87, Ito88, JMM87, MKD87a, MKD87b, MTA85, MS82a, Sch83, SS88b, Uim84, Wu85, Wu86]. magnetically [Bal87]. magnetism [FT88, PT87]. Magnetization [BPP84, LW88, Syl83, AF86a, ACCN88, AM84, BL86, DD81, De 87, GD85, Geo81, Lin87, LM88, Mon80, Pea81, Siu85, Sla83, Wre83]. magnetizations [KP89a]. magnetoelastic [BTZ85]. Magnetotransport [OS84]. magnets [BCSY84, KTB86]. majority [DvE89]. majority-spin [DvE89]. Making [Shl88]. Mandelbrot [Nau87]. Manhattan [Dup87, DD88]. MANIAC [And86]. manifold [KW83, vK86]. manifolds [FR81, MT87]. mantle [Bau85]. many [AM83, Bak82, CC82, CWD82, CWD83, DPV86, FOT85, Gri88b, GOSS86, Neg86, OP87, PPP84, Wag89, vEF89]. many-body [AM83, CWD82, CWD83, Gri88b, GOSS86, Wag89]. many-fermion [Bak82, OP87]. many-particle [Neg86, vEF89]. map [CM89, DH87, Doi89, Dör85a, FCK83, JJ86, MR87, MR88a, Str87, Tél83, YMS83, de 88]. Mapping [Sch86a, FR81, HK85, KH86]. mappings [BZ87, GS81]. maps [BGH84, BN88, Boy87, CEK81, FKP88, GS84b, Hel88, Lop88, MS88b, NZ86, Okn88, PV89, SMYO83]. Mark [Kas87, Fei85]. Markoffian [Zwa82]. Markov [DFGW89, FZ85, Gat88, Gid85, HCM85, KMST86a, KMST86b, Lév86, MT84, NNN89, SS84a, Str88, Van83b, Zeg86]. Markovian [CH88b, Els84]. mass [AS89, FO89, Joh89, ORD86, Sha86, ST88, VKK85]. masses [MM83]. massive [MK83b]. massless [BLM87]. Master [Els84, Fai81, FL85, Ken83, LN84, RPSS85, CH88a, Con82, DG83, Gzy81, NMM80, Vig82]. materials [OS81]. mathematical [Fis88, GBD^{+}86, Kra81]. Mathematics [Ano81o, Ano85s, Lax86, SS84b]. Mather [VT89]. Mathieu [CD89b]. Matrices [Rav89, CK87a, CPV88, Gro83c, Gut86, KW81, Sha82b]. matrix [AKS89, ABCT86, Boy88, Con88, DZVS84, DS83b, Gie89, Hub83, JMF88, Kho87, Lam89, MPK88, PS89, SD86, SC85b, VR85, Zol86]. matter [BH83, Kie89, Red89, Sca86, Shl85]. Mattis [Nis81]. maximal [Lop88, Zeg86]. Maximum [MS83b, Had81, Lan84b, Opp80]. Maxwell [CG80, CG81a, DF80, Fis88, HN82]. Maxwellian [Dic85, EP88]. Maxwellians [PG83]. Mayer [Ben85b, Bry86, BK87, BW88b, Lav82c, Lav82a, Lav82b, Rai81, Spe86, WB99]. McKean [KST89, Sch87c]. Mean [Ala87a, Ala87b, ACH88, Ken84, Pea81, Sim80, Sok82, THFT87, TB88, TH84, VB80, Bra89, Daw83, DGK81, FVV82b, FT88, GPS86, GP89a, LFtH88, Mac83a, PHB81, PWvH84, Ros84, SW85, Sla83, Smi88, Smi89, Tho82b, ZS82, LRS89, Van83a]. Mean-field [Ala87a, Pea81, Sok82, Ala87b, Daw83, FT88, GPS86, SW85]. meanfield [Bon87]. meaning [Oji89]. means [DST83, Lav82a, Lav82b]. Measure [Len88, BK88b, Cha82, FLS86, Lop88, MS88b, Meu84, Süt89, VTB86, WW88a]. measure-theoretical [Meu84]. Measurement [Nos84, Lan89]. Measurements [GRS80, LR88b]. measures [BPP87, Fri86, GBP88b, Gut84, Man84b, MR83, Zeg86]. Meblin [BGM84]. mechanical [Bau87, Ben84, Ben80, Ben82, BW88a, BS81, CDK89, CD89a, GKI85, Jan85, KW88, KD83, Par87a, PW88a, PW88b, Row85, Ruj87, SLM89, Wre83, ’t 88]. Mechanics [Ano80o, Ano80n, Ano81o, Ano81p, Ano82q, Ano82n, Ano82o, Ano82p, Ano83s, Ano83t, Ano84u, Ano85a, Ano85q, Ano85r, Ano85s, Ano86q, Ano87o, Ano87p, Ano87a, Ano88b, Ano88a, Ano88o, Ano88p, Ano89b, Ano89o, Leb86b, Leb86c, Leb88d, Leb89b, Leb89c, Leb89d, BFA86, BH83, BCF86, Bav87, Bav89, BEF86, Chu85b, Dem83, Der87a, Der87b, Dup89, FS81, Gol87, Gre85, Gre89, Gri84, Høy87, Kie89, KW81, Kog84, Kra81, KD82b, LG87, LPW81, LLE^{+}83, Leb86a, LRS88, Leb89a, LRS89, Leu87, Lin84, Lin88, Lin89b, LPRV87a, MS82b, Nak86, Nel89b, Omn88a, Omn88b, Omn88c, Omn89, Par85, PS88b, PS88c, Ram86, Rei84, Req82, Rue86b, Sch86c, SG81, Spo86a, WP82, Zwa81]. Mechanism [ZR87, KI83, PGDT86]. mechanisms [BGWY85, Bro81]. media [EEEF84, EMDvB84, FM87, GP85a, Gro81b, Gro81a, Gro81c, HZ86, HHO86, LHFW88, Le 84, MEvBD84, MG88, Ram84b, RHHG88, Sta84, Tor86b, Vid87, dH89b]. medical [PP83]. Medium [GP83, DS86b, Gar87, HWW82, JMF88, Nau85, Ser86, ST81c]. Medium-dependent [GP83]. Meeting [Ano81o, Ano82m, Ano83p, Ano83q, Ano84q, Ano84v, Ano85s, Ano87o, Ano88o, Ano88p, SS84b, Ano80o, Ano80n, Ano82n, Ano82o, Ano82p, Ano83s, Ano83t, Ano84u, Ano85q, Ano85r, Leb86b, Leb86c, Leb88d, Leb89b, Leb89c, Leb89d]. Meetings [Ano87p]. Melbourne [Ano85s]. Melnikov [Jau86]. Melting [Ima86, Kho87, Fis84b]. membranes [ADFJ89, Kle89]. Memory [CH88a, BGM86, Ern88, Leb81, ME87, VH89]. meniscus [DdGM89]. Mermin [BPK82, MM84]. mesh [Lav82c]. Mesoscopic [Nak81, Bur86, Jan89b, Jan89c]. metal [KD84a]. metal-insulator [KD84a]. metallic [HGR87]. metals [HGR87, KC85, Muk82, OS84, Pes85].

Metastability [Dav82, RS84, Sch85b, HCKS84, Jan87, PS82a, THFT87]. Metastable [ABvT87, CGOV84, DGK81, Han86b, LO85, MKB86, NS80]. methane [dSvLH89]. method [AN87, An88, Bax82, BWdF88, BS85, Boy88, CC86, Che87, Cho88a, CFS84, ER81, ED81, Fre87, GOSS86, GWU86, HW82, HR85, Loh86, LS89b, MS88a, MR88a, MO82, Mon88, Mor84, NG87, PWvH84, PS82b, Sch88b, Suz88, Tou85, WKUK88, WKC83, WT89b, YS85, vBD80a, vK87a]. Methods [BTZ85, Rei84, Bak82, BB80, Ber86, BBAH83, Fed88a, FCD86, IVMA83, KSF89, Koo86, LM86, MY85, Rap87, Rap88, Sch84, Suz86a, Wei84a, Wei86c, bA88a, Wes86a]. metric [Aur87a, Aur87b]. micellar [SZ88]. micelles [BGBS88]. Micro [HLC88]. Micro-hydrodynamics [HLC88]. microcanonical [DHB88, RZ88]. Microdynamics [GHT80]. microinclusions [Kho87]. micromorphic [OS81]. Microscopic [BCD^{+}86, DdGM89, Els84, Key83, LS82b, LPS88, PB83, PS87, Rei80a, Rei80b, SKS82, Ben80, DFV89, DGL83, Tit88, Wag89, dCD87]. Microstructure [Tor86b]. microstructures [DS81]. Migdal [Lar86, TS81]. migration [Han83a]. Mikhailovich [Kag85]. minority [ABL87]. Misra [dlL82]. mixed [AB87a, GW87a, Kam86, Uey80a]. mixed-mode [GW87a]. Mixing [NT89]. mixmaster [Eis87]. mixture [Blu80, GDx82, MM83, Woo89]. mixtures [FGS85, HP85, Joh89, Mei88a, VD89b]. Mobility [HM88, KD83, Rod89]. Mode [DML82, DR83, dSvLH89, vB84, vB85, BF81a, DH87, FT81, FTZ84, GW87a, Sch83, TL82, Yip89]. Mode-coupling [dSvLH89, vB84, Sch83, Yip89]. Model [MTA85, Spo83, AS86, AKS89, ACS89, AB87a, AM84, ABF84, AB86, AB87b, ABCT86, AMR82, AB83, ABF^{+}85, APSJ85, BF81a, BK81, BB82, BD82, BD88, BBP87, Bax81, BA86, Bax88, Bax89, BZ87, BG82, Ben80, BT89a, BDS86, BNHF83, BBF86, BS87a, Bil88, Bin81, BC89b, BM84b, BM88, BN87, BOP87, BK88b, Bov89, BT88, Bra89, BD85, BS84, BK88c, BMS87, Bur87, BL90, CDK89, CD89a, CCP87, CGVG83, CH88b, Cer86, Cer87, CCS87, CkM82, Cho88a, CS86b, CK87b, Chu85b, Cla83, CHM83, CFGK87, Cor80, Cor85, CP88, CMM86, CW81, DD81, DV87a, DP87, DDdG83, Dav82, Daw83, DFV89, DL82, Der83, DdVG89, DB80, DS83b, Dot84, DP89, EE88, EFN83, Els84, EF87, Fer88, Fer81, FdF86b, FdF86a, FFS84, FLO88, FB85b]. model [For86a, Fra80, FCD86, Fri85, Fri87, FH85, GLM87, GK82, GKT84, GK84, GP86a, GS85, GR87, GM84, GW87b, GP89a, GOSS86, Haa89, HW82, Ham87, HBB88, HMOV87, HSU80, HR82, Hio83, Hir86, Hög88, HD87, HM84, HR86, Igl89, ITH83, Ito84a, Ito88, Jan87, Kad85, KP89a, KT85, Ker88a, Ker86, KPS89, KSF89, KMP82, Kir84b, KLS89, Ko88a, Ko88b, KP89b, KU86, KMS87, KST89, Kur83a, KM87, Kür88, Lam88, Lar86, LS84, LPT87, Leu87, LR88a, LGP85, Lin87, LW88, LM88, Lin89a, LMR88, Luc86a, Mal82, MV87, MV88b, MKH81, MO82, MB81, MKB86, MPK88, MZ89, MS84, MJ83, MJ85, MH81, MB84, Mou82, Nag85, NOR89, NR82, NE85, Nis81, NKO^{+}89, O’C84a, OB84, OS84, OY87, OP87, OB87, PF85, Pan89, Par87a, Pas82]. model [Pas83, PFvdMZ89, Pel88, PB83, PHB81, PWvH84, PW86, Pfi82, Pfi87, Pic83, Pic84, PW84, PG83, Pri81a, Rag88, RT86, Rod89, RS84, RLB86, RHG89, RZ88, RG86b, Rui89, Ruj82b, Sal88, SW85, SS88a, Sch88a, Sch87c, ST81b, SH89a, SKO87, SR86, Sha88, SS86, Sim81a, Sim81b, Sim85, Smi88, Smi89, SD85, Spo87, Sta88a, SN84, Ste89, Sti80, Stö88, SH89c, Str88, TS81, TL82, Tej87, TF86, Tho82a, Tra86, TGN87, Tra88, Tru86, Uch88, Uim84, VB80, Wei87a, WK84, Wu80, Wu85, Wu86, WW88b, Wu88, YHM80, Yeu89, ZP89, ZS82, ZL81, Zol86, dOB89, vBS83, vD88, Bro82, Pel89]. modeling [CDPP89, MS81a, MR88b, Ros87]. Models [LR88b, vHZ88, ABvT87, AN84, AF86a, ABL87, ABF87, ACCN88, ADFJ89, ABCN81, BPP84, Bau83, Bax82, Ben82, BM84a, Ber85, Bon87, Bra88, BW82, BL87, BS86, BS87b, Cal83b, CG85, Cor87, Cor88, Cor89a, CG81b, Dav87, DMMSR88, DS80, DDI83, DFP86, Dic89, Dre87, Dup87, EEEF84, EW82, EvB81, FVV82b, FE87, FM88, GMP89, GT83, GL89a, GLR86, GRS80, Gra84, GT84a, GT84b, GG83, HCKS84, HN82, Her86, HKZ88, HS87, HD83, HR85, Ito82, JS83, Jau85, JKHV85, KW82, Kal81, KIM86, KLS84, Kat85, KTP88, KK88, KU86, Kor86, KSS87, KY88b, Kur83b, LPS88, LWS88, Mon85, Mon88, MS86b, NNK89, NS81, Nie84, OMS86, Opp88a, OB87, PYC85, Pea81, Per86, Per88, PT86, PS88b, PS88c, Red81, Roe81, Ruj87, SG88]. models [Sch86a, Sch88c, SS82a, SVBO88, Sel89, Sha82b, Siu85, ST88, ST89b, Sor88, Sta89, SD81, SYG83, SPA88, TG88, Tas87, WKUK88, WL88, WZK84, Wei86e, WC89, Wre83, Zah87, dFSZ87, vBK87, vE87b]. moderately [MK85a, NOR89]. modern [Lin88]. modes [Bis82, Bur86, Joh89, Per89, Rei80a, Rei80b, Sch89b, dSC82]. modification [OP87, Sch86c]. modified [Pol89b, Ros86, SH89a]. modified-hypernetted-chain [Ros86]. Modular [TGN87]. Modulated [FT88, JW89, Maz81, SG88]. moduli [BCF86, Bav86, Tej89]. Molecular [Ben85a, BDL81, BGWY85, Bon85, EW81, EW84, EW85, Per84, RBHL86, SIvL^{+}88, CT80, GG87, HEH89, HMH84, HK86, KTB86, KL83, KH87, LH85, MK87, MK85b, PAF84, PAF85, Rap85, VH87, VH89]. molecule [CK88a, GG86]. molecules [Abd87, Dic85, Har84, Mei88a, MS81b, SM86, YS85]. Moment [Bak82, WT89b]. moments [BMS89, LSSW80b, NH88, SKW82, WHM89]. momentum [SH82]. momentum-shell [SH82]. monolayer [YS85]. monomer [Jer87]. monomer-dimer [Jer87]. monomode [Kod85]. monotonic [Boy87, PL87]. Monte [Ber86, BS85, BK88a, BK80, BW89, dAB87, BF88, BSF89, BBAH83, CC86, CLWH82, CDM86, CG85, Cep86, CS86b, DL82, DC89, FOU86, Gla88, GWU86, Gut84, HW82, HW86, HGR87, HHB87, HR82, HMV84, JS83, JMS83, KW82, KTB86, KL83, Koo86, Kun86, Loh86, MS87, MS88a, Mei83, MGJ89, MS86b, Mut83, Neg86, PSZ87, PGS87, PA88, Rap89, RBHL86, SS89, SM86, SF89, Sug86, STA88b, Suz86a, Swe84, TWB86, WZK84, WK84, Zha89a, Rap87]. Montroll [Opp86b]. morphogenesis [KHM^{+}83]. Morphology [Cho88b]. Morse [AP89]. Mossotti [FFC83]. motility [Sta88a]. Motion [Lan88, PPK84, BH88b, CDK89, CH88b, CK88a, Cuk82, Gee82, Ken83, kM81, MK84, Maz81, McL81, Oka86, Pet87, Rav88, RV85, Rou87, RB88, SS80, SB88, ST87, Wei86a, Wei87b, Wes84b, WT89b, vK86]. motions [BG86b, DFGW89]. MSA [HHO86]. multichannel [TW85]. multicomponent [GP85a, Sim80]. multiconnected [ZP89]. Multicritical [Opp86a, Sen86]. multidimensional [CCRG80, WJS84]. Multifluxon [LKMJ85]. multifractal [GV89a]. Multifractality [HR88]. multifractals [Blu89]. Multiparticle [Vos84]. multiphase [FM87]. Multiple [EE88, Gro81b, Gro81a, Gro81c, CF88c, DD88, Joh89, KHNA83, MGP^{+}83, Nak86, Wer86a]. multiple-trapping [MGP^{+}83]. multiplicative [KFB89, RL86, bARC89]. multiprocessor [Che86]. multisite [BS89d]. multispecies [CWD82, CWD83]. multispin [PGS87, WZK84, WK84, Zha89a]. multistage [LBF83]. multistate [RS85a, RS85b]. multivariate [NMM80]. Must [Rue81].

N [HW86]. Narrow [De 85b, Alb89b, Lor84]. narrow-band [Alb89b]. nascent [LTL87]. National [Ano87o]. natural [Chu85a, RFH88, Wei87a, Wei84a, Wei86c]. nature [Coo83, Rap85, VTB86]. Navier [Ala88, Ala89, BF81a, Che82, Cor89b, FT81, FTZ84, FGN88, TL82, Vis85]. near [AMC89, AG85, BG86b, DT89a, GKN84, GS84a, Igl89, Jan82a, Jan82b, KC83, LW85, NL87, SMYO83, SYOM84, Tit84, vBC88]. nearest [BD82, Coo83, Mür87, Per82, Pol89a, Ros87]. nearest-neighbor [Per82, Pol89a]. Néel [KLS88a]. Negative [Rot89, For86b]. Negative-viscosity [Rot89]. Negentropy [Nak81]. neighbor [BD82, Huc83, Mür87, Per82, Pol89a, Ros87]. neighborhood [Coo83]. nematic [KFB89]. nervous [Kür88]. NESS [Key82]. Network [BvH89, BSZ89, KP89b, LBF83, Rei88a, ZP89, dAC87]. networks [BMV85, Blu89, DZVS84, DN87, Dom88, Dup89, KM88, MHV89, NG87, PGDT86, Sta87b, dH84b, vHGHK88a, vHGHK88b]. Neural [Dom88, DN87, KP89b, Mas89, MHV89, vHGHK88a, vHGHK88b]. neurons [PGDT86]. neutral [LW85, PFvdMZ89]. neutrality [GLM80]. neutralizing [GP89a]. neutron [Sen88]. Newton [Shl85]. Newtonian [HEH89]. next [BD82, Ros87]. next-nearest-neighbor [BD82, Ros87]. nitric [MV86]. Noise [Dei85, EH89, Fed88a, HS89, MT83a, MT83b, Mat84, MV88c, Ryt87, Wes85a, AN87, BMV82, BSTK87, BDE89, CL89, DMM^{+}89, Dei89, DM87, Doi89, FFGM88, FFG^{+}88, FCK83, Fri86, GT84a, GT84b, Gra89, Gro83b, HDL89, HJM89, HL89, HS88b, HGPRS89, Jau85, Jau86, KFB89, KDMS89, Lan89, LTL87, LMN84, LV88, Luc86a, Luc86b, Luc87, MW88, MKH81, MKK89, MS83b, NE85, RL86, RPSS85, SS80, SS84a, Sch89a, SH89a, SH89b, Tou89, Val85, Van83b, VMB82, ZL81, vBC88, vK89]. Noise-induced [EH89, MT83a, MT83b, Mat84, Ryt87, Doi89, FCK83]. Noise-sustained [Dei85]. noises [CS89]. Noisy [Wie85, Kai82, Lan88, NZ86]. Non [CH88b, DvE89, GK84, Spo85a, Zwa82, CS89, CH88a, HEH89, KP85, KLS88b, Opp88b, Str88, Zeg86]. Non-equilibrium [Spo85a, Opp88b]. non-FKG [Zeg86]. Non-Gaussian [GK84, CS89]. Non-Markoffian [Zwa82]. non-Markov [Str88]. Non-Markovian [CH88b]. non-Newtonian [HEH89]. non-self-adjoint [KP85]. non-Wiener [KLS88b]. nonadditive [Gri82]. Noncommutative [HS88a, HR83b]. Nonconcavity [Mon80]. nonconstant [Fra86, SSD82]. nonconvex [Cho88a, SG88]. noncorrelated [Sta89]. nondegenerate [BBF86]. nondiagonal [Con88]. nondifferentiable [Jau86, Okn88]. Nonequilibrium [Bal87, BSTK87, DT89a, DBW81, Dic89, EC81, HNMK80, HL82, KLS84, KSB85, Leb86a, MLSK85, MV88b, Opp82, VM86, VM87, VM88, Bon87, CT80, CWD82, CWD83, Der87a, Der87b, Eva89, FLM88, Fri84, GLM87, GP81, Gra81a, HMH84, HK86, KH87, LH85, MTG87, MK87, Nak81, Oji89, Opp88a, Sch89b, VtLZ89, Var81, WL88, Wes88, ZWLV88, Col86a, Wes84a]. Nonequilibriurn [MV87]. Nonergodicity [MS87, DD81]. nonexpanding [GBP88a]. Nonexponential [RNRT83]. nonextensivity [DDL87]. nongradient [Wic89]. nonhomogeneous [AV81]. nonhyperbolic [GBP88b]. noninteger [BB82, GG83, WHM89]. nonintegrable [Hel88]. noninteracting [Bau83, DV87b]. Nonintersecting [PW86]. Nonisotropic [BG86a, HN82]. Nonlinear [AC84, AM83, Cor85, Hoo86, Lin89b, Meu85, Spi85, SW88, Suz87, Wes83a, deV85, vHGHK88a, vHGHK88b, AN87, BMV82, Bar88b, BTZ85, BGWY85, Blu89, BCM87, BCM88, BDE89, BLS82, BT89b, CP87, CIPS88, Chu85b, CD89c, DS86b, DR83, ER81, EW88, Ern84, FFG^{+}88, FSW86, GH82, Git89, GM88b, GM96, GHT80, GP89b, HP81, HN82, HD87, Ito84b, Iye86, LRS88, LRS89, LPRV87a, MCDK85, MT82, NZ86, NC85, OHI88, PC86, Pol88a, Pol89b, Ram85, RL86, Ram86, Saw84, Sch89a, Str88, TC84, TM89, VMA83, VMB82, WRL83, Wie85, YKS81, ZR87, vK81c, KY85]. nonlocal [FFC83, Fer81]. Nonperturbative [LSM89]. nonphase [Pol84]. nonphysical [RHM86]. nonpolar [CF88b]. nonpotential [CCR82, Gar83]. nonrecurrence [KB86]. Nonstationary [Gid85, Gri88b]. nonuniform [Per86, Per88, RV81, Wat80]. Nonunique [EHH82]. nonzero [Wu85]. normal [DH86b, Maj87, vBD80a]. normality [KS81, Sta88a]. norms [vEF89]. Note [EC86, SF88, Wid88a, An88, DG83, Gee82, Leb88a, Leb88b, Lév86, Pom85, PS85, Wag83]. nuclear [Neg86]. Nucleation [BM88, HCKS84, HW82, Hee82, NC85, PB83, SKO87, SMK88]. nucleotide [Kre80]. number [BS89c, Cia85, DF87, DH86a, DH86b, For89b, Gee82, Kai82, LG87, LD87, Lan84c, MO83, Nis81, PGS87, Sar84, WS82, dH84a]. numbers [AA88, CF80, FHT88, LH88b]. Numerical [BL85, Fox89, GMLC87, GP86b, GSH88, Lin85, ORD86, vVE87, De 85b, FRY87, Haa89, ST81b, ST84a, Sen88, ZBKS89, WC81].

O [BWdF88]. objections [Van88]. objects [Hol88, WBKS83]. observable [Hio83]. Observables [Nak86]. obstacles [FOT85]. obtained [Lav82b]. occupancy [BDKW88, MV82]. occupation [LH88b]. occurrence [LS87, vDE86]. occurring [Cha86]. October [Ano88o]. odd [CF80]. off [Bov89]. off-diagonal [Bov89]. Oguchi [Siu85]. One [BZ87, BDS83, Bur87, DPST86, DPST88, ERH85, kM83, Per82, Red81, SS86, Tej87, Zol87, Agm86, AF81, ACCN88, Ala87b, ADM83, ADG88, AV81, ABG^{+}89, Ban87, BG82, BS83, BDL81, Bis82, BW88a, BPK82, BOP87, BK88b, Bov89, Boy88, Bra88, CLWH82, Cal83a, CCO83, CkM82, CFG80, CFS83, Cho88a, Cia85, CJB88, DF85, DL82, DLS85, DL89, Der83, DFP86, DB80, Dic85, Dom89, EP88, EF87, FF86, Fel86, FH87, FJS83, Fri85, Fri87, GMO89, Gar87, GD85, GP86a, GM88b, GM96, GS84b, HW85, HMV84, Hög88, IKY88, Jan84, Jan85, JLM85, JMM87, JLM95, Jen86b, Jon86, KC85, KD84b, KLPS83a, KLPS83b, Kle86, KY88a, Kur83b, KvB85, LPT87, LW83, LMR88, LP86, Mac83a, MP82, Mar81, MM83, MO82, MSW82, Mür87, Nau85, OP89]. one [PMBB^{+}84, PPK84, PL87, PTB^{+}89, Pol89a, PS85, PW88a, PW88b, Rad84, RR82, Rez81, RW84, RT81, Ros84, Ros87, Rub84, RG86b, Sch88a, SKA86, SS88b, SKO87, SL80, Sha88, Sim81a, Sin82, Sin87, Sin83, ST84b, SWYM87, TG88, TCB89, VB86, VD89b, Vie85, VB89, Wei81, Zwa82, vBS83, vEvH85, vE87b]. one- [BPK82, DLS85]. one-component [ADM83, CLWH82, CFG80, CFS83, CJB88, FJS83, HW85, HMV84, Jan84, Jan85, JLM85, JMM87, JLM95, LW83, Mac83a, PL87, Ros84, SKA86, SS88b, Sin83, Vie85, VB89]. One-dimensional [BZ87, BDS83, Bur87, DPST86, DPST88, ERH85, kM83, Per82, Red81, SS86, Tej87, Zol87, Agm86, AF81, ACCN88, ADG88, AV81, ABG^{+}89, Ban87, BG82, BS83, BDL81, Bis82, BW88a, BOP87, BK88b, Bov89, Boy88, Bra88, Cal83a, CCO83, CkM82, Cho88a, Cia85, DF85, DL82, DL89, Der83, DFP86, DB80, Dic85, Dom89, EP88, EF87, Fel86, FH87, Fri85, Fri87, GMO89, Gar87, GD85, GP86a, GS84b, Hög88, IKY88, Jen86b, Jon86, KC85, KD84b, Kle86, KY88a, Kur83b, KvB85, LPT87, LMR88, LP86, MP82, Mar81, MM83, MO82, MSW82, Mür87, OP89, PMBB^{+}84, PTB^{+}89, Pol89a, PS85, PW88a, PW88b, RR82, Rez81, RT81, Rub84, Sch88a, SKO87, Sha88, Sim81a, Sin82, Sin87, ST84b, SWYM87, TG88, TCB89, VB86, VD89b, Wei81, Zwa82, vBS83]. one-dimensional [vEvH85, vE87b]. one-site [GM88b, GM96]. one-state [RG86b]. Onsager [CF80, GVV89, HT87b, HT88, Shl85, Tit88]. onset [AMC89, Lib84, SYOM84]. onto [Sch86a]. Open [AAF89, FGP80, AM83, GG82, Len88, Nak81, Pol88b]. operator [AV81, Dör85a, FFG^{+}88, FOT85, FHT88, Gol85, MR88a, McL82b, Rez81, Sin87]. operators [CP82, CD89b, CdS87a, CdS87b, CFS84, Del85, DE89b, KP85, Mai87, Mez87, SH82, Sza89, Zag82]. Optical [Kod85, HNMK80, LTL87, Sch89a]. optics [Gri88a, Wes87]. Optimal [BSZ89, FM87, Ito84b, ER81]. Optimization [Kir84a, RG86a, Ebe86, LG87, Mas89, SW88, dO89]. optimum [KR85]. Orbach [PSZ87]. orbits [DST83, De 85a, GN83, GW87a, GS84a, Rie85, SV80]. Order [Git88, MK87, Ruj82a, Ruj84, WSL83, AJ81, Bau80, Bre87, DP87, Doi89, FJ86, GKK87, Gro81d, KT85, KLS88a, LSSW80b, MT83a, MT83b, Mat84, MBG82, MH81, NR82, NKO^{+}89, Par80, PLM^{+}81, PLM^{+}84, PS82c, PF83, RT81, RS87b, RHG89, Ruj82b, Tar87, VM87, VM88]. order-by-order [Tar87]. ordered [Kad83, SB84, Sar84]. ordering [BWdF88, KS85]. organization [HD83, dH84b]. organized [TB88]. Orientational [LW85, LPT87, LvB89b]. oriented [DST89, DT89b]. origin [CS83, Dei89]. originating [DGL83]. Ornstein [Blu80, CW81, WC81]. Orsay [JKHV85]. oscillate [Kir89]. oscillating [POD87, ZR87]. oscillations [GW87a, GR89, KY85, SH89b, Spi85, Wes86b, YKS81]. oscillator [BLS82, FLO88, FRY87, FGMZ85, KWL87a, KN87, LSSW80b, Maz81, Spi85]. Oscillators [ACT82, BCM87, BCM88, EW88, GRS80, SHKM86, Str87, Tél88]. Oscillatory [Ste89, PL87, Vid87]. Ostwald [Voo85]. other [Dav87, LW86, Man84c, Red81]. Outlines [HL82]. overdamped [BBLT81]. Overhauser [DP89]. Overlap [Kra82b]. overlapping [Fal87]. Overview [Wei86e, Bau87, Dom88, Lom85]. oxidation [MV86]. Oxygen [BWdF88]. Oxygen-vacancy [BWdF88].

packed [Kho87]. Padé [DLT88, Gib81]. Pair [Zha88, Cia85, GP83, HV85, LvB89b, LPdS84, YNS86]. pair-composite [GP83]. pair-potentials [LPdS84]. pairing [KP86]. paper [PS85, Wag83, dH85a]. parabola [TG81]. paradox [BS81, Lan84a, WW82]. Parallel [Ano89o, HH85, Ano88o, Chr86, Stö88]. parameter [AO87, GKN84]. parameters [GP85a, GG83]. parametric [Spi85]. parametrically [McL81]. parametrization [RT86]. parastatistics [dSVT87]. parking [Hem89]. part [SC85b]. Partial [GMAK84, dCD87, Bro82, Ito84b, SLW84]. Partial-dimensional [GMAK84]. partially [Wei86d, WF88]. participants [Ano84p, Ano86p]. Particle [AV03, BW84, AV87, BG80a, BG80b, BS84, BH88b, CPPV89, CDPP89, DFL86, DFL87, DK86b, DV87b, FFC82b, Fri84, FvdM89, Gro83b, Hug85, Ker88b, KST89, ME82, Mag87, MD83, MK84, MK85b, MK83b, Neg86, NLSV89, NR88, Pia84, Pol84, Req83, Sch87c, SG87, Wag89, Wer88, Wic85, Wic89, vEF89, KST89, ORD86]. Particles [BCD89, Alb89b, AAF89, BK88a, CC82, Cia85, CR81, DPV86, EP88, Heu86, KD82a, KLPS83a, KLPS83b, MS82a, NOR89, Nos84, Pul82, SSD82, Skr84, SG81, Spo87, WMvBL88]. particulates [JMF88]. Partition [Gro83b, DD88, Ken82, Lév86, Mon83]. partitioning [LG87]. partitions [FZ85, NNN89]. parts [Gib81]. passage [BDKW88, CCD86, FHW88, KMST86b, LSSW80a, ME87, MW88, THFT87, Wei81, WSL83, Wei84b, WDG86, WHM89]. Passive [FKP88, KK88]. past [Rub86]. Path [Ber86, TWB86, FCD86, Gre82a, Gro83c, Gro84, SIK86, WKUK88, bA88a]. Path-integral [TWB86, bA88a]. pathways [MJ83]. pathwise [CGOV84]. Pattern [Ebe86, WKL89, Jen86b, KMD89, WC89, YAR89]. pattern-recognizing [Jen86b]. Patterns [Har84, Cho88b, Fif85, LHFW88, Mad88, VV89, Vid87]. Pearson [Hio83]. pedestrian [Kad85]. Peierls [BD85]. pendulum [McL81]. Penrose [AO87, GLO86, Kat85]. Penrose-type [Kat85]. percolating [Heu86, LMJ87]. Percolation [AA88, Aha84, ABL87, BLM87, FH85, HBAM84, KD84a, LB87, SS82a, AN84, BCSY84, BFL82, CO84, CG81b, Dom89, DST89, DT89b, GMAK84, GM84, HR88, HKLW87, HHB87, KGA84, Kes81, KZ87, KL82, Kra88b, Lam88, Lam89, NH88, NS81, New86, New87, Ngu87a, Ngu87b, Ngu88, NM84, NHH88, PSMZ84, PSZ87, RK88a, RG86b, Sha82a, WB87, Web84, ZBS86, vE87a, vVE87]. Percus [Var81]. perfect [DMMSR88, PPK84, Vie85]. performance [BSZ89]. Period [CEK81, McL81, GM88b, GM96, mMH87]. Period-doubling [McL81]. Periodic [BLP81b, NL87, SHKM86, Ala87a, Ala87b, ACH88, Alb89b, BL85, BLS82, CJB88, Cos80, DST83, De 85a, Der83, DH85b, For86a, FGN88, GGL88, HMH84, HK86, Iye86, KS86, Mez87, MHB87, RS85b, SS84b, TG81, TG87, VtLZ89, GM85]. periodically [DGP84]. perpendicular [FE87]. Perron [Dör85a]. persistence [CV87]. Persistent [ST84b]. perspectives [Dre87]. Perturbation [Bov89, Bre87, Chu85b, GKT84, GGL88, KMST86b, MPPT78, MPPT80, Wer84b]. perturbations [BT89a, Boy87, LFtH88, LOS88, MOS89]. Perturbative [Enz81]. perturbed [BJLS89, Ito84a, MNT85]. perturber [HR83a]. Petersburg [PMBB^{+}84]. Phase [Bar89, BCSY84, BS89c, Cho88a, FILS80, GS88, Huc83, Kir84b, KY88a, KP86, KU86, Kor86, Kur83a, LPS85, Mag87, MTA85, Pen86, PP83, PW81, Pfi82, Pfi87, Ram89, Süt80, SGL89, WL88, Ade86, ABF87, AG85, BD88, BT85, Bon87, BN87, BF86, BLP81a, BKL83, Cal83b, DP87, DPT80, DBW81, DP89, EE88, GN85, GG82, GKK87, GLO86, GKT^{+}86, Gro81d, Gun84, Haa89, HHB87, Hög88, Kam86, Kat85, KT85, Kra88b, KN87, MP89, MLSK85, MV87, MV88b, NR82, PF85, Par80, PLM^{+}81, PLM^{+}84, PS82c, RS87a, Req82, Req83, RW86, Req88, RW84, SKA86, Sch85d, Sin82, SWYM87, Tar87, Tél83, Tor86b, Uim84, VM86, VM87, VM88, VA81, Wei87a, Wic85, Wid88a, Wid88b, Zah87, dS88a, vEvH85, vHZ88, BC89b, Pen87]. phase-separated [Wid88a, Wid88b]. phase-separating [Req83]. phases [DdGM89, FT88, Nar82, NR82]. Phenomena [HL82, bA88b, Bal87, Bar89, Bin81, DZ87, Fam84, FCK83, GKN84, Kes88, MS83b, Opp86a, Par87b, RNRT83, RFH88, Sen86, Suz88, TB88, Wei87b, Wes83a, Opp88a]. Phenomenological [BW89, VA81]. phenomenon [BPK82]. philosopher [Fis88]. Phonon [LP86, Gro89a, Gro89b, Gro89c, JW89]. phonon-modulated [JW89]. photocounts [CLRT83]. Photon [Nie88, Nos86b]. Phys [AV03]. Physical [Dor83, Bar85, Ben80, BGM84, Wes84a, Oji89]. physicist [Fis88, WTB89]. Physics [Ano81o, Han83a, Fis88, FOU86, Fox87, GBD^{+}86, HBB^{+}84, Leb86a, Lin89b, Man84b, Omn89, Rap89, RW86, RBHL86, Shl85, Wei83, Wei84a, Wei86c, Wes83a, Wes85a, Wes85b, Wes88, bA88a, dST86, Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano82l, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l, Ano83m, Ano83n, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano85b, Ano85c, Ano85d, Ano85e]. Physics [Ano85f, Ano85g, Ano85h, Ano85i, Ano85j, Ano85k, Ano85l, Ano85m, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano86m, Ano86b, Ano87h, Ano87c, Ano87d, Ano87i, Ano87e, Ano87f, Ano87j, Ano87k, Ano87l, Ano87m, Ano87n, Ano87g, Ano88e, Ano88c, Ano88d, Ano88f, Ano88g, Ano88h, Ano88i, Ano88j, Ano88k, Ano88l, Ano88m, Ano88n, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k, Ano89l, Ano89m, Ano89n, Ket84]. picture [dFSZ87]. Piecewise [ER81, Boy87, DH87, Rue81, Tél83]. Pierre [Leb81, PD81]. piston [Dri81, Dri82, MS81a]. pitchfork [SH89a]. pivot [MS88a]. planar [Cor87, HC88, KIM86]. Planck [BP83, BT81b, BT81a, Eng88, FFGM88, FFG^{+}88, Fed88b, FFdlM87, Gra81a, GT84a, GT84b, HR85, Jau85, MGL87, MK83b, MS83a, RV85, RSRD83, TC84, Val83, VR85]. plane [BWdF88, FT81, GPS86, GL89b, Jan82a, Jan82b, KIM86, Pic83, Pic84, WBKS83, ZT85]. plasma [ADM83, BS86, CLWH82, CFS83, CJB88, FJS83, For84, HW85, HMV84, Jan84, Jan85, JLM85, JMM87, LW83, LM86, Mac83a, Meu85, PL87, Ros84, SS88b, Sin83, SSM85, Vie85, VB89, Wat80, Wes85b, JLM95]. plasmas [Bal87, GDx82]. plus [KS86, Mez87]. PMC [Mar87]. Podolsky [Omn88b]. Point [FF87, Abr85, AG85, Coo83, DDdG83, FTZ84, Fri85, Gib81, GDx82, Gra82a, Mar81, Meu84, PS85, SYOM84]. points [BW89, Fei88, GS86, Han83b, Tél88, Vis81, WS82, dHK85]. poisons [Pan84]. Poisson [HGPRS89, Pel88, Pel89, Pom81, RPSS85, ZS80]. Pol [Spi85]. polar [SIK86, YS85]. polarizable [CF88b, FFC82a, HHO86]. polarization [Ram84b, YS85]. Pólya [WZ89]. polyacetylene [Par87a]. polyatomic [CL81, Har84]. polycrystalline [FM87]. polydisperse [XS89]. polymer [Dup89, FK83, Guj82, Gut84, HR82, KW88, KS85, Nag85, Pri86, Spo86b, Sta84, Wid88a, Wid88b, bA88a]. polymerization [Spo83, SW87, Wer86b, Zif80, vDE84]. Polymers [CD89c, DS88b, Spo85b, Ben84, CDM86, DC89, IS88, Mut83, Sal86, Zha89b]. Polynomially [DS86b]. pools [Kre80]. Poor [RW86, Req82]. population [Dem83, Wei82]. populations [BCM87, BCM88]. porous [LHFW88, Sta84]. position [FF87, SS80]. Positive [For86b, BN88, Cor88, LPdS84, SSM85]. positivity [FLM88, FILS80, OHI88, PW81]. possibility [Req83]. Possible [Tsa88]. Posters [Ano84r]. potential [Agm86, AJ84, Ala87a, Ala87b, ACH88, Alb89b, CCRG80, CCR81, FH87, GGL88, GB84, Gou86, Gro81c, IVMA83, LM84, LLS88, Lop88, PC87, PA88, RR82, RV85, RV81, Ryt87, Sin87, VR85, WJS84, WMvBL88]. potentials [Bar88a, Bau80, Cho88a, DT89a, FHT88, HZ82, Jau86, KS86, KLS89, LPdS84, Mez87, Ros84, SS84b, Spe88, Süt80]. Potsdam [Ano88o]. Potts [ACCN88, Bax88, Bax89, Bin81, BL87, DP87, DMMSR88, Dot84, Dup87, EW82, HBB88, HC88, MS84, PW86, PT86, RHG89, RG86b, Rui89, Sch88c, Sel89, TCB89, Wu80, Wu88]. Power [Mac83b, MR88b, Rav89, YAR89, ZWLV88, Bra89, Kes81, SMYO83, YMS83]. Powerlike [CG81a]. powers [HS88b]. Precise [NHH88]. Precision [ZBS86]. precursors [Wie85]. prediction [Wes84a]. Preface [Ano84s, DKS89, GL88, Gub86, Leb87, LN87, Leb88c, Mei88b, Wal84, Leb85a, Leb85b, LP89, vBEL88]. preferred [Pan84]. Preliminaries [CGVG83]. prescribed [Ala89, CS83]. presence [BHR^{+}88, DF85, Fri86, MTA85, PPK84, SS84a, Str82, YKS81]. Presentation [Fei88]. preserving [FKP88]. pressure [Bar88a, BGN89, BPTV88, Mei83, MF83, Vai89, WW86b]. Prigogine [MT87, Pet87, dlL82]. principle [BCD^{+}86, DFGW89, Fer88, Had81, ST81a, Saw84]. Principles [Wei89, KB85]. priori [Fri87]. probabilistic [GL89a, SYG83]. probabilities [KY88b, Par89, ZS82]. Probability [CLRT83, HKLW87, Ala88, Ben84, Bon87, DMM^{+}89, GM88b, GM96, Luc86a, LBF83, Sta87a, Sta89, Ull81, WKUK88, WHB85]. Probes [TM89]. problem [AKvL88, ABKS89, ADG88, BM89, BFA86, BDKW88, CF89, FFdlM87, Hem89, Hug85, IT88, LG87, LO85, Luc86b, Luc87, Mac83b, MST84, MSW82, PMBB^{+}84, Pol88b, Pom81, RG86a, ST89a, TW85, VA81, WHM89, ZL81]. Problems [DDL87, HS82a, CP87, CIPS88, CO84, FHW88, FLNO88, GBD^{+}86, LG87, MW88, PC86, Sch88b, Spo86a, Wei81, Wes84a, WT89b, dO89]. procedure [MS83a]. procedures [DK86a, Suz87]. process [BCFN89, CH88a, Cha82, DM81, DIP82, DFIP82, DPV86, GMO89, Gat88, Go83, MM86, MM88, Sch85b, Sch86b, Shl84, Van83b, Wic89, WE82]. processes [BH88a, BGWY85, BGM86, CS89, Cha86, Che87, DF84, DFGW89, DL89, DGL83, EEEF84, FF87, GHT80, HMS83, KK84, KDMS89, KMST86a, KMST86b, Kop86, KHM^{+}83, KEKS83, LHFW88, LS82a, Le 84, MT84, Mar87, MW88, MK89, MC81, Ngu87a, Ngu87b, Ngu88, NG87, Ram85, SKO87, SL80, Spo85b, Suz87, SLW84, VKK85, Wei82, Wei83, WSL83, Opp88b]. processing [MHV89, vHGHK88b]. processor [FOU86, Sug86, SIvL^{+}88]. produce [Doi89]. Product [Rav89]. production [KMD89, Len88, OHI88, Oji89]. products [CPV88]. profile [BP85a, Cia85, Tit84]. Program [Ano80o, Ano80n, Ano81o, Ano81p, Ano82q, Ano82n, Ano82o, Ano82p, Ano83r, Ano83s, Ano83t, Ano84u, Ano84v, Ano84t, Ano85q, Ano85r, Ano87o, Ano88o, Leb86b, Leb86c, Leb88d, Leb89b, Leb89c, Leb89d, Ano84w, DZVS84, Red82]. Programs [Ano87p]. Progress [Kog86]. projection [FFGM88, FFG^{+}88]. Projectors [Omn89]. Proof [DS84, Nag85, vE87a, EW87, MG83, Sok81b]. propagating [FC89, Ker86, LCG85, LCG86]. Propagation [BT89b, Dee85, KN85, RHG89, Sch85d, Web84, Bra88, DCW85, FF86, Fre87, Gla85, Ker88b, Kos86, deV85]. propagators [Gro81c]. Properties [GS84b, HKW^{+}85, HEH89, Lee83, WHM89, AB83, ABG^{+}89, AGL88, Aur87a, Aur87b, BMV82, BCT86, Bax89, BG84, Ber85, Blo81, BPK82, Bor87, BL85, CPRV89, DL82, FH87, GH82, GLM80, HMOV87, Han83b, HW85, Hio83, HT88, Jan84, Jen86a, Jen86b, JJ86, Jon86, JE89, Kad86, KCH88, Kes81, KW88, KI83, LMJ87, LPT87, tL88, LW83, May85, MT82, NT89, NH88, O’C84a, OB84, OMS86, PG83, Rav88, RS85a, RS85b, RRO88, Row86, SS86, Sto86, Sza89, Tak89, Tho82a, Tor86b, TGN87, Tru86, VMB82, dSVT87, WWH89, YHOF85, Zab88, Zag82, vBS83, vVE87]. property [Zeg86]. Protein [Go83]. proteins [MJ83]. proton [STA88b]. Pseudo [SH89b]. Pseudo-regular [SH89b]. pseudobimolecular [GG82]. pulse [HS88b, HJ87]. pure [BF88]. Purely [CD89b, Del85, vB84, vB85]. purpose [HBB^{+}84].

QCD [Sha86, dF86a]. quadratic [BGH84, DF87, GDx82, Luc86a, Luc86b, Luc87, Tou89]. quadratically [Sar84]. Qualitative [Ala87b]. quantal [CH88b]. Quantitative [Kir84a, AKvL88]. quantity [Sch87a, Sch87b]. Quantization [Hel88]. quantized [HHO86]. Quantum [Ano80o, Ben82, CDM86, DZ87, HS88b, Igl89, Kun86, LS89c, MY85, Par85, RSK86, Row85, SD85, STA88b, Suz86a, AM83, APSJ85, BR80, Ben80, BCD89, Cha86, CCR85, Che87, cC87, DB89, FK87, FP80, GM87, Gib81, GVV89, Gol87, GKT^{+}86, Gri84, Gri88a, GOSS86, HW86, Hio83, Ima86, Jan85, KF85, KLT88, KD83, Kna88, Kog84, Kos86, LO85, Len88, LN85, Loh86, LS89b, LOS88, Maa84, McL82a, McL82b, MPK88, Nak86, NTW88, NT89, NLSV89, Nis81, OHI88, Oji89, Omn88a, Omn88b, Omn88c, Omn89, OMS86, Par82, Pel88, Pel89, PW81, PS82b, RBHL86, Rub82, Rue86b, Sca86, Sch89a, Sch86a, ST81b, SG87, Sch86c, SLM89, Sto86, Str87, SS83, Suz86b, WO86, Wre83, ZT85, ’t 88, vHBW84, Ruj82a]. Quantum-statistical [LS89c]. quasi [BG86b, DLS85, FGP80, Nos84, Pel88, Pel89]. quasi-elastic [Nos84]. quasi-energies [Pel88, Pel89]. quasi-free [FGP80]. quasi-integrable [BG86b]. quasi-one-dimensional [DLS85]. quasiaverage [AZ85]. Quasicrystal [AF86b]. quasicrystalline [Bur87]. quasicrystals [Bur88, LP86]. quasilattice [LB87]. quasiperiodic [AGL88, EH89, GG88, Hio83, LOS88, Ser86, Sin87, Spe88]. Quasiperiodicity [GL89b]. Quasiprobabilities [HW88]. Quasistationary [PP87, Pol88b]. quenched [FH87, FLNO88, NE85, VTH86]. quenched-state [FH87]. quenches [DV87a]. quenching [Sch88a].

radial [WW86a]. radiant [ACS89, BG82, CGVG83]. radiation [Ser86, VD89b]. radical [EN85, Pan84]. Radin [Wag83]. radiophysics [Wei89]. radius [LT86]. Raman [DCW85, WCD85]. Ramanujan [ABF84, Bax81, FB85b]. Ramanujan-type [ABF84]. ramified [HBAM84, VV89]. Random [AS89, Bar85, BEF86, CV87, CS83, EvV86, Ern88, EN85, Hal83, HGPRS89, HS82b, Imr84, KdH83, KI83, Lec85, MS88b, Nau85, Pap83, Ram84a, Rav89, Ric83, SKW82, SN84, ST81c, WC89, WS83, Wes82, bARC89, dH84a, dHK85, dH89b, Agm86, Ala89, AKN84, AV81, AAK84, ABG^{+}89, AGL88, BR86, BZ87, BDKW88, BS83, Ber85, BCSY84, BZ83, BC89a, BK88c, BLS82, BHR^{+}88, BLP89, CCP87, CK87a, CkM82, CEGM84, CFS84, CD89c, CPV88, DPV86, DFGW89, Del85, DZVS84, DS86b, DH85b, Dom83, Dom89, EEEF84, EMDvB84, Ern87, EH84, Far87, FFS84, FH87, For89b, GMP89, GR87, GN87, Gro81b, Gro81a, Gro81c, Gun84, Han83b, HZ86, Hem89, Hey82, HWW82, HFJ86, Hio83, HMS82, IS88, Iye86, Jan87, JM86, KK84, Kes88, Key82, KB85]. random [KS86, KBZ84, KHNA83, KLS88b, Law88, LMN84, LSSW80a, LN88, Mac83c, MEvBD84, MG88, MV88a, Mai87, MOS89, MPK88, Mez87, MH81, MV82, Nea80, Nie88, NL85, NLC^{+}86, NL87, NG87, PMBB^{+}84, Pan84, PSMZ84, Par89, PP83, PWvH84, Pic83, Pic84, PGS87, Ram84b, Rav88, Rez81, RS85a, RS85b, Rub86, SW85, Sch85c, SVBO88, SS84b, ST89b, Spe88, Spi85, Spo85b, Sta87a, Sta88a, Sta89, Sta87b, SLW84, ST84b, Tas87, Tas89, TF86, Tor86a, VH89, WR80, Wei81, WS82, Wei84b, WH84, WHB85, WDG86, WW88a, WWH89, Whi89, Zol87, Zwa83, dAC87, dFKP86, dFKP87, dlSLW88, vEvH83, vEvH85]. random-field [Jan87]. random-matrix [MPK88]. random-number [PGS87]. random-random [ABG^{+}89]. Random-walk [SN84, ST89b]. random-walking [AS89]. randomly [Geo81, Ito84a, Sta84, VD89b, dH84a]. randomness [GL89b, GMC81, NB87]. range [AM85, Bau80, BP83, BZ83, CCO83, CFG80, DF84, FHT88, FILS80, Gra86, GLM80, HMV84, Ito82, Ker88a, KP86, MBG82, NKD84, NKO^{+}89, Pen86, Pen87, Pia81b, Pic83, Pic84, RK88a, RT81, SS85, TCB89, Tor86a, VB89, Zeg87, vEvH83, vEvH85, vE85, vE87b]. rare [LS87]. rarefied [Ban87]. Rate [Kop86, ACS89, EK85, Fai81, Gat88, KDMS89, KLSZ86, MST84, Sim81b]. Rates [TG87, dlL82, GHKM89a, GHKM89b, Hyn86, Rei88b, THFT87, VM86, Wei86e, vBK87]. ratio [Lan84c]. rational [Lop88, MS88b]. ratios [CR86]. Rayleigh [Dri81, Dri82, EH89, MS81a, ST87, VA81, vBC88]. Reacting [AKN84, Uey80a, Vid87]. Reaction [DFL86, DFL87, BM84b, BM88, BCD^{+}86, Bur86, EEEF84, Ebe86, Els84, EK85, GHKM89a, GHKM89b, Hyn86, Ker88b, KBZ84, KHM^{+}83, NOR89, NMM80, Pol88b, RDRS87, SBS88, TG87, Wei86e, ZR87, vD88, vD89a]. Reaction-diffusion [DFL86, DFL87, NOR89, NMM80, vD88]. reaction-limited [vD89a]. reactions [BA87, Blo81, BGM86, Cuk83, Cuk86, Jan89a, Jan89b, Jan89c, MY85, POD87, Pan84, PP87, WKL89, vK87b]. reactive [MC81]. reagent [Fif85]. real [BGH84, Han83b, KSB85, MO82, YHM80, Ber83]. real-space [YHM80, Ber83]. realistically [d’E89a]. rearrangement [Bro81]. reciprocity [Tit88]. recognition [HH85]. recognizing [Jen86b]. recombination [EN85]. reconsidered [RG86b]. Rectangular [vBK87]. Recurrence [DGP84, Man84a]. recursion [SH82]. red [Row85, WP82]. reduce [Pea85]. Reduced [Kra88a]. reducibility [Mor84]. reducible [Jan89b]. Reduction [WB99, BW88b, Mon88]. refined [MM84]. Reflection [KB85, Fel86, FILS80, PW81]. Reformulation [WKUK88, Omn88a, Omn88b, Omn88c, Omn89]. regime [EH89, LO85, Req82]. regimes [SHKM86]. region [Gro81d, Ima86, Wat80, WS88]. regions [FOT85]. Regression [Del84a]. Regular [Wes84b, HM84, MR86, MH81, SH89b]. Regularity [KLS89]. related [DFV89, PT86, dSEC81]. Relation [PS89, Bax82, BA86, HZ86, KF85, Nau87, Rod89, ST81b, Van83b, VKK85, dCD87]. Relations [Sha82a, dFSZ87, Dav87, DD85, DdGM89, GVV89, HT87b, Mor84, Rei80b, SH82, Sha88, Str88, Tit88, ’t 88]. relationship [BS87b, Kei83, RV81]. relative [Don87]. relativistic [Dud89, KLK^{+}85]. Relaxation [DM87, DC89, Luc86b, Luc87, RRO88, Ben84, BS87b, CS89, Der87a, DB80, GRS80, GR89, Gzy82a, Ken83, LAA87, Mac87, MR87, MR88a, MPK88, MB84, RNRT83, Shl84, Vig82, VKK85]. relaxing [Shi81]. relevance [RW86]. relevant [FvdM89]. remark [CF88a, DF84, HT80, PS82a, vEF89]. Remarks [ADM83, Fri84, PG83]. renormalizable [BMV85]. Renormalization [CL89, ED81, Fam88, FK83, Kad86, LS84, Mac83c, MGJ89, VT89, Vis85, dOB89, ADFJ89, AMV85, AO87, BK81, BB80, BS87a, DGK81, DS80, Del84a, FO89, GWU86, JMS83, Kle89, Lar86, LSM89, MO82, Mis85, MS86b, Mut83, RGJ85, SF88, SD81, Swe84, TS81, YHM80, Ber83]. Renormalization-group [Vis85, DGK81]. Renormalized [CF88c]. Rényi [Les82]. repeated [CK88a, MK83a, MK83b]. repeated-ring [MK83a]. replacement [KLW86, KWL87b]. replica [CFS84, PWvH84]. Representation [Gzy82b, Ala89, Ker88b, Kra82b, WRL83]. representations [Ram89]. represented [Ben80]. repulsion [LM87]. repulsive [dS86a]. Résibois [PD81, Leb81, Pet87]. Residual [Sch88a, GS88, Nak81, SGL89]. resistance [FF86, SN84]. Resistivity [LH88a, PMBB^{+}84]. resistor [Blu89, DZVS84, NE85, dAC87]. resolution [BS81]. resolvent [PS82b]. resonances [BG86b, Rue86a]. Response [BL90, Kei84, LV80, GDx82, KC85, LOS88, Muk82, OHI88, Pan89, Van88]. Restoration [PS88d]. restricted [CW81, KY88b, PS82a, Gro81b]. Results [Mon88, ABCT86, ABG^{+}89, BN87, CJB88, CS83, FdF86a, FP87, For86b, For89a, GR87, GP86b, hGH86, Jan84, KH82, KP89b, Leb89a, MV82, PS88d, Red81, Req88, Ruj84, SK82, SPA88, WC81, Wu80, ZS88]. Resummation [LM86]. retrovirus [Pan89]. Return [HR86, Maa84, Maj89, Fei87a]. reunions [FG88, FG89]. reversals [Ito88]. Reversible [EK85, DFGW89, Tak89, vDE84]. review [Agm88, Bak82, Bar85, Bar89, Ben85a, Ber83, BN87, Bru88, CS86a, Cha84, Col86a, Col86b, Cuk82, Dar82, Dom87, Dor83, Dor85b, DZ87, Fam88, Fis88, Fox83, Git88, Git89, Gli83, Gre82a, Gre82b, Gre85, Gre86, Gre89, Gri88a, Kes88, KY85, Leb86a, Leb89a, Lec85, Lee83, Lin84, LL86, Lin88, Lin89b, Mas89, Nos84, Nos86b, Nos86a, Opp80, OP81, OI82, Opp82, Opp86a, Opp86b, Opp88a, Opp88b, Opp89, Pap83, Par87b, Per84, Rap87, Rap88, Rap89, Red89, Rei82, Rei84, Rei88a, Sen82, Sen86, Shl85, Shl88, Sil81, Sil82, Spo85a, Spo86a, Ste87, Wei82, Wei83, Wei84a, Wei85a, Wei86a, Wei86b, Wei86c, WW87, Wei87b, Wei89, WC89, Wes83a, Wes83b, Wes84a, Wes84b, Wes85a, Wes85b, Wes86a, Wes86b, WO86, Wes87, Wes88, Wes89, Wol82, Zwa81, bA88a, bA88b]. review [dST86, dH89b, vK81a]. Reviews [AM87a, MBAWP87, WP87, BL84, PM82]. revised [Pia87]. revisited [HT87b]. Rigid [HKZ88, CT84]. Rigorous [DK86a, Leb89a, SS81, SS82b, Sok81b, BN87, DST83, ERH85, Har87, HT87a, KK88, KP89b, Req88, dHK85]. ring [CK88a, Dre87, MK83a, MK83b, Spo86b, YAR89]. ripening [Voo85]. RISM [MS81b]. rod [BDS83, Pia84, PS88d]. rod-to-coil [PS88d]. rods [For86a, For86b, For88, MM83, Pia83]. Rogers [ABF84, Bax81, FB85b]. role [DV87a, MT87, VKK85]. Roles [Cha86]. Rosen [Omn88b]. rotating [Sek83, Ste89, vBD82]. rotating-wave [Sek83]. rotation [DF87, Lan84c]. Rotational [Cuk82, MK85b]. Rotations [Kiv88]. rotator [CT84, DGP84, Pel88, Pic83, Pic84, Pel89]. rotators [Dun85, FGS85]. rotors [BC89a]. rough [CL88]. Roughening [BEF86, KMS87, Abr84, BFL82, Gra84]. rouleaux [WP82]. round [OB87]. rounding [BEF86]. row [GS85]. Ruelle [Mon83, Zag82]. Rule [Zab88, GV89b, PSZ87]. rules [GLM80, hGH86, Vie85, VB89]. runaway [Pia81a]. Rupture [RHHG88]. Rutgers [Ano81o].

S [dH85a]. salesman [AKvL88, BM89, BFA86, RG86a]. Salsburg [Zag82]. salts [CW81]. sandbox [WTB89]. sausage [BMS89]. Scalar [Alb89b]. scalars [FKP88]. scale [Che86, ED81, KD84a, NT81, Reb86]. scale-similar [NT81]. scaled [BG80a]. scaled-particle-like [BG80a]. Scaling [AGL88, DT89b, Fei87a, GBP88b, LR88b, May85, NNK89, PAF84, PAF85, Wid88b, bA88b, de 88, vDE88, BW89, Bra89, DHB88, DS83a, DS86c, Fei88, GK84, GS81, GGP86, Ham87, HBB88, LB87, mMH87, MHB86, MS83b, PS88d, Rav89, RGJ85, TW85, VM87, VM88, WF88]. scalling [BT88]. scan [BK88a]. scattered [CLRT83]. scatterers [FF86, FC89]. Scattering [FvdM89, Ser86, BK88a, CF88c, DCW85, Gro81b, Gro81a, Gro81c, HM88, Kiv88, Koh88, MGTM88, MK89, Nos84, Nos86b, Okn88, SS85, SSM85, VD89b, WCD85, WS84, dSEC81]. scheme [Kor86]. schemes [HT80, Pol88b]. Schlögl [BM84b]. schrödinger [CFS84, Del85, CC86, Chu85b, DF86b, DS86b, FHT88, LRS88, LRS89, Mai87, Rez81, Sin87]. Schwarzschild [Ale85]. Schwinger [WKC83]. science [Shl88, bA88a]. sciences [Bar85, Wei84a, Wei86c, Wes84a]. Scientific [And86, Ano84w, Che86]. screening [BS86, GN85, GLM80, KMD89, Vie85, Yan87]. search [BDWK87, WS88]. Second [DPST88, Gro81d, Huc83, LSSW80b, Par80, VM87, VM88]. second-neighbor [Huc83]. second-order [Gro81d, LSSW80b, Par80, VM87, VM88]. sections [CG80]. Sedimentation [GM88a]. segment [Doi89]. selected [Ano85p]. selection [BCD^{+}86, Chu85a, DML82, Wei87a]. Selectively [BT81a]. Self [BCM87, DF85, HGR87, KLPS83a, KLPS83b, KK85, Pia81b, Whi83, WMvBL88, YKS81, dFKP86, AH89, ABG^{+}89, BS85, BBAH83, CG85, Dom83, FC89, Gla88, HCM85, HD83, Kie89, KP85, Kle80, Law88, LS82b, LS82c, LCY83, Mac83a, MS87, Mad88, MS88a, Nel89a, NCF88, NM84, SD86, TB88, Tél88, WE82, Yip89, Zol87, dH84b, dFKP87, BCM88]. self-averaging [ABG^{+}89]. Self-avoiding [Whi83, dFKP86, AH89, BS85, BBAH83, CG85, Dom83, Gla88, Kle80, Law88, MS87, Mad88, MS88a, NCF88, SD86, dFKP87]. Self-consistent [HGR87, KK85, Mac83a, NM84, Yip89]. self-correlation [LCY83]. Self-diffusion [DF85, KLPS83a, KLPS83b, Pia81b, WMvBL88, LS82b, LS82c, WE82]. self-energy [FC89]. self-gravitating [Kie89]. self-interaction [Zol87]. self-organization [HD83, dH84b]. self-organized [TB88]. self-similar [HCM85]. self-similarity [Nel89a]. self-sustained [Tél88]. Self-synchronization [BCM87, YKS81, BCM88]. semi [ABCN81]. semi-infinite [ABCN81]. Semiannual [Leb89d]. Semiclassical [Koh88, Omn89]. semiconductors [KR85]. semigroup [Maj89]. semigroups [AM83, HN80, Maj87]. seminar [Ano80o]. semiperiodic [CFS83]. separated [Wid88a, Wid88b]. separating [Req83]. separation [BLP81a, HHB87, PF85, RW84, SKA86, Suz87]. separations [CF80, Kur83a]. September [JKHV85]. sequence [Fra80]. sequences [GMAK84, MT82]. sequential [HFJ86, Stö88]. Series [Dic89, Bre87, GKT84, HS88a, Pol89a, Vai89]. session [Kes88]. set [Nau87, Süt89, Wei86d]. sets [BGH84, BGM84, BPTV88, Fei87b, GV89a, Lav82a, LSSW80a, VT89]. seven [FT81, TL82]. seven-mode [FT81, TL82]. Seventh [Ano81p]. shape [AG85, MM86, MM88]. shapes [ATZ83, Zia86]. sharp [ABF87]. sharpness [Ker88a, PIP89, Rui89]. Shear [MIH86, EW81, LH85, MD83, RSRD83, Str82, SKS82, XS89, dSvLH89]. shearing [LCG85, LCG86]. sheet [KM87]. sheets [Man84c]. shell [SH82]. Sherman [Gra82b]. shield [WS88]. Shielded [Per80]. shielding [MG83]. shift [Mar87, NR82]. shifted [VtLZ89]. shifts [MT84]. Shil’nikov [ACT82]. Shlosman [RS87a]. Shock [Abd87, BCFN89, Cor88, FFdlM87]. short [Ern88, FILS80, KP86, LvB89a, VB89, dSEC81, dSC82]. short-range [FILS80, KP86, VB89]. shortening [KV82]. shot [HS88b, MW88, Van83b]. shoulder [WMvBL88]. sided [SS89]. Siegel [LT86]. similar [HCM85, NT81]. similarity [Cor86, Nel89a]. Simple [GT83, Sti80, Abd87, AF81, AM84, APSJ85, Bax89, BCFN89, BGM86, BSF89, CD89a, Fed88a, FCD86, Hio83, Huc83, KM88, KK88, KVGZ87, LH88b, MTG87, MGTM88, MKB86, NKD84, OS84, Rav88, Ros86, SM89, Vil87, Opp88a]. simplified [SW88]. simulate [BDS86]. simulated [AKvL88, Han83b, Kir84a, RG86a, Mas89]. Simulating [Kad85, Kog86]. Simulation [MLSK85, SIvL^{+}88, BK80, BCFN89, Bon85, CG85, CS86b, FRY87, FGMZ85, GG87, Her86, JS86, JE89, LCY83, Lin85, NNK89, Pan84, Pan89, PB83, PS87, PA88, Rap88, Rap89, RDRS87, SM89, SLM89, TC84, WBL89, WZK84, WK84].

Simulations [Hir86, Sca86, BW89, BDL81, CPRV89, DBK^{+}86, Fox89, FKRB88, FKRB89, GOSS86, HGR87, JS83, KTB86, KL83, KHNA83, Kop86, MK87, PLM^{+}84, RSK86, SS89, SKA86, SF89, TWB86, Zha89a, dAHK^{+}89]. simulator [SMM89]. Sinai [BL84, Lév86]. sine [BBLT81, LSM89]. sine-Gordon [BBLT81]. Single [AAK84, BH88b, KHNA83, RHM86, GOSS86, HJ87, Ull81, WHB85, WRL83]. single-atom [HJ87]. Single-lane [RHM86]. Single-vacancy [BH88b]. Singular [Boy87, NL85, Süt89, CG80, Del85, KLS89, KMST86b, Pet87, TM88]. singularities [Coo83, Mez87, Mou82, NLC^{+}86, NL87, SS83]. Singularity [MS88b, Kur83b, Pol84]. sinks [Cuk83, MC81]. Site [CO84, BF88, BSF89, GM88b, GM96, Sha82a]. Site-bond [CO84]. site-diluted [BF88, BSF89]. sites [BMS87, Cor89a, MPH87, Wer86a]. six [Cla83, DdVG89]. six-vertex [Cla83, DdVG89]. Size [IKY88, ADS80, BW89, BT88, Bra89, DHB88, Fro89, Gut84, HW82, Ham87, HBB88, HW85, Lu87, MHB86, Ngu88, Oli88, Pes85, PF83, RGJ85, Spo83, SPA88, VM87, VM88]. skeleton [HKW^{+}85]. slab [IT88]. Slawny [Mon88]. Slip [Cor89b]. slit [BK88a]. slit-scan [BK88a]. Slow [Sch88a, GG88, ME87, Rub86, SC85b]. slowing [AFV85]. slowly [Kle86]. Small [Cer88, DIP82, DFIP82, MOS89, BM84b, DH86a, DH86b, FOT85, Fri86, GP86b, GG87, Jau85, Lav87, LS82a, WT89a, Gro89c, MIH86]. Small-B [Gro89c]. Smoluchowski [CD89a, SS80, vDE86, vDE88]. smooth [SMK88]. Sobolev [HS87]. soft [BC89b, CW81, HEH89, Rad81, Ros84]. soft-core [CW81]. soft-sphere [HEH89]. soft-spin [BC89b]. softcore [WC81]. solar [Pet87, SC88]. solid [ANFdL^{+}83, Bar88a, Ben84, Fer88, KY88b, PS88c, SIvL^{+}88, SPA88, PS88c]. solid-fluid [SIvL^{+}88]. solid-on-solid [Fer88, KY88b, SPA88, PS88c]. solidification [DML82]. solids [RW86]. solitary [KN85]. Soliton [WCD85, BBLT81, DCW85, UP84]. Solitons [Lom85, Kod85, LCG85, LCG86]. soluble [Luc86a, MPK88, OS84, SS88a, vD88]. Solution [Blu80, BS84, Cer86, CW81, DS83b, EP88, FFdlM87, HN82, SW87, Val83, WC81, AKS89, Ala87a, Ala87b, Bar88b, DGK81, DH86a, EFN83, FBG86, GvD89, GP89a, HZ82, HS86, IVMA83, Maa84, PHB81, Pol88a, Sch83, Sha82b, SS83, TC84, Tou83, TW85, Wu85, vBD80a, vDG89]. Solutions [CG80, VR85, BGBS88, BG86a, Bro82, BMS82, CH88a, Cer87, Che82, CK88a, CG81a, Cor85, Cor86, Cor87, Cor88, DH86b, Eng88, EHH82, Ern84, FB85b, Fre87, GM85, Gzy82b, HM88, Ito84b, KP85, KEKS83, Pol88a, SZ88, Spo86b, TM88, Vig82, Wid88a, Wid88b, vDE88]. Solvable [For88, HR85, Roe81, Bax88, BD85, Cla83, DFP86, EK85, For84, For86a, KT85, KMP82, KY88b, MB81, Per86, Pom81, Pri81a, RLB86]. solvation [SIK86]. solved [AS86, Bax82]. solvent [BGM86, Hyn86, SW87]. solvents [SIK86]. Some [Ade86, Ber85, Blo81, BR85, Cha84, Fei87b, Fri84, GS81, KY89, KP89b, LMJ87, MW88, New86, Per86, Per87, PG83, Req88, Bax82, CLP87, CL89, Fal87, Gzy81, Hio83, JE89, KLS88a, KZ87, LPT87, Man84c, PS82a, Row86, Sch85a, SVBO88, Sha82b, Wei84b, Sen88, dHK85]. SOS [ABF84, FB85b]. Sound [Sch89b, Joh89, Sch85d]. space [Ala89, Ark83, CDK89, CIPS88, CS83, Eis87, EC81, Mag87, MR87, MR88a, MO82, Ram89, SS80, YHM80, Ber83]. space-homogeneous [Ark83]. space-time [Ala89, Eis87]. span [WR80]. span-constrained [WR80]. spanning [Lav82b]. spans [WDG86]. Spatial [vD89a, BB82, BDE89, LLS88, Req83, SKW82, YAR89]. Spatially [Cor85, SG88]. Spatiotemporal [MKK89]. Special [HBB^{+}84, NLC^{+}86, Sch83, NL87]. Specific [CT84, Fer81]. spectoscopy [Nos86b]. spectra [BMV85, Ben85a, EW88, Fei87a, Har84, LP86, SMYO83, SYOM84, TG81, YMS83, YAR89]. Spectral [BSV87, BLS82, LGP85, Zag82, CK87a, CP82, HR83a, HR83b, Lor84, Man84a, Sen88]. spectral-line [HR83a, HR83b]. Spectroscopy [Gee82, HJ87, Reb86]. Spectrum [AP89, Dör85a, KP85, AKS89, Alb89a, CD89b, CLP87, Del85, Dri81, Dri82, LMR88, PS89, Rez81, Sha86, Süt89, Zol86, vdB83]. speed [Che86]. sphere [ANFdL^{+}83, DE89b, EW84, EW85, EC86, HEH89, Kir89, KH87, LAA87, LA89, LvB89a, LvB89b, PP82, WW86a, WW86b, WT89a, XS89, ZS88, dSEC81, dSCKP89, vBD80a, vBD80b, vBD82]. spheres [CL88, EW81, Gib81, Han83b, Huc83, Kra81, KD82b, Kra88b, LW85, Pom85, Rub86]. Spherical [Kat85, BT88, Bra89, FFC83, GP89a, Mac83a, Pas82, Pas83, PHB81, PWvH84, Phi82, Ros84, SR86, Smi88, Smi89, TF86, VB80]. spherical-model [TF86]. spherically [CWD82, CWD83]. Spin [AM87b, Ko88a, Ko88b, Nis81, You84, Zeg87, vHBW84, AB87a, ABCT86, ARS81, BM89, BK81, BM84a, BBF86, BC89b, BF86, BD85, BLP81b, CEGM84, DD85, DS80, DS88b, DvE89, FLM88, FE87, GD85, GL89a, Gro89a, Gro89b, Gro89c, Har87, HT87a, Kam86, KLS88a, LS87, LR88a, Lin89a, Mal82, MO82, Nie84, O’C84b, Oli88, PPKP89, Pea81, Rag88, RT81, SD81, Sto86, SYG83, Suz86a, Suz86b, TS81, Tas85, Tas89, TCB89, Tho82b, Tru86, Wat89, Wre83, Wu86, Zah87, Zha89a, vE85, vE87b]. spin- [KLS88a, TS81]. spin-boson [LR88a, Rag88]. spin-exchange [MO82, Zha89a]. spin-glass [vE87b]. spin-isospin [PPKP89]. spin-phonon [Gro89a, Gro89b, Gro89c]. Spin-spin [Ko88a, Ko88b]. spin-systems [Sto86]. spinodal [DV87a, EF87, KL83, MV87, MV88b, NC85, SS88a]. spins [ABL87, BG84, CF80, IP86, Loh86, Nea80, Ruj82a, Zha88]. spiral [MPH87]. Spitzer [Fri85]. splitting [AAF89, SMYO83]. Spontaneous [Geo81, KR85, LN85, Lin87, LM88, Sek83, KLS81, NB87, Wre83, YS85]. spreading [dSH88]. square [BET80, Guj82, Haa89, KTP88, LM88, Mei83, MHB86, Ros87, WMvBL88]. square-lattice [KTP88]. square-shoulder [WMvBL88]. square-well [WMvBL88]. squares [PS88a, RS87a]. squeezed [HW88]. Squig [Man84b, Man84c]. ST [Sug86]. ST-100 [Sug86]. Stability [BG86b, DP85, Gri88b, KD82a, LFtH88, Wes83b, BT89b, Maj89]. stabilize [BvH89]. Stable [Bon87, FR81, Ben84, DGK81, FN83, MT87, MB84]. staircases [Uim84]. Stan [Fei85]. Star [HJ87, Lav82c, Sha88, BA86]. star-mesh [Lav82c]. star-triangle [Sha88]. Star-unitary [HJ87]. Stat [AV03]. state [AL81, ABF^{+}85, ARS81, Bar88a, BBF86, Bin81, Bre87, Büt89, CFG80, Coh88, Dee85, EW84, EW85, FH87, FN83, Gro89a, Gro89b, Gro89c, Haa89, HBB88, Han86b, HN80, HR80, KLT88, KRS89, KP86, KY88b, LO85, LS82c, MCDK85, Mei88a, MHV89, Nis81, Pea85, Pia83, Rad81, RHG89, RG86b, SS84a, Sek83, SM89, SS83, Val85, Yeu89, ZT85]. state-dependent [Büt89, SM89]. States [AF81, ABvT87, Ala89, AZ85, AP89, BGBS88, BPT83, BLP81b, BS89c, Bru88, Bur88, CT80, CFS84, CFGK87, DGK81, DG83, Dic89, DP85, Eva89, FGS84, GLM87, GKI85, GM88b, GM96, Gra81a, Gri88b, Gro85, HW88, JK89, KIM86, KD84b, KLS84, Key82, KLS81, Kle86, KLS89, KMS87, Lan88, Mai87, MB81, McL82a, McL82b, MKB86, NL85, Par87a, Pra83, Pul82, Rad84, Rad86, SS89, WL88]. Static [Ber86, Bin81, FJ86, YHOF85, Req82, Sch86a]. Static- [Ber86]. stationarity [Oji89]. Stationary [GLM87, GP86a, GKI85, IT88, WJS84, ABF^{+}85, BPT83, BC89b, CIPS88, Cuk83, DM81, EMDvB84, Fed88a, Fri86, HZ86, HN80, MEvBD84, MGL87, MB81, MC81, Pia83, SC85a, SC85b, SC87, Tél88, ZWLV88]. Statistical [Ale85, Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano80o, Ano80n, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano81o, Ano81p, Ano82l, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano82q, Ano82n, Ano82o, Ano82p, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l, Ano83m, Ano83n, Ano83s, Ano83t, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano84u, Ano84v, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano85i, Ano85j, Ano85k, Ano85l]. Statistical [Ano85m, Ano85q, Ano85r, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano86m, Ano86b, Ano86q, Ano87h, Ano87c, Ano87d, Ano87i, Ano87e, Ano87f, Ano87j, Ano87k, Ano87l, Ano87m, Ano87n, Ano87g, Ano87o, Ano87p, Ano87a, Ano88e, Ano88c, Ano88d, Ano88f, Ano88g, Ano88h, Ano88i, Ano88j, Ano88k, Ano88l, Ano88m, Ano88n, Ano88b, Ano88a, Ano88o, Ano88p, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k, Ano89l, Ano89m, Ano89n, Ano89b, Ano89o, BH83, Bau87, BCF86, Bav87, Bav89, Dem83, DPT80, Dup89, Gre85, Gre89, Høy87, JJ86, Ket84, KW88, KLW86, KN87, Leb86a, Leb86b, Leb86c, Leb88d, LRS88, Leb89a, Leb89b, Leb89c, Leb89d, LRS89]. Statistical [Leu87, MS82b, Nel89b, RP85, Rei84, Shl85, Wes85b, Wes87, WP82, BFA86, BCG^{+}80, BSZ89, BS81, BEF86, Cep86, Che82, Chu85b, Col86a, DK86a, Der87a, Der87b, DK86b, FS81, Hio83, Ito82, Jen81, Kie89, KW81, Kog84, KWL87b, Kra81, KD82b, LG87, LPW81, LLE^{+}83, Lev88, Lin84, Lin88, LPRV87a, LS89b, LS89c, LWS88, LV80, MB89, Nak86, Opp82, Par85, Par87a, PS88b, PS88c, Ram86, Rap89, Red81, Req82, Rub82, Ruj87, Sch86c, SG81, Spo86a, Suz86a, Wes86a, Wes88, Zwa81, Ano85s, BT85, Wer84a, Opp88b]. Statistics [Gee82, HK85, Bax82, FP80, HR83a, Par85, PP83, Pom81, Pri81a, Pri81b, Pri86, PS88b, Ram84a, SL80, Sta84, Tsa88, WSL83, Whi89]. Statphys [Dom87]. stay [WZ89]. Steady [LS82c, Del82a, Del82b, Dic89, Gra81a, KLS84, Key82, ME87, WL88, Yeu89]. stellar [Gla85, Wei85b]. step [PS89, WS82]. stepped [WM83]. steps [BZ83, dH84a]. stereochemistry [Bro81]. sticky [HR80]. stiffness [PS89]. Stillinger [MG83]. stimulated [DCW85, WCD85]. Stochastic [BNMT83, DK86b, DGL83, Fox87, FL85, FN83, FCK83, Gol87, GR89, GNKW89, HR83a, HR83b, Ito88, JS86, KHM^{+}83, LN84, LLS88, ME82, MV86, MR88b, Nak86, Nos83, PW84, Wei83, AMC89, ABKS89, BZ87, Blo81, BM84b, BS87b, CDPP89, CS89, CGOV84, DM81, DLT88, ER81, ED81, Fed88a, Fox89, GHKM89a, GHKM89b, Go83, GKI85, GP86b, GG87, GHT80, GSH88, HS88a, HS87, HMS83, Ito84b, KLS84, KLPS83a, KLPS83b, KWL87a, LS87, LSSW80b, Lin89b, MLSK85, Mar85, Mas89, PC86, PW88a, PW88b, Rei88b, Ruj84, Sch86a, SS82a, Shl84, ST88, Spi85, TC84, Uey80b, VM86, VM87, VM88, VMA83, WL88, Wei84a, Wei86a, Wei86c, Wes84b, ZS82, vBS83]. Stochastically [BJLS89]. stochasticity [BG82, KLK^{+}85, PPP84]. stochastics [Opp86b]. Stokes [Ala88, Ala89, BF81a, Che82, Cor89b, FT81, FTZ84, FGN88, TL82, Vis85]. straight [Fed88b]. Straley [vE87a]. Strange [MR83, WBKS83, BPTV88, Bro89, Fei87b, HK85, Kad86, KWL87a]. stratified [Str82]. Stratonovich [vK81b]. stress [Eva80, Wag89]. stress-stress [Eva80]. Strict [KZ87, AJ81]. string [Bor87, PW86]. strings [Jen86b]. strip [BBP87, CFS83, FJS83, KLS89, MO83]. strips [Kle80]. Strong [BCG^{+}80, MKD87a, MKD87b, Meu85]. strongly [BLM87, DMM^{+}89, HJM89, LM86]. Structural [LAA87, Tru86, LGP85, NH88, Sch85d, vHZ88]. Structure [KFB89, Sin83, AM84, AGL88, BK88a, BLP81a, Dei85, Dei89, DDI83, GLO86, GV89b, Haa89, LW85, Mag87, MS82a, Req88]. structures [BTZ85, BDWK87, HH85, KK85, Orb84, Pro84, Ram84a, SB84, SG88, UP84, YMS83]. studied [BWdF88]. Studies [Rei88a, Zwa81, BD82, CDM86, Cha84, Gut84, HW86, Kir84a, MS86b, Neg86, PAF84, PAF85, SSM85, Sug86, dOB89]. Study [BBAH83, Fye86, HD83, IM86, SIK86, AKvL88, dAB87, BL85, BF88, CLWH82, DMMSR88, DST83, De 85a, FCK83, FCD86, Gla88, Haa89, HHO86, IVMA83, JM86, Ker86, KBZ84, Lam89, LS84, Mei83, Pol89a, SMYO83, STA88b, Vid87, YMS83, ZBKS89]. studying [Sca86]. sub [Con82]. Subdominant [Mou82]. subdynamics [GM87]. subgraphs [Lav82a]. subject [Gro83b]. subset [WS82]. substitutionally [KFS83]. such [LW86]. Sum [Vie85, VB89, GLM80]. summability [GKT84]. Summary [BH83]. Summertop [ZAT88]. Super [Suz88]. Super-effective-field [Suz88]. superconducting [Abr85]. superconductors [Nel89b]. Supercritical [Kad83, Sar84, GMO89]. superfluid [HS89, Rei80a, Rei80b, deV85].

Superintegrable [Bax89]. Superionic [Die85]. supersaturated [BK80]. Superstability [BMPP80, LPdS84]. superstable [CCO83, Par85]. supersymmetric [Con88]. superuniversality [Blu89]. Surface [BT89a, BFL82, BKL83, CPRV89, Jan84, Jan85, KP89a, LW83, MS82a, MJ85, BW82, HW85, KVGZ87, MK89, PAF84, PAF85, PS89, SK82, SBS88, Smi89, WW86b, Wid88b, WM83]. Surfaces [Abr84, Bor87, BEF86, Gla88, Koh88, Sch85c, SMK88, WM83]. surfactant [BGBS88]. survival [WHB85]. susceptibilities [MH81]. susceptibility [BG84, BSF89, CPR86, CPR87, DD85, FE87]. suspended [Nos84]. suspension [FFC82a, FFC83]. suspensions [CF88b, Coh88, GM88a]. sustained [Dei85, Tél88]. swarms [FvdM89, PFvdMZ89]. Symmetric [CdS87a, CdS87b, CWD82, CWD83, KK84, Ken82, MS81b, PC87, dHK85]. Symmetries [Kal81, RT86, FM84, KD82a]. Symmetry [BF81a, GH82, AP89, BPZ84, BT85, BC89a, HT88, KLS81, KR85, LN84, LLS88, LN85, NB87, Pic83, Sim81a, ST89c, Sza89, WW88a]. symmetry-breaking [LN84, LLS88, NB87, ST89c]. symplectic [LPRV87b]. Symposium [Lee83, Ano87q]. synchronization [BCM87, YKS81, BCM88]. System [MS82a, AJ81, BvH89, Ben80, BW88a, BLP81a, Bur86, CDPP89, CF88a, Che86, Cor89b, DM87, EK85, FC89, For88, FN83, GG82, GLR86, GKI85, GG87, Gro89a, Gro89b, Gro89c, GNKW89, HS86, Iye86, Ken82, KK85, KD82b, Kür88, tLC86a, tLC86b, tL88, LSM89, Lu87, LOS88, MTG87, MP82, Mar81, MLSK85, Mür87, NMM80, OY87, Pan89, PPKP89, PIP89, Pet87, PS85, PW88a, PW88b, Rei88b, RR82, RLB86, Rub84, Sch83, SKA86, Sek83, SM89, Skr84, Wic85, ZS80, vB85]. Systematic [NMM80]. Systems [AV03, bA88b, AF81, AL81, AJ84, AM83, AFV85, AN87, AMV85, AV87, ABCN81, AAF89, Bak82, BMV85, BH88a, Bau83, BG86b, BMPP80, BS83, BDL81, Bis82, BPT83, BPK82, BC89a, Boy88, BDE89, BLP81b, BKL83, BLM87, BS89c, Bur86, CPPV89, CCO83, CCR82, CGVG83, CCR85, CFG80, CPR86, CPR87, CPRV89, CF88c, CLP87, CL89, CFGK87, CWD82, CWD83, DD85, De 85a, DFL86, DFL87, DMM^{+}89, Dei89, DLS85, DK86a, DG83, DV87b, Ebe86, EJT87, ED81, FVV82a, Fed88a, Fei87a, FLM88, FF87, FM84, Fox83, FGP80, Fri84, Fri85, Fri86, FP80, FILS80, FSW86, Gar83, GKN84, GW87a, GM87, GP83, GG88, GMC81, GP81, GP86b, GP85b, Gri88b, GLM80, GS88, HW82, HL89, Har87, HT87a, HNMK80, HW85, Hug85, IKY88]. systems [Jan82a, Jan82b, Jer87, KF85, KCH88, Ken84, Ker88b, Kes88, KK88, KRS89, KW83, KP86, KY85, KLW86, KWL87b, KS81, KLSZ86, KN87, LH88b, LLE^{+}83, LM84, LS87, LPT87, Len88, LN85, Lin89b, LPRV87a, LN88, LV80, Mac87, MGTM88, MPPT78, MY80, Mar85, MOS89, MR83, MPK88, MBG82, MGP^{+}83, Muk83, Nak81, NTW88, NT89, Neg86, NB87, NM84, O’C84b, OHI88, Oji89, Oli88, PSMZ84, Par80, Par82, PW81, Pol84, PS89, Pul82, RL86, RHM86, RPSS85, RT81, RRO88, RSK86, Rub82, Rue86a, Ruj82a, Ruj82b, Ruj84, SSD82, Saw84, Sca86, Sch89a, Sch85a, Sch86a, ST89a, Shi83, SV80, Spe86, Spi85, Ste81, SLM89, Sto86, SS83, SGL89, Suz86a, Suz86b, Tar87, Tas85, Tas89, TM89, TG87, VM86, VM87, VM88, VtLZ89, Vig82]. systems [Vil87, Wag89, WBL89, Wei86a, Wes83b, WF88, YHOF85, Yan87, ’t 88, vB84, vDE87, vD87, vD88, vEvH83, vEvH85, MPPT80, WO86]. systems- [GP86b].

table [ATZ83]. Tagged [MD83, CK88a, MK84]. Tail [KV82, Eva80, SC88]. tailed [WWH89]. tails [EMDvB84, EvV89, HZ86, KS86, LMN84, LN88, MEvBD84, Mac86, MK85a, MK83a, MK85b, MS83b, Sim85, Sim87, Woo89]. takes [Pea85]. tale [MS83b]. talks [Ano87q]. Tangential [Kle89]. target [Vid87]. tasks [BSZ89]. Taylor [BMO^{+}84]. technique [KW82]. Techniques [Gre82a, Ken83, Pra83]. Teller [Igl89, Pfi82, Sal88, dOB89]. Temperature [BFG86, MNT85, ACH88, BP89, CkM82, DV87a, Deu83, DS86c, GD85, GPS86, Gut84, Joh89, Jon86, Mon85, O’C84a, OB84, Pra83, RS87b, SG87, Sim80, Sim81b, Sok81b, SGL89, Tar87, Uim84, Zah87]. temperatures [GLM87, HKZ88, KD83, Lav87, MZ89, Yan87, Zeg87]. Temperley [DDdG83]. temporal [SKW82]. tension [Bor87, BFL82, BKL83, KVGZ87, MJ85, Wid88a, Wid88b]. tensor [Wag89]. tent [Dör85a, SMYO83, YMS83]. Tenth [Ano84v]. term [kM83]. terms [VR85]. test [Pia84]. Testing [CR86, DP87]. Tests [FRY87, MS86b]. their [Bar85, BA87, PSMZ84, RW86]. theme [Kas87]. theorem [ACT82, Bav86, BS89b, CC82, CF88a, CL81, Del84b, FHT88, Gol85, GP89b, LPW81, Mar81, Pia87, Pla86, PG83, Rez81, RL80, Wer88]. theorems [AC84, Col86b, CG81b, Mon83]. theoretic [Par80]. Theoretical [Ano81o, Meu84, Wei86e, vDG89]. theories [Bau87, Bor87, DP87, DBK^{+}86, PYC85, SI87, d’E89a]. Theory [Ano80o, BK80, cC87, Fai81, FGMZ85, GVV89, HS89, Kop86, KVGZ87, MGP^{+}83, SS84a, Wol82, YS85, Ala87a, Ala87b, ACH88, AF86b, BK81, BNMT83, BR80, BW89, BS89b, BP89, BM88, BF86, BLS82, CR86, Cam81, CK87a, CGVG83, CP87, CL88, CIPS88, CC84, qC81, qCeY81, Chr86, Chu85b, CF89, CdS87a, CdS87b, Cuk82, DS81, DGK81, DF87, Der87a, Di 84, DS83a, DZ87, Eis87, ERH85, EB88, Fam88, Fei88, Fer81, FCB86, FT88, FvdM89, GHKM89a, GHKM89b, Gol87, GP86b, Gre86, Gro85, GGP86, GV89b, Hee82, HR83a, HR83b, JMF88, JK89, KGA84, Ken84, Kes81, Key82, KL82, Key83, KS85, KC83, Kog84, Kog86, Kos86, KMS87, tLC86a, tLC86b, LYAA83, LCY83, LV80, Mag87, MK85a, MKD87a, MKD87b, MT87, MY85, MHB86, MS86a]. theory [NGH82, OHI88, Oka86, PS88a, PB83, PLM^{+}84, Per84, PT87, RV81, Ros86, SC85a, SS88b, ST81b, Sen88, Shi83, Shl85, Sla83, Spe84, SKS82, Suz88, ST87, TB88, Van88, Voo85, Wat89, Wer84b, Wes85a, WT89b, Wol86, Wu88, dlL82, vBD80a, vBD80b, vBD82, vB84, vB85, vHGHK88a, EMDvB84, RS85a, Kes88]. there [d’E89a]. Thermal [SVBO88, Fri86, GP86b, GG87, LS82a, LS89a, Maa84, SF89, WP82, vBC88]. thermally [FGS84]. Thermo [Suz86b]. thermochemical [TvdB88, Van83a]. Thermodynamic [Bax89, LS82a, RHM86, dSVT87, BS81, Bon87, BCM87, BCM88, DL82, FP80, IP86, Kei84, Kra88b, PV89, PPKP89, Rue81, Saw84, SBS88, Tak89, Wes83b, vEvH83, Wer84b]. Thermodynamics [Ano88a, Chu85a, DP89, Hug85, JW89, VB80, AL81, Bal87, Ban87, CCP87, Del84b, DPT80, DDL87, Eu84, GCF82, GCdHR^{+}84, Gre86, HNMK80, Kei83, Lan84a, LR88a, Lin84, MK89, Nak81, Opp82, Opp88b, Opp89, Rei88a, Sin83, WW82, Wer84a, BMS82, Col86a]. Thermophoresis [Key83]. Thermophysical [Lee83]. thick [HP87]. thin [Bab86, KD84a]. Third [Gib81, AL81]. thirds [cCM89]. Thouless [BP89, Dun85, MP89]. Three [Bau85, BBF86, CF89, EN85, Lin89a, McK87, AB87a, BNHF83, Bil88, CS86b, CK87b, CMM86, DdGM89, EK85, FKP88, FG88, FG89, GDx82, HCKS84, HK86, HS86, KD82a, LS84, Law88, LW83, kM83, Mal82, MLSK85, MHV89, MJ85, NR88, PSMZ84, PSZ87, Par80, Pri81a, Pri81b, RHG89, SL80, dFKP87, vVE87]. three- [NR88]. Three-body [CF89, kM83]. three-component [HS86]. Three-dimensional [Bau85, EN85, BNHF83, Bil88, CS86b, CK87b, CMM86, FKP88, HCKS84, LS84, LW83, MLSK85, MJ85, PSZ87, Par80, Pri81a, Pri81b]. three-point [GDx82]. Three-spin [BBF86, Lin89a, AB87a, Mal82]. three-state [MHV89, RHG89]. threshold [ED81, KL82, PSMZ84]. thresholds [Sha82a]. Thue [AP89]. TI [SH89a]. TII [SH89a]. Tiling [Sch88b]. tilings [GL89b]. Time [APSJ85, CC82, Ern87, JLM85, JMM87, MP82, MO83, MSW82, PS85, Vig82, Ala89, ADG88, BDKW88, Ber86, CCR82, CCD86, Der87a, Der87b, Eis87, EMDvB84, EvV89, EW81, Eva80, FHW88, Gar83, Gar87, Gzy82a, HZ86, HR83a, KB85, KWL87b, KSB85, Kür88, LMN84, LSSW80b, LS89b, LN88, LOS88, Mac87, MEvBD84, Mac86, MV88a, MK85a, MW88, MK83a, MK84, MK85b, MR83, MR87, MR88a, NM84, Pea85, PS82b, Pul82, Rue86b, SS88b, SG87, Shl84, Skr84, SF89, Vai89, Wei81, WDG86, WHM89, WWH89, Wes84a, Woo89, YAA82, Zwa83, dSCKP89, vBLLS80, vVE87, JLM95]. time-correlated [KWL87b]. time-correlation [Ber86, EW81]. Time-dependent [Ern87, JLM85, JMM87, FHW88, LOS88, vVE87, JLM95]. time-local [Der87a, Der87b, NM84]. times [BR86, CS89, Fei87a, HDL89, KMST86b, LS87, LvB89a, LvB89b, LSSW80a, MOS89, MV82, THFT87, VB86, WSL83, Wei84b, dSEC81]. Titles [Ano87q]. Tjon [Cor85]. Tolman [Hee82]. tool [HJ87, Sca86]. tools [VV89]. top [Rav89]. topological [BKLP89]. topology [Dup89]. torichaos [Rie85]. torque [vBD82]. torus [Sal88]. Tracer [Spo87, vBK87, BH88b, KvB85, dH89a]. traffic [RHM86]. trajectories [EC86, NZ86, Rap85, Sar84]. Trajectory [GRS80, Kad83]. Transfer [AKS89, Le 84, SD86, BZ83, Con88, DZVS84, Gie89, Gut86, KW81, Lam89, MR88a, PS89, Zol86]. transform [BFG86, PA88]. transformation [GM88b, Hol88, Lav82c, SF88, GM96]. Transformations [HC88, Cos80, DvE89, GBP88a, HJ87, MT82]. transforms [BGM84, LWS88]. Transient [Agm86, CPPV89, FN83]. Transition [BMV82, BG82, NC85, Rei88b, ZS82, AS86, ANFdL^{+}83, ABF87, AG85, BP89, BD85, Cal83b, DS86c, Dun85, FCK83, GN85, GG82, GKT^{+}86, Gra84, HHB87, Hög88, Huc83, Kam86, Kat85, KY88a, KMS87, Kra88b, KN87, LD87, MP89, MLSK85, MST84, NR82, PLM^{+}81, PLM^{+}84, PP83, Pol84, PS88d, SH89a, Sim80, SMM89, Süt80, SWYM87, TL82, Tou89, VA81, Wei87a, Wic85, dS88a, vDE86, vHZ88]. transition1 [Per87]. Transitions [Rie85, BD88, Bar89, BSTK87, BCSY84, BT85, Bon87, BS89c, DP87, DBW81, DH85b, DP89, EE88, FILS80, GKK87, Gro81d, Gun84, KT85, Kir84b, KP86, KU86, Kor86, MV87, MV88b, Par80, Pen86, Pen87, PW81, Pfi82, PS82c, PF83, Req82, Ryt87, Sch85d, SMYO83, Sin82, Uim84, VM86, VM87, VM88, WL88, Wes85a, vEvH85]. Transitivity [NTW88]. translation [LN85]. Translational [MK85b]. Transmission [Fel86, DS86b]. Transport [Büt89, NM84, Rub84, WWH89, vBS83, Bal87, Bur86, CP82, CH87, Del84a, Der87a, FvdM89, Hor85, KK88, KFS83, KEKS83, MKD87a, MKD87b, MGP^{+}83, Ram86, Sen88, Sta84, VD89b, vVE87]. transverse [DD81, FE87, HGR87, Kir84b]. trap [PPK84, WHB85]. Trapping [Hub83, WH84, ADG88, KBZ84, KHNA83, Mac86, MSW82, MGP^{+}83, Web84, Wei86d, dH84a]. traps [Agm86, Hal83, SLW84, Wei86d, dH84a]. traveling [AKvL88, BM89, BFA86, DS88b, RG86a, Vid87]. traversal [Gar87]. Treatment [CC86, Val85, AMV85, Bau85, CCRG80, DH87, LV88, RHM86, VD89b]. Trebon [Ano87q]. Tree [AN84, GvD89, HKW^{+}85, HKLW87, HM84, ITH83, MTA85, MH81, Pri86, Tho82a]. trees [CG85, DS88b, Lav82b]. triangle [BA86, Ngu87b, Sha88]. Triangular [EW82, KIM86, KU86, Mal82, Wu85, Zia86]. triangular/honeycomb [Zia86]. tribute [Opp86b]. Tricritical [VA81, dAB87]. Trinomial [AB87b]. Triple [McL82b]. Triple-collision [McL82b]. Triviality [GK82]. trotter [Fye86, Val83]. trotter-like [Fye86]. Truncated [TL82, BF81a, FTZ84, Gra82a]. truncation [FT81, HT80]. truncations [FGN88]. tubes [Bab86, Kle80]. Tunneling [PMBB^{+}84, JW89, SS85, SD85]. Turbulence [Bro89, Ala88, BMO^{+}84, HS89, Hor85, Kra88a, LLE^{+}83, Lin85, LFtH88, Nel89a, OY87, Pro84, Str82, Yeu89, ZBKS89, Wes86b]. turbulent [Ala89, qC81, qCeY81, DP85, Tou89]. Turing [Ben80, Ben82, BDWK87]. Twelfth [Ano86q]. Two [BM84b, BL84, Cor89a, Deu83, FFC82b, Gib81, HV85, Jer87, Joh89, Mal82, PW85, SI87, SS89, UP84, Wu86, AJ81, AJ84, ACH88, ADM83, AB87a, ADFJ89, BM89, BD82, BPP84, BD88, Bax82, BPZ84, BPP87, BT89a, BS87a, Bin81, BPK82, dAB87, BLP81a, BH88b, BR85, Bur88, CR86, CLWH82, CG85, CCD86, CCS87, CFS83, Cor88, CJ87, CJB88, DP87, Dic85, Dup87, ED81, Fed88b, Fer81, FdF86b, FLNO88, FJS83, For84, For88, For89a, FM87, GN85, GKN84, Gat88, GS86, Gra82a, Gro81b, Gut84, Hoo86, Ima86, Jan84, KP89a, KTB86, KLT88, Ker88b, KZ87, KRS89, KLS81, Ko88a, Ko88b, KU86, Kor86, KLS88b, LMJ87, LH88a, Lav87, Law88, LOS88, MP89, MK87, MM84, MS88b, MIH86, MPH87, NNK89, Ngu88, Nie84, NR88]. two [NHH88, OMS86, PV89, PSMZ84, Par80, PS82b, PL87, PS88c, Req83, Req88, SB84, Sal86, SD86, SS84a, Sch83, Sch85a, Sch88c, Sek83, Sel89, Ser86, SL80, Sor88, Spe86, Sta87a, Sta88a, Stö88, SPA88, TCB89, Tor86b, VM87, VM88, VA81, Wag83, Yan87, dAC87, dHK85, dFSZ87, vVE87]. Two- [Mal82, AB87a, Par80]. Two-body [HV85, Gro81b, Hoo86, MIH86]. Two-cell [BM84b]. two-component [Dic85, For84, For88, For89a, Sch83, Sor88, VA81, dAC87]. Two-dimensional [Cor89a, Deu83, Jer87, PW85, SI87, UP84, Wu86, AJ81, AJ84, ADM83, ADFJ89, BM89, BD82, BPP84, BD88, Bax82, BPP87, BT89a, BS87a, Bin81, BPK82, BLP81a, BH88b, Bur88, CLWH82, CCD86, CCS87, CFS83, CJ87, CJB88, DP87, Fed88b, Fer81, FdF86b, FLNO88, FJS83, FM87, GN85, Gat88, HV85, Ima86, Jan84, KP89a, KTB86, KLT88, KZ87, Ko88a, Ko88b, KU86, Kor86, KLS88b, LH88a, Lav87, MP89, MK87, MM84, NNK89, Ngu88, Nie84, NR88, NHH88, OMS86, PV89, PL87, SB84, Sal86, SD86, Sch85a, Sel89, Ser86, Spe86, Sta87a, Sta88a, Stö88, Yan87, dFSZ87]. two-level [LOS88, Sek83]. two-parameter [GKN84]. Two-particle [FFC82b, Ker88b, Req83]. two-phase [Req88, Tor86b]. Two-point [Gib81, Gra82a]. Two-sided [SS89]. two-spin [TCB89]. two-state [KRS89, SS84a]. Two-temperature [Joh89]. two-time [PS82b]. type [ABF87, ABF84, ABCN81, Fed88a, Kat85, KEKS83, LPS85, LPdS84, Meu84, OB87, Ste89, VMA83, ZT85]. type-I [Meu84]. Typical [Ngu88].

Uhlenbeck [Row86]. Ulam [Fei85]. ultrarelativistic [Cer86, Fre87]. ultrasound [deV85]. unbounded [BG84]. undecidability [Sch88b]. underdamped [RV85]. Understanding [Fif85]. uniaxial [FdF86a]. unicity [Man84b]. Uniform [MST84, BSF89, JMM87, KvB85, MD83, RSRD83, Tej89]. uniformly [KU86, Kor86]. unique [FN83, KLT88]. Uniqueness [Kle86, MN84, CP87, Gro81d, Par87a, Pra83, RS87a]. unit [Cos80]. unitary [HJ87]. universal [BL85, GS81]. Universality [Blu89, Fam84, Stö88, Tra88, MJ85, Pol84, PS88d, WL88]. universe [Eis87]. University [Ano85s, Ano87o, Ano88o, Ano81o]. Unpredictability [MOS89]. Unstable [Val85, AMV85, DPV86, Dee85, Dei89, DM87, FR81]. until [dH84a]. updating [Stö88]. Upon [Phi82, CC86]. Upper [Gro83c, Lav82b, Mon85, Pic84, vE87b, KK88, Lav82a, Sim80, Siu85, Tas89]. use [De 85a, SV80]. uses [And86]. using [BB80, Cho88a, cCM89, DBK^{+}86, Gro83c, Loh86, Mon83, PA88, RR82, Sha82b, SIK86, Val83, Zha89a]. utility [Phi82].

vacancy [BWdF88, BH88b, KI83]. vacua [AF81]. Vacuum [GM87]. validity [CC84, vK81c]. value [CP87, IT88, ST89a]. Vanishing [Wre83, Hug85]. vapor [BK80, HGR87, HP87]. variable [SS85]. variable-range [SS85]. variables [CD89c, GHT80, Jan89a, Jan89b, Jan89c, Shi81, YS85, vK87a]. Variance [Tor86a, Sch87a, Sch87b]. variant [ABF^{+}85]. variation [An88, BWdF88, Mor84, Sch84, Sch88b]. Variational [Mal82, ST81a, BS89b, CK88a, CWD82, CWD83, PC86, PC87, Ros86, Saw84, Sch84, Sch85a, WKC83, Gro89a]. Variations [Kas87]. various [Mon85]. Vasserstein [KRS89]. VBS [Kna88]. VBS-quantum [Kna88]. vector [AZ85, Bau83, Kal81, KY85, Kur83b, Pea85, WZK84, vE85]. Vectorized [BS89d, Zha89a]. vectors [GM88b, GM96]. Velocity [Der83, Kos86, MPPT78, CK88b, Cor87, CWD82, CWD83, GP86a, Kir89, MK85a, PP82, SC88, VKK85, WS83, dH89a, MPPT80, vBS83]. versions [CR86]. versus [BW84, GNKW89, LH88b, MV88a, vK81b]. vertex [ABF84, BBP87, Cla83, DdVG89, DS83b, FB85b, Ham87, NNK89, Sha82b, Tru86]. Very [dSC82, BS84, KW82, Lor84]. Very-short-wavelength [dSC82]. vessels [GM88a]. VI [Spo86a]. via [HMH84, HK86, KI83, KH87, LH85, MN84, VH87]. Vibrational [KCH88]. Vicious [For89b, FG88, FG89]. view [Der87a, Der87b, Kog84, Meu84, Rub84]. viewpoint [GP83]. VIII [NGH82]. violations [GG83]. virial [ANFdL^{+}83, Gib81, Kra82a, Kra85, Rai81]. viscoelasticity [Bav87]. viscosities [SKS82, XS89, dSvLH89]. viscosity [EW81, LH85, MIH86, Rot89]. viscous [Rub86]. visited [WS82]. visual [BGWY85]. Vlasov [Ala87a]. Vogel [BS88]. volatile [DPT80]. voltage [NT81]. Volume [Rap87, Wes84a, BMS89, FKP88, FK83]. volume-preserving [FKP88]. Voronoi [Han83b]. vortex [BMO^{+}84, CF88a]. vortices [MO83].

W [Opp86b]. Waals [DS80, NGH82, SD81, vL89]. Wagner [BPK82]. Wainwright [Oka86]. waiting [WWH89]. waiting-time [WWH89]. walk [AV81, BDKW88, BS85, BZ83, BMS87, BLP89, CV87, GN87, HGPRS89, HWW82, Hio83, KBZ84, KI83, KLS88b, Law88, MS87, MS88a, Nau85, NG87, PP83, Ram84a, Ric83, ST89b, Sta87a, Sta88a, Sta89, SN84, ST81c, Tor86a, WS82, WHB85, WDG86, WWH89, WS83, Zol87, bARC89, dFKP87]. walker [AAK84]. walkers [AKN84, FG88, FG89, For89b, LSSW80a, Pri86]. walking [AS89]. Walks [Fis84b, AH89, ABG^{+}89, Bar85, BR86, BBAH83, BHR^{+}88, Dom83, Dup87, DD88, EvV86, Ern87, Ern88, Hal83, Hey82, HMS82, HS82b, JM86, KdH83, KB85, Kle80, KHNA83, Lec85, LMN84, LSSW80a, Mac83c, Mad88, MV88a, MV82, NCF88, Par89, PS88d, RS85a, RS85b, SD86, SKW82, SLW84, ST84b, WR80, Wei81, Wei84b, WH84, WW88a, Wes82, Whi83, ZBKS89, Zwa83, dFKP86, dlSLW88, dH84a, dHK85]. Wall [Req88, BP85a, Bav86, Jan82a, Jan82b, Key83, LW85, Per80, Tit84]. wall-bounded [Per80]. walls [Fis84b]. Ward [FB85a]. water [TWB86]. Watts [MB84, Shl84]. Wave [Ste81, Cor88, Far87, FF86, FC89, FFdlM87, Lev88, Sek83, Sor89]. wavelength [dSC82]. wavelet [GV89a, Hol88]. waves [BCFN89, DS88b, Fel86, KN85, Meu85, MPH87, Vid87, WS84, Wes86b, vHBW84]. Weak [BK88b, ZD88, AN87, BMV82, GT84a, GT84b, Gra89, HMC87, Jau86, LFtH88, Pia81b, VMB82, WS84]. weak-noise [GT84a, GT84b, Jau86]. weakening [KSB85]. weakly [CCR82, CWD82, CWD83, HDL89, HW85, MKDC85, RRO88]. weakly-coupled [HW85]. Weiss [dH85a, AZ85, EE88]. well [RV85, Ryt87, VR85, WMvBL88]. West [Ano81p, Ano82q, Ano85a, Ano86q, Ano87a, Ano88b, Ano89b, Ano84v]. Wetting [BL87, AS86, BEF86, DMMSR88, Fis84b, FLNO88, Per87, SIvL^{+}88, dCD87]. where [DdGM89]. which [DIP82, DFIP82, KW88]. white [HGPRS89, RPSS85, Van83b]. width [AKS89]. Wiener [BMS89, cCM89, Gro83a, KLS88b, Nak86, SL80]. Wigner [Ima86, NR88]. Wild [Bar88b]. Williams [MB84, Shl84]. Wilson [Mis85]. wind [SC88]. winding [Sar84]. wire [MB81]. within [Bab86, Git88]. without [Bro89, DM81, Mie89]. WKB [CCRG80]. wonderful [Opp86b]. Works [GP85b, Ano85p]. workshop [BH83, HL82, LLE^{+}83, dAHK^{+}89, JKHV85]. world [Loh86, Opp86b]. Wu [Cor80].

XXZ [NKO^{+}89]. XY [AB83, HR86, KU86, Kor86, SS86]. XYZ [OMS86].

Yang [Dav87, De 87, Kur83b, MS84, VTB86]. YBa [BWdF88]. Yevick [Var81]. York [Ano88o]. Young [dCD87]. Yukawa [BR80, Ben85b, Blu80, CW81, WC81].

Zernike [Blu80, CW81, WC81]. Zero [GD85, Tej89, BBF86, DV87a, De 87, DF84, EF87, FHT88, Süt89]. zero-degree [EF87]. Zero-frequency [Tej89]. zero-range [DF84, FHT88]. zero-temperature [DV87a]. Zeros [Mon83, DDI83, GPS86, MS84]. zeta [BSV87]. Zhabotinskii [RDRS87]. Zipf [NNN89].

References

Adler:1988:PCN

[AA88] Joan Adler and Amnon Aharony. Percolation cluster numbers. Journal of Statistical Physics, 52(1–2):509–517, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

            tronic). URL http://link.springer.com/article/10.1007/
            BF01016433.

Antonets:1989:OSS

[AAF89] V. A. Antonets, M. A. Antonets, and V. A. Farfel’. Open systems of splitting particles. Journal of Statistical Physics, 55(5– 6):1053–1064, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

            com/article/10.1007/BF01041078.

Argyrakis:1984:SRW

[AAK84] P. Argyrakis, L. W. Anacker, and R. Kopelman. Single random walker on disordered lattices. Journal of Statistical Physics, 36(5– 6):579–589, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012923.

Araki:1983:DEP

[AB83] Huzihiro Araki and Eytan Barouch. On the dynamics and ergodic properties of the XY model. Journal of Statistical Physics, 31(2):327–345, May 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011585.

Andrews:1986:LGG

[AB86] George E. Andrews and R. J. Baxter. Lattice gas generalization of the hard hexagon model. II. The local densities as elliptic functions. Journal of Statistical Physics, 44(5–6):713–728, September 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011904.

Alcaraz:1987:CBI

[AB87a] Francisco C. Alcaraz and Michael N. Barber. On the critical behavior of the Ising model with mixed two- and three-spin interactions. Journal of Statistical Physics, 46(3–4):435–453, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013367.

Andrews:1987:LGG

[AB87b] George E. Andrews and R. J. Baxter. Lattice gas generalization of the hard hexagon model. III. q-Trinomial coefficients. Journal of Statistical Physics, 47(3–4):297–330, May 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01007513.

Angelescu:1981:CBK

[ABCN81] N. Angelescu, M. Bundaru, G. Costache, and G. Nenciu. The critical behavior of Kac-type models for semi-infinite systems. Journal of Statistical Physics, 24(3):529–552, March 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012819.

Angelescu:1986:ERG

[ABCT86] N. Angelescu, M. Bundaru, G. Costache, and C. J. Thompson. Exact results for a generalized classical O(n) matrix spin model. Journal of Statistical Physics, 43(1–2):33–49, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010571.

Abdelazim:1987:SCS

[Abd87] M. S. Abdelazim. Shock compression of simple molecules. Journal of Statistical Physics, 46(3–4):635–658, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01013378.

Andrews:1984:EVS

[ABF84] George E. Andrews, R. J. Baxter, and P. J. Forrester. Eightvertex SOS model and generalized Rogers–Ramanujan-type identities. Journal of Statistical Physics, 35(3–4):193–266, May 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014383.

Artuso:1985:SSH

[ABF^{+}85] R. Artuso, V. Benza, A. Frigerio, V. Gorini, and E. Montaldi. The stationary state and the heat equation for a variant of Davies’ model of heat conduction. Journal of Statistical Physics, 38(5– 6):1051–1070, March 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010429.

Aizenman:1987:PTG

[ABF87] M. Aizenman, D. J. Barsky, and R. Fernández. The phase transition in a general class of Ising-type models is sharp. Journal of Statistical Physics, 47(3–4):343–374, May 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01007515.

Aslangul:1989:ERS

[ABG^{+}89] Claude Aslangul, Jean-Philippe Bouchaud, Antoine Georges, Noëlle Pottier, and Daniel Saint-James. Exact results and self-averaging properties for random-random walks on a onedimensional infinite lattice. Journal of Statistical Physics, 55 (1–2):461–468, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01042612.

Albeverio:1989:IPS

[ABKS89] S. Albeverio, Ph. Blanchard, S. Kusuoka, and L. Streit. An inverse problem for stochastic differential equations. Journal of Statistical Physics, 57(1–2):347–356, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01023648.

Aizenman:1987:PMS

[ABL87] M. Aizenman, J. Bricmont, and J. L. Lebowitz. Percolation of the minority spins in high-dimensional Ising models. Journal of Statistical Physics, 49(3–4):859–865, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009363.

Abraham:1984:SR

[Abr84] D. B. Abraham. Surfaces and roughening. Journal of Statistical Physics, 34(5–6):793–801, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009441.

Abrahams:1985:ILS

[Abr85] Elihu Abrahams. Interplay of localization and superconducting fluctuations above the critical point. Journal of Statistical Physics, 38(1–2):89–95, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01017850.

Achilles:1987:MSH

[ABvT87] M. Achilles, J. Bendisch, and H. von Trotha. Metastable states in homogeneous Ising models. Journal of Statistical Physics, 47 (1–2):257–263, April 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009045.

Alberti:1984:NEE

[AC84] Peter M. Alberti and Bernd Crell. Nonlinear evolution equations and H theorems. Journal of Statistical Physics, 35(1–2):131–149, April 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017370.

Aizenman:1988:DMO

[ACCN88] M. Aizenman, J. T. Chayes, L. Chayes, and C. M. Newman. Discontinuity of the magnetization in one-dimensional 1/|x−y|^{2} Ising and Potts models. Journal of Statistical Physics, 50(1–2):1–40, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01022985.

Alastuey:1988:MFK

[ACH88] Angel Alastuey, Jean Clérouin, and Jean-Pierre Hansen. Mean field kinetic theory of a classical electron gas in a periodic potential. III. The high-temperature limit in two dimensions. Journal of Statistical Physics, 52(1–2):395–421, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01016422.

Alabiso:1989:ERE

[ACS89] C. Alabiso, M. Casartelli, and S. Sello. Equipartition and rate of energy exchanges in a model of a radiant cavity. Journal of Statistical Physics, 54(1–2):361–377, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01023484.

Arneodo:1982:OCB

[ACT82] A. Arneodo, P. Coullet, and C. Tresser. Oscillators with chaotic behavior: An illustration of a theorem by shil’nikov. Journal of Statistical Physics, 27(1):171–182, January 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011745.

Adelman:1986:SCC

[Ade86] Steven A. Adelman. Some concepts in condensed phase chemical kinetics. Journal of Statistical Physics, 42(1–2):37–48, January 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010839.

Ambjorn:1989:RGA

[ADFJ89] Jan Ambjørn, Bergfinnur Durhuus, Jürg Fröhlich, and Thórdur Jónsson. A renormalization group analysis of lattice models of two-dimensional membranes. Journal of Statistical Physics, 55 (1–2):29–85, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01042591.

Aldea:1988:LTA

[ADG88] A. Aldea, M. Dulea, and P. Gartner. Long-time asymptotics in the one-dimensional trapping problem with large bias. Journal of Statistical Physics, 52(3–4):1061–1068, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019739.

Albeverio:1983:RIF

[ADM83] S. Albeverio, D. Dürr, and D. Merlini. Remarks on the independence of the free energy from crystalline boundary conditions in the two-dimensional one-component plasma. Journal of Statistical Physics, 31(2):389–407, May 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011589.

Aizenman:1980:LBC

[ADS80] Michael Aizenman, FranÇois Delyon, and Bernard Souillard. Lower bounds on the cluster size distribution. Journal of Statistical Physics, 23(3):267–280, September 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011369.

Aizenman:1981:SOD

[AF81] Michael Aizenman and Jürg Fröhlich. States of one-dimensional Coulomb systems as simple examples of θ vacua and confinement. Journal of Statistical Physics, 26(2):347–364, October 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013176.

Aizenman:1986:CBM

[AF86a] M. Aizenman and R. Fernández. On the critical behavior of the magnetization in high-dimensional Ising models. Journal of Statistical Physics, 44(3–4):393–454, August 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011304.

Allouche:1986:QIC

[AF86b] Jean-Paul Allouche and Michel Mendes France. Quasicrystal Ising chain and automata theory. Journal of Statistical Physics, 42 (5–6):809–821, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010446.

Alicki:1985:DBC

[AFV85] R. Alicki, M. Fannes, and A. Verbeure. Detailed balance and critical slowing down for classical lattice systems. Journal of Statistical Physics, 41(1–2):263–272, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020612.

Andreev:1985:CSN

[AG85] A. F. Andreev and I. A. Grishchyuk. Crystal shape near the faceting phase transition point. Journal of Statistical Physics, 38 (1–2):183–190, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01017856.

Aharony:1984:MCF

[AGK84] Amnon Aharony, Yuval Gefen, and Yacov Kantor. Magnetic correlations on fractals. Journal of Statistical Physics, 36(5–6): 795–805, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012939.

Aubry:1988:SPS

[AGL88] S. Aubry, C. Godrèche, and J. M. Luck. Scaling properties of a structure intermediate between quasiperiodic and random. Journal of Statistical Physics, 51(5–6):1033–1075, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014899.

Agmon:1986:DRT

[Agm86] Noam Agmon. Diffusion with random traps: Transient onedimensional kinetics in a linear potential. Journal of Statistical Physics, 43(3–4):537–559, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020652.

Agmon:1988:BRF

[Agm88] Noam Agmon. Book review: The fluctuating enzyme. Journal of Statistical Physics, 52(1–2):519–520, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01016434.

Arneodo:1988:FDH

[AH88] A. Arneodo and M. Holschneider. Fractal dimensions and homeomorphic conjugacies. Journal of Statistical Physics, 50(5–6): 995–1020, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019150.

Aharony:1989:FAS

[AH89] Amnon Aharony and A. Brooks Harris. Flory approximant for self-avoiding walks on fractals. Journal of Statistical Physics, 54(3–4):1091–1097, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01019789.

Aharony:1984:PFA

[Aha84] Amnon Aharony. Percolation, fractals, and anomalous diffusion. Journal of Statistical Physics, 34(5–6):931–939, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009449.

Alastuey:1981:ASC

[AJ81] A. Alastuey and B. Jancovici. Absence of strict crystalline order in a two-dimensional electron system. Journal of Statistical Physics, 24(3):443–449, March 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012815.

Alastuey:1984:PFF

[AJ84] A. Alastuey and B. Jancovici. On potential and field fluctuations in two-dimensional classical charged systems. Journal of Statistical Physics, 34(3–4):557–569, February 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01018558.

Anacker:1984:FCK

[AKN84] L. W. Anacker, R. Kopelman, and J. S. Newhouse. Fractal chemical kinetics: Reacting random walkers. Journal of Statistical Physics, 36(5–6):591–602, September 1984. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012924.

Abraham:1989:TMS

[AKS89] D. B. Abraham, L. F. Ko, and N. M. Svrakić. Transfer matrix spectrum for the finite-width Ising model with adjustable boundary conditions: Exact solution. Journal of Statistical Physics, 56(5–6):563–587, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01016767.

Aarts:1988:QAS

[AKvL88] Emile H. L. Aarts, Jan H. M. Korst, and Peter J. M. van Laarhoven. A quantitative analysis of the simulated annealing algorithm: A case study for the traveling salesman problem. Journal of Statistical Physics, 50(1–2):187–206, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01022991.

Aizenman:1981:TLT

[AL81] Michael Aizenman and Elliott H. Lieb. The third law of thermodynamics and the degeneracy of the ground state for lattice systems. Journal of Statistical Physics, 24(1):279–297, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01007649.

Alastuey:1987:MFKa

[Ala87a] Angel Alastuey. Mean-field kinetic theory of a classical electron gas in a periodic potential. I. Formal solution of the Vlasov equation. Journal of Statistical Physics, 48(3–4):839–871, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01019699.

Alastuey:1987:MFKb

[Ala87b] Angel Alastuey. Mean field kinetic theory of a classical electron gas in a periodic potential. II. Qualitative analysis of the meanfield solution in one dimension. Journal of Statistical Physics, 49(3–4):685–724, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009353.

Alankus:1988:GFP

[Ala88] Taylan Alankus. The generating functional for the probability density functions of Navier–Stokes turbulence. Journal of Statistical Physics, 53(5–6):1261–1271, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01023868.

Alankus:1989:ERS

[Ala89] Taylan Alankus. An exact representation of the space-time characteristic functional of turbulent Navier–Stokes flows with prescribed random initial states and driving forces. Journal of Statistical Physics, 54(3–4):859–872, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019778.

Albanese:1989:SHH

[Alb89a] Claudio Albanese. On the spectrum of the Heisenberg Hamiltonian. Journal of Statistical Physics, 55(1–2):297–309, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01042601.

Albanese:1989:SPN

[Alb89b] Claudio Albanese. Scalar particles in a narrow-band periodic potential. Journal of Statistical Physics, 55(1–2):279–295, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01042600.

Alexanian:1985:SES

[Ale85] Moorad Alexanian. Statistical entropy of a Schwarzschild black hole. Journal of Statistical Physics, 41(3–4):709–717, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009029.

Alicki:1983:NQD

[AM83] R. Alicki and J. Messer. Nonlinear quantum dynamical semigroups for many-body open systems. Journal of Statistical Physics, 32(2):299–312, August 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012712.

Anderson:1984:CSA

[AM84] Scott R. Anderson and Gene F. Mazenko. Construction of simple approximants for the structure factor and magnetization for the simple cubic Ising model. Journal of Statistical Physics, 35(1–2):1–18, April 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01017361.

Alastuey:1985:DCC

[AM85] A. Alastuey and Ph. A. Martin. Decay of correlations in classical fluids with long-range forces. Journal of Statistical Physics, 39 (3–4):405–426, May 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01018670.

Adler:1987:BR

[AM87a] Robert J. Adler and Frank Moss. Book reviews. Journal of Statistical Physics, 47(1–2):289–292, April 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009047.

Aizenstadt:1987:SII

[AM87b] V. V. Aizenstadt and V. A. Malyshev. Spin interaction with an ideal Fermi gas. Journal of Statistical Physics, 48(1–2):51–68, July 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010399.

Ahlers:1989:DSE

[AMC89] Guenter Ahlers, Christopher W. Meyer, and David S. Cannell. Deterministic and stochastic effects near the convective onset. Journal of Statistical Physics, 54(5–6):1121–1131, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01044706.

AnglesdAuriac:1982:CDF

[AMR82] J. C. Anglès d’Auriac, R. Maynard, and R. Rammal. Critical dynamics of finite Ising model. Journal of Statistical Physics, 28(2):307–323, June 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012608.

Anderson:1985:GRG

[AMV85] Scott R. Anderson, Gene F. Mazenko, and Oriol T. Valls. Generalized renormalization group treatment of the growth kinetics of unstable systems. Journal of Statistical Physics, 41(1–2):17–36, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020602.

Aizenman:1984:TGI

[AN84] Michael Aizenman and Charles M. Newman. Tree graph inequalities and critical behavior in percolation models. Journal of Statistical Physics, 36(1–2):107–143, July 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01015729.

Altares:1987:NMA

[AN87] V. Altares and G. Nicolis. A new method of analysis of the effect of weak colored noise in nonlinear dynamical systems. Journal of Statistical Physics, 46(1–2):191–205, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010340.

An:1988:NCV

[An88] Guozhong An. A note on the cluster variation method. Journal of Statistical Physics, 52(3–4):727–734, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019726.

Anderson:1986:SUM

[And86] H. L. Anderson. Scientific uses of the MANIAC. Journal of Statistical Physics, 43(5–6):731–748, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02628301.

Aguilera-Navarro:1983:HSF

[ANFdL^{+}83] V. C. Aguilera-Navarro, M. Fortes, M. de Llano, A. Piastino, J. del Río, and O. Rojo. Hard-sphere fluid-to-solid transition and the virial expansion. Journal of Statistical Physics, 32(1):95–104, July 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009422.

Anonymous:1980:FCBa

[Ano80a] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 22(1):121, January 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01007994.

Anonymous:1980:FCBb

[Ano80b] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 22(2):279, February 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008052.

Anonymous:1980:FCBc

[Ano80c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 22(3):407, March 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014650.

Anonymous:1980:FCBd

[Ano80d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 22(4):513, April 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012870.

Anonymous:1980:FCBe

[Ano80e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 22(5):625, May 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011342.

Anonymous:1980:FCBf

[Ano80f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 22(6):745, June 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013940.

Anonymous:1980:FCBg

[Ano80g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 23(1):129, July 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014433.

Anonymous:1980:FCBh

[Ano80h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 23(2):265, August 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012595.

Anonymous:1980:FCBi

[Ano80i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 23(3):403, September 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011374.

Anonymous:1980:FCBj

[Ano80j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 23(4):537, October 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011580.

Anonymous:1980:FCBk

[Ano80k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 23(5):653, November 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011735.

Anonymous:1980:FCBl

[Ano80l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 23(6):783–784, December 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008520.

Anonymous:1980:IA

[Ano80m] Anonymous. Important announcement. Journal of Statistical Physics, 23(6):655, December 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008514.

Anonymous:1980:PSM

[Ano80n] Anonymous. Program of the 42nd Statistical Mechanics Meeting. Journal of Statistical Physics, 22(1):123–129, January 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01007995.

Anonymous:1980:PFA

[Ano80o] Anonymous. Program of the French American seminar on Statistical Mechanics and Quantum Field Theory and the 43rd Statistical Mechanics Meeting. Journal of Statistical Physics, 22(6):747– 754, June 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013941.

Anonymous:1981:A

[Ano81a] Anonymous. Announcements. Journal of Statistical Physics, 26(1):i, September 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01106795.

Anonymous:1981:FCBa

[Ano81b] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 24(1):299, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01007650.

Anonymous:1981:FCBb

[Ano81c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 24(2):403, February 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013309.

Anonymous:1981:FCBc

[Ano81d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 24(3):553, March 1981.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012820.

Anonymous:1981:FCBd

[Ano81e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 24(4):705, April 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011381.

Anonymous:1981:FCBe

[Ano81f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 25(1):181, May 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008485.

Anonymous:1981:FCBf

[Ano81g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 25(2):367–368, June 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01022192.

Anonymous:1981:FCBg

[Ano81h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 25(3):583, July 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010805.

Anonymous:1981:FCBh

[Ano81i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 25(4):775, August 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01022367.

Anonymous:1981:FCBi

[Ano81j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 26(1):219–220, September 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01106794.

Anonymous:1981:FCBj

[Ano81k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 26(2):425–426, October 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013183.

Anonymous:1981:FCBk

[Ano81l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 26(3):633–634, November 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011440.

Anonymous:1981:FCBl

[Ano81m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 26(4):849–850, December 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010947.

Anonymous:1981:IA

[Ano81n] Anonymous. Important announcement. Journal of Statistical Physics, 24(1):1, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01007628.

Anonymous:1981:PSM

[Ano81o] Anonymous. Program of the 44th Statistical Mechanics Meeting and the 19th Eastern Theoretical Physics Conference Department of Mathematics and Physics Rutgers University. Journal of Statistical Physics, 24(4):707–712, April 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011382.

Anonymous:1981:PSW

[Ano81p] Anonymous. Program of the Seventh West Coast Statistical Mechanics Conference. Journal of Statistical Physics, 26(3):631–632, November 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011439.

Anonymous:1982:FCBa

[Ano82a] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 27(1):221, January 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011749.

Anonymous:1982:FCBb

[Ano82b] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 27(2):423, February 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008948.

Anonymous:1982:FCBc

[Ano82c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 27(3):629–630, March 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011095.

Anonymous:1982:FCBd

[Ano82d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 27(4):841–842, April 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013455.

Anonymous:1982:FCBe

[Ano82e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 28(1):233–234, May 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011636.

Anonymous:1982:FCBf

[Ano82f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 28(2):411–412, June 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012616.

Anonymous:1982:FCBg

[Ano82g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 28(3):619–620, July 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008331.

Anonymous:1982:FCBh

[Ano82h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 28(4):843, August 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011885.

Anonymous:1982:FCBi

[Ano82i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 29(1):157–158, September 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008256.

Anonymous:1982:FCBj

[Ano82j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 29(2):399–400, October 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020796.

Anonymous:1982:FCBk

[Ano82k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 29(3):647–648, November 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01342196.

Anonymous:1982:FCB

[Ano82l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 29(4):885–886, December 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011799.

Anonymous:1982:MA

[Ano82m] Anonymous. Meeting announcements. Journal of Statistical Physics, 29(2):397, October 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020795.

Anonymous:1982:PSMa

[Ano82n] Anonymous. Program of the 45th Statistical Mechanics Meeting. Journal of Statistical Physics, 27(4):829–833, April 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013453.

Anonymous:1982:PSMb

[Ano82o] Anonymous. Program of the 46th Statistical Mechanics Meeting. Journal of Statistical Physics, 27(4):835–839, April 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013454.

Anonymous:1982:PSMc

[Ano82p] Anonymous. Program of the 47th Statistical Mechanics Meeting. Journal of Statistical Physics, 28(4):837–842, August 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011884.

Anonymous:1982:PEW

[Ano82q] Anonymous. Program of the Eighth West Coast Statistical Mechanics Conference. Journal of Statistical Physics, 29(4):879–881, December 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011797.

Anonymous:1983:A

[Ano83a] Anonymous. Announcements. Journal of Statistical Physics, 33(1):i, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009762.

Anonymous:1983:CA

[Ano83b] Anonymous. Conference announcements. Journal of Statistical Physics, 31(2):429, May 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011593.

Anonymous:1983:FCBa

[Ano83c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 30(1):245–247, January 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010878.

Anonymous:1983:FCBb

[Ano83d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 30(2):561–562, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012329.

Anonymous:1983:FCBc

[Ano83e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 30(3):793, March 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009689.

Anonymous:1983:FCBd

[Ano83f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 31(1):229–230, April 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010930.

Anonymous:1983:FCBe

[Ano83g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 31(2):431–432, May 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011594.

Anonymous:1983:FCBf

[Ano83h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 31(3):713–714, June 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01019508.

Anonymous:1983:FCBg

[Ano83i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 32(1):207–208, July 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009431.

Anonymous:1983:FCBh

[Ano83j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 32(2):441, August 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012721.

Anonymous:1983:FCBi

[Ano83k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 32(3):657–658, September 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008963.

Anonymous:1983:FCBj

[Ano83l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 33(1):239–240, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009761.

Anonymous:1983:FCBk

[Ano83m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 33(2):491–492, November 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009809.

Anonymous:1983:FCBl

[Ano83n] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 33(3):759–760, December 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01018843.

Anonymous:1983:I

[Ano83o] Anonymous. Introduction. Journal of Statistical Physics, 30(2): 249–250, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012298.

Anonymous:1983:MAa

[Ano83p] Anonymous. Meeting announcements. Journal of Statistical Physics, 30(3):785, March 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009687.

Anonymous:1983:MAb

[Ano83q] Anonymous. Meeting announcements. Journal of Statistical Physics, 32(3):649, September 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008961.

Anonymous:1983:P

[Ano83r] Anonymous. Program. Journal of Statistical Physics, 30(2): 251–253, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012299.

Anonymous:1983:PSMa

[Ano83s] Anonymous. Program of the 48th Statistical Mechanics Meeting. Journal of Statistical Physics, 30(3):787–792, March 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009688.

Anonymous:1983:PSMb

[Ano83t] Anonymous. Program of the 49th Statistical Mechanics Meeting. Journal of Statistical Physics, 32(3):651–656, September 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008962.

Anonymous:1984:A

[Ano84a] Anonymous. Announcement. Journal of Statistical Physics, 34 (5–6):I, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01009461;   http://link.
springer.com/article/10.1007/BF01009461.

Anonymous:1984:C

[Ano84b] Anonymous. Committees. Journal of Statistical Physics, 34(5–6): 663, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01009434;   http://link.
springer.com/article/10.1007/BF01009434.

Anonymous:1984:D

[Ano84c] Anonymous. Dedication. Journal of Statistical Physics, 36(5–6): ii, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01012916;   http://link.
springer.com/article/10.1007/BF01012916.

Anonymous:1984:E

[Ano84d] Anonymous. Erratum. Journal of Statistical Physics, 34(1– 2):345, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01770364.

Anonymous:1984:FCBa

[Ano84e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 34(1–2):359–360, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01770369.

Anonymous:1984:FCBb

[Ano84f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 34(3–4):657–658, February 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01018565.

Anonymous:1984:FCBc

[Ano84g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 34(5–6):1075–1076, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009460.

Anonymous:1984:FCBd

[Ano84h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 35(1–2):191–192, April 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017375.

Anonymous:1984:FCBe

[Ano84i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 35(3–4):471–472, May 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014397.

Anonymous:1984:FCBf

[Ano84j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 35(5–6):775–776, June 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010835.

Anonymous:1984:FCBg

[Ano84k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 36(1–2):291, July 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01015737.

Anonymous:1984:FCBh

[Ano84l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 36(3–4):517–518, August 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010994.

Anonymous:1984:FCBi

[Ano84m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 36(5–6):923–924, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01012952;   http://link.
springer.com/article/10.1007/BF01012952.

Anonymous:1984:FCBj

[Ano84n] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 37(3–4):495–496, November 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011848.

Anonymous:1984:FCBk

[Ano84o] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 37(5–6):715–716, December 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010506.

Anonymous:1984:LP

[Ano84p] Anonymous. List of participants. Journal of Statistical Physics, 34(5–6):1065–1074, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009459.

Anonymous:1984:MA

[Ano84q] Anonymous. Meeting announcements. Journal of Statistical Physics, 34(3–4):i, February 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01018566.

Anonymous:1984:Pc

[Ano84r] Anonymous. Posters. Journal of Statistical Physics, 36(5–6):921, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01012951;   http://link.
springer.com/article/10.1007/BF01012951.

Anonymous:1984:Pb

[Ano84s] Anonymous. Preface. Journal of Statistical Physics, 36(5–6):iii, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01012917;   http://link.
springer.com/article/10.1007/BF01012917.

Anonymous:1984:Pa

[Ano84t] Anonymous. Program. Journal of Statistical Physics, 34(1–2): 353–355, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01770367.

Anonymous:1984:PSM

[Ano84u] Anonymous. Program of the 50th Statistical Mechanics Meeting. Journal of Statistical Physics, 35(5–6):767–773, June 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010834.

Anonymous:1984:PTW

[Ano84v] Anonymous. Program of the tenth West Coast statistical meeting. Journal of Statistical Physics, 37(5–6):713–714, December 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010505.

Anonymous:1984:SP

[Ano84w] Anonymous. Scientific program. Journal of Statistical Physics, 36(5–6):917–919, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012950.

Anonymous:1985:EWC

[Ano85a] Anonymous. Eleventh West Coast Statistical Mechanics Conference. Journal of Statistical Physics, 41(5–6):1047–1049, December 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010016.

Anonymous:1985:FCBa

[Ano85b] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 38(1–2):433–434, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017872.

Anonymous:1985:FCBb

[Ano85c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 38(3–4):807–808, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010492.

Anonymous:1985:FCBc

[Ano85d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 38(5–6):1103–1104, March 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010432.

Anonymous:1985:FCBd

[Ano85e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 39(1–2):265, April 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01007984.

Anonymous:1985:FCBe

[Ano85f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 39(3–4):451–452, May 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01018675.

Anonymous:1985:FCBf

[Ano85g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 39(5–6):817–818, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01008367;   http://link.
springer.com/article/10.1007/BF01008367.

Anonymous:1985:FCBg

[Ano85h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 40(1–2):369, July 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010542.

Anonymous:1985:FCBh

[Ano85i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 40(3–4):611, August 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017190.

Anonymous:1985:FCBi

[Ano85j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 40(5–6):803–804, September 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009904.

Anonymous:1985:FCBj

[Ano85k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 41(1–2):351–352, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020620.

Anonymous:1985:FCBk

[Ano85l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 41(3–4):731–732, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009032.

Anonymous:1985:FCBl

[Ano85m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 41(5–6):1157, December 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010017.

Anonymous:1985:IAa

[Ano85n] Anonymous. Important announcement. Journal of Statistical Physics, 41(3–4):i, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009012.

Anonymous:1985:IAb

[Ano85o] Anonymous. Important announcement. Journal of Statistical Physics, 41(5–6):i, December 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010000.

Anonymous:1985:LSW

[Ano85p] Anonymous. List of selected works of I. M. Lifshitz. Journal of Statistical Physics, 38(1–2):11–24, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01017845.

Anonymous:1985:PSMa

[Ano85q] Anonymous. Program of the 51st Statistical Mechanics Meeting. Journal of Statistical Physics, 40(5–6):789–794, September 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009902.

Anonymous:1985:PSMb

[Ano85r] Anonymous. Program of the 52nd Statistical Mechanics Meeting. Journal of Statistical Physics, 40(5–6):795–802, September 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009903.

Anonymous:1985:SMM

[Ano85s] Anonymous. Statistical Mechanics Meeting Mathematics Department, University of Melbourne. Journal of Statistical Physics, 41 (3–4):729, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009031.

Anonymous:1986:A

[Ano86a] Anonymous. Announcement. Journal of Statistical Physics, 42 (5–6):1139, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010469.

Anonymous:1986:FCBl

[Ano86b] Anonymous. Future contribution to Journal of Statistical Physics. Journal of Statistical Physics, 45(5–6):1123–1124, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020595.

Anonymous:1986:FCBa

[Ano86c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 42(1–2):245–246, January 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010848.

Anonymous:1986:FCBb

[Ano86d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 42(3–4):725–726, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01127737.

Anonymous:1986:FCBc

[Ano86e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 42(5–6):1141–1142, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010470.

Anonymous:1986:FCBd

[Ano86f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 43(1–2):393–394, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010590.

Anonymous:1986:FCBe

[Ano86g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 43(3–4):727–728, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020665.

Anonymous:1986:FCBf

[Ano86h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 43(5–6):1243–1244, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02628343.

Anonymous:1986:FCBg

[Ano86i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 44(1–2):279–280, July 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010920.

Anonymous:1986:FCBh

[Ano86j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 44(3–4):711–712, August 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011318.

Anonymous:1986:FCBi

[Ano86k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 44(5–6):1069– 1070, September 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011922.

Anonymous:1986:FCBj

[Ano86l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 45(1–2):357–358, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01033097.

Anonymous:1986:FCBk

[Ano86m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 45(3–4):775–776, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01021097.

Anonymous:1986:IA

[Ano86n] Anonymous. Important announcement. Journal of Statistical Physics, 42(1–2):i, January 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010836.

Anonymous:1986:I

[Ano86o] Anonymous. Introduction. Journal of Statistical Physics, 42 (1–2):1–2, January 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010837.

Anonymous:1986:LP

[Ano86p] Anonymous. List of participants. Journal of Statistical Physics, 43(5–6):1239–1242, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02628342.

Anonymous:1986:TWC

[Ano86q] Anonymous. Twelfth West Coast Statistical Mechanics Conference. Journal of Statistical Physics, 45(5–6):1121–1122, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020594.

Anonymous:1987:WCS

[Ano87a] Anonymous. 13th West Coast Statistical Mechanics Conference. Journal of Statistical Physics, 49(3–4):875–876, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009368.

Anonymous:1987:A

[Ano87b] Anonymous. Announcement. Journal of Statistical Physics, 46 (3–4):797, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013388.

Anonymous:1987:FCBb

[Ano87c] Anonymous. Future contribution to Journal of Statistical Physics. Journal of Statistical Physics, 46(3–4):799–800, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013389.

Anonymous:1987:FCBc

[Ano87d] Anonymous. Future contribution to Journal of Statistical Physics. Journal of Statistical Physics, 46(5–6):1281–1282, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011164.

Anonymous:1987:FCBe

[Ano87e] Anonymous. Future contribution to Journal of Statistical Physics. Journal of Statistical Physics, 47(3–4):605–606, May 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01007530.

Anonymous:1987:FCBf

[Ano87f] Anonymous. Future contribution to Journal of Statistical Physics. Journal of Statistical Physics, 47(5–6):987–988, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01206174.

Anonymous:1987:FCBl

[Ano87g] Anonymous. Future contribution to Journal of Statistical Physics. Journal of Statistical Physics, 49(5–6):1283–1284, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017574.

Anonymous:1987:FCBa

[Ano87h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 46(1–2):433–434, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010354.

Anonymous:1987:FCBd

[Ano87i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 47(1–2):295–296, April 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009049.

Anonymous:1987:FCBg

[Ano87j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 48(1–2):363–364, July 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010415.

Anonymous:1987:FCBh

[Ano87k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 48(3–4):955–956, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01019708.

Anonymous:1987:FCBi

[Ano87l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 48(5–6):1373– 1374, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009551.

Anonymous:1987:FCBj

[Ano87m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 49(1–2):409–410, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009973.

Anonymous:1987:FCBk

[Ano87n] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 49(3–4):877–878, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009369.

Anonymous:1987:PAS

[Ano87o] Anonymous. Program of Australian Statistical Mechanics Meeting Australian National University Canberra. Journal of Statistical Physics, 46(3–4):795–796, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013387.

Anonymous:1987:PSM

[Ano87p] Anonymous. Programs of the 56th and 57th statistical mechanics meetings. Journal of Statistical Physics, 49(1–2):395–407, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009972.

Anonymous:1987:TAT

[Ano87q] Anonymous. Titles of additional talks given at the Trebon symposium. Journal of Statistical Physics, 47(5–6):985–986, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01206173.

Anonymous:1988:IIC

[Ano88a] Anonymous. 17th IUPAP international conference on thermodynamics and statistical mechanics. Journal of Statistical Physics, 53(3–4):1015, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014241.

Anonymous:1988:FWC

[Ano88b] Anonymous. Fourteenth West Coast Statistical Mechanics Conference. Journal of Statistical Physics, 53(3–4):1011–1012, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014239.

Anonymous:1988:FCBb

[Ano88c] Anonymous. Future contribution to Journal of Statistical Physics. Journal of Statistical Physics, 50(3–4):851–852, February 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01026504.

Anonymous:1988:FCBc

[Ano88d] Anonymous. Future contribution to Journal of Statistical Physics. Journal of Statistical Physics, 50(5–6):1291–1292, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01019173.

Anonymous:1988:FCBa

[Ano88e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 50(1–2):473–474, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023008.

Anonymous:1988:FCBd

[Ano88f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 51(1–2):325–326, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01015337.

Anonymous:1988:FCBe

[Ano88g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 51(3–4):739–740, May 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01028482.

Anonymous:1988:FCBf

[Ano88h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 51(5–6):1127–1128, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014903.

Anonymous:1988:FCBg

[Ano88i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 52(1–2):525–526, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016436.

Anonymous:1988:FCBh

[Ano88j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 52(3–4):1139–1140, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019748.

Anonymous:1988:FCBi

[Ano88k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 52(5–6):1479– 1480, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011661.

Anonymous:1988:FCBj

[Ano88l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 53(1–2):549–550, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011573.

Anonymous:1988:FCBk

[Ano88m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 53(3–4):1017–1018, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014242.

Anonymous:1988:FCBl

[Ano88n] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 53(5–6):1345–1346, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023875.

Anonymous:1988:PFM

[Ano88o] Anonymous. Program of the First Meeting on the ‘Statistical Mechanics at the 45th parallel” 17 October 1987, Clarkson University, Potsdam, New York. Journal of Statistical Physics, 50(5– 6):1289–1290, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019172.

Anonymous:1988:SMM

[Ano88p] Anonymous. Statistical mechanics meeting. Journal of Statistical Physics, 53(3–4):1013, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014240.

Anonymous:1989:A

[Ano89a] Anonymous. Announcement. Journal of Statistical Physics, 54 (1–2):561, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023495.

Anonymous:1989:FWC

[Ano89b] Anonymous. Fifteenth West Coast Statistical Mechanics Conference. Journal of Statistical Physics, 57(5–6):1141–1142, December 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020055.

Anonymous:1989:FCBa

[Ano89c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 54(1–2):565–566, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023497.

Anonymous:1989:FCBb

[Ano89d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 54(3–4):1109–1110, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019793.

Anonymous:1989:FCBc

[Ano89e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 54(5–6):1539–1540, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01044734.

Anonymous:1989:FCBd

[Ano89f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 55(1–2):473–474, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01042616.

Anonymous:1989:FCBe

[Ano89g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 55(3–4):857–858, May 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01041609.

Anonymous:1989:FCBf

[Ano89h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 55(5–6):1371–1372, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01041098.

Anonymous:1989:FCBg

[Ano89i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 56(1–2):245–246, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01044245.

Anonymous:1989:FCBh

[Ano89j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 56(3–4):561–562, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01044456.

Anonymous:1989:FCBi

[Ano89k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 56(5–6):985–986, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01016793.

Anonymous:1989:FCBj

[Ano89l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 57(1–2):429–430, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023657.

Anonymous:1989:FCBk

[Ano89m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 57(3–4):965, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01022843.

Anonymous:1989:FCBl

[Ano89n] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 57(5–6):1165–1166, December 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020056.

Anonymous:1989:SMP

[Ano89o] Anonymous. Statistical mechanics on the 45th parallel II. Journal of Statistical Physics, 56(1–2):243–244, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01044244.

Aoyama:1987:EPR

[AO87] Hideaki Aoyama and Takashi Odagaki. Eight-parameter renormalization group for Penrose lattices. Journal of Statistical Physics, 48(3–4):503–511, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01019684.

Axel:1989:SES

[AP89] Françoise Axel and Jacques Peyrière. Spectrum and extended states in a harmonic chain with controlled disorder: Effects of the Thue–Morse symmetry. Journal of Statistical Physics, 57(5–6): 1013–1047, December 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020046.

Aslangul:1985:TBC

[APSJ85] C. Aslangul, N. Pottier, and D. Saint-James. Time behavior of the correlation functions in a simple dissipative quantum model. Journal of Statistical Physics, 40(1–2):167–189, July 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010531.

Arkeryd:1983:ESH

[Ark83] Leif Arkeryd. L^{∞} estimates for the space-homogeneous Boltzmann equation. Journal of Statistical Physics, 31(2):347–361, May 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011586.

Avron:1981:GSD

[ARS81] J. E. Avron, G. Roepstorff, and L. S. Schulman. Ground state degeneracy and ferromagnetism in a spin glass. Journal of Statistical Physics, 26(1):25–36, September 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01106783.

Abraham:1986:ESM

[AS86] D. B. Abraham and E. R. Smith. An exactly solved model with a wetting transition. Journal of Statistical Physics, 43(3–4):621– 643, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020656.

Akuezue:1989:RAR

[AS89] H. C. Akuezue and J. Stringer. Random aggregation and randomwalking center of mass. Journal of Statistical Physics, 56(3– 4):461–470, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01044446.

Avron:1983:ESC

[ATZ83] J. E. Avron, J. E. Taylor, and R. K. P. Zia. Equilibrium shapes of crystals in a gravitational field: Crystals on a table. Journal of Statistical Physics, 33(3):493–522, December 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01018830.

Aurell:1987:MPFa

[Aur87a] Erik Aurell. On the metric properties of the Feigenbaum attractor. Journal of Statistical Physics, 47(3–4):439–458, May 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01007519.

Aurell:1987:MPFb

[Aur87b] Erik Aurell. On the metric properties of the Feigenbaum attractor. Journal of Statistical Physics, 49(5–6):1261, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017573.

Anshelevich:1981:LOR

[AV81] V. V. Anshelevich and A. V. Vologodskii. Laplace operator and random walk on one-dimensional nonhomogeneous lattice. Journal of Statistical Physics, 25(3):419–430, July 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010797.

Andjel:1987:HEA

[AV87] Enrique Daniel Andjel and Maria Eulália Vares. Hydrodynamic equations for attractive particle systems on Z. Journal of Statistical Physics, 47(1–2):265–288, April 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01009046. See cor-

rection [AV03].

Andjel:2003:CHE

[AV03] Enrique Daniel Andjel and Maria Eulália Vares. Correction to ‘Hydrodynamic Equations for Attractive Particle Systems on Z,’ J. Stat. Phys. 47:265 (1987). Journal of Statistical Physics, 113 (1–2):379–380, October 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1025739208508. See [AV87].

Angelescu:1985:GQA

[AZ85] N. Angelescu and V. A. Zagrebnov. A generalized quasiaverage approach to the description of the limit states of the n-vector Curie–Weiss ferromagnet. Journal of Statistical Physics, 41(1– 2):323–334, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020617.

Baxter:1986:LGG

[BA86] R. J. Baxter and George E. Andrews. Lattice gas generalization of the hard hexagon model. I. Star–triangle relation and local densities. Journal of Statistical Physics, 44(1–2):249–271, July 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010916.

Ben-Avraham:1987:DFT

[BA87] Daniel Ben-Avraham. Discrete fluctuations and their influence on kinetics of reactions. Journal of Statistical Physics, 48(1–2):315– 328, July 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010411.

ben-Avraham:1988:BRB

[bA88a] Daniel ben Avraham. Book review: Introduction to path-integral methods in physics and polymer science. Journal of Statistical Physics, 52(3–4):1135–1136, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019746.

ben-Avraham:1988:BRS

[bA88b] Daniel ben Avraham. Book review: Scaling phenomena in disordered systems. Journal of Statistical Physics, 53(3–4):1005, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014237.

Babovsky:1986:KFW

[Bab86] Hans Babovsky. On Knudsen flows within thin tubes. Journal of Statistical Physics, 44(5–6):865–878, September 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011911.

Baker:1982:BRM

[Bak82] George A. Baker, Jr. Book review: Moment methods in manyfermion systems. Journal of Statistical Physics, 27(1):219–220, January 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011748.

Balescu:1987:NTT

[Bal87] R. Balescu. Nonequilibrium thermodynamics and the transport phenomena in magnetically confined plasmas. Journal of Statistical Physics, 48(5–6):1091–1107, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009535.

Banach:1987:FET

[Ban87] Zbigniew Banach. On the fundamentals of extended thermodynamics (ET) of a one-dimensional rarefied gas. Journal of Statistical Physics, 48(3–4):813–837, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019698.

Barakat:1985:BRR

[Bar85] Richard Barakat. Book review: Random walks and their applications in the physical and biological sciences. Journal of Statistical Physics, 40(1–2):365–368, July 1985. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010541.

Barker:1988:HPE

[Bar88a] J. A. Barker. High-pressure equation of state for solid krypton from interatomic potentials. Journal of Statistical Physics, 52(5–6):1359–1365, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011653.

Barrachina:1988:WSN

[Bar88b] R. O. Barrachina. Wild’s solution of the nonlinear Boltzmann equation. Journal of Statistical Physics, 52(1–2):357–368, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016419.

Bartelt:1989:BRP

[Bar89] Maria C. T. P. C. Bartelt. Book review: Phase transitions and critical phenomena. Journal of Statistical Physics, 55(5– 6):1331–1332, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01041095.

ben-Avraham:1989:RWR

[bARC89] D. ben Avraham, S. Redner, and Z. Cheng. Random walk in a random multiplicative environment. Journal of Statistical Physics, 56(3–4):437–459, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01044445.

Baus:1980:ALR

[Bau80] Marc Baus. Absence of long-range order with long-range potentials. Journal of Statistical Physics, 22(1):111–119, January 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01007993.

Baumgartner:1983:GIN

[Bau83] Bernhard Baumgartner. Griffiths inequalities for noninteracting n-vector (classical Heisenberg) models and applications to interacting systems. Journal of Statistical Physics, 32(3):615– 625, September 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008959.

Baumgardner:1985:TDT

[Bau85] John R. Baumgardner. Three-dimensional treatment of convective flow in the earth’s mantle. Journal of Statistical Physics, 39 (5–6):501–511, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008348.

Baus:1987:SMT

[Bau87] M. Baus. Statistical mechanical theories of freezing: An overview. Journal of Statistical Physics, 48(5–6):1129–1146, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009537.

Bavaud:1986:WTE

[Bav86] F. Bavaud. The wall theorem for elastic moduli. Journal of Statistical Physics, 45(1–2):171–181, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01033085.

Bavaud:1987:SMV

[Bav87] François Bavaud. Statistical mechanics of viscoelasticity. Journal of Statistical Physics, 46(3–4):753–775, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01013384.

Bavaud:1989:SMC

[Bav89] F. Bavaud. Statistical mechanics of convex bodies. Journal of Statistical Physics, 57(5–6):1059–1068, December 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020048.

Baxter:1981:RRI

[Bax81] R. J. Baxter. Rogers–Ramanujan identities in the hard hexagon model. Journal of Statistical Physics, 26(3):427–452, November 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011427.

Baxter:1982:IRM

[Bax82] R. J. Baxter. The inversion relation method for some twodimensional exactly solved models in lattice statistics. Journal of Statistical Physics, 28(1):1–41, May 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011621.

Baxter:1988:FES

[Bax88] R. J. Baxter. Free energy of the solvable chiral Potts model. Journal of Statistical Physics, 52(3–4):639–667, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019722.

Baxter:1989:SCP

[Bax89] R. J. Baxter. Superintegrable chiral Potts model: Thermodynamic properties, an ‘inverse’ model, and a simple associated Hamiltonian. Journal of Statistical Physics, 57(1–2):1–39, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023632.

Barber:1980:BFE

[BB80] Michael N. Barber and Peter T. Binns. Bounds for free energy derivatives using renormalization group methods. Journal of Statistical Physics, 22(1):91–96, January 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01007991.

Baker:1982:FIM

[BB82] George A. Baker, Jr. and L. P. Benofy. On the ferromagnetic Ising model in noninteger spatial dimension. Journal of Statistical Physics, 29(4):699–716, December 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011786.

Brender:1983:SMC

[BBAH83] C. Brender, D. Ben-Avraham, and S. Havlin. Study of Monte Carlo methods for generating self-avoiding walks. Journal of Statistical Physics, 31(3):661–670, June 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019503.

Bidaux:1986:TSI

[BBF86] R. Bidaux, N. Boccara, and G. Forgàcs. Three-spin interaction Ising model with a nondegenerate ground state at zero applied field. Journal of Statistical Physics, 45(1–2):113–134, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01033081.

Bennett:1981:KFO

[BBLT81] C. H. Bennett, M. Büttiker, R. Landauer, and H. Thomas. Kinematics of the forced and overdamped sine-Gordon soliton gas. Journal of Statistical Physics, 24(3):419–442, March 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012814.

Batchelor:1987:BAC

[BBP87] Murray T. Batchelor, Michael N. Barber, and Paul A. Pearce. Bethe ansatz calculations for the eight-vertex model on a finite strip. Journal of Statistical Physics, 49(5–6):1117–1163, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017563.

Boldrighini:1983:BEL

[BBS83] C. Boldrighini, L. A. Bunimovich, and Ya. G. Sinai. On the Boltzmann equation for the Lorentz gas. Journal of Statistical Physics, 32(3):477–501, September 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008951.

Bonato:1989:ASB

[BC89a] Cezar A. Bonato and Massimo Campanino. Absence of symmetry breaking for systems of rotors with random interactions. Journal of Statistical Physics, 54(1–2):81–88, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01023474.

Bonilla:1989:DSS

[BC89b] Luis L. Bonilla and José M. Casado. Dynamics of a soft-spin van Hemmen model. I. Phase and bifurcation diagrams for stationary distributions. Journal of Statistical Physics, 56(1–2):113– 125, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01044235.

Bramson:1986:MSP

[BCD^{+}86] M. Bramson, P. Calderoni, A. De Masi, P. Ferrari, J. Lebowitz, and R. H. Schonmann. Microscopic selection principle for a diffusion-reaction equation. Journal of Statistical Physics, 45(5– 6):905–920, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020581.

Blanchard:1989:PBQ

[BCD89] Ph. Blanchard, E. A. Carlen, and G. F. Dell’Antonio. Particles and ‘bumps’ in quantum field configurations. Journal of Statistical Physics, 55(3–4):769–786, May 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01041607.

Bavaud:1986:SME

[BCF86] F. Bavaud, Ph. Choquard, and J.-R. Fontaine. Statistical mechanics of elastic moduli. Journal of Statistical Physics, 42(3– 4):621–646, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01127732.

Boldrighini:1989:CSS

[BCFN89] C. Boldrighini, G. Cosimi, S. Frigio, and M. Grasso Nunes. Computer simulation of shock waves in the completely asymmetric simple exclusion process. Journal of Statistical Physics, 55 (3–4):611–623, May 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01041600.

Bender:1980:SCE

[BCG^{+}80] Carl M. Bender, Fred Cooper, Gerald Guralnik, Harvey A. Rose, and David H. Sharp. Strong coupling expansion for classical statistical dynamics. Journal of Statistical Physics, 22(6):647–660, June 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013934.

Bonilla:1987:SSP

[BCM87] Luis L. Bonilla, José M. Casado, and Manuel Morillo. Selfsynchronization of populations of nonlinear oscillators in the thermodynamic limit. Journal of Statistical Physics, 48(3–4):571–591, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/

10.1007/BF01019689. See corrigendum [BCM88].

Bonilla:1988:CSS

[BCM88] L. L. Bonilla, J. M. Casado, and M. Morillo. Corrigendum to Selfsynchronization of populations of nonlinear oscillators in the thermodynamic limit. Journal of Statistical Physics, 50(3–4):849–850, February 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026503. See [BCM87].

Birgeneau:1984:PTD

[BCSY84] R. J. Birgeneau, R. A. Cowley, G. Shirane, and H. Yoshizawa. Phase transitions in diluted magnets: Critical behavior, percolation, and random fields. Journal of Statistical Physics, 34(5– 6):817–848, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009443.

Baras:1986:CAP

[BCT86] F. Baras, P. H. Coullet, and E. Tirapegui. Comment on asymptotic properties of coupled Langevin equations. Journal of Statistical Physics, 45(3–4):745–752, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01021093.

Barber:1982:HST

[BD82] Michael N. Barber and Phillip M. Duxbury. Hamiltonian studies of the two-dimensional axial next-nearest-neighbor Ising (ANNNI) model. Journal of Statistical Physics, 29(3):427–432, November 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01342182.

Brazovsky:1985:ESM

[BD85] S. Brazovsky and I. Dzyaloshinsky. Exactly solvable XY model of the spin Peierls transition. Journal of Statistical Physics, 38 (1–2):115–124, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01017852.

Barber:1988:DPT

[BD88] Michael N. Barber and Bernard Derrida. Dynamical phase transitions in the two-dimensional ANNNI model. Journal of Statistical Physics, 51(5–6):877–891, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014890.

Brand:1989:ENI

[BDE89] Helmut R. Brand, Charles R. Doering, and Robert E. Ecke. External noise and its interaction with spatial degrees of freedom in nonlinear dissipative systems. Journal of Statistical Physics, 54 (5–6):1111–1119, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01044705.

Bendler:1988:FPT

[BDKW88] John T. Bendler, Menachem Dishon, Harry Kesten, and George H. Weiss. A first passage time problem for random walk occupancy. Journal of Statistical Physics, 50(5–6):1069–1087, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019153.

Bishop:1981:MDS

[BDL81] Marvin Bishop, Marie Derosa, and Joanne Lalli. Molecular dynamics simulations of one-dimensional Lennard-Jones systems. Journal of Statistical Physics, 25(2):229–235, June 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01022184.

Boldrighini:1983:ODH

[BDS83] C. Boldrighini, R. L. Dobrushin, and Yu. M. Sukhov. Onedimensional hard rod caricature of hydrodynamics. Journal of Statistical Physics, 31(3):577–616, June 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019499.

Bhanot:1986:FAC

[BDS86] Gyan Bhanot, Dennis Duke, and Román Salvador. A fast algorithm for the cyber 205 to simulate the 3D Ising model. Journal of Statistical Physics, 44(5–6):985–1002, September 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011918.

Beetem:1986:G

[BDW86] J. Beetem, M. Denneau, and D. Weingarten. GF11. Journal of Statistical Physics, 43(5–6):1171–1183, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02628339.

Borckmans:1987:STS

[BDWK87] P. Borckmans, G. Dewel, D. Walgraef, and Y. Katayama. The search for Turing structures. Journal of Statistical Physics, 48 (5–6):1031–1044, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009531.

Bricmont:1986:RSS

[BEF86] J. Bricmont, A. El Mellouki, and J. Fröhlich. Random surfaces in statistical mechanics: Roughening, rounding, wetting, . . .. Journal of Statistical Physics, 42(5–6):743–798, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010444.

Benioff:1980:CPS

[Ben80] Paul Benioff. The computer as a physical system: A microscopic quantum mechanical Hamiltonian model of computers as represented by Turing machines. Journal of Statistical Physics, 22(5):563–591, May 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011339.

Benioff:1982:QMH

[Ben82] Paul Benioff. Quantum mechanical Hamiltonian models of Turing machines. Journal of Statistical Physics, 29(3):515–546, November 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01342185.

Bendler:1984:LSP

[Ben84] John T. Bendler. Levy (stable) probability densities and mechanical relaxation in solid polymers. Journal of Statistical Physics, 36(5–6):625–637, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012927.

Bendler:1985:BRM

[Ben85a] John T. Bendler. Book review: Molecular diffusion and spectra. Journal of Statistical Physics, 40(3–4):607–608, August 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017188.

Benfatto:1985:IME

[Ben85b] G. Benfatto. An iterated Mayer expansion for the Yukawa gas. Journal of Statistical Physics, 41(3–4):671–684, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009027.

Bers:1983:BRB

[Ber83] Helene Cooper Bers. Book review: Real-Space Renormalization. Journal of Statistical Physics, 33(2):487–490, November 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009808.

Berretti:1985:SPR

[Ber85] Alberto Berretti. Some properties of random Ising models. Journal of Statistical Physics, 38(3–4):483–496, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010473.

Berne:1986:PIM

[Ber86] Bruce J. Berne. Path integral Monte Carlo methods: Staticand time-correlation functions. Journal of Statistical Physics, 43 (5–6):911–929, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02628319.

Baxter:1980:HSL

[BET80] R. J. Baxter, I. G. Enting, and S. K. Tsang. Hard-square lattice gas. Journal of Statistical Physics, 22(4):465–489, April 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012867.

Baive:1981:SBM

[BF81a] Daniela Baivé and Valter Franceschini. Symmetry breaking on a model of five-mode truncated Navier–Stokes equations. Journal of Statistical Physics, 26(3):471–484, November 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011429.

Bricmont:1981:CIC

[BF81b] Jean Bricmont and Jean-Raymond Fontaine. Correlation inequalities and contour estimates. Journal of Statistical Physics, 26(4): 745–753, December 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010936.

Bovier:1986:HTS

[BF86] Anton Bovier and Jürg Fröhlich. A heuristic theory of the spin glass phase. Journal of Statistical Physics, 44(3–4):347–391, August 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011303.

Braun:1988:MCS

[BF88] P. Braun and M. Fähnle. Monte Carlo study of the critical behavior of pure and site-diluted Ising ferro-and ferrimagnets. Journal of Statistical Physics, 52(3–4):775–794, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019729.

Baskaran:1986:SMT

[BFA86] G. Baskaran, Yaotian Fu, and P. W. Anderson. On the statistical mechanics of the traveling salesman problem. Journal of Statistical Physics, 45(1–2):1–25, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01033073.

Barrachina:1986:TTB

[BFG86] R. O. Barrachina, D. H. Fujii, and C. R. Garibotti. Temperature transform of the Boltzmann equation. Journal of Statistical Physics, 45(3–4):647–668, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01021089.

Bricmont:1982:STP

[BFL82] Jean Bricmont, Jean-Raymond Fontaine, and Joel L. Lebowitz. Surface tension, percolation, and roughening. Journal of Statistical Physics, 29(2):193–203, October 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020782.

Barboy:1980:HPFa

[BG80a] Boris Barboy and William M. Gelbart. Hard-particle fluids. I. General scaled-particle-like descriptions. Journal of Statistical Physics, 22(6):685–708, June 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01013937.

Barboy:1980:HPFb

[BG80b] Boris Barboy and William M. Gelbart. Hard-particle fluids. II. General y-expansion-like descriptions. Journal of Statistical Physics, 22(6):709–742, June 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01013938.

Benettin:1982:TSO

[BG82] Giancarlo Benettin and Luigi Galgani. Transition to stochasticity in a one-dimensional model of a radiant cavity. Journal of Statistical Physics, 27(1):153–169, January 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011744.

Benfatto:1984:SCP

[BG84] G. Benfatto and Ch. Gruber. On the susceptibility and clustering properties of unbounded spins. Journal of Statistical Physics, 37 (1–2):237–256, October 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012913.

Barrachina:1986:NSB

[BG86a] R. O. Barrachina and C. R. Garibotti. Nonisotropic solutions of the Boltzmann equation. Journal of Statistical Physics, 45(3–4): 541–560, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01021085.

Benettin:1986:SMN

[BG86b] Giancarlo Benettin and Giovanni Gallavotti. Stability of motions near resonances in quasi-integrable Hamiltonian systems. Journal of Statistical Physics, 44(3–4):293–338, August 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011301.

Bagdassarian:1988:LCS

[BGBS88] Carey Bagdassarian, William M. Gelbart, and Avinoam Ben-Shaul. Liquid crystalline states of surfactant solutions of isotropie micelles. Journal of Statistical Physics, 52(5–6):1307– 1313, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011648.

Barnsley:1984:GCJ

[BGH84] M. F. Barnsley, J. S. Geronimo, and A. N. Harrington. Geometry and combinatorics of Julia sets of real quadratic maps. Journal of Statistical Physics, 37(1–2):51–92, October 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012905.

Bessis:1984:MTA

[BGM84] D. Bessis, J. S. Geronimo, and P. Moussa. Meblin transforms associated with Julia sets and physical applications. Journal of Statistical Physics, 34(1–2):75–110, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01770350.

Borgis:1986:CCP

[BGM86] Daniel Borgis, Bernard Gaveau, and Michel Moreau. A class of collision processes with memory and application to simple chemical reactions in a solvent. Journal of Statistical Physics, 45(1– 2):319–347, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01033094.

Benfatto:1989:APH

[BGN89] G. Benfatto, G. Gallavotti, and F. Nicolo. On the analyticity of the pressure in the hierarchical dipole gas. Journal of Statistical Physics, 55(3–4):739–744, May 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01041605.

Bitensky:1985:MMV

[BGWY85] M. W. Bitensky, J. S. George, M. Whalen, and A. Yamazaki. Molecular mechanisms of visual excitation: A concatenation of nonlinear cellular processes. Journal of Statistical Physics, 39 (5–6):513–541, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008349.

Baus:1983:SMI

[BH83] M. Baus and J. P. Hansen. Statistical mechanics of ionic matter: Summary of a workshop. Journal of Statistical Physics, 31(2):409–422, May 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011590.

Batrouni:1988:FAI

[BH88a] Ghassan George Batrouni and Alex Hansen. Fourier acceleration of iterative processes in disordered systems. Journal of Statistical Physics, 52(3–4):747–773, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01019728.

Brummelhuis:1988:SVI

[BH88b] M. J. A. M. Brummelhuis and H. J. Hilhorst. Single-vacancy induced motion of a tracer particle in a two-dimensional lattice gas. Journal of Statistical Physics, 53(1–2):249–278, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011556.

Bunde:1988:FDR

[BHR^{+}88] A. Bunde, S. Havlin, H. E. Roman, G. Schildt, and H. E. Stanley. On the field dependence of random walks in the presence of random fields. Journal of Statistical Physics, 50(5–6):1271–1276, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019166.

Bilalbegovic:1988:IDT

[Bil88] G. Bilalbegović. Interface delocalization in the three-dimensional Ising model. Journal of Statistical Physics, 50(5–6):1131–1140, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019157.

Binder:1981:SDC

[Bin81] K. Binder. Static and dynamic critical phenomena of the twodimensional q-state Potts model. Journal of Statistical Physics, 24(1):69–86, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01007636.

Bishop:1982:CMO

[Bis82] Marvin Bishop. Collective modes of one-dimensional Lennard-Jones systems. Journal of Statistical Physics, 29(3):623–629, November 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01342191.

Benzi:1989:SPL

[BJLS89] Roberto Benzi, Giovanni Jona-Lasinio, and Alfonso Sutera. Stochastically perturbed Landau–Ginzburg equations. Journal of Statistical Physics, 55(3–4):505–522, May 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01041596.

Binder:1980:CCS

[BK80] K. Binder and M. H. Kalos. ‘Critical clusters’ in a supersaturated vapor: Theory and Monte Carlo simulation. Journal of Statistical Physics, 22(3):363–396, March 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014648.

Baker:1981:CSI

[BK81] George A. Baker, Jr. and John M. Kincaid. The continuousspin Ising model, g_{0}: ϕ^{4}:_{d} field theory, and the renormalization group. Journal of Statistical Physics, 24(3):469–528, March 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012818.

Brydges:1987:MEH

[BK87] D. C. Brydges and T. Kennedy. Mayer expansions and the Hamilton–Jacobi equation. Journal of Statistical Physics, 48 (1–2):19–49, July 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010398.

Bhandari:1988:MCA

[BK88a] R. Bhandari and M. Kerker. Monte Carlo analysis of the internal structure of light scattering particles with slit-scan illumination. Journal of Statistical Physics, 52(5–6):1263–1283, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011646.

Bovier:1988:WDE

[BK88b] Anton Bovier and Abel Klein. Weak disorder expansion of the invariant measure for the one-dimensional Anderson model. Journal of Statistical Physics, 51(3–4):501–517, May 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01028469.

Bricmont:1988:HRF

[BK88c] J. Bricmont and A. Kupiainen. The hierarchical random field Ising model. Journal of Statistical Physics, 51(5–6):1021–1032, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014898.

Bricmont:1983:STP

[BKL83] Jean Bricmont, Koji Kuroda, and Joel L. Lebowitz. Surface tension and phase coexistence for general lattice systems. Journal of Statistical Physics, 33(1):59–75, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009748.

Block:1989:IAC

[BKLP89] Louis Block, James Keesling, Shihai Li, and Kevin Peterson. An improved algorithm for computing topological entropy. Journal of Statistical Physics, 55(5–6):929–939, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01041072.

Bricmont:1984:BRT

[BL84] J. Bricmont and Joel L. Lebowitz. Book reviews: Two from Sinai. Journal of Statistical Physics, 34(3–4):651–656, February 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01018564.

Bouchaud:1985:NSD

[BL85] Jean-Philippe Bouchaud and Pierre Le Doussal. Numerical study of a d-dimensional periodic Lorentz gas with universal properties. Journal of Statistical Physics, 41(1–2):225–248, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020610.

Bricmont:1986:CME

[BL86] J. Bricmont and J. L. Lebowitz. On the continuity of the magnetization and energy in Ising ferromagnets. Journal of Statistical Physics, 42(5–6):861–869, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010449.

Bricmont:1987:WPB

[BL87] J. Bricmont and J. L. Lebowitz. Wetting in Potts and Blume– Capel models. Journal of Statistical Physics, 46(5–6):1015–1029, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011154.

Bellomo:1988:AEB

[BL88] N. Bellomo and M. Lachowicz. On the asymptotic equivalence between the Enskog and the Boltzmann equations. Journal of Statistical Physics, 51(1–2):233–247, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01015329.

Buttiker:1990:RBL

[BL90] M. Büttiker and R. Landauer. Response to ‘The Búttiker– Landauer model generalized’. Journal of Statistical Physics, 58 (1–2):371–373, January 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020298. See [SH89c].

Bricmont:1987:PSC

[BLM87] Jean Bricmont, Joel L. Lebowitz, and Christian Maes. Percolation in strongly correlated systems: The massless Gaussian field. Journal of Statistical Physics, 48(5–6):1249–1268, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009544.

Blomberg:1981:SPS

[Blo81] C. Blomberg. Some properties of stochastic equations for coupled chemical reactions far from equilibrium. Journal of Statistical Physics, 25(1):73–109, May 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008480.

Bricmont:1981:LSP

[BLP81a] J. Bricmont, J. L. Lebowitz, and C. E. Pfister. On the local structure of the phase separation line in the two-dimensional Ising system. Journal of Statistical Physics, 26(2):313–332, October 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013174.

Bricmont:1981:PGS

[BLP81b] Jean Bricmont, Joel L. Lebowitz, and Charles E. Pfister. Periodic Gibbs states of ferromagnetic spin systems. Journal of Statistical Physics, 24(1):269–277, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01007648.

Burdzy:1989:CER

[BLP89] Krzysztof Burdzy, Gregory F. Lawler, and Thomas Polaski. On the critical exponent for random walk intersections. Journal of Statistical Physics, 56(1–2):1–12, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01044226.

Bulsara:1982:SAN

[BLS82] Adi R. Bulsara, Katja Lindenberg, and K. E. Shuler. Spectral analysis of a nonlinear oscillator driven by random and periodic forces. I. Linearized theory. Journal of Statistical Physics, 27 (4):787–808, April 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013448.

Blum:1980:SOZ

[Blu80] Lesser Blum. Solution of the Ornstein–Zernike equation for a mixture of hard ions and Yukawa closure. Journal of Statistical Physics, 22(6):661–672, June 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01013935.

Blumenfeld:1989:USM

[Blu89] Raphael Blumenfeld. Universality and superuniversality of multifractals in nonlinear resistor networks. Journal of Statistical Physics, 56(1–2):233–241, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01044243.

Berker:1984:HMC

[BM84a] A. Nihat Berker and Susan R. McKay. Hierarchical models and chaotic spin glasses. Journal of Statistical Physics, 36(5– 6):787–793, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012938.

Borgis:1984:TCS

[BM84b] D. Borgis and M. Moreau. Two-cell stochastic model of the Schlögl reaction with small diffusional coupling. Journal of Statistical Physics, 37(5–6):631–651, December 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010499.

Borgis:1988:NTM

[BM88] D. Borgis and M. Moreau. Nucleation theory for a model bistable chemical reaction. Journal of Statistical Physics, 50(5– 6):935–962, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019148.

Bacci:1989:TSP

[BM89] S. Bacci and E. N. Miranda. The traveling salesman problem and its analogy with two-dimensional spin glasses. Journal of Statistical Physics, 56(3–4):547–551, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01044452.

Brachet:1984:TGV

[BMO^{+}84] M. E. Brachet, D. Meiron, S. Orszag, B. Nickel, R. Morf, and U. Frisch. The Taylor–Green vortex and fully developed turbulence. Journal of Statistical Physics, 34(5–6):1049–1063, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009458.

Benfatto:1980:SEA

[BMPP80] G. Benfatto, C. Marchioro, E. Presutti, and M. Pulvirenti. Superstability estimates for anharmonic systems. Journal of Statistical Physics, 22(3):349–362, March 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014647.

Brostow:1982:MCSb

[BMS82] Witold Brostow, M. Antonieta Macip, and Jerzy S. Sochanski. Macromolecular conformations in solutions. II. Thermodynamics of interactions. Journal of Statistical Physics, 29(4):865–878, December 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011796.

Bunde:1987:EMK

[BMS87] Armin Bunde, Sasuke Miyazima, and H. Eugene Stanley. From the Eden model to the kinetic growth walk: A generalized growth model with a finite lifetime of growth sites. Journal of Statistical Physics, 47(1–2):1–16, April 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009033.

Berezhkovskii:1989:WSV

[BMS89] A. M. Berezhkovskii, Yu. A. Makhnovskii, and R. A. Suris. Wiener sausage volume moments. Journal of Statistical Physics, 57(1–2):333–346, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01023647.

Baras:1982:APC

[BMV82] F. Baras, M. Malek Mansour, and C. Van den Broeck. Asymptotic properties of coupled nonlinear Langevin equations in the limit of weak noise. II: Transition to a limit cycle. Journal of Statistical Physics, 28(3):577–587, July 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008325.

Barnsley:1985:INS

[BMV85] M. F. Barnsley, T. D. Morley, and E. R. Vrscay. Iterated networks and the spectra of renormalizable electromechanical systems. Journal of Statistical Physics, 40(1–2):39–67, July 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010526.

Balescu:1981:F

[BN81] R. Balescu and G. Nicolis. Foreword. Journal of Statistical Physics, 24(1):5, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01007630.

Borgs:1987:PDA

[BN87] Christian Borgs and Florian Nill. The phase diagram of the Abelian lattice Higgs model. A review of rigorous results. Journal of Statistical Physics, 47(5–6):877–904, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01206163.

Benatti:1988:EDC

[BN88] F. Benatti and H. Narnhofer. Entropic dimension for completely positive maps. Journal of Statistical Physics, 53(5–6):1273–1298, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023869.

Bhattacharjee:1983:CBT

[BNHF83] Somendra M. Bhattacharjee, John F. Nagle, David A. Huse, and Michael E. Fisher. Critical behavior of a three-dimensional dimer model. Journal of Statistical Physics, 32(2):361–374, August 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012715.

Baras:1983:STA

[BNMT83] F. Baras, G. Nicolis, M. Malek Mansour, and J. W. Turner. Stochastic theory of adiabatic explosion. Journal of Statistical Physics, 32(1):1–23, July 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009416.

Brown:1987:GEA

[BOG87] Reggie Brown, Edward Ott, and Celso Grebogi. The goodness of ergodic adiabatic invariants. Journal of Statistical Physics, 49(3– 4):511–550, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009347.

Bonomi:1985:MDS

[Bon85] Ernesto Bonomi. Molecular dynamical simulation of the canonical ensemble. Journal of Statistical Physics, 39(1–2):167–180, April 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01007978.

Bonilla:1987:SNP

[Bon87] Luis L. Bonilla. Stable nonequilibrium probability densities and phase transitions for meanfield models in the thermodynamic limit. Journal of Statistical Physics, 46(3–4):659–678, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013379.

Borzi:1987:DCF

[BOP87] C. Borzi, G. Ord, and J. K. Percus. The direct correlation function of a one-dimensional Ising model. Journal of Statistical Physics, 46(1–2):51–66, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010330.

Borgs:1987:CSA

[Bor87] Christian Borgs. Charged surfaces and the analyticity properties of the string tension in lattice gauge theories. Journal of Statistical Physics, 47(5–6):867–876, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01206162.

Bovier:1989:PEO

[Bov89] Anton Bovier. Perturbation expansion for a one-dimensional Anderson model with off-diagonal disorder. Journal of Statistical Physics, 56(5–6):645–668, September 1989. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01016772.

Boyarsky:1987:SPP

[Boy87] Abraham Boyarsky. Singular perturbations of piecewise monotonic maps of the interval. Journal of Statistical Physics, 48(3– 4):561–569, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019688.

Boyarsky:1988:MME

[Boy88] Abraham Boyarsky. A matrix method for estimating the Liapunov exponent of one-dimensional systems. Journal of Statistical Physics, 50(1–2):213–229, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01022993.

Beals:1983:HRC

[BP83] R. Beals and V. Protopopescu. Half-range completeness for the Fokker–Planck equation. Journal of Statistical Physics, 32(3): 565–584, September 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008957.

Ballone:1985:ADP

[BP85a] P. Ballone and G. Pastore. Asymptotic density profile of a classical fluid against a hard wall. Journal of Statistical Physics, 38 (5–6):861–865, March 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010419.

Buffet:1985:HCB

[BP85b] E. Buffet and J. V. Pulé. A hard core Bose gas. Journal of Statistical Physics, 40(5–6):631–653, September 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009894.

Bonato:1989:KTT

[BP89] C. A. Bonato and J. Fernando Perez. Kosterlitz–Thouless transition for the finite-temperature d = 2 + 1, U(1) Hamiltonian lattice gauge theory. Journal of Statistical Physics, 56(1–2):13– 22, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01044227.

Bonato:1982:MWP

[BPK82] C. A. Bonato, J. Fernando Perez, and Abel Klein. The Mermin– Wagner phenomenon and cluster properties of one- and twodimensional systems. Journal of Statistical Physics, 29(2): 159–175, October 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020779.

Barber:1984:MCT

[BPP84] Michael N. Barber, Ingo Peschel, and Paul A. Pearce. Magnetization at corners in two-dimensional Ising models. Journal of Statistical Physics, 37(5–6):497–527, December 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010493.

Benfatto:1987:IMT

[BPP87] G. Benfatto, P. Picco, and M. Pulvirenti. On the invariant measures for the two-dimensional Euler flow. Journal of Statistical Physics, 46(3–4):729–742, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013382.

Boldrighini:1983:CSS

[BPT83] Carlo Boldrighini, Alessandro Pellegrinotti, and Livio Triolo. Convergence to stationary states for infinite harmonic systems. Journal of Statistical Physics, 30(1):123–155, January 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010871.

Bessis:1988:GDE

[BPTV88] D. Bessis, G. Paladin, G. Turchetti, and S. Vaienti. Generalized dimensions, entropies, and Liapunov exponents from the pressure function for strange sets. Journal of Statistical Physics, 51 (1–2):109–134, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01015323.

Bolshov:1985:IEC

[BPU85] L. A. Bol’shov, V. L. Pokrovsky, and G. V. Uimin. On the interface at the equilibrium crystallization. Journal of Statistical Physics, 38(1–2):191–205, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01017857.

Belavin:1984:ICS

[BPZ84] A. A. Belavin, A. M. Polyakov, and A. B. Zamolodchikov. Infinite conformal symmetry of critical fluctuations in two dimensions. Journal of Statistical Physics, 34(5–6):763–774, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009438.

Battle:1980:FIY

[BR80] Guy A. Battle and Lon Rosen. The FKG inequality for the Yukawa_{2} quantum field theory. Journal of Statistical Physics, 22 (2):123–192, February 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008048.

Buff:1985:SCF

[BR85] Frank P. Buff and Mark A. Robert. Some considerations of fluid interfaces in two dimensions. Journal of Statistical Physics, 41(5–6):1037–1044, December 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010014.

Barma:1986:ETI

[BR86] Mustansir Barma and Ramakrishna Ramaswamy. Escape times in interacting biased random walks. Journal of Statistical Physics, 43(3–4):561–570, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020653.

Bramson:1988:FPC

[Bra88] Maury Bramson. Front propagation in certain one-dimensional exclusion models. Journal of Statistical Physics, 51(5–6):863–869, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014888.

Brankov:1989:FSS

[Bra89] Jordan G. Brankov. Finite-size scaling for the mean spherical model with inverse power law interaction. Journal of Statistical Physics, 56(3–4):309–330, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01044439.

Breen:1987:LOE

[Bre87] Stephen Breen. Large-order estimates for ground-state energy perturbation series. Journal of Statistical Physics, 46(5–6): 1233–1280, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011163.

Brocas:1981:CRM

[Bro81] J. Brocas. Classification of rearrangement mechanisms and dynamic stereochemistry. Journal of Statistical Physics, 24(1): 199–214, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01007644.

Brostow:1982:MCSa

[Bro82] Witold Brostow. Macromolecular conformations in solutions. I. Model for chains with partial flexibility. Journal of Statistical Physics, 29(4):849–864, December 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011795.

Brosa:1989:TSA

[Bro89] U. Brosa. Turbulence without strange attractor. Journal of Statistical Physics, 55(5–6):1303–1312, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01041090.

Brush:1988:BRC

[Bru88] Stephen G. Brush. Book review: On the continuity of the gaseous and liquid states. Journal of Statistical Physics, 53(5–6): 1337–1339, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023873.

Brydges:1986:CME

[Bry86] David C. Brydges. Convergence of Mayer expansions. Journal of Statistical Physics, 42(3–4):425–435, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01127719.

Boley:1981:SMR

[BS81] C. D. Boley and Marlan O. Scully. The statistical mechanical resolution of a thermodynamic ‘paradox’. Journal of Statistical Physics, 24(1):159–174, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01007641.

Bernasconi:1983:DRO

[BS83] J. Bernasconi and W. R. Schneider. Diffusion in random onedimensional systems. Journal of Statistical Physics, 30(2): 355–362, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012309.

Brey:1984:SBM

[BS84] J. J. Brey and A. Santos. Solution of the BGK model kinetic equation for very hard particle interaction. Journal of Statistical Physics, 37(1–2):123–149, October 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012908.

Berretti:1985:NMC

[BS85] Alberto Berretti and Alan D. Sokal. New Monte Carlo method for the self-avoiding walk. Journal of Statistical Physics, 40(3– 4):483–531, August 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01017183.

Brydges:1986:ASC

[BS86] D. C. Brydges and E. Seiler. Absence of screening in certain lattice gauge and plasma models. Journal of Statistical Physics, 42(3–4):405–424, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01127718.

Bilalbegovic:1987:IFE

[BS87a] G. Bilalbegović and N. M. Svrakić. Interfacial free energy of the two-dimensional Ising model from the renormalization group. Journal of Statistical Physics, 46(3–4):777–787, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013385.

Budimir:1987:RBS

[BS87b] J. Budimir and J. L. Skinner. On the relationship between T_{1} and T_{2} for stochastic relaxation models. Journal of Statistical Physics, 49(5–6):1029–1042, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01017558.

Bendler:1988:GVL

[BS88] John T. Bendler and Michael F. Shlesinger. Generalized Vogel law for glass-forming liquids. Journal of Statistical Physics, 53 (1–2):531–541, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011571.

Benzi:1989:BLG

[BS89a] Roberto Benzi and Sauro Succi. Bifurcations of a lattice gas flow under external forcing. Journal of Statistical Physics, 56(1–2):69– 81, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01044232.

Blawzdziewicz:1989:LTK

[BS89b] Jerzy Bławzdziewicz and George Stell. Local H-theorem for a kinetic variational theory. Journal of Statistical Physics, 56(5–6): 821–840, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016781.

Bricmont:1989:PTS

[BS89c] J. Bricmont and J. Slawny. Phase transitions in systems with a finite number of dominant ground states. Journal of Statistical Physics, 54(1–2):89–161, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01023475.

Brosa:1989:VMC

[BS89d] U. Brosa and D. Stauffer. Vectorized multisite coding for hydrodynamic cellular automata. Journal of Statistical Physics, 57(1– 2):399–403, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023652.

Braun:1989:MCE

[BSF89] P. Braun, U. Staaden, and M. Fähnle. Monte Carlo evidence for a cusp in the uniform susceptibility of the site-diluted simple cubic Ising antiferromagnet. Journal of Statistical Physics, 57(1– 2):411–414, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023654.

Behn:1987:NTD

[BSTK87] U. Behn, K. Schiele, A. Teubel, and A. Kühnel. Nonequilibrium transitions driven by external dichotomous noise. Journal of Statistical Physics, 47(5–6):947, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01206168.

Balazs:1987:SFZ

[BSV87] N. L. Balazs, C. Schmit, and A. Voros. Spectral fluctuations and zeta functions. Journal of Statistical Physics, 46(5–6):1067–1090, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011157.

Bialek:1989:OPF

[BSZ89] W. Bialek, R. Scalettar, and A. Zee. Optimal performance of a feed-forward network at statistical discrimination tasks. Journal of Statistical Physics, 57(1–2):141–156, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01023637.

Burschka:1981:KBLb

[BT81a] M. A. Burschka and U. M. Titulaer. The kinetic boundary layer for the Fokker–Planck equation: Selectively absorbing boundaries. Journal of Statistical Physics, 26(1):59–71, September 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01106786.

Burschka:1981:KBLa

[BT81b] M. A. Burschka and U. M. Titulaer. The kinetic boundary layer for the Fokker–Planck equation with absorbing boundary. Journal of Statistical Physics, 25(3):569–582, July 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010804.

Birman:1985:SSA

[BT85] Joseph L. Birman and H.-R. Trebin. ‘Statistical’ symmetry with applications to phase transitions. Journal of Statistical Physics, 38(1–2):371–391, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01017868.

Brankov:1988:FSS

[BT88] Jordan G. Brankov and Nicholai S. Tonchev. On the finite-size scalling equation for the spherical model. Journal of Statistical Physics, 52(1–2):143–159, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01016408.

Berche:1989:CIP

[BT89a] Bertrand Berche and Loïc Turban. Conformal invariance and perturbations in the two-dimensional Ising model: Surface defects. Journal of Statistical Physics, 56(5–6):589–607, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016768.

Buttiker:1989:PSK

[BT89b] M. Büttiker and H. Thomas. Propagation and stability of kinks in driven and damped nonlinear Klein–Gordon chains. Journal of Statistical Physics, 54(5–6):1427, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01044726.

Belobrov:1985:MND

[BTZ85] P. I. Belobrov, A. G. Tret’yakov, and G. M. Zaslavsky. Methods of nonlinear dynamics and equilibrium structures of magnetoelastic chains. Journal of Statistical Physics, 38(1–2):393–404, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017869.

Burschka:1986:FDR

[Bur86] Martin A. Burschka. Fluctuations in diffusion reaction systems. I: Adiabatic elimination of transport modes from a mesoscopic nbody system and the Ω-expansion. Journal of Statistical Physics, 45(3–4):715–744, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01021092.

Burkov:1987:ODM

[Bur87] S. E. Burkov. One-dimensional model of the quasicrystalline alloy. Journal of Statistical Physics, 47(3–4):409–438, May 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01007518.

Burkov:1988:GST

[Bur88] S. E. Burkov. Ground states of two-dimensional quasicrystals. Journal of Statistical Physics, 52(1–2):453–461, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016426.

Buttiker:1989:TCS

[Büt89] M. Büttiker. Transport as a consequence of state-dependent diffusion. Journal of Statistical Physics, 54(5–6):1425, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01044725.

Behn:1989:NDI

[BvH89] Ulrich Behn and J. Leo van Hemmen. Network description of the immune system: Dormant b cells stabilize cycles. Journal of Statistical Physics, 56(3–4):533–545, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01044451.

Brandenberger:1982:DCS

[BW82] R. Brandenberger and C. E. Wayne. Decay of correlations in surface models. Journal of Statistical Physics, 27(3):425–440, March 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011084.

Ball:1984:PAV

[BW84] R. C. Ball and T. A. Witten. Particle aggregation versus cluster aggregation in high dimensions. Journal of Statistical Physics, 36(5–6):873–879, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012946.

Boldrighini:1988:FOD

[BW88a] C. Boldrighini and W. David Wick. Fluctuations in a onedimensional mechanical system. I. The Euler limit. Journal of Statistical Physics, 52(3–4):1069–1095, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019740.

Brydges:1988:MEH

[BW88b] D. C. Brydges and J. D. Wright. Mayer expansions and the Hamilton–Jacobi equation. II. Fermions, dimensional reduction formulas. Journal of Statistical Physics, 51(3–4):435–456, May 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01028465. See erratum [BW88b].

Binder:1989:FSE

[BW89] Kurt Binder and Jian-Sheng Wang. Finite-size effects at critical points with anisotropic correlations: Phenomenological scaling theory and Monte Carlo simulations. Journal of Statistical Physics, 55(1–2):87–126, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01042592.

Berera:1988:OVO

[BWdF88] A. Berera, L. T. Wille, and D. de Fontaine. Oxygen-vacancy ordering in the YBa_{2} cu_{3} o_{z} basal plane studied by the cluster variation method. Journal of Statistical Physics, 50(5–6):1245–1253, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019163.

Blumen:1983:ETR

[BZ83] A. Blumen and G. Zumofen. Energy transfer as a random walk with long-range steps. Journal of Statistical Physics, 30 (2):487–495, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012322.

Behn:1987:ODR

[BZ87] U. Behn and V. A. Zagrebnov. One-dimensional random field Ising model and discrete stochastic mappings. Journal of Statistical Physics, 47(5–6):939–946, June 1987. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01206167.

Calef:1983:IDO

[Cal83a] Daniel F. Calef. Impurity dynamics in a one-dimensional chain. Journal of Statistical Physics, 32(1):81–94, July 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009421.

Calheiros:1983:EPT

[Cal83b] Francisco J. L. C. Calheiros. The existence of a phase transition in classical antiferromagnetic models. Journal of Statistical Physics, 31(1):141–154, April 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010927.

Cammarota:1981:ILF

[Cam81] C. Cammarota. Ising-like field theory. Journal of Statistical Physics, 25(1):143–152, May 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008483.

Calderoni:1982:TEI

[CC82] P. Calderoni and S. Caprino. Time evolution of infinitely many particles: An existence theorem. Journal of Statistical Physics, 28(4):815–833, August 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011882.

Chayes:1984:VIC

[CC84] J. T. Chayes and L. Chayes. On the validity of the inverse conjecture in classical density functional theory. Journal of Statistical Physics, 36(3–4):471–488, August 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010992.

Caffarel:1986:TSE

[CC86] Michel Caffarel and Pierre Claverie. Treatment of the Schrödinger equation through a Monte Carlo method based upon the generalized Feynman–Kac formula. Journal of Statistical Physics, 43 (5–6):797–801, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02628305.

Chen:1987:TQD

[cC87] Yong cong Chen. Theory of quantum dynamics in fermionic environment: An influence functional approach. Journal of Statistical Physics, 47(1–2):17–55, April 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009034.

Chayes:1986:CBT

[CCD86] J. T. Chayes, L. Chayes, and R. Durrett. Critical behavior of the two-dimensional first passage time. Journal of Statistical Physics, 45(5–6):933–951, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020583.

Chung:1989:DFT

[cCM89] Tung chen Chung and William C. Meecham. Defect of the five-thirds law using the Wiener–Hermite expansion. Journal of Statistical Physics, 55(5–6):1089–1107, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01041080.

Campanino:1983:AOD

[CCO83] M. Campanino, D. Capocaccia, and E. Olivieri. Analyticity for one-dimensional systems with long-range superstable interactions. Journal of Statistical Physics, 33(2):437–476, November 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009805.

Capocaccia:1987:ETG

[CCP87] D. Capocaccia, M. Cassandro, and P. Picco. On the existence of thermodynamics for the generalized random energy model. Journal of Statistical Physics, 46(3–4):493–505, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013370.

Caroli:1981:DBP

[CCR81] B. Caroli, C. Caroli, and B. Roulet. Diffusion in a bistable potential: The functional integral approach. Journal of Statistical Physics, 26(1):83–111, September 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01106788.

Caroli:1982:KTW

[CCR82] B. Caroli, C. Caroli, and B. Roulet. Kramers time in weakly nonpotential systems. Journal of Statistical Physics, 28(4):757–767, August 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011879.

Chayes:1985:DFA

[CCR85] J. T. Chayes, L. Chayes, and Mary Beth Ruskai. Density functional approach to quantum lattice systems. Journal of Statistical Physics, 38(3–4):497–518, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010474.

Caroli:1980:WTD

[CCRG80] B. Caroli, C. Caroli, B. Roulet, and J. F. Gouyet. A WKB treatment of diffusion in a multidimensional bistable potential. Journal of Statistical Physics, 22(5):515–536, May 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011336.

Chayes:1987:EDC

[CCS87] J. T. Chayes, L. Chayes, and R. H. Schonmann. Exponential decay of connectivities in the two-dimensional Ising model. Journal of Statistical Physics, 49(3–4):433–445, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009344.

Calderoni:1989:SLS

[CD89a] Paola Calderoni and Detlef Dürr. The Smoluchowski limit for a simple mechanical model. Journal of Statistical Physics, 55 (3–4):695–738, May 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01041604.

Chulaevsky:1989:PAC

[CD89b] Victor Chulaevsky and François Delyon. Purely absolutely continuous spectrum for almost Mathieu operators. Journal of Statistical Physics, 55(5–6):1279–1284, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01041087.

Cook:1989:PDH

[CD89c] J. Cook and B. Derrida. Polymers on disordered hierarchical lattices: A nonlinear combination of random variables. Journal of Statistical Physics, 57(1–2):89–139, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01023636.

Calderoni:1989:MMB

[CDK89] Paola Calderoni, Detlef Dürr, and Shigeo Kusuoka. A mechanical model of Brownian motion in half-space. Journal of Statistical Physics, 55(3–4):649–693, May 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01041603.

Campbell:1986:QMC

[CDM86] D. K. Campbell, T. A. DeGrand, and S. Mazumdar. Quantum Monte Carlo studies of electron-electron interaction effects in conducting polymers. Journal of Statistical Physics, 43(5–6):803– 814, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02628306.

Caprino:1989:SPS

[CDPP89] S. Caprino, A. De Masi, E. Presutti, and M. Pulvirenti. A stochastic particle system modeling the Carleman equation. Journal of Statistical Physics, 55(3–4):625–638, May 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01041601.

Cohen:1987:SLCa

[CdS87a] E. G. D. Cohen and I. M. de Schepper. Symmetric linear collision operators in kinetic theory. Journal of Statistical Physics, 46 (5–6):949–969, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011151.

Cohen:1987:SLCb

[CdS87b] E. G. D. Cohen and I. M. de Schepper. Symmetric linear collision operators in kinetic theory. Journal of Statistical Physics, 49 (3–4):869, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009365.

Collet:1984:SGR

[CEGM84] P. Collet, J.-P. Eckmann, V. Glaser, and A. Martin. A spin glass with random couplings. Journal of Statistical Physics, 36 (1–2):89–106, July 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01015728.

Collet:1981:PDB

[CEK81] P. Collet, J.-P. Eckmann, and H. Koch. Period doubling bifurcations for families of maps on R^{n}. Journal of Statistical Physics, 25(1):1–14, May 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008475.

Ceperley:1986:SEG

[Cep86] D. M. Ceperley. The statistical error of Green’s function Monte Carlo. Journal of Statistical Physics, 43(5–6):815–826, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02628307.

Cercignani:1986:SLK

[Cer86] Carlo Cercignani. Solution of a linearized kinetic model for an ultrarelativistic gas. Journal of Statistical Physics, 42(3–4): 601–620, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01127731.

Cercignani:1987:GES

[Cer87] C. Cercignani. Global existence of solutions for a model Boltzmann equation. Journal of Statistical Physics, 49(5–6):1083– 1091, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01017561.

Cercignani:1988:SDE

[Cer88] Carlo Cercignani. Small data existence for the Enskog equation in L^{1}. Journal of Statistical Physics, 51(1–2):291–297, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01015332.

Chubb:1980:CFO

[CF80] Scott R. Chubb and David Fox. Correlation functions of odd numbers of spins with finite separations on the Onsager–Ising lattice. Journal of Statistical Physics, 23(5):627–638, November 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011733.

Celletti:1988:RKT

[CF88a] Alessandra Celletti and Corrado Falcolini. A remark on the KAM theorem applied to a four-vortex system. Journal of Statistical Physics, 52(1–2):471–477, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01016428.

Cichocki:1988:DCP

[CF88b] B. Cichocki and B. U. Felderhof. Dielectric constant of polarizable, nonpolar fluids and suspensions. Journal of Statistical Physics, 53(1–2):499–521, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011569.

Cichocki:1988:RCE

[CF88c] B. Cichocki and B. U. Felderhof. Renormalized cluster expansion for multiple scattering in disordered systems. Journal of Statistical Physics, 51(1–2):57–76, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01015320.

Cichocki:1989:TBP

[CF89] B. Cichocki and B. U. Felderhof. Three-body problem in the theory of the dielectric constant. Journal of Statistical Physics, 57(3– 4):871–885, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01022838.

Choquard:1980:ESC

[CFG80] Ph. Choquard, P. Favre, and Ch. Gruber. On the equation of state of classical one-component systems with long-range forces. Journal of Statistical Physics, 23(4):405–442, October 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011574.

Constantinescu:1987:ADS

[CFGK87] F. Constantinescu, G. Felder, K. Gawedzki, and A. Kupiainen. Analyticity of density of states in a gauge-invariant model for disordered electronic systems. Journal of Statistical Physics, 48 (3–4):365–391, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019679.

Choquard:1983:TDO

[CFS83] Ph. Choquard, P. J. Forrester, and E. R. Smith. The twodimensional one-component plasma at Γ = 2: The semiperiodic strip. Journal of Statistical Physics, 33(1):13–22, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009744.

Constantinescu:1984:ADS

[CFS84] F. Constantinescu, J. Fröhlich, and T. Spencer. Analyticity of the density of states and replica method for random schrödinger operators on a lattice. Journal of Statistical Physics, 34(3–4):571–596, February 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01018559.

Cornille:1980:SBE

[CG80] H. Cornille and A. Gervois. Solutions of the Boltzmann equation for Maxwell interactions and singular angle-dependent cross sections. Journal of Statistical Physics, 23(2):167–202, August 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012590.

Cornille:1981:PDS

[CG81a] H. Cornille and A. Gervois. Powerlike decreasing solutions of the Boltzmann equation for a Maxwell gas. Journal of Statistical Physics, 26(1):181–217, September 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01106793.

Cox:1981:CLT

[CG81b] J. Theodore Cox and Geoffrey Grimmett. Central limit theorems for percolation models. Journal of Statistical Physics, 25(2):237– 251, June 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01022185.

Caracciolo:1985:NMC

[CG85] Sergio Caracciolo and Ueli Glaus. A new Monte Carlo simulation for two models of self-avoiding lattice trees in two dimensions. Journal of Statistical Physics, 41(1–2):95–114, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020605.

Cassandro:1984:MBS

[CGOV84] Marzio Cassandro, Antonio Galves, Enzo Olivieri, and Maria Eulália Vares. Metastable behavior of stochastic dynamics: A pathwise approach. Journal of Statistical Physics, 35(5–6):603–634, June 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010826.

Casati:1983:PET

[CGVG83] Giulio Casati, Italo Guarneri, and Fausto Valz-Gris. Preliminaries to the ergodic theory of infinite-dimensional systems: A model of radiant cavity. Journal of Statistical Physics, 30(1):195–217, January 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010875.

Costa:1987:ETC

[CH87] U. M. S. Costa and H. J. Herrmann. Energy transport in a cellular automaton. Journal of Statistical Physics, 47(3–4):597–602, May 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01007528.

Cataldo:1988:MLP

[CH88a] H. M. Cataldo and E. S. Hernández. Memory loss process and non–Gibbsian equilibrium solutions of master equations. Journal of Statistical Physics, 53(3–4):673–688, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014219.

Cataldo:1988:NMQ

[CH88b] H. M. Cataldo and E. S. Hernández. Non-Markovian quantal Brownian motion model. Journal of Statistical Physics, 50(1– 2):383–403, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023000.

Chan:1988:IDC

[CH88c] Derek Y. C. Chan and Barry D. Hughes. Ion diffusion in a Coulombic field. Journal of Statistical Physics, 52(1–2):383–394, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01016421.

Chang:1982:EMB

[Cha82] Mou-Hsiung Chang. Equilibrium measure and Brownian escape process. Journal of Statistical Physics, 28(1):189–191, May 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011632.

Chandrasekhar:1984:SRS

[Cha84] S. Chandrasekhar. Some recent studies of liquid crystals: A review. Journal of Statistical Physics, 34(5–6):883–893, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009447.

Chandler:1986:RCD

[Cha86] David Chandler. Roles of classical dynamics and quantum dynamics on activated processes occurring in liquids. Journal of Statistical Physics, 42(1–2):49–67, January 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010840.

Chen:1982:CGH

[Che82] Su Shing Chen. Crystallographic groups and homogeneous statistical solutions of Navier–Stokes equations. Journal of Statistical Physics, 29(3):579–589, November 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01342188.

Chen:1986:ALS

[Che86] Steve S. Chen. Advanced large-scale and high-speed multiprocessor system for scientific applications. Journal of Statistical Physics, 43(5–6):1185–1208, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02628340.

Chen:1987:NMQ

[Che87] Yong-Cong Chen. A new method for quantum processes in fermionic heat baths. Journal of Statistical Physics, 49(3–4): 811–826, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009357.

Cohen:1983:GEC

[CHM83] Joel Cohen, Loren Haskins, and Jean-Paul Marchand. Geometry of equilibrium configurations in the Ising model. Journal of Statistical Physics, 31(3):671–678, June 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019504.

Chou:1988:PDO

[Cho88a] Weiren Chou. Phase diagram of a one-dimensional model with nonconvex interactions, using the method of effective potentials. Journal of Statistical Physics, 50(1–2):207–212, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01022992.

Chowdhury:1988:MGI

[Cho88b] Debashish Chowdhury. Morphology of growing interfacial patterns. Journal of Statistical Physics, 50(5–6):1111–1130, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01019156.

Christ:1986:LGT

[Chr86] N. H. Christ. Lattice gauge theory with a fast highly parallel computer. Journal of Statistical Physics, 43(5–6):1061–1072, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02628330.

Chudnovsky:1985:TNS

[Chu85a] E. M. Chudnovsky. Thermodynamics of natural selection. Journal of Statistical Physics, 41(5–6):877–897, December 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010008.

Chung:1985:FPT

[Chu85b] S. G. Chung. Fermionic perturbation theory for the statistical mechanics of the nonlinear Schrödinger model. Journal of Statistical Physics, 40(1–2):303–328, July 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010538.

Ciach:1985:ECN

[Cia85] A. Ciach. The effect of a constant number of particles on the pair correlation function and the density profile in a one-dimensional lattice gas. Journal of Statistical Physics, 40(3–4):593–606, August 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017187.

Cercignani:1988:NSH

[CIPS88] Carlo Cercignani, Reinhard Illner, Mario Pulvirenti, and Marvin Shinbrot. On nonlinear stationary half-space problems in discrete kinetic theory. Journal of Statistical Physics, 52(3–4):885–896, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019733.

Cornu:1987:TDC

[CJ87] Françoise Cornu and Bernard Jancovici. On the two-dimensional Coulomb gas. Journal of Statistical Physics, 49(1–2):33–56, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009953.

Cornu:1988:TDO

[CJB88] Françoise Cornu, Bernard Jancovici, and Lesser Blum. The twodimensional one-component plasma in a doubly periodic background: Exact results. Journal of Statistical Physics, 50(5–6): 1221–1243, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019162.

Carmona:1987:IST

[CK87a] René Carmona and Shinichi Kotani. Inverse spectral theory for random Jacobi matrices. Journal of Statistical Physics, 46(5– 6):1091–1114, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011158.

Chowdhury:1987:DGT

[CK87b] Debashish Chowdhury and S. Kumar. Domain growth in the three-dimensional dilute Ising model. Journal of Statistical Physics, 49(3–4):855–858, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009362.

Cole:1988:ERR

[CK88a] R. G. Cole and T. Keyes. The Enskog repeated ring equations for tagged molecule motion and variational solutions. Journal of Statistical Physics, 51(1–2):249–273, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01015330.

Cole:1988:VCF

[CK88b] R. G. Cole and T. Keyes. The velocity correlation function for the Lorentz gas. Journal of Statistical Physics, 51(1–2):275–289, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01015331.

Chen:1982:LTB

[CkM82] Hsin-Hsiung Chen and Shang keng Ma. Low-temperature behavior of a one-dimensional random Ising model. Journal of Statistical Physics, 29(4):717–746, December 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011787.

Cercignani:1981:HTP

[CL81] Carlo Cercignani and Maria Lampis. On the h-theorem for polyatomic gases. Journal of Statistical Physics, 26(4):795–801, December 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010940.

Cercignani:1988:KTD

[CL88] Carlo Cercignani and Maria Lampis. On the kinetic theory of a dense gas of rough spheres. Journal of Statistical Physics, 53(3– 4):655–672, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014218.

Collet:1989:RGA

[CL89] P. Collet and A. Lesne. Renormalization group analysis of some dynamical systems with noise. Journal of Statistical Physics, 57(5–6):967–992, December 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020044.

Classer:1983:SSV

[Cla83] M. L. Classer. A solvable six-vertex model with defects. Journal of Statistical Physics, 33(3):753–755, December 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01018841.

Collet:1987:DSS

[CLP87] P. Collet, J. L. Lebowitz, and A. Porzio. The dimension spectrum of some dynamical systems. Journal of Statistical Physics, 47 (5–6):609–644, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01206149.

Chen:1983:PDP

[CLRT83] S. H. Chen, C. C. Lai, J. Rouch, and P. Tartaglia. Probability distribution of photocounts of the light scattered by critical fluctuations. Journal of Statistical Physics, 30(3):699–707, March 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009684.

Caillol:1982:MCS

[CLWH82] J. M. Caillol, D. Levesque, J. J. Weis, and J. P. Hansen. A Monte Carlo study of the classical two-dimensional one-component plasma. Journal of Statistical Physics, 28(2):325–349, June 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012609.

Chate:1989:CML

[CM89] Hugues Chaté and Paul Manneville. Coupled map lattices as cellular automata. Journal of Statistical Physics, 56(3–4):357–370, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01044441.

Creutz:1986:CIT

[CMM86] Michael Creutz, P. Mitra, and K. J. M. Moriarty. Computer investigations of the three-dimensional Ising model. Journal of Statistical Physics, 42(5–6):823–832, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010447.

Chang:1984:SBP

[CO84] K. C. Chang and T. Odagaki. Site-bond percolation problems. Journal of Statistical Physics, 35(5–6):507–516, June 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010823.

Cohen:1988:ESA

[Coh88] Ruben D. Cohen. Equilibrium state of aggregation in suspensions comprising linear clusters. Journal of Statistical Physics, 51(3–4):559–567, May 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01028472.

Colin:1986:BRB

[Col86a] Leopolde Garcia Colin. Book review: Nonequilibrium statistical thermodynamics. Journal of Statistical Physics, 44(1–2):275–276, July 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010918.

Collet:1986:BRE

[Col86b] Pierre Collet. Book review: Ergodic theorems. Journal of Statistical Physics, 43(3–4):713, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020662.

Conlan:1982:SME

[Con82] Francis J. Conlan. The sub master equation. Journal of Statistical Physics, 27(4):767–785, April 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01013447.

Constantinescu:1988:STM

[Con88] F. Constantinescu. The supersymmetric transfer matrix for linear chains with nondiagonal disorder. Journal of Statistical Physics, 50(5–6):1167–1177, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019159.

Coopersmith:1983:NNS

[Coo83] Michael Coopersmith. On the nature of the nearest singularities of the free energy in the neighborhood of a critical point. Journal of Statistical Physics, 33(1):43–58, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009747.

Cornille:1980:KWM

[Cor80] H. Cornille. On the Krook–Wu model of the Boltzmann equation. Journal of Statistical Physics, 23(2):149–166, August 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012589.

Cornille:1985:NKM

[Cor85] H. Cornille. Nonlinear Kac model: Spatially homogeneous solutions and the Tjon effect. Journal of Statistical Physics, 39 (1–2):181–213, April 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01007979.

Cornille:1986:ISS

[Cor86] Henri Cornille. Inhomogeneous similarity solutions of the Boltzmann equation with confining external forces. Journal of Statistical Physics, 45(3–4):611–646, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01021088.

Cornille:1987:EDS

[Cor87] Henri Cornille. Exact (1+1)-dimensional solutions of discrete planar velocity Boltzmann models. Journal of Statistical Physics, 48 (3–4):789–811, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019697.

Cornille:1988:CPE

[Cor88] Henri Cornille. Construction of positive exact (2+1)-dimensional shock wave solutions for two discrete Boltzmann models. Journal of Statistical Physics, 52(3–4):897–949, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019734.

Cornu:1989:TDM

[Cor89a] Françoise Cornu. Two-dimensional models for an electrode with adsorption sites. Journal of Statistical Physics, 54(3–4):681–706, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019771.

Coron:1989:DSB

[Cor89b] François Coron. Derivation of slip boundary conditions for the Navier–Stokes system from the Boltzmann equation. Journal of Statistical Physics, 54(3–4):829–857, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019777.

Coste:1980:ITU

[Cos80] Jean Coste. Iterations of transformations on the unit interval: Approach to a periodic attractor. Journal of Statistical Physics, 23(4):521–536, October 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011579.

Cercignani:1982:SDC

[CP82] C. Cercignani and V. Protopopescu. The spectral decomposition for a class of linear transport operators. Journal of Statistical Physics, 29(2):363–374, October 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020792.

Cercignani:1987:EUN

[CP87] Carlo Cercignani and Andrzej Palczewski. Existence and uniqueness for nonlinear boundary value problems in kinetic theory. Journal of Statistical Physics, 46(1–2):273–281, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010346.

Corsten:1988:IDK

[CP88] Martin Corsten and Peter Poole. Initiation of damage in the Kauffman model. Journal of Statistical Physics, 50(1–2):461–463, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023004.

Calderoni:1989:TBI

[CPPV89] P. Calderoni, A. Pellegrinotti, E. Presutti, and M. E. Vares. Transient bimodality in interacting particle systems. Journal of Statistical Physics, 55(3–4):523–577, May 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01041597.

Choquard:1986:DSC

[CPR86] Ph. Choquard, B. Piller, and R. Rentsch. On the dielectric susceptibility of classical Coulomb systems. Journal of Statistical Physics, 43(1–2):197–205, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010577.

Choquard:1987:DSC

[CPR87] Ph. Choquard, B. Piller, and R. Rentsch. On the dielectric susceptibility of classical Coulomb systems. II. Journal of Statistical Physics, 46(3–4):599–633, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013377.

Choquard:1989:SPF

[CPRV89] Ph. Choquard, B. Piller, R. Rentsch, and P. Vieillefosse. Surface properties of finite classical Coulomb systems: Debye-Hückel approximation and computer simulations. Journal of Statistical Physics, 55(5–6):1185–1262, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01041085.

Crisanti:1988:GLE

[CPV88] Andrea Crisanti, Giovanni Paladin, and Angelo Vulpiani. Generalized Lyapunov exponents in high-dimensional chaotic dynamics and products of large random matrices. Journal of Statistical Physics, 53(3–4):583–601, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01014215.

Ciccotti:1981:DGL

[CR81] G. Ciccotti and J.-P. Ryckaert. On the derivation of the generalized Langevin equation for interacting Brownian particles. Journal of Statistical Physics, 26(1):73–82, September 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01106787.

Cahill:1986:TTV

[CR86] Kevin Cahill and Randolph Reeder. Testing two versions of lattice gauge theory: Creutz ratios in U (1)_{3}. Journal of Statistical Physics, 43(5–6):1043–1059, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02628329.

Cummings:1983:RFE

[CS83] P. T. Cummings and G. Stell. Random flights in Euclidean space. I. General analysis and results for flights with prescribed hit expectance density about the origin. Journal of Statistical Physics, 33(3):709–751, December 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01018840.

Cawley:1986:BR

[CS86a] Robert Cawley and Francis Sullivan. Book review. Journal of Statistical Physics, 45(3–4):769–773, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01021096.

Chowdhury:1986:MCS

[CS86b] Debashish Chowdhury and Dietrich Stauffer. Monte Carlo simulation of three-dimensional dilute Ising model. Journal of Statistical Physics, 44(1–2):203–210, July 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010912.

Casademunt:1989:LRT

[CS89] J. Casademunt and J. M. Sancho. Linear relaxation times of stochastic processes driven by non-Gaussian noises. Journal of Statistical Physics, 56(5–6):911–929, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01016785.

Ciccotti:1980:CEN

[CT80] G. Ciccotti and A. Tenenbaum. Canonical ensemble and nonequilibrium states by molecular dynamics. Journal of Statistical Physics, 23(6):767–772, December 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008518.

Caride:1984:SHA

[CT84] Anibal O. Caride and Constantino Tsallis. Specific heat of the anisotropic rigid rotator. Journal of Statistical Physics, 35(1– 2):187–190, April 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01017374.

Cukier:1982:BRR

[Cuk82] Robert I. Cukier. Book review: Rotational Brownian motion and dielectric theory. Journal of Statistical Physics, 28(2):407–408, June 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012614.

Cukier:1983:DCR

[Cuk83] Robert I. Cukier. Diffusion-controlled reactions among stationary sinks. Journal of Statistical Physics, 30(2):383–389, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012312.

Cukier:1986:DIR

[Cuk86] R. I. Cukier. Diffusion-influenced reactions. Journal of Statistical Physics, 42(1–2):69–82, January 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010841.

Curry:1981:CEH

[Cur81] James H. Curry. On computing the entropy of the Henon attractor. Journal of Statistical Physics, 26(4):683–695, December 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010933.

Claes:1987:RWP

[CV87] I. Claes and C. Van den Broeck. Random walk with persistence. Journal of Statistical Physics, 49(1–2):383–392, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009970.

Cummings:1981:SOZ

[CW81] P. T. Cummings and C. C. Wright. Solution of the Ornstein– Zernike equation for a soft-core Yukawa fluid. III. A restricted model for electrolytes and fused salts. Journal of Statistical Physics, 24(3):405–412, March 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012812.

Cuperman:1982:VDV

[CWD82] S. Cuperman, I. Weiss, and M. Dryer. A variational derivation of the velocity distribution functions for nonequilibrium, multispecies, weakly interacting, spherically symmetric many-body systems. Journal of Statistical Physics, 29(4):803–812, December 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011793.

Cuperman:1983:VDV

[CWD83] S. Cuperman, I. Weiss, and M. Dryer. A variational derivation of the velocity distribution functions for nonequilibrium, multispecies, weakly interacting, spherically symmetric many-body systems. Journal of Statistical Physics, 32(2):439, August 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012720.

Bonfim:1987:MCS

[dAB87] O. F. de Alcantara Bonfim. Monte Carlo study of tricritical dynamics in two dimensions. Journal of Statistical Physics, 48(3– 4):919–924, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019702.

deArcangelis:1987:IHE

[dAC87] Lucilla de Arcangelis and Antonio Coniglio. Infinite hierarchy of exponents in a two-component random resistor network. Journal of Statistical Physics, 48(3–4):935–942, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019704.

deArcangelis:1989:ACW

[dAHK^{+}89] L. de Arcangelis, H. J. Herrmann, M. Kolb, René Thomas, C. Amitrano, L. Peliti, M. Saber, G. Weisbuch, Sara A. Solla, F. Bagnoli, A. Francescato, R. Livi, S. Ruffo, Pablo Tamayo, Michel Droz, and et al. Abstracts from the CECAM workshop on computer simulations of cellular automata. Journal of Statistical Physics, 55(5–6):1333–1359, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01041096.

Darvey:1982:BR

[Dar82] Ivan G. Darvey. Book review. Journal of Statistical Physics, 28 (4):835–836, August 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011883.

Davies:1982:MIM

[Dav82] E. B. Davies. Metastability and the Ising model. Journal of Statistical Physics, 27(4):657–675, April 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013440.

Davies:1987:YBO

[Dav87] B. Davies. Yang–Baxter and other relations for free-fermion and Ising models. Journal of Statistical Physics, 46(3–4):535–557, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013374.

Dawson:1983:CDF

[Daw83] Donald A. Dawson. Critical dynamics and fluctuations for a mean-field model of cooperative behavior. Journal of Statistical Physics, 31(1):29–85, April 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010922.

Dhar:1980:EDR

[DB80] Deepak Dhar and Mustansir Barma. Effect of disorder on relaxation in the one-dimensional Glauber model. Journal of Statistical Physics, 22(2):259–277, February 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008051.

Dufty:1989:CQK

[DB89] James W. Dufty and David B. Boercker. Classical and quantum kinetic equations with exact conservation laws. Journal of Statistical Physics, 57(3–4):827–839, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01022835.

Davies:1986:LSL

[DBK^{+}86] C. Davies, G. Batrouni, G. Katz, A. Kronfeld, P. Lepage, P. Rossi, B. Svetitsky, and K. Wilson. Langevin simulations of lattice field theories using Fourier acceleration. Journal of Statistical Physics, 43(5–6):1073–1075, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02628331.

Dewel:1981:NPT

[DBW81] G. Dewel, P. Borckmans, and D. Walgraef. Nonequilibrium phase transitions and chemical instabilities. Journal of Statistical Physics, 24(1):119–137, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01007639.

Duarte:1989:RMC

[DC89] J. A. M. S. Duarte and Antonio M. R. Cadilhe. Relaxation Monte Carlo for 3D branched polymers: The leading confluence exponent. Journal of Statistical Physics, 56(5–6):951–955, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016788.

deConinck:1987:PCW

[dCD87] Joël de Coninck and François Dunlop. Partial to complete wetting: A microscopic derivation of the Young relation. Journal of Statistical Physics, 47(5–6):827–849, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01206160.

Druhl:1985:ASP

[DCW85] K. J. Drühl, J. L. Carlsten, and R. G. Wenzel. Aspects of soliton propagation in stimulated Raman scattering. Journal of Statistical Physics, 39(5–6):615–620, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008355.

Dagonnier:1981:NTM

[DD81] R. Dagonnier and Martine Dumont. On the nonergodicity of the transverse magnetization in the transverse Ising model. Journal of Statistical Physics, 24(1):59–68, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01007635.

DeConinck:1985:FSR

[DD85] Joël De Coninck and François Dunlop. Fluctuation susceptibility relations for classical spin systems. Journal of Statistical Physics, 40(1–2):241–248, July 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010535.

Duplantier:1988:EPF

[DD88] Bertrand Duplantier and François David. Exact partition functions and correlation functions of multiple Hamiltonian walks on the Manhattan lattice. Journal of Statistical Physics, 51 (3–4):327–434, May 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01028464.

Datoussaid:1983:CPL

[DDdG83] H. Datoussaïd, J. De Coninck, and Ph. de Gottal. Critical-point limit law for Temperley’s continuous model. Journal of Statistical Physics, 31(2):379–388, May 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011588.

DeConinck:1984:NII

[DdG84] J. De Coninck and Ph. de Gottal. New inequalities for Ising ferromagnets. Journal of Statistical Physics, 36(1–2):181–188, July 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01015732.

DeConinck:1989:MWT

[DdGM89] Joël De Coninck, Philippe de Gottal, and Frédéric Menu. A meniscus where three phases coexist at equilibrium: Microscopic derivation of the Herring relations. Journal of Statistical Physics, 56(1–2):23–32, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01044228.

Derrida:1983:FSZ

[DDI83] B. Derrida, L. De Seze, and C. Itzykson. Fractal structure of zeros in hierarchical models. Journal of Statistical Physics, 33 (3):559–569, December 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01018834.

Dunning-Davies:1987:PNH

[DDL87] J. Dunning-Davies and P. T. Landsberg. Problems of nonextensivity in hadron thermodynamics. Journal of Statistical Physics, 46(1–2):87–98, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010332.

Destri:1989:SVM

[DdVG89] C. Destri, H. J. de Vega, and H. J. Giacomini. The six-vertex model eigenvectors as critical limit of the eight-vertex model Bethe ansatz. Journal of Statistical Physics, 56(3–4):291–308, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01044438.

DeGregorio:1985:SPO

[De 85a] S. De Gregorio. The study of periodic orbits of dynamical systems. The use of a computer. Journal of Statistical Physics, 38 (5–6):947–972, March 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010424.

DeGregorio:1985:NBN

[De 85b] Salvatore De Gregorio. Narrow bounds for numerical integration of differential equations. Journal of Statistical Physics, 41(5–6): 865–876, December 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010007.

DeConinck:1987:GFM

[De 87] Joël De Coninck. Gaussian fluctuations for the magnetization of Lee–Yang ferromagnets at zero external field. Journal of Statistical Physics, 47(3–4):397–407, May 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01007517.

deSousaVieira:1988:SFA

[de 88] M. C. de Sousa Vieira. Scaling factors associated with Mfurcations of the 1 − μ||x||^{z} map. Journal of Statistical Physics, 53(5–6):1315–1325, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01023871.

dEspagnat:1989:TRI

[d’E89a] B. d’Espagnat. Are there realistically interpretable local theories? Journal of Statistical Physics, 56(5–6):747–766, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016778.

Dorfman:1989:HSB

[DE89b] J. R. Dorfman and M. H. Ernst. Hard-sphere binary-collision operators. Journal of Statistical Physics, 57(3–4):581–593, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01022823.

Dee:1985:PUS

[Dee85] Gregory Dee. Propagation into an unstable state. Journal of Statistical Physics, 39(5–6):705–717, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008361.

Deissler:1985:NSS

[Dei85] Robert J. Deissler. Noise-sustained structure, intermittency, and the Ginzburg–Landau equation. Journal of Statistical Physics, 40 (3–4):371–395, August 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01017180.

Deissler:1989:ENO

[Dei89] Robert J. Deissler. External noise and the origin and dynamics of structure in convectively unstable systems. Journal of Statistical Physics, 54(5–6):1459–1488, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01044729.

Delale:1982:EHE

[Del82a] Can F. Delale. Erratum: The Hilbert expansion to the Boltzmann equation for steady flow. Journal of Statistical Physics, 29(4):883, December 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011798. See [Del82b].

Delale:1982:HEB

[Del82b] Can F. Delale. The Hilbert expansion to the Boltzmann equation for steady flow. Journal of Statistical Physics, 28(3):589–602, July 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/

10.1007/BF01008326. See erratum [Del82a].

DelRio:1984:RLR

[Del84a] J. L. Del Rio. Regression law and the renormalization of the transport coefficients. Journal of Statistical Physics, 34(1–2):329–343, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01770363.

Delale:1984:TIT

[Del84b] Can F. Delale. The H theorem and irreversible thermodynamics. Journal of Statistical Physics, 37(3–4):451–463, November 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011844.

Delyon:1985:APS

[Del85] F. Delyon. Appearance of a purely singular continuous spectrum in a class of random schrödinger operators. Journal of Statistical Physics, 40(5–6):621–630, September 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009893.

Demetrius:1983:SMP

[Dem83] Lloyd Demetrius. Statistical mechanics and population biology. Journal of Statistical Physics, 30(3):709–753, March 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009685.

Derrida:1983:VDC

[Der83] Bernard Derrida. Velocity and diffusion constant of a periodic one-dimensional hopping model. Journal of Statistical Physics, 31(3):433–450, June 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019492;   http://www.lps.ens.
fr/~derrida/PAPIERS/1983/jsp1983-vitesse.pdf.

Der:1987:TLVa

[Der87a] R. Der. The time-local view of nonequilibrium statistical mechanics. I. Linear theory of transport and relaxation. Journal of Statistical Physics, 46(1–2):349–389, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010350.

Der:1987:TLVb

[Der87b] R. Der. The time-local view of nonequilibrium statistical mechanics. II. Generalized Langevin equations. Journal of Statistical Physics, 46(1–2):391–424, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010351.

Deutsch:1983:TDE

[Deu83] Claude Deutsch. Two-dimensional electron fluid at high temperature. Journal of Statistical Physics, 32(1):115–122, July 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009424.

deVegvar:1985:NPU

[deV85] P. G. N. deVegvar. Nonlinear propagation of ultrasound in superfluid ^{3}He. Journal of Statistical Physics, 39(5–6):675–685, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008359.

Davis:1980:MB

[DF80] Joel L. Davis and Ronald Forrest Fox. Maxwell →_{t→∞} Boltzmann. Journal of Statistical Physics, 22(6):627–645, June 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013933.

DeMasi:1984:RHZ

[DF84] A. De Masi and P. Ferrari. A remark on the hydrodynamics of the zero-range processes. Journal of Statistical Physics, 36(1–2):81– 87, July 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01015727.

DeMasi:1985:SDO

[DF85] A. De Masi and P. A. Ferrari. Self-diffusion in one-dimensional lattice gases in the presence of an external field. Journal of Statistical Physics, 38(3–4):603–613, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010480.

deForcrand:1986:QCF

[dF86a] Philippe de Forcrand. QCD from Chippewa Falls. Journal of Statistical Physics, 43(5–6):1077–1094, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02628332.

Delyon:1986:AIA

[DF86b] François Delyon and Patrick Foulon. Adiabatic invariants and asymptotic behavior of Lyapunov exponents of the Schrödinger equation. Journal of Statistical Physics, 45(1–2):41–47, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01033075.

Delyon:1987:ATL

[DF87] François Delyon and Patrick Foulon. Adiabatic theory, Liapunov exponents, and rotation number for quadratic Hamiltonians. Journal of Statistical Physics, 49(3–4):829–840, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009359.

DeMasi:1989:IPR

[DFGW89] A. De Masi, P. A. Ferrari, S. Goldstein, and W. D. Wick. An invariance principle for reversible Markov processes. Applications to random motions in random environments. Journal of Statistical Physics, 55(3–4):787–855, May 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01041608.

DeMasi:1982:SDLb

[DFIP82] A. De Masi, P. Ferrari, N. Ianiro, and E. Presutti. Small deviations from local equilibrium for a process which exhibits hydrodynamical behavior. II. Journal of Statistical Physics, 29 (1):81–93, September 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008249.

deForcrand:1986:SAR

[dFKP86] Ph. de Forcrand, F. Koukiou, and D. Petritis. Self-avoiding random walks on the hexagonal lattice. Journal of Statistical Physics, 45(3–4):459–470, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01021082.

deForcrand:1987:CES

[dFKP87] Ph. de Forcrand, F. Koukiou, and D. Petritis. Critical exponents for the self-avoiding random walk in three dimensions. Journal of Statistical Physics, 49(1–2):223–234, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009959.

DeMasi:1986:RDE

[DFL86] A. De Masi, P. A. Ferrari, and J. L. Lebowitz. Reaction-diffusion equations for interacting particle systems. Journal of Statistical Physics, 44(3–4):589–644, August 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011311.

DeMasi:1987:RDE

[DFL87] A. De Masi, P. A. Ferrari, and J. L. Lebowitz. Reaction-diffusion equations for interacting particle systems. Journal of Statistical Physics, 47(1–2):293, April 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009048.

Derrida:1986:ESO

[DFP86] B. Derrida, M. Mendès France, and J. Peyrière. Exactly solvable one-dimensional inhomogeneous models. Journal of Statistical Physics, 45(3–4):439–449, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01021080.

diFrancesco:1987:RBC

[dFSZ87] P. di Francesco, H. Saleur, and J. B. Zuber. Relations between the Coulomb gas picture and conformal invariance of two-dimensional critical models. Journal of Statistical Physics, 49(1–2):57–79, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009954.

DeMasi:1989:MMI

[DFV89] A. De Masi, P. A. Ferrari, and M. E. Vares. A microscopic model of interface related to the Burgers equation. Journal of Statistical Physics, 55(3–4):601–609, May 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01041599.

DiPrisco:1983:NAM

[DG83] A. Di Prisco and G. Gutiérrez. A note on the asymptotic master equations for systems with bound states. Journal of Statistical Physics, 31(3):617–621, June 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01019500.

Dee:1981:ESR

[DGK81] Gregory Dee, James D. Gunton, and Kyozi Kawasaki. An exact solution of the renormalization-group equations for the mean field theory of stable and metastable states. Journal of Statistical Physics, 24(1):87–107, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01007637.

Durr:1983:SPO

[DGL83] Detlef Dürr, Sheldon Goldstein, and Joel L. Lebowitz. Stochastic processes originating in deterministic microscopic dynamics. Journal of Statistical Physics, 30(2):519–526, February 1983.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012325.

Dorizzi:1984:PKR

[DGP84] B. Dorizzi, B. Grammaticos, and Y. Pomeau. The periodically kicked rotator: Recurrence and/or energy growth. Journal of Statistical Physics, 37(1–2):93–108, October 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012906.

denHollander:1984:RWL

[dH84a] W. Th. F. den Hollander. Random walks on lattices with randomly distributed traps I. The average number of steps until trapping. Journal of Statistical Physics, 37(3–4):331–367, November 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011838.

dHumieres:1984:DSO

[dH84b] D. d’Humières and B. A. Huberman. Dynamics of selforganization in complex adaptive networks. Journal of Statistical Physics, 34(3–4):361–379, February 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01018550.

denHollander:1985:CPG

[dH85a] W. Th. F. den Hollander. Comment on a paper by G. H. Weiss, S. Havlin, and A. Bunde. Journal of Statistical Physics, 40 (1–2):201–204, July 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010533. See [WHB85].

Doering:1985:CBT

[DH85b] C. R. Doering and W. Horsthemke. A comparison between transitions induced by random and periodic fluctuations. Journal of Statistical Physics, 38(3–4):763–783, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010489.

Ding:1986:ILS

[DH86a] E. J. Ding and Z. Q. Huang. On the initial layer solution of the Boltzmann equation with small Knudsen number. Journal of Statistical Physics, 45(3–4):589–610, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01021087.

Ding:1986:NSB

[DH86b] E. J. Ding and Z. Q. Huang. On the normal solutions of the Boltzmann equation with small Knudsen number. Journal of Statistical Physics, 45(3–4):561–588, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01021086.

Ding:1987:ETM

[DH87] E. J. Ding and P. C. Hemmer. Exact treatment of mode locking for a piecewise linear map. Journal of Statistical Physics, 46(1– 2):99–110, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010333.

deHaro:1989:CEV

[dH89a] M. López de Haro. Chapman–Enskog velocity distribution for tracer diffusion. Journal of Statistical Physics, 57(3–4):907–920, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01022840.

denHollander:1989:BRR

[dH89b] Frank den Hollander. Book review: Random media. Journal of Statistical Physics, 55(5–6):1327–1330, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01041094.

Desai:1988:FSS

[DHB88] Rashmi C. Desai, Dieter W. Heermann, and K. Binder. Finitesize scaling in a microcanonical ensemble. Journal of Statistical Physics, 53(3–4):795–823, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01014226.

denHollander:1985:RWL

[dHK85] W. Th. F. den Hollander and P. W. Kasteleyn. Random walks on lattices with points of two colors. II. Some rigorous inequalities for symmetric random walks. Journal of Statistical Physics, 39(1–2):15–52, April 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01007973.

DiMarzio:1984:FTC

[Di 84] Edmund A. Di Marzio. Field theory, curdling, limit cycles, and cellular automata. Journal of Statistical Physics, 36(5–6): 897–907, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012948.

Dickman:1985:AEO

[Dic85] Ronald Dickman. Approach to equilibrium in a one-dimensional, two-component gas of Maxwellian molecules. Journal of Statistical Physics, 41(3–4):607–619, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009024.

Dickman:1989:NLM

[Dic89] Ronald Dickman. Nonequilibrium lattice models: Series analysis of steady states. Journal of Statistical Physics, 55(5–6):997–1026, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01041076.

Dieterich:1985:SC

[Die85] W. Dieterich. Superionic conductors. Journal of Statistical Physics, 39(5–6):583–596, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008353.

DeMasi:1982:SDLa

[DIP82] A. De Masi, N. Ianiro, and E. Presutti. Small deviations from local equilibrium for a process which exhibits hydrodynamical behavior. I. Journal of Statistical Physics, 29(1):57–79, September 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008248.

Denker:1986:RSP

[DK86a] Manfred Denker and Gerhard Keller. Rigorous statistical procedures for data from dynamical systems. Journal of Statistical Physics, 44(1–2):67–93, July 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010905.

Dimits:1986:SPA

[DK86b] Andris M. Dimits and John A. Krommes. Stochastic particle acceleration and statistical closures. Journal of Statistical Physics, 44(5–6):879–906, September 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011912.

Dorfman:1989:P

[DKS89] J. R. Dorfman, T. R. Kirkpatrick, and J. V. Sengers. Preface. Journal of Statistical Physics, 57(3–4):430–432, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01022814.

DeRaedt:1982:MCC

[DL82] Hans De Raedt and Ad Lagendijk. Monte Carlo calculation of the thermodynamic properties of a one-dimensional fermion lattice model. Journal of Statistical Physics, 27(4):731–744, April 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013445.

Delyon:1989:BGO

[DL89] François Delyon and Jean-François Luciani. Behavior of general one-dimensional diffusion processes. Journal of Statistical Physics, 54(3–4):1065–1080, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019787.

delaLlave:1982:RCE

[dlL82] R. de la Llave. Rates of convergence to equilibrium in the Prigogine–Misra–Courbage theory of irreversibility. Journal of Statistical Physics, 29(1):17–31, September 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008245.

Delyon:1985:ALO

[DLS85] François Delyon, Yves Lévy, and Bernard Souillard. Anderson localization for one- and quasi-one-dimensional systems. Journal of Statistical Physics, 41(3–4):375–388, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009014.

delaSelva:1988:CRW

[dlSLW88] S. M. T. de la Selva, Katja Lindenberg, and Bruce J. West. Correlated random walks. Journal of Statistical Physics, 53(1– 2):203–219, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011553.

Dammers:1988:SLE

[DLT88] A. J. Dammers, Y. K. Levine, and J. A. Tjon. The stochastic Liouville equation and Padé approximants. Journal of Statistical Physics, 53(1–2):109–124, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011548.

Dambrine:1981:SDS

[DM81] S. Dambrine and M. Moreau. On the stationary distribution of a stochastic chemical process without detailed balance. Journal of Statistical Physics, 26(1):137–148, September 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01106790.

Dhara:1987:RUS

[DM87] A. K. Dhara and S. V. G. Menon. Relaxation of an unstable system driven by a colored noise. Journal of Statistical Physics, 46(3– 4):743–751, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013383.

Datye:1982:MSC

[DML82] V. Datye, R. Mathur, and J. S. Langer. Mode selection in a caricature of eutectic solidification. Journal of Statistical Physics, 29(1):1–16, September 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008244.

Debnath:1989:HPD

[DMM^{+}89] G. Debnath, Frank Moss, F. Marchesoni, Th. Leiber, and H. Risken. Holes in the probability density of strongly colored noise driven systems. Journal of Statistical Physics, 54(5–6): 1381–1382, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01044721.

DeConinck:1988:SPW

[DMMSR88] Joël De Coninck, Alain Messager, Salvador Miracle-Sole, and Jean Ruiz. A study of perfect wetting for Potts and Blume– Capel models with correlation inequalities. Journal of Statistical Physics, 52(1–2):45–60, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01016403.

Derrida:1987:LFA

[DN87] B. Derrida and J. P. Nadal. Learning and forgetting on asymmetric, diluted neural networks. Journal of Statistical Physics, 49(5– 6):993–1009, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01017556.

deOliveira:1988:ACE

[dO88] César R. de Oliveira. Attractors and characteristic exponents. Journal of Statistical Physics, 53(3–4):603–612, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014216.

deOliveira:1989:GOP

[dO89] Mário J. de Oliveira. Graph optimization problems on a Bethe lattice. Journal of Statistical Physics, 54(1–2):477–493, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023490.

deOliveira:1989:RGS

[dOB89] Paulo Murilo C. de Oliveira and F. C. Sá Barreto. Renormalization group studies of the Ashkin–Teller model. Journal of Statistical Physics, 57(1–2):53–63, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01023634.

Doi:1989:CMF

[Doi89] Shinji Doi. A chaotic map with a flat segment can produce a noise-induced order. Journal of Statistical Physics, 55(5–6):941– 964, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01041073.

Domb:1983:RSA

[Dom83] Cyril Domb. From random to self-avoiding walks. Journal of Statistical Physics, 30(2):425–436, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012316.

Domb:1987:BRS

[Dom87] C. Domb. Book review: Statphys 16. Journal of Statistical Physics, 49(5–6):1257, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01017571.

Domany:1988:NNB

[Dom88] Eytan Domany. Neural networks: A biased overview. Journal of Statistical Physics, 51(5–6):743–775, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014882.

Domb:1989:CRI

[Dom89] Cyril Domb. Covering by random intervals and one-dimensional continuum percolation. Journal of Statistical Physics, 55(1– 2):441–460, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01042611.

Donald:1987:FER

[Don87] Matthew J. Donald. Free energy and the relative entropy. Journal of Statistical Physics, 49(1–2):81–87, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009955.

Doolen:1985:F

[Doo85] Gary Doolen. Foreword. Journal of Statistical Physics, 39(5–6): 453, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01008344;   http://link.
springer.com/article/10.1007/BF01008344.

Dorfman:1983:BRP

[Dor83] J. R. Dorfman. Book review: Physical kinetics. Journal of Statistical Physics, 33(1):237–238, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009760.

Dorfle:1985:SEF

[Dör85a] M. Dörfle. Spectrum and eigenfunctions of the Frobenius–Perron operator of the tent map. Journal of Statistical Physics, 40 (1–2):93–132, July 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010528.

Dorfman:1985:BR

[Dor85b] J. R. Dorfman. Book review. Journal of Statistical Physics, 39 (3–4):445–447, May 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01018673.

Dotsenko:1984:CBA

[Dot84] Vl. S. Dotsenko. Critical behavior and associated conformal algebra of the Z_{3} Potts model. Journal of Statistical Physics, 34 (5–6):781–791, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009440.

DuBois:1985:SIA

[DP85] D. F. DuBois and D. Pesme. Stability of ion acoustic turbulent states. Journal of Statistical Physics, 39(5–6):783–816, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008366.

Dasgupta:1987:TAT

[DP87] Chandan Dasgupta and Rahul Pandit. Testing approximate theories of first-order phase transitions on the two-dimensional Potts model. Journal of Statistical Physics, 47(3–4):375–396, May 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01007516.

Duffield:1989:TPT

[DP89] N. G. Duffield and J. V. Pulé. Thermodynamics and phase transitions in the Overhauser model. Journal of Statistical Physics, 54 (1–2):449–475, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023489.

Dobrushin:1986:ODH

[DPST86] R. L. Dobrushin, A. Pellegrinotti, Yu. M. Suhov, and L. Triolo. One-dimensional harmonic lattice caricature of hydrodynamics. Journal of Statistical Physics, 43(3–4):571–607, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020654.

Dobrushin:1988:ODH

[DPST88] R. L. Dobrushin, A. Pellegrinotti, Yu. M. Suhov, and L. Triolo. One-dimensional harmonic lattice caricature of hydrodynamics: Second approximation. Journal of Statistical Physics, 52(1–2):423–439, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016423.

Derjaguin:1980:STF

[DPT80] B. V. Derjaguin, A. V. Prokhorov, and N. N. Tunitskij. Statistical thermodynamics of the formation of a new phase in the boiling up of volatile liquids. Journal of Statistical Physics, 23(4):495–520, October 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011578.

DeMasi:1986:EUE

[DPV86] A. De Masi, E. Presutti, and M. E. Vares. Escape from the unstable equilibrium in a random process with infinitely many interacting particles. Journal of Statistical Physics, 44(3–4):645–696, August 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011312.

Dufty:1983:MCL

[DR83] James W. Dufty and Rosalío F. Rodríguez. Mode coupling from linear and nonlinear kinetic equations. Journal of Statistical Physics, 33(2):261–286, November 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009797.

Dresden:1987:NPK

[Dre87] M. Dresden. New perspectives on Kac ring models. Journal of Statistical Physics, 46(5–6):829–842, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011144.

Driessler:1981:SRP

[Dri81] Wulf Driessler. On the spectrum of the Rayleigh piston. Journal of Statistical Physics, 24(4):595–606, April 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011377.

Driessler:1982:SRP

[Dri82] W. Driessler. On the spectrum of the Rayleigh piston II. Journal of Statistical Physics, 29(1):95–112, September 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008250.

Dekeyser:1980:DRV

[DS80] Raf Dekeyser and Attilio Stella. Differential renormalization of van der Waals spin models. Journal of Statistical Physics, 23(5): 587–607, November 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011731.

Davis:1981:GTF

[DS81] H. T. Davis and L. E. Scriven. Gradient theory of fluid microstructures. Journal of Statistical Physics, 24(1):243–268, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01007647.

Dickman:1983:STC

[DS83a] Ronald Dickman and William C. Schieve. A scaling theory of clusters in the lattice gas. Journal of Statistical Physics, 33(3): 527–547, December 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01018832.

Doria:1983:SFM

[DS83b] M. Doria and R. Shankar. Solution of a factorizable S-matrix and an asymmetric eight-vertex model. Journal of Statistical Physics, 33(3):695–708, December 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01018839.

Dickman:1984:PEC

[DS84] Ronald Dickman and William C. Schieve. Proof of the existence of the cluster free energy. Journal of Statistical Physics, 36(3– 4):435–446, August 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010990.

deSmedt:1986:ERI

[dS86a] Philip de Smedt. The effect of repulsive interactions on Bose– Einstein condensation. Journal of Statistical Physics, 45(1–2): 201–213, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01033087.

Devillard:1986:PDT

[DS86b] Pierre Devillard and Bernard Souillard. Polynomially decaying transmission for the nonlinear Schrödinger equation in a random medium. Journal of Statistical Physics, 43(3–4):423–439, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020646.

Dickman:1986:CTA

[DS86c] Ronald Dickman and William C. Schieve. Collapse transition and asymptotic scaling behavior of lattice animals: Low-temperature expansion. Journal of Statistical Physics, 44(3–4):465–489, August 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011306.

Dobrushin:1987:CAI

[DS87] R. L. Dobrushin and S. B. Shlosman. Completely analytical interactions: Constructive description. Journal of Statistical Physics, 46(5–6):983–1014, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011153.

daSilva:1988:FDP

[dS88a] L. R. da Silva. Fractal dimension at the phase transition of inhomogeneous cellular automata. Journal of Statistical Physics, 53(3–4):985–990, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014234.

Derrida:1988:PDT

[DS88b] B. Derrida and H. Spohn. Polymers on disordered trees, spin glasses, and traveling waves. Journal of Statistical Physics, 51 (5–6):817–840, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014886.

deSchepper:1982:VSW

[dSC82] I. M. de Schepper and E. G. D. Cohen. Very-short-wavelength collective modes in fluids. Journal of Statistical Physics, 27(2): 223–281, February 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008939.

deSchepper:1989:CET

[dSCKP89] I. M. de Schepper, E. G. D. Cohen, and B. Kamgar-Parsi. On the calculation of equilibrium time correlation functions in hardsphere fluids. Journal of Statistical Physics, 54(1–2):273–313, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023481.

deSchepper:1981:ISF

[dSEC81] I. M. de Schepper, M. H. Ernst, and E. G. D. Cohen. The incoherent scattering function and related correlation functions in hard sphere fluids at short times. Journal of Statistical Physics, 25(2):321–360, June 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01022190.

daSilva:1988:DSG

[dSH88] L. R. da Silva and H. J. Herrmann. Damage spreading in a gradient. Journal of Statistical Physics, 52(1–2):463–470, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016427.

DeGregorio:1983:RSP

[DST83] S. De Gregorio, E. Scoppola, and B. Tirozzi. A rigorous study of periodic orbits by means of a computer. Journal of Statistical Physics, 32(1):25–33, July 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009417.

deSmedt:1986:BRC

[dST86] Ph. de Smedt and J. Talbot. Book review: Computational physics. Journal of Statistical Physics, 44(3–4):709, August 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011317.

Durrett:1989:CLO

[DST89] Richard Durrett, Roberto H. Schonmann, and Nelson I. Tanaka. Correlation lengths for oriented percolation. Journal of Statistical Physics, 55(5–6):965–979, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01041074.

deSchepper:1989:MCE

[dSvLH89] I. M. de Schepper, J. J. van Loef, and A. F. E. M. Haffmans. Mode-coupling effects in the density dependence of the shear viscosities of dense argon and methane. Journal of Statistical Physics, 57(3–4):631–645, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01022825.

Vieira:1987:DIG

[dSVT87] M. C. de Sousa Vieira and C. Tsallis. d-dimensional ideal gas in parastatistics: Thermodynamic properties. Journal of Statistical Physics, 48(1–2):97–120, July 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010402.

Descalzi:1989:NPN

[DT89a] Orazio Descalzi and Enrique Tirapegui. Nonequilibrium potentials near instabilities. Journal of Statistical Physics, 57(5–6): 993–1012, December 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020045.

Durrett:1989:SIO

[DT89b] Richard Durrett and Nelson I. Tanaka. Scaling inequalities for oriented percolation. Journal of Statistical Physics, 55(5–6):981– 995, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01041075.

Dudynski:1989:LRB

[Dud89] Marek Dudyński. On the linearized relativistic Boltzmann equation. II. Existence of hydrodynamics. Journal of Statistical Physics, 57(1–2):199–245, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01023641.

Dunlop:1985:CIK

[Dun85] François Dunlop. Correlation inequalities and the Kosterlitz– Thouless transition for anisotropic rotators. Journal of Statistical Physics, 41(5–6):733–743, December 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010001.

Duplantier:1987:CEM

[Dup87] Bertrand Duplantier. Critical exponents of Manhattan Hamiltonian walks in two dimensions, from Potts and O(n) models. Journal of Statistical Physics, 49(3–4):411–431, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009343.

Duplantier:1989:SMP

[Dup89] Bertrand Duplantier. Statistical mechanics of polymer networks of any topology. Journal of Statistical Physics, 54(3–4):581–680, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019770.

Dale:1987:RDK

[DV87a] Andrew F. Dale and Oriol T. Valls. The role of dimensionality in the kinetic Ising model of spinodal decomposition: Evidence from zero-temperature quenches. Journal of Statistical Physics, 49(3– 4):661–667, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009351.

Donsker:1987:LDN

[DV87b] M. D. Donsker and S. R. S. Varadhan. Large deviations for noninteracting infinite-particle systems. Journal of Statistical Physics, 46(5–6):1195–1232, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011162.

Dorlas:1989:NGL

[DvE89] T. C. Dorlas and A. C. D. van Enter. Non–Gibbsian limit for large-block majority-spin transformations. Journal of Statistical Physics, 55(1–2):171–181, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01042596.

Durr:1987:BRQ

[DZ87] Detlef Dürr and Nino Zanghi. Book review: Quantum theory of collective phenomena. Journal of Statistical Physics, 47(3–4):603– 604, May 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01007529.

Derrida:1984:TMP

[DZVS84] B. Derrida, J. G. Zabolitzky, J. Vannimenus, and D. Stauffer. A transfer matrix program to calculate the conductivity of random resistor networks. Journal of Statistical Physics, 36(1–2):31–42, July 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01015724.

Ernst:1988:LLG

[EB88] M. H. Ernst and P. M. Binder. Lorentz lattice gases: Basic theory. Journal of Statistical Physics, 51(5–6):981–990, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014895.

Ebeling:1986:PDO

[Ebe86] Werner Ebeling. Pattern dynamics and optimization by reaction diffusion systems. Journal of Statistical Physics, 45(5–6):891–903, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020580.

Ernst:1981:NFS

[EC81] M. H. Ernst and E. G. D. Cohen. Nonequilibrium fluctuations in μ space. Journal of Statistical Physics, 25(1):153–180, May 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008484.

Erpenbeck:1986:NFD

[EC86] J. J. Erpenbeck and E. G. D. Cohen. Note on the fractal dimension of hard sphere trajectories. Journal of Statistical Physics, 43(1–2):343–347, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010586.

Escande:1981:RMC

[ED81] D. F. Escande and F. Doveil. Renormalization method for computing the threshold of the large-scale stochastic instability in two degrees of freedom Hamiltonian systems. Journal of Statistical Physics, 26(2):257–284, October 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013171.

Eisele:1988:MPT

[EE88] Theodor Eisele and Richard S. Ellis. Multiple phase transitions in the generalized Curie–Weiss model. Journal of Statistical Physics, 52(1–2):161–202, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016409.

Ebeling:1984:DRR

[EEEF84] W. Ebeling, A. Engel, B. Esser, and R. Feistel. Diffusion and reaction in random media and models of evolution processes. Journal of Statistical Physics, 37(3–4):369–384, November 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011839.

Elskens:1987:AKO

[EF87] Yves Elskens and Harry L. Frisch. Aggregation kinetics for a one-dimensional zero-degree Kelvin model of spinodal decomposition. Journal of Statistical Physics, 48(5–6):1243–1248, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009543.

Elskens:1983:ESD

[EFN83] Yves Elskens, Harry L. Frisch, and Grégoire Nicolis. Exact solution of a deterministic model for isomerization kinetics. Journal of Statistical Physics, 33(2):317–339, November 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009800.

Evans:1984:EKA

[EH84] J. W. Evans and D. K. Hoffman. Exact kinetics for ‘almost random’ irreversible filling of lattices. Journal of Statistical Physics, 36(1–2):65–80, July 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01015726.

Ecke:1989:NII

[EH89] Robert Ecke and Hans Haucke. Noise-induced intermittency in the quasiperiodic regime of Rayleigh–Bénard convection. Journal of Statistical Physics, 54(5–6):1153–1172, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01044708.

Ernst:1982:NSK

[EHH82] M. H. Ernst, K. Hellesoe, and E. H. Hauge. Nonunique solutions of kinetic equations. Journal of Statistical Physics, 27(4):677–691, April 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013441.

Eiskens:1987:ETM

[Eis87] Yves Eiskens. Ergodic theory of the mixmaster universe in higher space-time dimensions. II. Journal of Statistical Physics, 48(5–6): 1269–1282, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009545.

Elphick:1987:CBF

[EJT87] C. Elphick, M. Jeanneret, and E. Tirapegui. Comment on bifurcations in fluctuating systems. Journal of Statistical Physics, 48 (3–4):925–933, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019703.

Elskens:1985:RDM

[EK85] Yves Elskens and Raymond Kapral. Reversible dynamics and the macroscopic rate law for a solvable Kolmogorov system: The three bakers’ reaction. Journal of Statistical Physics, 38(5–6): 1027–1049, March 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010428.

Elskens:1984:MDM

[Els84] Yves Elskens. Microscopic derivation of a Markovian master equation in a deterministic model of chemical reaction. Journal of Statistical Physics, 37(5–6):673–695, December 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010501.

Ernst:1984:LTT

[EMDvB84] M. H. Ernst, J. Machta, J. R. Dorfman, and H. van Beijeren. Long time tails in stationary random media. I. Theory. Journal of Statistical Physics, 34(3–4):477–495, February 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01018555.

Evans:1985:RDF

[EN85] J. W. Evans and R. S. Nord. Random dimer filling of lattices: Three-dimensional application to free radical recombination kinetics. Journal of Statistical Physics, 38(3–4):681–705, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010485.

Englefield:1988:ESF

[Eng88] M. J. Englefield. Exact solutions of a Fokker–Planck equation. Journal of Statistical Physics, 52(1–2):369–381, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016420.

Ellis:1981:GIL

[ENO81] Richard S. Ellis, Charles M. Newman, and Michael R. O’Connell. The GHS inequality for a large external field. Journal of Statistical Physics, 26(1):37–52, September 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01106784.

Enz:1981:PAC

[Enz81] C. P. Enz. Perturbative analysis of the convection instability. Journal of Statistical Physics, 24(1):109–117, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01007638.

Eder:1988:SOD

[EP88] Otto J. Eder and Maximilian Posch. Solution of the onedimensional linear Boltzmann equation for charged Maxwellian particles in an external field. Journal of Statistical Physics, 52(3– 4):1031–1060, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019738.

Eaves:1981:POL

[ER81] John O. Eaves and William P. Reinhardt. Piecewise optimal linearization method for nonlinear stochastic differential equations. Journal of Statistical Physics, 25(1):127–141, May 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008482.

Elkoshi:1985:ODR

[ERH85] Zeev Elkoshi, Howard Reiss, and Audrey Dell Hammerich. Onedimensional rigorous hole theory of fluids: Internally constrained ensembles. Journal of Statistical Physics, 41(3–4):685–708, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009028.

Ernst:1984:ESN

[Ern84] Matthieu H. Ernst. Exact solutions of the nonlinear Boltzmann equation. Journal of Statistical Physics, 34(5–6):1001–1017, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009454.

Ernst:1987:TDC

[Ern87] M. H. Ernst. Time-dependent correlation functions for random walks on bond disordered cubic lattices. Journal of Statistical Physics, 48(3–4):645–676, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01019691.

Ernst:1988:RWS

[Ern88] Matthieu H. Ernst. Random walks with short memory. Journal of Statistical Physics, 53(1–2):191–201, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011552.

Eu:1984:CKC

[Eu84] Byung Chan Eu. Comment on Keizer’s critique on extended irreversible thermodynamics. Journal of Statistical Physics, 37(3–4): 485–490, November 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011846. See [Kei83, Kei84].

Evans:1980:ELT

[Eva80] Denis J. Evans. Enhanced t^{−3/2} long-time tail for the stressstress time correlation function. Journal of Statistical Physics, 22(1):81–90, January 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01007990.

Evans:1989:ENS

[Eva89] Denis J. Evans. On the entropy of nonequilibrium states. Journal of Statistical Physics, 57(3–4):745–758, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01022830.

Ernst:1981:CAA

[EvB81] M. H. Ernst and H. van Beijeren. On a conjecture of Alley and Alder for fluids and Lorentz models. Journal of Statistical Physics, 26(1):1–24, September 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01106782.

Ernst:1986:RWC

[EvV86] M. H. Ernst and P. F. J. van Velthoven. Random walks on cubic lattices with bond disorder. Journal of Statistical Physics, 45(5–6):1001–1030, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020586.

Ernst:1989:LTT

[EvV89] M. H. Ernst and G. A. van Velzen. Long-time tails in lattice Lorentz gases. Journal of Statistical Physics, 57(3–4):455–471, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01022816.

Erpenbeck:1981:MDC

[EW81] Jerome J. Erpenbeck and William W. Wood. Molecular dynamics calculations of shear viscosity time-correlation functions for hard spheres. Journal of Statistical Physics, 24(3):455–468, March 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012817.

Enting:1982:TLP

[EW82] I. G. Enting and F. Y. Wu. Triangular lattice Potts models. Journal of Statistical Physics, 28(2):351–373, June 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012610.

Erpenbeck:1984:MDC

[EW84] Jerome J. Erpenbeck and William W. Wood. Molecular dynamics calculations of the hard-sphere equation of state. Journal of Statistical Physics, 35(3–4):321–340, May 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01014387.

Erpenbeck:1985:EMD

[EW85] Jerome J. Erpenbeck and William W. Wood. Erratum: Molecular dynamics calculations of the hard sphere equation of state. Journal of Statistical Physics, 40(5–6):787, September 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009901.

Eckmann:1987:CPF

[EW87] Jean-Pierre Eckmann and Peter Wittwer. A complete proof of the Feigenbaum conjectures. Journal of Statistical Physics, 46(3– 4):455–475, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013368.

Eckmann:1988:LSI

[EW88] Jean-Pierre Eckmann and C. Eugene Wayne. Liapunov spectra for infinite chains of nonlinear oscillators. Journal of Statistical Physics, 50(5–6):853–878, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01019144.

Fain:1981:TRC

[Fai81] Benjamin Fain. Theory of rate constants: Master equation approach. Journal of Statistical Physics, 25(3):475–489, July 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010801.

Falconer:1987:HDS

[Fal87] K. J. Falconer. The Hausdorff dimension of some fractals and attractors of overlapping construction. Journal of Statistical Physics, 47(1–2):123–132, April 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009037.

Family:1984:FDG

[Fam84] Fereydoon Family. Fractal dimension and grand universality of critical phenomena. Journal of Statistical Physics, 36(5–6): 881–896, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012947.

Family:1988:BRR

[Fam88] Fereydoon Family. Book review: Renormalization group theory of macromolecules. Journal of Statistical Physics, 52(1–2):521–524, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01016435.

Faris:1987:LER

[Far87] William G. Faris. Localization estimates for a random discrete wave equation at high frequency. Journal of Statistical Physics, 46(3–4):477–491, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013369.

Ferrell:1985:WIC

[FB85a] Richard A. Ferrell and Jayanta K. Bhattacharhjee. Ward’s identity in critical dynamics. Journal of Statistical Physics, 41(5–6): 899–914, December 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010009.

Forrester:1985:FES

[FB85b] P. J. Forrester and R. J. Baxter. Further exact solutions of the eight-vertex SOS model and generalizations of the Rogers– Ramanujan identities. Journal of Statistical Physics, 38(3–4): 435–472, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010471.

Fujii:1986:ISB

[FBG86] D. H. Fujii, R. O. Barrachina, and C. R. Garibotti. Iterative solution of the Boltzmann equation. Journal of Statistical Physics, 44(1–2):95–106, July 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010906.

Felderhof:1989:GFL

[FC89] B. U. Felderhof and B. Cichocki. Generalization of the Foldy– Lax formula for the self-energy of a wave propagating in a disordered system of scatterers. Journal of Statistical Physics, 55 (5–6):1157–1168, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01041083.

Fleming:1986:ABC

[FCB86] Graham R. Fleming, Scott H. Courtney, and Michael W. Balk. Activated barrier crossing: Comparison of experiment and theory. Journal of Statistical Physics, 42(1–2):83–104, January 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010842.

Freeman:1986:FPI

[FCD86] D. L. Freeman, R. D. Coalson, and J. D. Doll. Fourier path integral methods: A model study of simple fluids. Journal of Statistical Physics, 43(5–6):931–934, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02628320.

Fraser:1983:SDC

[FCK83] Simon Fraser, Edward Celarier, and Raymond Kapral. Stochastic dynamics of the cubic map: A study of noise-induced transition phenomena. Journal of Statistical Physics, 33(2):341–370, November 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009801.

Finel:1986:UBM

[FdF86a] Alphonse Finel and D. de Fontaine. The uniaxial brickwork model, exact results, and CVM approximation. Journal of Statistical Physics, 43(3–4):663–670, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020658.

Finel:1986:TDA

[FdF86b] Alphonse Finel and Didier de Fontaine. The two-dimensional ANNNI model in the CVM approximation. Journal of Statistical Physics, 43(3–4):645–661, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020657.

Frank:1987:IPI

[FE87] B. Frank and M. Elofer. Initial perpendicular isothermal susceptibility formulas for the transverse general-spin Ising and Blume– Capel models. Journal of Statistical Physics, 46(3–4):559–565, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013375.

Fedchenia:1988:LMC

[Fed88a] I. I. Fedchenia. Local methods for constructing stationary distribution functions of systems of stochastic differential Langevintype equations: Noise influence on simple bifurcation. Journal of Statistical Physics, 50(5–6):1043–1068, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019152.

Fedchenia:1988:TDF

[Fed88b] I. I. Fedchenia. A two-dimensional Fokker–Planck equation degenerating on a straight line. Journal of Statistical Physics, 52(3– 4):1005–1029, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019737.

Feigenbaum:1985:ISU

[Fei85] Mitchell Feigenbaum. An interview with Stan Ulam and Mark Kac. Journal of Statistical Physics, 39(5–6):455–476, June 1985.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01008345. Reprinted from Los Alamos Science, Fall, 1982.

Feigenbaum:1987:SSR

[Fei87a] Mitchell J. Feigenbaum. Scaling spectra and return times of dynamical systems. Journal of Statistical Physics, 46(5–6):925–932, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011149.

Feigenbaum:1987:SCS

[Fei87b] Mitchell J. Feigenbaum. Some characterizations of strange sets. Journal of Statistical Physics, 46(5–6):919–924, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011148.

Feigenbaum:1988:PFF

[Fei88] Mitchell J. Feigenbaum. Presentation functions, fixed points, and a theory of scaling function dynamics. Journal of Statistical Physics, 52(3–4):527–569, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019716.

Felderhof:1986:TRW

[Fel86] B. U. Felderhof. Transmission and reflection of waves in a onedimensional disordered array. Journal of Statistical Physics, 43 (1–2):267–279, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010581.

Ferrell:1981:FTT

[Fer81] Richard A. Ferrell. Field theory of the two-dimensional Ising model: II. Nonlocal specific heat. Journal of Statistical Physics, 25(2):361–366, June 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01022191.

Ferrari:1988:IPS

[Fer88] Pablo A. Ferrari. Invariance principle for a solid-on-solid interface model. Journal of Statistical Physics, 51(5–6):1077–1090, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014900.

Felderhof:1986:GRD

[FF86] B. U. Felderhof and G. W. Ford. Growth of resistance with density of scatterers in one dimensional wave propagation. Journal of Statistical Physics, 45(3–4):695–714, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01021091.

Fichtner:1987:PPP

[FF87] K.-H. Fichtner and W. Freudenberg. Point processes and the position distribution of infinite boson systems. Journal of Statistical Physics, 47(5–6):959–978, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01206171.

Felderhof:1982:CED

[FFC82a] B. U. Felderhof, G. W. Ford, and E. G. D. Cohen. Cluster expansion for the dielectric constant of a polarizable suspension. Journal of Statistical Physics, 28(1):135–164, May 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011628.

Felderhof:1982:TPC

[FFC82b] B. U. Felderhof, G. W. Ford, and E. G. D. Cohen. Two-particle cluster integral in the expansion of the dielectric constant. Journal of Statistical Physics, 28(4):649–672, August 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011874.

Felderhof:1983:CMF

[FFC83] B. U. Felderhof, G. W. Ford, and E. G. D. Cohen. The Clausius– Mossotti formula and its nonlocal generalization for a dielectric suspension of spherical inclusions. Journal of Statistical Physics, 33(2):241–260, November 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009796.

Fernandez-Feria:1987:SFP

[FFdlM87] R. Fernandez-Feria and J. Fernandez de la Mora. Solution of the Fokker–Planck equation for the shock wave problem. Journal of Statistical Physics, 48(3–4):901–917, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019701.

Faetti:1988:POA

[FFG^{+}88] Sandro Faetti, Leone Fronzoni, Paolo Grigolini, Vincenzo Palleschi, and Girolamo Tropiano. The projection operator approach to the Fokker–Planck equation. II. Dichotomic and nonlinear Gaussian noise. Journal of Statistical Physics, 52(3–4): 979–1003, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019736.

Faetti:1988:PAF

[FFGM88] Sandro Faetti, Leone Fronzoni, Paolo Grigolini, and Riccardo Mannella. The projection approach to the Fokker–Planck equation. I. Colored Gaussian noise. Journal of Statistical Physics, 52 (3–4):951–978, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019735.

Fisher:1984:IMR

[FFS84] Daniel S. Fisher, Jürg Fröhlich, and Thomas Spencer. The Ising model in a random magnetic field. Journal of Statistical Physics, 34(5–6):863–870, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009445.

Fisher:1988:RTD

[FG88] Michael E. Fisher and Martin P. Gelfand. The reunions of three dissimilar vicious walkers. Journal of Statistical Physics, 53(1– 2):175–189, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011551.

Fisher:1989:RTD

[FG89] M. E. Fisher and M. P. Gelfand. The reunions of three dissimilar vicious walkers. Journal of Statistical Physics, 55(1–2):472, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01042615.

Fronzoni:1985:DOL

[FGMZ85] Leone Fronzoni, Paolo Grigolini, Riccardo Mannella, and Bruno Zambon. The Duffing oscillator in the low-friction limit: Theory and analog simulation. Journal of Statistical Physics, 41(3–4): 553–579, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009022.

Franceschini:1988:CPB

[FGN88] Valter Franceschini, Claudio Giberti, and Marco Nicolini. Common periodic behavior in larger and larger truncations of the Navier–Stokes equations. Journal of Statistical Physics, 50(5– 6):879–896, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019145.

Frigerio:1980:OQF

[FGP80] Alberto Frigerio, Vittorio Gorini, and Joe V. Pulé. Open quasifree systems. Journal of Statistical Physics, 22(4):409–433, April 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012865.

Farmer:1984:IST

[FGS84] J. Farmer, S. Goldstein, and E. R. Speer. Invariant states of a thermally conducting barrier. Journal of Statistical Physics, 34 (1–2):263–277, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01770358.

Fontaine:1985:EDC

[FGS85] J.-R. Fontaine, Ch. Gruber, and A. Sbaï. Exact decay of correlations for mixtures and rotators. Journal of Statistical Physics, 38 (5–6):891–899, March 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010422.

Frohlich:1985:PHC

[FH85] Jürg Fröhlich and Dale A. Huckaby. Percolation in hard-core lattice gases and a model ferrofluid. Journal of Statistical Physics, 38(5–6):809–821, March 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010416.

Fonk:1987:GSQ

[FH87] Y. Fonk and H. J. Hilhorst. Ground-state and quenched-state properties of a one-dimensional interacting lattice gas in a random potential. Journal of Statistical Physics, 49(5–6):1235–1254, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017569.

Figari:1988:LLN

[FHT88] R. Figari, H. Holden, and A. Tetab. A law of large numbers and a central limit theorem for the Schrödinger operator with zerorange potentials. Journal of Statistical Physics, 51(1–2):205–214, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01015327.

Fletcher:1988:FPT

[FHW88] John E. Fletcher, Shlomo Havlin, and George H. Weiss. First passage time problems in time-dependent fields. Journal of Statistical Physics, 51(1–2):215–232, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01015328.

Fife:1985:UPB

[Fif85] Paul C. Fife. Understanding the patterns in the BZ reagent. Journal of Statistical Physics, 39(5–6):687–703, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008360.

Frohlich:1980:PTR

[FILS80] Jürg Fröhlich, Robert B. Israel, Elliott H. Lieb, and Barry Simon. Phase transitions and reflection positivity. II. Lattice systems with short-range and Coulomb interactions. Journal of Statistical Physics, 22(3):297–347, March 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01014646.

Fisher:1984:BA

[Fis84a] Michael E. Fisher. The Boltzmann Award. Journal of Statistical Physics, 34(5–6):664–666, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009435.

Fisher:1984:WWW

[Fis84b] Michael E. Fisher. Walks, walls, wetting, and melting. Journal of Statistical Physics, 34(5–6):667–729, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009436.

Fisch:1988:ERP

[Fis88] Menachem Fisch. Essay review: A physicists’ philosopher — James Clerk Maxwell on mathematical physics. Journal of Statistical Physics, 51(1–2):309–319, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01015334.

Frobose:1986:SCH

[FJ86] K. Froböse and J. Jäckle. Static correlations of higher order and diffusion in an interacting lattice gas. Journal of Statistical Physics, 42(3–4):551–566, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01127727.

Forrester:1983:TDO

[FJS83] P. J. Forrester, B. Jancovici, and E. R. Smith. The twodimensional one-component plasma at Γ = 2: Behavior of correlation functions in strip geometry. Journal of Statistical Physics, 31(1):129–140, April 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010926.

Freed:1983:RGD

[FK83] Karl F. Freed and A. L. Kholodenko. Renormalization group description of polymer excluded volume. Journal of Statistical Physics, 30(2):437–447, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012317.

Ford:1987:QLE

[FK87] G. W. Ford and M. Kac. On the quantum Langevin equation. Journal of Statistical Physics, 46(5–6):803–810, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011142.

Feingold:1988:PST

[FKP88] Mario Feingold, Leo P. Kadanoff, and Oreste Piro. Passive scalars, three-dimensional volume-preserving maps, and chaos. Journal of Statistical Physics, 50(3–4):529–565, February 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01026490.

Fritter:1988:CSG

[FKRB88] Daniela Fritter, Charles M. Knobler, Didier Roux, and Daniel Beysens. Computer simulations of the growth of breath figures. Journal of Statistical Physics, 52(5–6):1447–1459, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011659.

Fritter:1989:CSG

[FKRB89] D. Fritter, C. M. Knobler, D. Roux, and D. Beysens. Computer simulations of the growth of breath figures. Journal of Statistical Physics, 54(1–2):563–564, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01023496.

Fraikin:1985:SAH

[FL85] A. Fraikin and H. Lemarchand. Stochastic analysis of a Hopf bifurcation: Master equation approach. Journal of Statistical Physics, 41(3–4):531–551, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009021.

Ferrari:1988:PCN

[FLM88] Pablo A. Ferrari, Joel L. Lebowitz, and Christian Maes. On the positivity of correlations in nonequilibrium spin systems. Journal of Statistical Physics, 53(1–2):295–305, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011558.

Forgacs:1988:ECB

[FLNO88] G. Forgacs, J. M. Luck, Th. M. Nieuwenhuizen, and H. Orland. Exact critical behavior of two-dimensional wetting problems with quenched disorder. Journal of Statistical Physics, 51(1–2):29–56, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01015319.

Ford:1988:IOM

[FLO88] G. W. Ford, J. T. Lewis, and R. F. O’Connell. Independent oscillator model of a heat bath: Exact diagonalization of the Hamiltonian. Journal of Statistical Physics, 53(1–2):439–455, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011565.

Feldmann:1986:GME

[FLS86] Tova Feldmann, R. D. Levine, and Peter Salamon. A geometrical measure for entropy changes. Journal of Statistical Physics, 42(5– 6):1127–1134, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010466.

Fontaine:1984:EES

[FM84] J. R. Fontaine and Ph. A. Martin. Equilibrium equations and symmetries of classical Coulomb systems. Journal of Statistical Physics, 36(1–2):163–179, July 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01015731.

Francfort:1987:OBC

[FM87] G. A. Francfort and G. W. Milton. Optimal bounds for conduction in two-dimensional, multiphase, polycrystalline media. Journal of Statistical Physics, 46(1–2):161–177, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010338.

Fritz:1988:DHE

[FM88] József Fritz and Christian Maes. Derivation of a hydrodynamic equation for Ginzburg–Landau models in an external field. Journal of Statistical Physics, 53(5–6):1179–1206, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023863.

Frankowicz:1983:TET

[FN83] M. Frankowicz and G. Nicolis. Transient evolution towards a unique stable state: Stochastic analysis of explosive behavior in a chemical system. Journal of Statistical Physics, 33(3):595–609, December 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01018836.

Ford:1989:CCM

[FO89] G. W. Ford and R. F. O’Connell. Canonical commutator and mass renormalization. Journal of Statistical Physics, 57(3–4): 803–810, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01022833.

Fontaine:1981:LFA

[Fon81] J. R. Fontaine. Low-fugacity asymptotic expansion for classical lattice dipole gases. Journal of Statistical Physics, 26(4):767–782, December 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010938.

Forrester:1984:IES

[For84] P. J. Forrester. Interpretation of an exactly solvable twocomponent plasma. Journal of Statistical Physics, 35(1–2):77–87, April 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017365.

Forrester:1986:CRP

[For86a] P. J. Forrester. Charged rods in a periodic background: A solvable model. Journal of Statistical Physics, 42(5–6):871–894, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010450.

Forrester:1986:PNC

[For86b] P. J. Forrester. Positive and negative charged rods alternating along a line: Exact results. Journal of Statistical Physics, 45(1– 2):153–169, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01033084.

Forrester:1988:SIT

[For88] P. J. Forrester. Solvable isotherms for a two-component system of charged rods on a line. Journal of Statistical Physics, 51(3–4):457–479, May 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01028466.

Forrester:1989:ERC

[For89a] P. J. Forrester. Exact results for correlations in a two-component log-gas. Journal of Statistical Physics, 54(1–2):57–79, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023473.

Forrester:1989:VRW

[For89b] P. J. Forrester. Vicious random walkers in the limit of a large number of walkers. Journal of Statistical Physics, 56(5–6): 767–782, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016779.

Figari:1985:LRM

[FOT85] R. Figari, E. Orlandi, and S. Teta. The Laplacian in regions with many small obstacles: Fluctuations around the limit operator. Journal of Statistical Physics, 41(3–4):465–487, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009018.

Fox:1986:MCP

[FOU86] G. C. Fox, S. W. Otto, and E. A. Umland. Monte Carlo physics on a concurrent processor. Journal of Statistical Physics, 43(5– 6):1209–1237, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02628341.

Fox:1983:BRI

[Fox83] Ronald F. Fox. Book review: Instabilities, bifurcations, and fluctuations in chemical systems. Journal of Statistical Physics, 32(1):205–206, July 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009430.

Fox:1987:SCP

[Fox87] Ronald F. Fox. Stochastic calculus in physics. Journal of Statistical Physics, 46(5–6):1145–1157, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011160.

Fox:1989:NSS

[Fox89] Ronald F. Fox. Numerical simulations of stochastic differential equations. Journal of Statistical Physics, 54(5–6):1353–1366, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01044719.

Frohlich:1980:CIT

[FP80] Jürg Fröhlich and Yong Moon Park. Correlation inequalities and the thermodynamic limit for classical and quantum continuous systems II. Bose–Einstein and Fermi–Dirac statistics. Journal of Statistical Physics, 23(6):701–753, December 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008516.

Forgacs:1987:DCL

[FP87] G. Forgacs and V. Privman. Directed compact lattice animals: Exact results. Journal of Statistical Physics, 49(5–6):1165–1180, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017564.

Franceschini:1981:SUM

[FR81] Valter Franceschini and Lucio Russo. Stable and unstable manifolds of the Hénon mapping. Journal of Statistical Physics, 25 (4):757–769, August 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01022365.

Franceschini:1980:FSB

[Fra80] Valter Franceschini. A Feigenbaum sequence of bifurcations in the Lorenz model. Journal of Statistical Physics, 22(3):397–406, March 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014649.

France:1986:ICN

[Fra86] M. Mendès France. The Ising chain with nonconstant external field. Journal of Statistical Physics, 45(1–2):89–97, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01033079.

Frezzotti:1987:PID

[Fre87] A. Frezzotti. The propagation of infinitesimal disturbances in an ultrarelativistic gas according to the method of elementary solutions. Journal of Statistical Physics, 46(1–2):255–272, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010345.

Fritz:1984:SRN

[Fri84] J. Fritz. Some remarks on nonequilibrium dynamics of infinite particle systems. Journal of Statistical Physics, 34(3–4):539–556, February 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01018557.

Fritz:1985:ABS

[Fri85] J. Fritz. On the asymptotic behaviour of Spitzer’s model for evolution of one-dimensional point systems. Journal of Statistical Physics, 38(3–4):615–645, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010481.

Fritz:1986:SMA

[Fri86] J. Fritz. On the stationary measures of anharmonic systems in the presence of a small thermal noise. Journal of Statistical Physics, 44(1–2):25–47, July 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010903.

Fritz:1987:HLO

[Fri87] J. Fritz. On the hydrodynamic limit of a one-dimensional Ginzburg–Landau lattice model. The a priori bounds. Journal of Statistical Physics, 47(3–4):551–572, May 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01007526.

Frobose:1989:FSE

[Fro89] K. Froböse. Finite-size effects in a cellular automaton for diffusion. Journal of Statistical Physics, 55(5–6):1285–1292, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01041088.

Fox:1987:TNS

[FRY87] Ronald F. Fox, Rajarshi Roy, and A. W. Yu. Tests of numerical simulation algorithms for the Kubo oscillator. Journal of Statistical Physics, 47(3–4):477–487, May 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01007521.

Frohlich:1981:SMC

[FS81] Jürg Fröhlich and Thomas Spencer. On the statistical mechanics of classical Coulomb and dipole gases. Journal of Statistical Physics, 24(4):617–701, April 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011379.

Frohlich:1986:LDN

[FSW86] Jürg Fröhlich, Thomas Spencer, and C. Eugene Wayne. Localization in disordered, nonlinear dynamical systems. Journal of Statistical Physics, 42(3–4):247–274, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01127712.

Franceschini:1981:SMT

[FT81] Valter Franceschini and Claudio Tebaldi. A seven-mode truncation of the plane incompressible Navier–Stokes equations. Journal of Statistical Physics, 25(3):397–417, July 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010796.

Forrester:1988:MPM

[FT88] P. J. Forrester and Colin J. Thompson. Modulated phases and the mean-field theory of magnetism with competing interactions. Journal of Statistical Physics, 51(3–4):519–536, May 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01028470.

Franceschini:1984:FPL

[FTZ84] Valter Franceschini, Claudio Tebaldi, and Fernando Zironi. Fixed point limit behavior of n-mode truncated Navier–Stokes equations as n increases. Journal of Statistical Physics, 35(3–4):387– 397, May 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014392.

Frosali:1989:STR

[FvdM89] Giovanni Frosali and Cornelis V. M. van der Mee. Scattering theory relevant to the linear transport of particle swarms. Journal of Statistical Physics, 56(1–2):139–148, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01044237.

Fannes:1982:EEI

[FVV82a] M. Fannes, P. Vanheuverzwijn, and A. Verbeure. Energy-entropy inequalities for classical lattice systems. Journal of Statistical Physics, 29(3):547–560, November 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01342186.

Fannes:1982:IMF

[FVV82b] M. Fannes, P. Vanheuverzwijn, and A. Verbeure. Inhomogeneous mean field models. Journal of Statistical Physics, 28(2):381–389, June 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012612.

Fye:1986:STL

[Fye86] R. M. Fye. Study of trotter-like approximations. Journal of Statistical Physics, 43(5–6):827–829, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02628308.

Franceschini:1985:CMP

[FZ85] Valter Franceschini and Fernando Zironi. On constructing Markov partitions by computer. Journal of Statistical Physics, 40(1–2):69–91, July 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010527.

Gardiner:1983:ETN

[Gar83] C. W. Gardiner. The escape time in nonpotential systems. Journal of Statistical Physics, 30(1):157–177, January 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010872.

Gardiner:1987:DTT

[Gar87] C. W. Gardiner. Diffusive traversal time of a one-dimensional medium. Journal of Statistical Physics, 49(3–4):827–828, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009358.

Gates:1988:GDT

[Gat88] D. J. Gates. Growth and decrescence of two-dimensional crystals: A Markov rate process. Journal of Statistical Physics, 52 (1–2):245–257, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016412.

Gouyet:1984:DBP

[GB84] J. F. Gouyet and A. Bunde. Diffusion in a bistable potential at intermediate and high friction. Journal of Statistical Physics, 36(1–2):43–64, July 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01015725.

Gallavotti:1986:AIC

[GBD^{+}86] G. Gallavotti, J. Thomas Beale, Ronald J. DiPerna, M. Dubois, J.-P. Eckmann, Valter Franceschini, Giovanni Gallavotti, M. Ghil, M. Giglio, N. Kopell, Oscar E. Lanford III, Joel L. Lebowitz, B. Legras, A. Libchaber, A. Mullhaupt, and et al. Abstracts from the international conference on mathematical problems from the physics of fluids. Journal of Statistical Physics, 44(5–6): 1007–1067, September 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011921.

Gora:1988:CAD

[GBP88a] P. Góra, A. Boyarsky, and H. Proppe. Constructive approximations to densities invariant under nonexpanding transformations. Journal of Statistical Physics, 51(1–2):179–194, April 1988.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01015325.

Grassberger:1988:SLI

[GBP88b] P. Grassberger, R. Badii, and A. Politi. Scaling laws for invariant measures on hyperbolic and nonhyperbolic attractors. Journal of Statistical Physics, 51(1–2):135–178, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01015324.

Garcia-Colin:1984:FEI

[GCdHR^{+}84] L. S. Garcia-Colin, M. López de Haro, R. F. Rodriguez, J. Casas-Vázquez, and D. Jou. On the foundations of extended irreversible thermodynamics. Journal of Statistical Physics, 37(3–4):465–484, November 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011845.

Garcia-Colin:1982:KDE

[GCF82] L. S. García-Colin and G. J. Fuentes y Martínez. A kinetic derivation of extended irreversible thermodynamics. Journal of Statistical Physics, 29(2):387–396, October 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020794.

Gardner:1985:ZTM

[GD85] E. Gardner and B. Derrida. Zero temperature magnetization of a one-dimensional spin glass. Journal of Statistical Physics, 39 (3–4):367–377, May 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01018668.

Golden:1982:DTP

[GDx82] Kenneth I. Golden and Lu De-xin. Dynamical three-point correlations and quadratic response functions in binary ionic mixture plasmas. Journal of Statistical Physics, 29(2):281–307, October 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020787.

Geerts:1982:NNF

[Gee82] H. Geerts. A note on number fluctuations: Statistics of fluorescence correlation spectroscopy as applied to Brownian motion.

Journal of Statistical Physics, 28(1):173–176, May 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011630.

Georgii:1981:SMR

[Geo81] Hans-Otto Georgii. Spontaneous magnetization of randomly dilute ferromagnets. Journal of Statistical Physics, 25(3):369–396, July 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010795.

Gaveau:1982:EPT

[GG82] Marc André Gaveau and Bernard Gaveau. An example of a phase transition in an open pseudobimolecular system. Journal of Statistical Physics, 28(4):745–756, August 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011878.

Griffiths:1983:CVN

[GG83] Robert B. Griffiths and P. D. Gujrati. Convexity violations for noninteger parameters in certain lattice models. Journal of Statistical Physics, 30(3):563–589, March 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009677.

Goldstein:1986:DFM

[GG86] S. Goldstein and J. Guetti. On the diffusion of a fast molecule. Journal of Statistical Physics, 43(1–2):321–341, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010585.

Gorecki:1987:ATE

[GG87] J. Gorecki and J. Gryko. The adiabatic thermal explosion in a small system: Comparison of the stochastic approach with the molecular dynamics simulation. Journal of Statistical Physics, 48(1–2):329–342, July 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010412.

Golden:1988:ASD

[GG88] K. Golden and S. Goldstein. Arbitrarily slow decay of correlations in quasiperiodic systems. Journal of Statistical Physics, 52(3–4):

1113–1118, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019742.

Golden:1988:DPP

[GGL88] K. Golden, S. Goldstein, and J. L. Lebowitz. Diffusion in a periodic potential with a local perturbation. Journal of Statistical Physics, 51(3–4):637–656, May 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01028476.

Goodrich:1986:FI

[GGM86] R. K. Goodrich, K. Gustafson, and B. Misra. On K-flows and irreversibility. Journal of Statistical Physics, 43(1–2):317–320, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010584.

Gupta:1986:FSI

[GGP86] Rajan Gupta, G. Guralnik, and A. Patel. The β function, scaling, and improved action for SU(3) lattice gauge theory. Journal of Statistical Physics, 43(5–6):1101–1104, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02628334.

Garrod:1982:SPN

[GH82] Claude Garrod and James P. Hurley. Symmetry properties of nonlinear barrier coefficients. Journal of Statistical Physics, 29 (1):33–37, September 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008246.

Gaveau:1989:STCa

[GHKM89a] Bernard Gaveau, James T. Hynes, Raymond Kapral, and Michel Moreau. A stochastic theory of chemical reaction rates. I. Formalism. Journal of Statistical Physics, 56(5–6):879–893, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016783.

Gaveau:1989:STCb

[GHKM89b] Bernard Gaveau, James T. Hynes, Raymond Kapral, and Michel Moreau. A stochastic theory of chemical reaction rates. II. Applications. Journal of Statistical Physics, 56(5–6):895–910, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01016784.

Grabert:1980:MNS

[GHT80] H. Grabert, P. Hänggi, and P. Talkner. Microdynamics and nonlinear stochastic processes of gross variables. Journal of Statistical Physics, 22(5):537–552, May 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011337.

Gibson:1981:TVC

[Gib81] W. G. Gibson. Third virial coefficient for quantum hard spheres: Two-point Padé approximants for direct and exchange parts. Journal of Statistical Physics, 26(2):333–346, October 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013175.

Gidas:1985:NMC

[Gid85] Basilis Gidas. Nonstationary Markov chains and convergence of the annealing algorithm. Journal of Statistical Physics, 39 (1–2):73–131, April 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01007975.

Gielerak:1989:ETM

[Gie89] Roman Gielerak. Existence of the transfer matrix formalism for a class of classical continuous gases. Journal of Statistical Physics, 55(1–2):183–201, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01042597.

Gitterman:1988:BRO

[Git88] M. Gitterman. Book review: Order within chaos. Journal of Statistical Physics, 50(5–6):1283–1284, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019169.

Gitterman:1989:BRC

[Git89] Moshe Gitterman. Book review: Chaos and integrability in nonlinear dynamics. Journal of Statistical Physics, 57(5–6):

1137–1138, December 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020053.

Gawcedzki:1982:TAH

[GK82] K. Gawcedzki and A. Kupiainen. Triviality of ϕ_{4}^{4} and all that in a hierarchical model approximation. Journal of Statistical Physics, 29(4):683–698, December 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011785.

Gawedzki:1984:NGS

[GK84] K. Gawedzki and A. Kupiainen. Non-Gaussian scaling limits. Hierarchical model approximation. Journal of Statistical Physics, 35(3–4):267–284, May 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014384.

Goldsterin:1985:SSM

[GKI85] S. Goldsterin, C. Kipnis, and N. Ianiro. Stationary states for a mechanical system with stochastic boundary conditions. Journal of Statistical Physics, 41(5–6):915–939, December 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010010.

Gawedzki:1987:CGA

[GKK87] K. Gawedzki, R. Kotecký, and A. Kupiainen. Coarse-graining approach to first-order phase transitions. Journal of Statistical Physics, 47(5–6):701–724, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01206154.

Gaspard:1984:BPN

[GKN84] P. Gaspard, R. Kapral, and G. Nicolis. Bifurcation phenomena near homoclinic systems: A two-parameter analysis. Journal of Statistical Physics, 35(5–6):697–727, June 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010829.

Gawedzki:1984:BSP

[GKT84] K. Gawedzki, A. Kupiainen, and B. Tirozzi. Borel summability of the perturbation series in a hierarchical λ(∇ϕ)^{4} model. Journal of Statistical Physics, 36(1–2):145–162, July 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01015730.

Gottlieb:1986:DPT

[GKT^{+}86] S. A. Gottlieb, J. Kuti, D. Toussaint, A. D. Kennedy, S. Meyer, B. J. Pendleton, and R. L. Sugar. The deconfining phase transition in lattice quantum chromodynamics. Journal of Statistical Physics, 43(5–6):1105–1115, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02628335.

Gelbart:1988:P

[GL88] William M. Gelbart and Joel L. Lebowitz. Preface. Journal of Statistical Physics, 52(5–6):1141–1142, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011637.

Georges:1989:ESM

[GL89a] Antoine Georges and Pierre Le Doussal. From equilibrium spin models to probabilistic cellular automata. Journal of Statistical Physics, 54(3–4):1011–1064, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019786.

Godreche:1989:QRT

[GL89b] C. Godrèche and J. M. Luck. Quasiperiodicity and randomness in tilings of the plane. Journal of Statistical Physics, 55(1–2):1–28, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01042590.

Glatzmaier:1985:MFP

[Gla85] Gary A. Glatzmaier. Magnetic field propagation in a stellar dynamo. Journal of Statistical Physics, 39(5–6):493–499, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008347.

Glaus:1988:MCS

[Gla88] U. Glaus. Monte Carlo study of self-avoiding surfaces. Journal of Statistical Physics, 50(5–6):1141–1166, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019158.

Glick:1983:BR

[Gli83] Arnold J. Glick. Book review. Journal of Statistical Physics, 30 (1):243–244, January 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010877.

Gruber:1980:EPC

[GLM80] Ch. Gruber, Ch. Lugrin, and Ph. A. Martin. Equilibrium properties of classical systems with long-range forces. BBGKY equation, neutrality, screening, and sum rules. Journal of Statistical Physics, 22(2):193–236, February 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008049.

Garrido:1987:SNS

[GLM87] P. L. Garrido, A. Labarta, and J. Marro. Stationary nonequilibrium states in the Ising model with locally competing temperatures. Journal of Statistical Physics, 49(3–4):551–568, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009348.

Godreche:1986:MPS

[GLO86] C. Godréche, J. M. Luck, and H. Orland. Magnetic phase structure on the Penrose lattice. Journal of Statistical Physics, 45(5– 6):777–800, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020574.

Goldstein:1986:AEM

[GLR86] S. Goldstein, J. L. Lebowitz, and K. Ravishankar. Approach to equilibrium in models of a system in contact with a heat bath. Journal of Statistical Physics, 43(1–2):303–315, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010583.

Gaveau:1983:EDD

[GM83] Bernard Gaveau and Alain Meritet. Example of diffusion in a disordered Lorentz gas. Journal of Statistical Physics, 32(3): 539–544, September 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008954.

Given:1984:NMP

[GM84] James A. Given and Benoît B. Mandelbrot. A new model of percolation clusters. Journal of Statistical Physics, 36(5–6):545, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01012920;   http://link.
springer.com/article/10.1007/BF01012920;   http://www.
springerlink.com/content/l745uv6w7n3m8835/.

Ghil:1985:BDE

[GM85] M. Ghil and A. Mullhaupt. Boolean delay equations. II. Periodic and aperiodic solutions. Journal of Statistical Physics, 41 (1–2):125–173, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020607.

George:1987:VSL

[GM87] C. George and F. Mayné. Vacuum subdynamics in large quantum systems. Journal of Statistical Physics, 48(5–6):1343– 1361, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009549.

Geigenmuller:1988:SHS

[GM88a] U. Geigenmüller and P. Mazur. Sedimentation of homogeneous suspensions in finite vessels. Journal of Statistical Physics, 53(1– 2):137–173, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011550.

Goles:1988:OSD

[GM88b] Eric Goles and Servet Martinez. The one-site distribution of Gibbs states on Bethe lattice are probability vectors of period ≤ 2 for a nonlinear transformation. Journal of Statistical Physics, 52(1–2):267–285, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016414. See erratum [GM96].

Goles:1996:EOS

[GM96] Eric Goles and Servet Martínez. Erratum: The one-site distribution of Gibbs states on Bethe lattice are probability vectors of period ≤ 2 for a nonlinear transformation. Journal of Statistical Physics, 82(1–2):451, January 1996. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02189240.

Gefen:1984:PDS

[GMAK84] Yuval Gefen, B. B. Mandelbrot, A. Aharony, and A. Kapitulnik. Partial-dimensional sequences and percolation. Journal of Statistical Physics, 36(5–6):827–830, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012942;   http://www.springerlink.com/content/j815gu7668r33v43/

.

Goldstein:1981:IRD

[GMC81] S. Goldstein, B. Misra, and M. Courtage. On intrinsic randomness of dynamical systems. Journal of Statistical Physics, 25(1):111– 126, May 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01008481.

Garcia:1987:NIF

[GMLC87] Alejandro L. Garcia, M. Malek Mansour, George C. Lie, and Enrico Cementi. Numerical integration of the fluctuating hydrodynamic equations. Journal of Statistical Physics, 47(1–2):209–228, April 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009043.

Galves:1989:LDF

[GMO89] Antonio Galves, Fabio Martinelli, and Enzo Olivieri. Largedensity fluctuations for the one-dimensional supercritical contact process. Journal of Statistical Physics, 55(3–4):639–648, May 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01041602.

Galves:1989:FDR

[GMP89] Antonio Galves, Servet Martinez, and Pierre Picco. Fluctuations in Derrida’s random energy and generalized random energy models. Journal of Statistical Physics, 54(1–2):515–529, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023492.

Gaspard:1983:WCW

[GN83] P. Gaspard and G. Nicolis. What can we learn from homoclinic orbits in chaotic dynamics? Journal of Statistical Physics, 31 (3):499–518, June 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019496.

Gallavotti:1985:SPT

[GN85] G. Gallavotti and F. Nicoló. The ‘screening phase transitions’ in the two-dimensional Coulomb gas. Journal of Statistical Physics, 39(1–2):133–156, April 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01007976.

Goldhirsch:1987:AFR

[GN87] I. Goldhirsch and S. H. Noskowicz. Anomalous fluctuations in random walk dynamics. Journal of Statistical Physics, 48 (1–2):291–314, July 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010410.

Gudowska-Nowak:1989:SVC

[GNKW89] E. Gudowska-Nowak, A. Kleczkowski, and G. O. Williams. Stochastic versus chaotic dynamics in a deterministic system. Journal of Statistical Physics, 54(1–2):539–560, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023494.

Go:1983:PFS

[Go83] Nobuhiro Go. Protein folding as a stochastic process. Journal of Statistical Physics, 30(2):413–423, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012315.

Goldstein:1985:LTD

[Gol85] M. S. Goldstein. Limit theorem for the distribution of eigenvalues of the operator of energy. Journal of Statistical Physics, 40 (1–2):329–360, July 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010539.

Goldstein:1987:SMQ

[Gol87] Sheldon Goldstein. Stochastic mechanics and quantum theory. Journal of Statistical Physics, 47(5–6):645–667, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01206150.

Gubernatis:1986:DMQ

[GOSS86] J. E. Gubernatis, T. C. Olson, D. J. Scalapino, and R. L. Sugar. The determinant method and quantum simulations of many-body effects in a single impurity Anderson model. Journal of Statistical Physics, 43(5–6):831–840, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02628309.

Gouyet:1986:DBP

[Gou86] J.-F. Gouyet. Diffusion in a bistable potential. General inverse friction expansion. Journal of Statistical Physics, 45(1– 2):267–293, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01033091.

Goldstein:1981:NED

[GP81] S. Goldstein and O. Penrose. A nonequilibrium entropy for dynamical systems. Journal of Statistical Physics, 24(2):325–343, February 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013304.

Goldberg:1983:MDD

[GP83] Abraham Goldberg and Robert D. Puff. Medium-dependent dynamics in Fermi systems from a pair-composite viewpoint. Journal of Statistical Physics, 33(3):571–594, December 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01018835.

Golden:1985:BEP

[GP85a] K. Golden and G. Papanicolaou. Bounds for effective parameters of multicomponent media by analytic continuation. Journal of Statistical Physics, 40(5–6):655–667, September 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009895.

Gredeskul:1985:WML

[GP85b] S. A. Gredeskul and L. A. Pastur. Works of I. M. Lifshitz on disordered systems. Journal of Statistical Physics, 38(1–2):25–36, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017846.

Gervois:1986:SVD

[GP86a] A. Gervois and J. Piasecki. Stationary velocity distribution in an external field: A one-dimensional model. Journal of Statistical Physics, 42(5–6):1091–1102, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010463.

Gorecki:1986:STA

[GP86b] J. Gorecki and J. Popielawski. On the stochastic theory of adiabatic thermal explosion in small systems- numerical results. Journal of Statistical Physics, 44(5–6):941–954, September 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011916.

Golovko:1989:ASM

[GP89a] M. F. Golovko and I. A. Protsykevich. Analytic solution of the mean spherical approximation for ion-dipole model in a neutralizing background. Journal of Statistical Physics, 54(3–4):707–733, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019772.

Greenberg:1989:GET

[GP89b] William Greenberg and Jacek Polewczak. A global existence theorem for the nonlinear BGK equation. Journal of Statistical Physics, 55(5–6):1313–1321, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01041091.

Glasser:1986:CTP

[GPS86] M. L. Glasser, V. Privman, and L. S. Schulman. Complex temperature plane zeros in the mean-field approximation. Journal of Statistical Physics, 45(3–4):451–457, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01021081.

Giacomini:1987:ERR

[GR87] H. J. Giacomini and J. A. Riera. Exact results for a random frustrated Ising model on the Kagomé lattice. Journal of Statistical Physics, 49(3–4):491–509, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009346.

Grasman:1989:SCR

[GR89] J. Grasman and J. B. T. M. Roerdink. Stochastic and chaotic relaxation oscillations. Journal of Statistical Physics, 54(3–4): 949–970, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019783.

Grabert:1981:FPE

[Gra81a] Hermann Grabert. Fokker–Planck equation approach to fluctuations about nonequilibrium steady states. Journal of Statistical Physics, 26(1):113–135, September 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01106789.

Grassberger:1981:HDF

[Gra81b] P. Grassberger. On the Hausdorff dimension of fractal attractors. Journal of Statistical Physics, 26(1):173–179, September 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01106792.

Graham:1982:CIT

[Gra82a] Ross Graham. Correlation inequalities for the truncated twopoint function of an Ising ferromagnet. Journal of Statistical Physics, 29(2):177–183, October 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020780.

Graham:1982:IGH

[Gra82b] Ross Graham. An improvement of the Griffiths–Hurst–Sherman inequality for the Ising ferromagnet. Journal of Statistical Physics, 29(2):185–191, October 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020781.

Graham:1984:GIR

[Gra84] Ross Graham. A geometric interpretation of the roughening transition in Ising models. Journal of Statistical Physics, 35(5–6):473– 488, June 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010821.

Grassberger:1986:LRE

[Gra86] Peter Grassberger. Long-range effects in an elementary cellular automaton. Journal of Statistical Physics, 45(1–2):27–39, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01033074.

Graham:1989:BUW

[Gra89] Robert Graham. Bifurcations under weak noise. Journal of Statistical Physics, 54(5–6):1207–1215, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01044712.

Greenberg:1982:BRT

[Gre82a] O. W. Greenberg. Book review: Techniques and applications of path integration. Journal of Statistical Physics, 27(4):827, April 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013452.

Greer:1982:BR

[Gre82b] Sandra C. Greer. Book review. Journal of Statistical Physics, 29 (3):645–646, November 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01342195.

Greer:1985:BRS

[Gre85] Sandra C. Greer. Book review: Statistical mechanics. Journal of Statistical Physics, 41(5–6):1045–1046, December 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010015.

Greer:1986:BRT

[Gre86] Sandra C. Greer. Book review: The theory of thermodynamics. Journal of Statistical Physics, 44(3–4):701–702, August 1986.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011314.

Greer:1989:BRS

[Gre89] Sandra C. Greer. Book review: Statistical mechanics. Journal of Statistical Physics, 56(3–4):557–558, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01044454.

Grigolini:1982:GLE

[Gri82] Paolo Grigolini. A generalized Langevin equation for dealing with nonadditive fluctuations. Journal of Statistical Physics, 27(2): 283–316, February 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008940.

Griffiths:1984:CHI

[Gri84] Robert B. Griffiths. Consistent histories and the interpretation of quantum mechanics. Journal of Statistical Physics, 36(1–2):219– 272, July 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01015734.

Grigolini:1988:BRI

[Gri88a] Paolo Grigolini. Book review: Instabilities and chaos in quantum optics. Journal of Statistical Physics, 52(3–4):1137, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01019747.

Grinstein:1988:SNS

[Gri88b] G. Grinstein. Stability of nonstationary states of classical, manybody dynamical systems. Journal of Statistical Physics, 51(5– 6):803–815, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014885.

Gross:1981:MSRb

[Gro81a] Eugene P. Gross. Multiple scattering in random media. Journal of Statistical Physics, 25(4):605–618, August 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01022356.

Gross:1981:MSRa

[Gro81b] Eugene P. Gross. Multiple scattering in random media I. Restricted two-body additive approximation. Journal of Statistical Physics, 25(4):585–604, August 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01022355.

Gross:1981:MSRc

[Gro81c] Eugene P. Gross. Multiple scattering in random media. III. Coherent potential propagators and fluctuations. Journal of Statistical Physics, 26(2):241–255, October 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013170.

Gross:1981:ASO

[Gro81d] Leonard Gross. Absence of second-order phase transitions in the Dobrushin uniqueness region. Journal of Statistical Physics, 25(1):57–72, May 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008479.

Gross:1983:LBW

[Gro83a] Eugene P. Gross. Lower bounds for Wiener integrals. Journal of Statistical Physics, 31(1):115–128, April 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010925.

Gross:1983:PFP

[Gro83b] Eugene P. Gross. Partition function of a particle subject to Gaussian noise. Journal of Statistical Physics, 33(1):107–132, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009752.

Gross:1983:UBD

[Gro83c] Eugene P. Gross. Upper bounds for density matrices using path integrals. Journal of Statistical Physics, 30(1):45–66, January 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010868.

Gross:1984:FEP

[Gro84] Eugene P. Gross. Functional equations for path integrals. Journal of Statistical Physics, 36(1–2):189–217, July 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01015733.

Grosberg:1985:TCS

[Gro85] A. Yu. Grosberg. On the theory of the condensed states of heteropolymers. Journal of Statistical Physics, 38(1–2):149–160, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017854.

Gross:1989:GSSa

[Gro89a] Eugene P. Gross. Ground state of a spin-phonon system. I. Variational estimates. Journal of Statistical Physics, 54(1–2):405–427, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023486.

Gross:1989:GSSb

[Gro89b] Eugene P. Gross. Ground state of a spin-phonon system. II. Adiabatic limit. Journal of Statistical Physics, 54(1–2):429–436, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023487.

Gross:1989:GSSc

[Gro89c] Eugene P. Gross. Ground state of a spin-phonon system. III. Small-b limit. Journal of Statistical Physics, 54(1–2):437–448, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023488.

Gollub:1980:TDC

[GRS80] J. P. Gollub, E. J. Romer, and J. E. Socolar. Trajectory divergence for coupled relaxation oscillators: Measurements and models. Journal of Statistical Physics, 23(3):321–333, September 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011372.

Grassberger:1981:SMU

[GS81] P. Grassberger and M. Scheunert. Some more universal scaling laws for critical mappings. Journal of Statistical Physics, 26(4): 697–717, December 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010934.

Glendinning:1984:LGB

[GS84a] Paul Glendinning and Colin Sparrow. Local and global behavior near homoclinic orbits. Journal of Statistical Physics, 35 (5–6):645–696, June 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010828.

Gyorgyi:1984:PFD

[GS84b] G. Györgyi and P. Szépfalusy. Properties of fully developed chaos in one-dimensional maps. Journal of Statistical Physics, 34(3– 4):451–475, February 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01018554.

Ghosh:1985:EER

[GS85] Ranjan K. Ghosh and Robert E. Shrock. Exact expressions for row correlation functions in the isotropic d = 2 Ising model. Journal of Statistical Physics, 38(3–4):473–482, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010472.

Glendinning:1986:PCT

[GS86] Paul Glendinning and Colin Sparrow. T -points: A codimension two heteroclinic bifurcation. Journal of Statistical Physics, 43 (3–4):479–488, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020649.

Gruber:1988:PDL

[GS88] Christian Gruber and András Sütö. Phase diagrams of lattice systems with residual entropy. Journal of Statistical Physics, 52 (1–2):113–142, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016407.

Greiner:1988:NIS

[GSH88] A. Greiner, W. Strittmatter, and J. Honerkamp. Numerical integration of stochastic differential equations. Journal of Statistical Physics, 51(1–2):95–108, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01015322;   http:
//www.springerlink.com/content/k655227v66316102/.

Gambaudo:1983:SMB

[GT83] J. M. Gambaudo and C. Tresser. Simple models for bifurcations creating horseshoes. Journal of Statistical Physics, 32(3): 455–476, September 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008950.

Graham:1984:WNLa

[GT84a] R. Graham and T. Tél. On the weak-noise limit of Fokker– Planck models. Journal of Statistical Physics, 35(5–6):729–748, June 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010830.

Graham:1984:WNLb

[GT84b] R. Graham and T. Tél. On the weak-noise limit of Fokker–Planck models. Journal of Statistical Physics, 37(5–6):709, December 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010503.

Gubernatis:1986:P

[Gub86] James E. Gubernatis. Preface. Journal of Statistical Physics, 43 (5–6):729–730, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02628300.

Gujrati:1982:LBE

[Guj82] P. D. Gujrati. Lower bounds on entropy for polymer chains on a square and a cubic lattice. Journal of Statistical Physics, 28(3):441–472, July 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008317.

Gunton:1984:DRI

[Gun84] J. D. Gunton. The dynamics of random interfaces in phase transitions. Journal of Statistical Physics, 34(5–6):1019–1037, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009455.

Guttman:1984:MCS

[Gut84] C. M. Guttman. Monte Carlo studies of two measures of polymer chain size as a function of temperature. Journal of Statistical Physics, 36(5–6):717–733, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012934.

Gutkin:1986:CBC

[Gut86] Eugene Gutkin. A comment on Baxter condition for commutativity of transfer matrices. Journal of Statistical Physics, 44 (1–2):193–202, July 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010911.

Ghez:1989:WAM

[GV89a] J.-M. Ghez and S. Vaienti. On the wavelet analysis for multifractal sets. Journal of Statistical Physics, 57(1–2):415–420, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023655.

Gutowitz:1989:LST

[GV89b] Howard A. Gutowitz and Jonathan D. Victor. Local structure theory: Calculation on hexagonal arrays, and interaction of rule and lattice. Journal of Statistical Physics, 54(1–2):495–514, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023491.

Geurts:1989:CBC

[GvD89] Bernard J. Geurts and Ruud van Damme. Complexes of block copolymers in solution: Tree approximation. Journal of Statistical Physics, 57(5–6):1069–1098, December 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020049.

Goderis:1989:TQF

[GVV89] D. Goderis, A. Verbeure, and P. Vets. Theory of quantum fluctuations and the Onsager relations. Journal of Statistical Physics, 56(5–6):721–746, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01016777.

Gaspard:1987:HOM

[GW87a] P. Gaspard and X. J. Wang. Homoclinic orbits and mixed-mode oscillations in far-from-equilibrium systems. Journal of Statistical Physics, 48(1–2):151–199, July 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010405.

Glaffig:1987:IDB

[GW87b] Clemens Glaffig and Ed Waymire. Infinite divisibility of a Bethe lattice Ising model. Journal of Statistical Physics, 47(1–2):185– 192, April 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009041.

Gupta:1986:IMC

[GWU86] Rajan Gupta, K. G. Wilson, and C. Umrigar. Improved Monte Carlo renormalization group method. Journal of Statistical Physics, 43(5–6):1095–1099, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02628333.

Gzyl:1981:ESE

[Gzy81] Henryk Gzyl. On the equivalence of some exact master equations. Journal of Statistical Physics, 26(4):803–805, December 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010941.

Gzyl:1982:IBE

[Gzy82a] Henryk Gzyl. Integration of the Boltzmann equation in the relaxation time approximation. Journal of Statistical Physics, 29 (3):617–622, November 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01342190.

Gzyl:1982:LBE

[Gzy82b] Henryk Gzyl. Linearized Boltzmann equations i: Representation of solutions. Journal of Statistical Physics, 28(1):165–171, May 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011629.

Haas:1989:NSP

[Haa89] Tobias M. Haas. A numerical study of the phase structure of the q-state generalization of the hard-square model. Journal of Statistical Physics, 54(1–2):201–219, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01023479.

Hadjisavvas:1981:MEP

[Had81] N. Hadjisavvas. The maximum entropy principle as a consequence of the principle of Laplace. Journal of Statistical Physics, 26(4): 807–815, December 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010942.

Halpern:1983:RWR

[Hal83] V. Halpern. Random walks on random lattices with traps. Journal of Statistical Physics, 30(2):373–381, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012311.

Hamer:1987:FSS

[Ham87] C. J. Hamer. Finite-size scaling corrections for the eight-vertex model. Journal of Statistical Physics, 47(3–4):331–342, May 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01007514.

Hanggi:1983:PLM

[Han83a] Peter Hanggi. Physics of ligand migration in biomolecules. Journal of Statistical Physics, 30(2):401–412, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012314.

Hanson:1983:VCP

[Han83b] Howard G. Hanson. Voronoi cell properties from simulated and real random spheres and points. Journal of Statistical Physics, 30(3):591–605, March 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009678.

Hanggi:1986:AE

[Han86a] Peter Hanggi. Addendum and erratum. Journal of Statistical Physics, 44(5–6):1003–1004, September 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011919.

Hanggi:1986:EMS

[Han86b] Peter Hanggi. Escape from a metastable state. Journal of Statistical Physics, 42(1–2):105–148, January 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010843.

Harter:1984:PEL

[Har84] William G. Harter. Patterns of energy levels and spectra for polyatomic molecules. Journal of Statistical Physics, 36(5–6): 749–777, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012936.

Hara:1987:RCLa

[Har87] Takashi Hara. A rigorous control of logarithmic corrections in four-dimensional φ^{4} spin systems. Journal of Statistical Physics, 47(1–2):57–98, April 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009035.

Havlin:1984:PIR

[HBAM84] S. Havlin, D. Ben-Avraham, and D. Movshovitz. Percolation on infinitely ramified fractals. Journal of Statistical Physics, 36(5– 6):831–841, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012943.

Hilhorst:1984:SPC

[HBB^{+}84] H. J. Hilhorst, A. F. Bakker, C. Bruin, A. Compagner, and A. Hoogland. Special purpose computers in physics. Journal of Statistical Physics, 34(5–6):987–1000, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009453.

Hamer:1988:LCF

[HBB88] C. J. Hamer, M. T. Batchelor, and Michael N. Barber. Logarithmic corrections to finite-size scaling in the four-state Potts model. Journal of Statistical Physics, 52(3–4):679–710, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01019724.

Hauser:1988:PFT

[HC88] Paulo R. Hauser and Evaldo M. F. Curado. Potts ferromagnet: Transformations and critical exponents in planar hierarchical lattices. Journal of Statistical Physics, 52(1–2):503–507, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016432.

Heermann:1984:NMT

[HCKS84] Dieter W. Heermann, Antonio Coniglio, W. Klein, and Dietrich Stauffer. Nucleation and metastability in three-dimensional Ising models. Journal of Statistical Physics, 36(3–4):447–470, August 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010991.

Hanson:1985:ADS

[HCM85] James D. Hanson, John R. Cary, and James D. Meiss. Algebraic decay in self-similar Markov chains. Journal of Statistical Physics, 39(3–4):327–345, May 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01018666.

Hongler:1983:SCM

[HD83] M.-O. Hongler and Rashmi C. Desai. Study of a class of models for self-organization: equilibrium analysis. Journal of Statistical Physics, 32(3):585–614, September 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008958.

Holloway:1987:DCN

[HD87] James Paul Holloway and J. J. Dorning. The dynamics of coupled nonlinear model Boltzmann equations. Journal of Statistical Physics, 49(3–4):607–660, November 1987. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009350.

Hagan:1989:DET

[HDL89] Patrick S. Hagan, Charles R. Doering, and C. David Levermore. the distribution of exit times for weakly colored noise. Journal of Statistical Physics, 54(5–6):1321–1352, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01044718.

Heermann:1982:CNT

[Hee82] D. W. Heermann. Classical nucleation theory with a Tolman correction. Journal of Statistical Physics, 29(3):631–640, November 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01342192.

Hood:1989:PSS

[HEH89] L. M. Hood, D. J. Evans, and H. J. M. Hanley. Properties of a soft-sphere liquid from non-Newtonian molecular dynamics. Journal of Statistical Physics, 57(3–4):729–743, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01022829.

Helleman:1988:QNM

[Hel88] Robert H. G. Helleman. Quantization of nonintegrable maps. Journal of Statistical Physics, 53(1–2):457–474, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011566.

Hemmer:1989:RPP

[Hem89] P. C. Hemmer. The random parking problem. Journal of Statistical Physics, 57(3–4):865–869, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01022837.

Herrmann:1986:FAS

[Her86] H. J. Herrmann. Fast algorithm for the simulation of Ising models. Journal of Statistical Physics, 45(1–2):145–151, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01033083.

Heupel:1986:DIP

[Heu86] Ludger Heupel. Diffusion of interacting particles on a percolating lattice. Journal of Statistical Physics, 42(3–4):541–549, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01127726.

Heyde:1982:ABRa

[Hey82] C. C. Heyde. On the asymptotic behavior of random walks on an anisotropic lattice. Journal of Statistical Physics, 27(4):721–730, April 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013444.

Hendriks:1983:CEG

[HEZ83] E. M. Hendriks, M. H. Ernst, and R. M. Ziff. Coagulation equations with gelation. Journal of Statistical Physics, 31(3):519–563, June 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019497.

Hinrichsen:1986:GRS

[HFJ86] Einar L. Hinrichsen, Jens Feder, and Torstein Jøssang. Geometry of random sequential adsorption. Journal of Statistical Physics, 44(5–6):793–827, September 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011908.

Guan:1986:ERD

[hGH86] Pu hua Guan and Yu He. Exact results for deterministic cellular automata with additive rules. Journal of Statistical Physics, 43 (3–4):463–478, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020648.

Hernandez-Garcia:1989:RWD

[HGPRS89] E. Hernández-García, L. Pesquera, M. A. Rodríguez, and M. San Miguel. Random walk in dynamically disordered chains: Poisson white noise disorder. Journal of Statistical Physics, 55(5– 6):1027–1052, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01041077.

Harris:1987:SCM

[HGR87] Jonathan G. Harris, Jan Gryko, and Stuart A. Rice. Selfconsistent Monte Carlo simulations of the electron and ion distributions of inhomogeneous liquid alkali metals. II. Longitudinal and transverse density distributions in the liquid-vapor interface of binary metallic alloys. Journal of Statistical Physics, 48(5–6): 1109–1128, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009536.

Hogg:1985:PCS

[HH85] T. Hogg and B. A. Huberman. Parallel computing structures capable of flexible associations and recognition of fuzzy inputs. Journal of Statistical Physics, 41(1–2):115–123, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020606.

Hayward:1987:DPT

[HHB87] S. Hayward, Dieter W. Heermann, and K. Binder. Dynamic percolation transition induced by phase separation: A Monte Carlo analysis. Journal of Statistical Physics, 49(5–6):1053–1081, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017560.

Hubner:1986:CSM

[HHO86] R. Hübner, J. S. Høye, and K. Olaussen. A case study of the MSA approach to quantized polarizable media. Journal of Statistical Physics, 42(3–4):523–539, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01127725.

Hioe:1983:PRW

[Hio83] F. T. Hioe. On Pearson’s random walk and some statistical properties of a quasiperiodic observable in a simple quantum model. Journal of Statistical Physics, 30(2):467–475, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012320.

Hirsch:1986:SHM

[Hir86] J. E. Hirsch. Simulations of the Hubbard model. Journal of Statistical Physics, 43(5–6):841–859, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02628310.

Henin:1987:SUT

[HJ87] F. Henin and J. Jeener. Star-unitary transformations as a tool in single-atom pulse spectroscopy. Journal of Statistical Physics, 48(5–6):1321–1342, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009548.

Hanggi:1989:EDS

[HJM89] Peter Hänggi, Peter Jung, and Fabio Marchesoni. Escape driven by strongly correlated noise. Journal of Statistical Physics, 54(5– 6):1367–1380, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01044720.

Hsu:1985:SSA

[HK85] C. S. Hsu and Myun C. Kim. Statistics of strange attractors by generalized cell mapping. Journal of Statistical Physics, 38(3– 4):735–761, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010488.

Hoover:1986:HCT

[HK86] William G. Hoover and Karl W. Kratky. Heat conductivity of three periodic hard disks via nonequilibrium molecular dynamics. Journal of Statistical Physics, 42(5–6):1103–1114, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010464.

Havlin:1987:PDP

[HKLW87] S. Havlin, J. E. Kiefer, F. Leyvraz, and G. H. Weiss. Probability distribution for percolation clusters generated on a Cayley tree at criticality. Journal of Statistical Physics, 47(1–2):173–184, April 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009040.

Havlin:1985:PSA

[HKW^{+}85] Shlomo Havlin, James E. Kiefer, George H. Weiss, Daniel Benavraham, and Yehoshua Glazer. Properties of the skeleton of aggregates grown on a Cayley tree. Journal of Statistical Physics, 41(3–4):489–496, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009019.

Holicky:1988:RIL

[HKZ88] P. Holický, R. Kotecký, and M. Zahradník. Rigid interfaces for lattice models at low temperatures. Journal of Statistical Physics, 50(3–4):755–812, February 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01026500.

Hohenberg:1982:NPO

[HL82] P. C. Hohenberg and J. S. Langer. Nonequilibrium phenomena: Outlines and bibliographies of a workshop. Journal of Statistical Physics, 28(1):193–226, May 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011633.

Hansel:1989:DED

[HL89] D. Hansel and J. F. Luciani. On diffusion equations for dynamical systems driven by noise. Journal of Statistical Physics, 54(3– 4):971–995, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019784.

Hannon:1988:MH

[HLC88] L. Hannon, G. C. Lie, and E. Clementi. Micro-hydrodynamics. Journal of Statistical Physics, 51(5–6):965–979, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014894.

Horiguchi:1984:BEF

[HM84] T. Horiguchi and T. Morita. Behavior of the effective fields for a regular Ising model on the Cayley tree. Journal of Statistical Physics, 35(3–4):355–374, May 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014389.

Hubbard:1988:MFE

[HM88] J. B. Hubbard and D. A. McQuarrie. Mobility fluctuations and electrophoretic light scattering from macromolecular solutions. Journal of Statistical Physics, 52(5–6):1247–1261, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011645.

Houghton:1987:LEW

[HMC87] A. Houghton, A. J. McKane, and Hilda A. Cerdeira. Localization: The effect of a weak magnetic field. Journal of Statistical Physics, 48(1–2):81–96, July 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010401.

Hoover:1984:HPH

[HMH84] William G. Hoover, Bill Moran, and James M. Haile. Homogeneous periodic heat flow via nonequilibrium molecular dynamics. Journal of Statistical Physics, 37(1–2):109–121, October 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012907.

Hansel:1987:APA

[HMOV87] D. Hansel, J. M. Maillard, J. Oitmaa, and M. J. Velgakis. Analytical properties of the anisotropic cubic Ising model. Journal of Statistical Physics, 48(1–2):69–80, July 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010400.

Hughes:1982:FRW

[HMS82] B. D. Hughes, E. W. Montroll, and M. F. Shlesinger. Fractal random walks. Journal of Statistical Physics, 28(1):111–126, May 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011626.

Hughes:1983:FLS

[HMS83] B. D. Hughes, E. W. Montroll, and M. F. Shlesinger. Fractal and lacunary stochastic processes. Journal of Statistical Physics, 30 (2):273–283, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012302.

Helfer:1984:FMC

[HMV84] H. Lawrence Helfer, Robert L. McCrory, and Hugh M. Van Horn. Further Monte Carlo calculations for the classical one-component plasma in the range 100 ≤ Γ ≤ 160: The FCC lattice. Journal of Statistical Physics, 37(5–6):577–588, December 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010496.

Hasegawa:1980:CSS

[HN80] Hiroshi Hasegawa and Teruaki Nakagomi. On the characterization of the stationary state of a class of dynamical semigroups. Journal of Statistical Physics, 23(5):639–652, November 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011734.

Hendriks:1982:SNB

[HN82] E. M. Hendriks and T. M. Nieuwenhuizen. Solution to the nonlinear Boltzmann equation for Maxwell models for nonisotropic initial conditions. Journal of Statistical Physics, 29(3):591–615, November 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01342189.

Hasegawa:1980:NTL

[HNMK80] H. Hasegawa, T. Nakagomi, M. Mabuchi, and K. Kondo. Nonequilibrium thermodynamics of lasing and bistable optical systems. Journal of Statistical Physics, 23(3):281–313, September 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011370.

Hoglund:1988:ODM

[Hög88] Thomas Höglund. A one-dimensional model with phase transition. Journal of Statistical Physics, 53(3–4):825–834, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014227.

Holschneider:1988:WTF

[Hol88] Matthias Holschneider. On the wavelet transformation of fractal objects. Journal of Statistical Physics, 50(5–6):963–993, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01019149.

Hoover:1986:NCE

[Hoo86] William G. Hoover. Nonlinear conductivity and entropy in the two-body Boltzmann gas. Journal of Statistical Physics, 42(3– 4):587–600, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01127730.

Horton:1985:IAT

[Hor85] W. Horton. Ion acoustic turbulence and anomalous transport. Journal of Statistical Physics, 39(5–6):739–754, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008363.

Hoye:1987:SMI

[Høy87] J. S. Høye. Statistical mechanics of interfaces. Journal of Statistical Physics, 49(1–2):297–317, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009965.

Hauge:1981:BAN

[HP81] E. H. Hauge and E. Præstgaard. The Bobylev approach to the nonlinear Boltzmann equation. Journal of Statistical Physics, 24 (1):21–38, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01007632.

Hsu:1985:EID

[HP85] Weichan Hsu and David Pines. Effective interactions in dilute mixtures of ^{3}He in ^{4}He. Journal of Statistical Physics, 38(1– 2):273–312, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01017863.

Heath:1987:HTL

[HP87] D. L. Heath and J. K. Percus. How thick is a liquid-vapor interface? Journal of Statistical Physics, 49(1–2):319–330, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009966.

Heitmann:1980:GSS

[HR80] Raymond C. Heitmann and Charles Radin. The ground state for sticky disks. Journal of Statistical Physics, 22(3):281–287, March 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014644.

Heilmann:1982:EMC

[HR82] O. J. Heilmann and J. Rotne. Exact and Monte Carlo computations on a lattice model for change of conformation of a polymer. Journal of Statistical Physics, 27(1):19–35, January 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011737.

Hegerfeldt:1983:SATa

[HR83a] Gerhard C. Hegerfeldt and Reinhard Reibold. Stochastic aspects in the theory of spectral-line broadening. I. Collision time statistics and n-perturber limit. Journal of Statistical Physics, 32 (2):313–335, August 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012713.

Hegerfeldt:1983:SATb

[HR83b] Gerhard C. Hegerfeldt and Reinhard Reibold. Stochastic aspects in the theory of spectral-line broadening. II. Noncommutative cluster expansions. Journal of Statistical Physics, 32(2):337–360, August 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012714.

Hron:1985:SMF

[HR85] M. Hron and M. Razavy. Solvable models of the Fokker–Planck equation: An approach based on the Gel’fand–Levitan method. Journal of Statistical Physics, 38(3–4):655–668, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010483.

Hume:1986:REX

[HR86] Leigh Hume and Derek W. Robinson. Return to equilibrium in the XY model. Journal of Statistical Physics, 44(5–6):829– 848, September 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011909.

Hansen:1988:MEP

[HR88] Alex Hansen and Stéphane Roux. Multifractality in elastic percolation. Journal of Statistical Physics, 53(3–4):759–771, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014224.

Hill:1982:PBB

[HS82a] Terrell L. Hill and J. S. Shiner. Problems in biochemistry and biophysics. Journal of Statistical Physics, 28(3):613–614, July 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008329.

Hughes:1982:RWB

[HS82b] Barry D. Hughes and Muhammad Sahimi. Random walks on the Bethe lattice. Journal of Statistical Physics, 29(4):781–794, December 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011791.

Huckaby:1986:EST

[HS86] Dale A. Huckaby and Masato Shinmi. Exact solution of a threecomponent system on the honeycomb lattice. Journal of Statistical Physics, 45(1–2):135–144, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01033082.

Holley:1987:LSI

[HS87] Richard Holley and Daniel Stroock. Logarithmic Sobolev inequalities and stochastic Ising models. Journal of Statistical Physics, 46 (5–6):1159–1194, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011161.

Hegerfeldt:1988:NCS

[HS88a] Gerhard C. Hegerfeldt and Henrik Schulze. Noncommutative cumulants for stochastic differential equations and for generalized Dyson series. Journal of Statistical Physics, 51(3–4):691–710, May 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01028479.

Hegerfeldt:1988:QSN

[HS88b] Gerhard C. Hegerfeldt and Henrik Schulze. Quantum shot noise: Expansions in powers of the pulse density. Journal of Statistical Physics, 51(3–4):711–728, May 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01028480.

Horsthemke:1989:NST

[HS89] Werner Horsthemke and Mark F. Schumaker. Noise and superfluid turbulence in He II: Theory. Journal of Statistical Physics, 54(5–6):1175–1188, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01044710.

Heidenreich:1980:LM

[HSU80] R. Heidenreich, R. Seiler, and D. A. Uhlenbrock. The Luttinger model. Journal of Statistical Physics, 22(1):27–57, January 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01007986.

Hanggi:1980:RTS

[HT80] P. Hänggi and P. Talkner. A remark on truncation schemes of cumulant hierarchies. Journal of Statistical Physics, 22(1):65–67, January 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01007988.

Hara:1987:RCLb

[HT87a] Takashi Hara and Hal Tasaki. A rigorous control of logarithmic corrections in four-dimensional φ^{4} spin systems. Journal of Statistical Physics, 47(1–2):99–121, April 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009036.

Hubmer:1987:OCR

[HT87b] G. F. Hubmer and U. M. Titulaer. The Onsager–Casimir relations revisited. Journal of Statistical Physics, 49(1–2):331–346, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009967.

Hubmer:1988:OCS

[HT88] G. F. Hubmer and U. M. Titulaer. Onsager–Casimir symmetry properties of the Burnett equations. Journal of Statistical Physics, 50(5–6):913–934, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01019147.

Huber:1983:TEA

[Hub83] D. L. Huber. Trapping of excitation in the average T -matrix approximation. Journal of Statistical Physics, 30(2):345–354, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012308.

Huckaby:1983:PTL

[Huc83] Dale A. Huckaby. Phase transition in a lattice gas of hard spheres with second-neighbor exclusions on the simple cubic lattice. Journal of Statistical Physics, 33(1):23–29, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009745.

Hugenholtz:1983:DBE

[Hug83] N. M. Hugenholtz. Derivation of the Boltzmann equation for a Fermi gas. Journal of Statistical Physics, 32(2):231–254, August 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012709.

Hughes:1985:TCS

[Hug85] Webster Hughes. Thermodynamics for Coulomb systems: A problem at vanishing particle densities. Journal of Statistical Physics, 41(5–6):975–1013, December 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010012.

Hansen:1985:TBC

[HV85] J. P. Hansen and P. Viot. Two-body correlations and pair formation in the two-dimensional Coulomb gas. Journal of Statistical Physics, 38(5–6):823–850, March 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010417.

Hale:1982:MCM

[HW82] Barbara N. Hale and Richard C. Ward. A Monte Carlo method for approximating critical cluster size in the nucleation of model systems. Journal of Statistical Physics, 28(3):487–495, July 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008320.

Hasegawa:1985:SPW

[HW85] M. Hasegawa and M. Watabe. On surface properties of a weaklycoupled classical one-component plasma: The size effect in finite systems. Journal of Statistical Physics, 41(1–2):281–297, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020614.

Hall:1986:GLA

[HW86] R. W. Hall and P. G. Wolynes. G. N. Lewis’ atom and quantum Monte Carlo studies of liquids. Journal of Statistical Physics, 43 (5–6):935–948, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02628321.

Haake:1988:QBS

[HW88] Fritz Haake and Martin Wilkens. Quasiprobabilities based on squeezed states. Journal of Statistical Physics, 53(1–2):345–360, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011561.

Heyde:1982:ABRb

[HWW82] C. C. Heyde, M. Westcott, and E. R. Williams. The asymptotic behavior of a random walk on a dual-medium lattice. Journal of Statistical Physics, 28(2):375–380, June 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012611.

Hynes:1986:CRR

[Hyn86] James T. Hynes. Chemical reaction rates and solvent friction. Journal of Statistical Physics, 42(1–2):149–168, January 1986.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010844.

Hongler:1982:ESD

[HZ82] M. O. Hongler and W. M. Zheng. Exact solution for the diffusion in bistable potentials. Journal of Statistical Physics, 29 (2):317–327, October 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020789.

Harrison:1986:RBF

[HZ86] Alan K. Harrison and Robert Zwanzig. On the relation between a fluctuating diffusion equation and long time tails in stationary random media. Journal of Statistical Physics, 42(5–6):935–940, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010454.

Igloi:1989:QAT

[Igl89] F. Iglói. Quantum Ashkin–Teller model near the decoupling limit. Journal of Statistical Physics, 55(1–2):311–330, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01042602.

Kondo:1987:MCI

[iK87] Kei ichi Kondo. More correlation inequalities for a class of even ferromagnets. Journal of Statistical Physics, 46(1–2):35–38, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010328.

Kondo:1985:CIC

[iKOS85] Kei ichi Kondo, Takeshi Otofuji, and Yüki Sugiyama. Correlation inequalities for a class of even ferromagnets. Journal of Statistical Physics, 40(3–4):563–575, August 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01017185.

Ito:1988:SDF

[IKY88] H. M. Ito, S. Kotani, and T. Yokota. Size distribution of fractured areas in one-dimensional systems. Journal of Statistical Physics, 51(3–4):569–584, May 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01028473.

Ikeda:1986:SHD

[IM86] Kensuke Ikeda and Kenji Matsumoto. Study of a highdimensional chaotic attractor. Journal of Statistical Physics, 44(5–6):955–983, September 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011917.

Imada:1986:MTD

[Ima86] M. Imada. Melting of the two-dimensional Wigner crystal in the quantum region. Journal of Statistical Physics, 43(5–6):861–863, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02628311.

Imry:1984:REF

[Imr84] Yoseph Imry. Random external fields. Journal of Statistical Physics, 34(5–6):849–862, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009444.

Issigoni:1986:IFS

[IP86] M. E. Issigoni and C. Papatriantafillou. The influence of fixed spins on the thermodynamic behavior of an Ising ferromagnet. Journal of Statistical Physics, 45(3–4):527–539, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01021084.

Imbrie:1988:DDP

[IS88] J. Z. Imbrie and T. Spencer. Diffusion of directed polymers in a random environment. Journal of Statistical Physics, 52(3– 4):609–626, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019720.

Ianiro:1988:SBE

[IT88] N. Ianiro and L. Triolo. Stationary Boltzmann equation for a degenerate gas in a slab: Boundary value problem and hydrodynamics. Journal of Statistical Physics, 51(3–4):677–690, May 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01028478.

Inawashiro:1983:IMC

[ITH83] Sakari Inawashiro, Colin J. Thompson, and Goshi Honda. Ising model with competing interactions on a Cayley tree. Journal of Statistical Physics, 33(2):419–436, November 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009804.

Ito:1982:CLD

[Ito82] K. R. Ito. Clustering in low-dimensional SO(N)-invariant statistical models with long-range interactions. Journal of Statistical Physics, 29(4):747–760, December 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011788.

Ito:1984:ERP

[Ito84a] H. M. Ito. Ergodicity of randomly perturbed Lorenz model. Journal of Statistical Physics, 35(1–2):151–158, April 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01017371.

Ito:1984:OGS

[Ito84b] H. M. Ito. Optimal Gaussian solutions of nonlinear stochastic partial differential equations. Journal of Statistical Physics, 37(5– 6):653–671, December 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010500.

Ito:1988:SDD

[Ito88] H. M. Ito. Stochastic disk dynamo as a model of reversals of the Earth’s magnetic field. Journal of Statistical Physics, 53 (1–2):19–39, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011542.

Indira:1983:DBP

[IVMA83] R. Indira, M. C. Valsakumar, K. P. N. Murthy, and G. Ananthakrishna. Diffusion in a bistable potential: A comparative study of different methods of solution. Journal of Statistical Physics, 33 (1):181–194, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009755.

Iyengar:1986:NSU

[Iye86] R. N. Iyengar. A nonlinear system under combined periodic and random excitation. Journal of Statistical Physics, 44(5–6): 907–920, September 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011913.

Jancovici:1982:CCSa

[Jan82a] B. Jancovici. Classical Coulomb systems near a plane wall. I. Journal of Statistical Physics, 28(1):43–65, May 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011622.

Jancovici:1982:CCSb

[Jan82b] B. Jancovici. Classical Coulomb systems near a plane wall. II. Journal of Statistical Physics, 29(2):263–280, October 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020786.

Jancovici:1984:SPC

[Jan84] B. Jancovici. Surface properties of a classical two-dimensional one-component plasma: Exact results. Journal of Statistical Physics, 34(5–6):803–815, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009442.

Jancovici:1985:SCQ

[Jan85] B. Jancovici. Surface correlations in a quantum mechanical onecomponent plasma. Journal of Statistical Physics, 39(3–4):427– 441, May 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01018671.

Janis:1987:DEM

[Jan87] V. Janis. Diagrammatic expansion and metastability in the random-field Ising model. Journal of Statistical Physics, 47(5– 6):931–938, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01206166.

Janssen:1989:EFVa

[Jan89a] J. A. M. Janssen. The elimination of fast variables in complex chemical reactions. I. Macroscopic level. Journal of Statistical Physics, 57(1–2):157–169, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01023638.

Janssen:1989:EFVb

[Jan89b] J. A. M. Janssen. The elimination of fast variables in complex chemical reactions. II. Mesoscopic level (reducible case). Journal of Statistical Physics, 57(1–2):171–185, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01023639.

Janssen:1989:EFVc

[Jan89c] J. A. M. Janssen. The elimination of fast variables in complex chemical reactions. III. Mesoscopic level (irreducible case). Journal of Statistical Physics, 57(1–2):187–198, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023640.

Jauslin:1985:CFP

[Jau85] H. R. Jauslin. A classification of Fokker–Planck models and the small and large noise asymptotics. Journal of Statistical Physics, 40(1–2):147–165, July 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010530.

Jauslin:1986:MCN

[Jau86] H. R. Jauslin. Melnikov’s criterion for nondifferentiable weaknoise potentials. Journal of Statistical Physics, 42(3–4):573–585, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01127729.

Joslin:1989:CSS

[JE89] C. G. Joslin and P. A. Egelstaff. Computer simulation of some dynamical properties of the Lorentz gas. Journal of Statistical Physics, 56(1–2):127–138, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01044236.

Jensen:1981:FIA

[Jen81] Roderick V. Jensen. Functional integral approach to classical statistical dynamics. Journal of Statistical Physics, 25(2):183– 210, June 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01022182.

Jen:1986:GPC

[Jen86a] Erica Jen. Global properties of cellular automata. Journal of Statistical Physics, 43(1–2):219–242, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010579.

Jen:1986:ISP

[Jen86b] Erica Jen. Invariant strings and pattern-recognizing properties of one-dimensional cellular automata. Journal of Statistical Physics, 43(1–2):243–265, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010580.

Jerrum:1987:TDM

[Jer87] Mark Jerrum. Two-dimensional monomer-dimer systems are computationally intractable. Journal of Statistical Physics, 48 (1–2):121–134, July 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010403.

Jensen:1986:SPC

[JJ86] R. V. Jensen and E. R. Jessup. Statistical properties of the circle map. Journal of Statistical Physics, 43(1–2):369–390, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010588.

Jones:1989:DFT

[JK89] Gerald L. Jones and Soon-Chul Kim. Density functional theory of homogeneous states. Journal of Statistical Physics, 56(5–6): 709–719, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016776.

Jullien:1985:CWK

[JKHV85] R. Jullien, M. Kolb, H. Herrmann, and J. Vannimenus. CECAM-Workshop on kinetic models for cluster formation: Orsay, September 17–28, 1984. Journal of Statistical Physics, 39(1– 2):241–258, April 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01007981.

Jancovici:1985:TDC

[JLM85] B. Jancovici, J. L. Lebowitz, and Ph. A. Martin. Time-dependent correlations in an inhomogeneous one-component plasma. Journal of Statistical Physics, 41(5–6):941–974, December 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010011.

Jancovici:1995:ETD

[JLM95] B. Jancovici, J. L. Lebowitz, and Ph. A. Martin. Erratum: Time-dependent correlations in an inhomogeneous onecomponent plasma. Journal of Statistical Physics, 79(3–4):789, May 1995. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02184885. See [JMM87].

John:1986:SCE

[JM86] T. M. John and K. P. N. Murthy. A study on correlated exponential random walks. Journal of Statistical Physics, 45(3–4): 753–763, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01021094.

Jhon:1988:EMT

[JMF88] Myung S. Jhon, Robert J. Metz, and Karl F. Freed. Effective medium theory for elastic matrix composites containing dispersed particulates. Journal of Statistical Physics, 52(5–6): 1325–1342, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011650.

Jancovici:1987:TDC

[JMM87] B. Jancovici, N. Macris, and Ph. A. Martin. Time-dependent correlations for a one-component plasma in a uniform magnetic field. Journal of Statistical Physics, 47(1–2):229–256, April 1987.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01009044. See erratum [JLM95].

Jan:1983:DMC

[JMS83] Naeem Jan, L. Leo Moseley, and Dietrich Stauffer. Dynamic Monte Carlo renormalization group. Journal of Statistical Physics, 33(1):1–11, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009743.

Johnson:1989:TTH

[Joh89] E. A. Johnson. Two-temperature hydrodynamics and multiple sound modes in disparate-mass gas mixtures. Journal of Statistical Physics, 57(3–4):647–664, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01022826.

Jones:1986:LTP

[Jon86] Gerald L. Jones. The low-temperature properties of the onedimensional fluid. Journal of Statistical Physics, 44(1–2):237– 248, July 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010915.

Jan:1983:CDB

[JS83] Naeem Jan and M. O. Steinitz. Comparison of different boundary conditions for Monte Carlo simulations of Ising models. Journal of Statistical Physics, 30(1):37–44, January 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010867.

Jauslin:1986:SSF

[JS86] H. R. Jauslin and T. Schneider. Stochastic simulation of fermions. Journal of Statistical Physics, 43(5–6):865–868, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02628312.

Junker:1989:TPM

[JW89] W. Junker and M. Wagner. Thermodynamics of phononmodulated tunneling centers. Journal of Statistical Physics, 56 (3–4):331–355, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01044440.

Kadanoff:1983:SBO

[Kad83] Leo P. Kadanoff. Supercritical behavior of an ordered trajectory. Journal of Statistical Physics, 31(1):1–27, April 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010921.

Kadanoff:1985:SHP

[Kad85] Leo P. Kadanoff. Simulating hydrodynamics: A pedestrian model. Journal of Statistical Physics, 39(3–4):267–283, May 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01018663.

Kadanoff:1986:RGA

[Kad86] Leo P. Kadanoff. Renormalization group analysis of the global properties of a strange attractor. Journal of Statistical Physics, 43(3–4):395–410, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020644.

Kaganov:1985:IML

[Kag85] M. I. Kaganov. Ilya Mikhailovich Lifshitz. Journal of Statistical Physics, 38(1–2):3–9, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01017844.

Kai:1982:LNN

[Kai82] Tohru Kai. Lyapunov number for a noisy 2^{n} cycle. Journal of Statistical Physics, 29(2):329–343, October 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020790.

Kalus:1981:SCI

[Kal81] Norbert Kalus. Symmetries and correlation inequalities for classical n-vector models. Journal of Statistical Physics, 25(4):635–643, August 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01022358.

Kamphorst:1986:MPT

[Kam86] S. Oliffson Kamphorst. A mixed phase transition for a hierarchical spin glass. Journal of Statistical Physics, 45(3–4):369–393, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01021077.

Kasteleyn:1987:VTM

[Kas87] P. W. Kasteleyn. Variations on a theme by mark Kac. Journal of Statistical Physics, 46(5–6):811–827, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011143.

Katz:1985:SMG

[Kat85] Talbot Michael Katz. Spherical models with a Gates–Penrosetype phase transition. Journal of Statistical Physics, 38(3–4): 589–602, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010479.

Khantha:1985:RPB

[KB85] M. Khantha and V. Balakrishnan. Reflection principles for biased random walks and application to escape time distributions. Journal of Statistical Physics, 41(5–6):811–824, December 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010005.

Keller:1986:IN

[KB86] Joseph B. Keller and Luis L. Bonilla. Irreversibility and nonrecurrence. Journal of Statistical Physics, 42(5–6):1115–1125, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010465.

Klafter:1984:FBT

[KBZ84] J. Klafter, A. Blumen, and G. Zumofen. Fractal behavior in trapping and reaction: A random walk study. Journal of Statistical Physics, 36(5–6):561–577, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012922.

Kirkpatrick:1983:KTF

[KC83] T. R. Kirkpatrick and E. G. D. Cohen. Kinetic theory of fluctuations near a convective instability. Journal of Statistical Physics, 33(3):639–694, December 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01018838.

Kaner:1985:LED

[KC85] É. A. Kaner and L. V. Chebotarev. Localization effects and dispersion of the dielectric response function in one-dimensional metals. Journal of Statistical Physics, 38(1–2):77–87, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017849.

Keirstead:1988:VPH

[KCH88] W. P. Keirstead, H. A. Ceccatto, and B. A. Huberman. Vibrational properties of hierarchical systems. Journal of Statistical Physics, 53(3–4):733–757, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01014223.

Katz:1982:SSE

[KD82a] A. Katz and M. Duneau. Stability of symmetries for equilibrium configurations of n particles in three dimensions. Journal of Statistical Physics, 29(3):475–498, November 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01342183.

Kratky:1982:IDS

[KD82b] Karl W. Kratky and Herbert Drexler. Intersecting disks (and spheres) and statistical mechanics. II. The hard-disk system. Journal of Statistical Physics, 29(3):499–513, November 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01342184.

Kirkpatrick:1983:EMG

[KD83] T. R. Kirkpatrick and J. R. Dorfman. Electron mobility in gases at low temperatures: The quantum mechanical Lorentz gas, I. Journal of Statistical Physics, 30(1):67–106, January 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010869.

Kapitulnik:1984:PSE

[KD84a] Aharon Kapitulnik and Guy Deutscher. Percolation scale effects in metal-insulator thin films. Journal of Statistical Physics, 36(5– 6):815–826, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012941.

Katz:1984:COD

[KD84b] A. Katz and M. Duneau. The convergence of the one-dimensional ground states to an infinite lattice. Journal of Statistical Physics, 37(1–2):257–268, October 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012914.

Kasteleyn:1983:RWI

[KdH83] P. W. Kasteleyn and W. Th. F. den Hollander. Random walks on inhomogeneous lattices. Journal of Statistical Physics, 30(2): 363–372, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012310.

Klosek-Dygas:1989:CNA

[KDMS89] M. M. Klosek-Dygas, B. J. Matkowsky, and Z. Schuss. Colored noise in activated rate processes. Journal of Statistical Physics, 54 (5–6):1309–1320, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01044717.

Kehr:1983:DCL

[Keh83] K. W. Kehr. Diffusion in concentrated lattice gases. Journal of Statistical Physics, 30(2):509–518, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012324.

Keizer:1983:RBF

[Kei83] Joel Keizer. On the relationship between fluctuating irreversible thermodynamics and ‘extended’ irreversible thermodynamics. Journal of Statistical Physics, 31(3):485–497, June 1983.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01019495. See comments [Eu84, Kei84].

Keizer:1984:RCK

[Kei84] Joel Keizer. Response to ‘Comments on Keizer’s critique on extended irreversible thermodynamics’ by B. C. Eu. Journal of Statistical Physics, 37(3–4):491–493, November 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01011847. See [Eu84].

Kremp:1983:HTA

[KEKS83] Dietrich Kremp, Werner Ebeling, Hartmut Krienke, and Rainer Sändig. HNC-type approximation for transport processes in electrolytic solutions. Journal of Statistical Physics, 33(1):99–106, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009751.

Kennedy:1982:LBP

[Ken82] Tom Kennedy. A lower bound on the partition function for a classical charge symmetric system. Journal of Statistical Physics, 28(4):633–638, August 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011872.

Kenkre:1983:MET

[Ken83] V. M. Kenkre. Master equation techniques for exciton motion, relaxation, capture, and annihilation. Journal of Statistical Physics, 30(2):293–303, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012304.

Kennedy:1984:MFT

[Ken84] Tom Kennedy. Mean field theory for Coulomb systems. Journal of Statistical Physics, 37(5–6):529–559, December 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010494.

Kerstein:1986:CSP

[Ker86] Alan R. Kerstein. Computational study of propagating fronts in a lattice-gas model. Journal of Statistical Physics, 45(5–6):

921–931, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020582.

Kerimov:1988:ISI

[Ker88a] Azer Kerimov. Interface sharpness in the Ising model with longrange interaction. Journal of Statistical Physics, 52(1–2):69–98, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01016405.

Kerstein:1988:TPR

[Ker88b] Alan R. Kerstein. A two-particle representation of front propagation in diffusion-reaction systems. Journal of Statistical Physics, 53(3–4):703–712, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014221.

Kesten:1981:APP

[Kes81] Harry Kesten. Analyticity properties and power law estimates of functions in percolation theory. Journal of Statistical Physics, 25 (4):717–756, August 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01022364.

Kesten:1988:BRB

[Kes88] Harry Kesten. Book review: Les Houches 1984, session 43, critical phenomena, random systems, gauge theory. Journal of Statistical Physics, 50(5–6):1277–1279, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019167.

Ketzer:1984:FCB

[Ket84] Joel Ketzer. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 37(1–2):269–270, October 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012915.

Keyes:1982:NTR

[Key82] T. Keyes. NESS theory of random steady states. Journal of Statistical Physics, 28(3):413–429, July 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008315.

Keyes:1983:MTL

[Key83] T. Keyes. Microscopic theory of the liquid-wall boundary conditions; thermophoresis. Journal of Statistical Physics, 33(2): 287–306, November 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009798.

Kaganov:1985:ERQ

[KF85] M. I. Kaganov and V. B. Fiks. Einstein relation for quantum systems. Journal of Statistical Physics, 38(1–2):329–345, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017865.

Kai:1989:SCI

[KFB89] Shoichi Kai, Hidehiko Fukunaga, and Helmut R. Brand. Structure changes induced by external multiplicative noise in the electrohydrodynamic instability of nematic liquid crystals. Journal of Statistical Physics, 54(5–6):1133–1152, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01044707.

Korzeniewski:1983:LET

[KFS83] G. Korzeniewski, R. Friesner, and R. Silbey. Localized excitation transport on substitutionally disordered lattices. Journal of Statistical Physics, 31(3):451–465, June 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019493.

Kapitulnik:1984:FDC

[KGA84] A. Kapitulnik, Y. Gefen, and A. Aharony. On the fractal dimension and correlations in percolation theory. Journal of Statistical Physics, 36(5–6):807–814, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012940.

Kjaer:1982:DGC

[KH82] K. H. Kjaer and H. J. Hilhorst. The discrete Gaussian chain with 1/r^{n} interactions: Exact results. Journal of Statistical Physics, 28(4):621–632, August 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011871.

Kim:1986:CLL

[KH86] Myun C. Kim and C. S. Hsu. Computation of the largest Lyapunov exponent by the generalized cell mapping. Journal of Statistical Physics, 45(1–2):49–61, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01033076.

Kratky:1987:HSH

[KH87] Karl W. Kratky and William G. Hoover. Hard-sphere heat conductivity via nonequilibrium molecular dynamics. Journal of Statistical Physics, 48(3–4):873–900, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019700.

Kozak:1983:SFR

[KHM^{+}83] John J. Kozak, Michael D. Hatlee, Matthew K. Musho, Philip A. Politowicz, and Cecilia A. Walsh. Stochastic flows, reaction– diffusion processes, and morphogenesis. Journal of Statistical Physics, 30(2):263–271, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012301.

Kopelman:1983:SMR

[KHNA83] R. Kopelman, J. Hoshen, J. S. Newhouse, and P. Argyrakis. Single and multiple random walks on random lattices: Excitation trapping and annihilation simulations. Journal of Statistical Physics, 30(2):335–343, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012307.

Kholopov:1987:MMC

[Kho87] Eugene V. Kholopov. Melting of microinclusions close packed in an elastic matrix. Journal of Statistical Physics, 48(1–2):215– 230, July 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010407.

Koiwa:1983:RWP

[KI83] Masahiro Koiwa and Shunya Ishioka. Random walk properties of lattices and correlation factors for diffusion via the vacancy mechanism in crystals. Journal of Statistical Physics, 30(2):477–485, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012321.

Kiessling:1989:ESM

[Kie89] Michael K.-H. Kiessling. On the equilibrium statistical mechanics of isothermal classical self-gravitating matter. Journal of Statistical Physics, 55(1–2):203–257, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01042598.

Katsura:1986:GSC

[KIM86] Shigetoshi Katsura, Tsugio Ide, and Tohru Morita. The ground states of the classical Heisenberg and planar models on the triangular and plane hexagonal lattices. Journal of Statistical Physics, 42(3–4):381–404, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01127717.

Kirkpatrick:1984:OSA

[Kir84a] Scott Kirkpatrick. Optimization by simulated annealing: Quantitative studies. Journal of Statistical Physics, 34(5–6):975–986, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009452.

Kirkwood:1984:PTI

[Kir84b] James R. Kirkwood. Phase transitions in the Ising model with transverse field. Journal of Statistical Physics, 37(3–4):407–417, November 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011841.

Kirkpatrick:1989:DVA

[Kir89] T. R. Kirkpatrick. Does the velocity autocorrelation function oscillate in a hard-sphere crystal? Journal of Statistical Physics, 57(3–4):483–496, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01022818.

Kivelson:1988:DLS

[Kiv88] Daniel Kivelson. Depolarized light scattering from liquids: Rotations, collisions, and hydrodynamics. Journal of Statistical Physics, 52(5–6):1285–1305, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011647.

Kawazu:1984:BDP

[KK84] Kiyoshi Kawazu and Harry Kesten. On birth and death processes in symmetric random environment. Journal of Statistical Physics, 37(5–6):561–576, December 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010495.

Kosevich:1985:SCD

[KK85] A. M. Kosevich and I. L. Kruglikov. Self-consistent diffusive kinetics and dissipative structures in a distributed cell system. Journal of Statistical Physics, 38(1–2):361–370, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01017867.

Kim:1988:IRU

[KK88] Chang-Bae Kim and John A. Krommes. Improved rigorous upper bounds for transport due to passive advection described by simple models of bounded systems. Journal of Statistical Physics, 53(5–6):1103–1137, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01023860.

Keyes:1982:TED

[KL82] T. Keyes and J. W. Lyklema. A theory of exciton dynamics with a percolation threshold. Journal of Statistical Physics, 27 (3):487–498, March 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011088.

Koch:1983:CMD

[KL83] S. W. Koch and Rainer Liebmann. Comparison of molecular dynamics and Monte Carlo computer simulations of spinodal decomposition. Journal of Statistical Physics, 33(1):31–41, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009746.

Klein:1980:ADS

[Kle80] Douglas J. Klein. Asymptotic distributions for self-avoiding walks constrained to strips, cylinders, and tubes. Journal of Statistical Physics, 23(5):561–586, November 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011730.

Klein:1986:UCO

[Kle86] David Klein. Uniqueness of continuum one-dimensional Gibbs states for slowly decaying interactions. Journal of Statistical Physics, 43(1–2):207–218, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010578.

Kleinen:1989:TFF

[Kle89] H. Kleinen. Tangential flow in fluid membranes. Absence of renormalization effects. Journal of Statistical Physics, 56(1–2):227– 232, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01044242.

Khalatnikov:1985:SRC

[KLK^{+}85] I. M. Khalatnikov, E. M. Lifshitz, K. M. Khanin, L. N. Shchur, and Ya. G. Sinai. On the stochasticity in relativistic cosmology. Journal of Statistical Physics, 38(1–2):97–114, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017851.

Kipnis:1983:SDPa

[KLPS83a] C. Kipnis, J. L. Lebowitz, E. Presutti, and H. Spohn. Selfdiffusion for particles with stochastic collisions in one dimension. Journal of Statistical Physics, 30(1):107–121, January 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010870.

Kipnis:1983:SDPb

[KLPS83b] C. Kipnis, J. L. Lebowitz, E. Presutti, and H. Spohn. Selfdiffusion for particles with stochastic collisions in one dimension. Journal of Statistical Physics, 31(3):711, June 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019507.

Klein:1981:ASB

[KLS81] Abel Klein, Lawrence J. Landau, and David S. Shucker. On the absence of spontaneous breakdown of continuous symmetry for equilibrium states in two dimensions. Journal of Statistical Physics, 26(3):505–512, November 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011431.

Katz:1984:NSS

[KLS84] Sheldon Katz, Joel L. Lebowitz, and Herbert Spohn. Nonequilibrium steady states of stochastic lattice gas models of fast ionic conductors. Journal of Statistical Physics, 34(3–4):497–537, February 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01018556.

Kennedy:1988:ENO

[KLS88a] Tom Kennedy, Elliott H. Lieb, and B. Sriram Shastry. Existence of Néel order in some spin-1/2 Heisenberg antiferromagnets. Journal of Statistical Physics, 53(5–6):1019–1030, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023854.

Kramli:1988:NWR

[KLS88b] A. Krámli, P. Lukács, and D. Szász. A non-Wiener random walk in a two-dimensional Bernoulli environment. Journal of Statistical Physics, 50(3–4):599–609, February 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01026492.

Klein:1989:RDS

[KLS89] Abel Klein, Jean Lacroix, and Athanasios Speis. Regularity of the density of states in the Anderson model on a strip for potentials with singular continuous distributions. Journal of Statistical Physics, 57(1–2):65–88, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01023635.

Krug:1986:FRL

[KLSZ86] J. Krug, J. L. Lebowitz, H. Spohn, and M. Q. Zhang. The fast rate limit of driven diffusive systems. Journal of Statistical Physics, 44 (3–4):535–565, August 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011309.

Kennedy:1988:TDI

[KLT88] Tom Kennedy, Elliott H. Lieb, and Hal Tasaki. A twodimensional isotropic quantum antiferromagnet with unique disordered ground state. Journal of Statistical Physics, 53(1–2): 383–415, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011563.

Kottalam:1986:SRS

[KLW86] J. Kottalam, Katja Lindenberg, and Bruce J. West. Statistical replacement for systems with delta-correlated fluctuations. Journal of Statistical Physics, 42(5–6):979–1008, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010458.

Ma:1981:CED

[kM81] Shang keng Ma. Calculation of entropy from data of motion. Journal of Statistical Physics, 26(2):221–240, October 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013169.

Ma:1983:ODB

[kM83] Shang keng Ma. One-dimensional Boltzmann equation with a three-body collision term. Journal of Statistical Physics, 31 (1):107–114, April 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010924.

Kuroda:1987:IIM

[KM87] Koji Kuroda and Hiroko Manaka. The interface of the Ising model and the Brownian sheet. Journal of Statistical Physics, 47(5–6):979–984, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01206172.

Kerszberg:1988:DSC

[KM88] Michel Kerszberg and David Mukamel. Dynamics of simple computer networks. Journal of Statistical Physics, 51(5–6):777–795, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014883.

Kaufman:1989:ESE

[KMD89] J. H. Kaufman, O. R. Melroy, and G. M. Dimino. The effect of screening on entropy production in pattern formation. Journal of Statistical Physics, 55(1–2):331–340, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01042603.

Kipnis:1982:HFE

[KMP82] C. Kipnis, C. Marchioro, and E. Presutti. Heat flow in an exactly solvable model. Journal of Statistical Physics, 27(1):65–74, January 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011740.

Kotecky:1987:RTI

[KMS87] R. Kotecký and S. Miracle-Sole. Roughening transition for the Ising model on a BCC lattice. A case in the theory of ground states. Journal of Statistical Physics, 47(5–6):773–799, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01206158.

Knessl:1986:BBD

[KMST86a] C. Knessl, B. J. Matkowsky, Z. Schuss, and C. Tier. Boundary behavior of diffusion approximations to Markov jump processes. Journal of Statistical Physics, 45(1–2):245–266, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01033090.

Knessl:1986:SPA

[KMST86b] C. Knessl, B. J. Matkowsky, Z. Schuss, and C. Tier. A singular perturbation approach to first passage times for Markov jump processes. Journal of Statistical Physics, 42(1–2):169–184, January 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010845.

Knickerbocker:1985:PSW

[KN85] C. J. Knickerbocker and Alan C. Newell. Propagation of solitary waves in channels of decreasing depth. Journal of Statistical Physics, 39(5–6):653–674, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008358.

Kuramoto:1987:SML

[KN87] Yoshiki Kuramoto and Ikuko Nishikawa. Statistical macrodynamics of large dynamical systems. Case of a phase transition in oscillator communities. Journal of Statistical Physics, 49(3–4): 569–605, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009349.

Knabe:1988:EGE

[Kna88] Stefan Knabe. Energy gaps and elementary excitations for certain VBS-quantum antiferromagnets. Journal of Statistical Physics, 52(3–4):627–638, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01019721.

Ko:1988:SSCa

[Ko88a] Lee-Fen Ko. Spin-spin correlation function in the two-dimensional Ising model with linear defects. I. T < T_{c}. Journal of Statistical Physics, 52(3–4):795–813, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01019730.

Ko:1988:SSCb

[Ko88b] Lee-Fen Ko. Spin-spin correlation function in the two-dimensional Ising model with linear defects. II. T > T_{c}. Journal of Statistical Physics, 52(3–4):815–827, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01019731.

Kodama:1985:OSM

[Kod85] Yuji Kodama. Optical solitons in a monomode fiber. Journal of Statistical Physics, 39(5–6):597–614, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008354.

Kogut:1984:SMV

[Kog84] John B. Kogut. A statistical mechanics view of quantum chromodynamics: Lattice gauge theory. Journal of Statistical Physics, 34 (5–6):941–961, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009450.

Kogut:1986:PBS

[Kog86] John B. Kogut. Progress and bottlenecks in simulating lattice gauge theory with fermions. Journal of Statistical Physics, 43 (5–6):771–795, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02628304.

Kohn:1988:SDI

[Koh88] Walter Kohn. Semiclassical description of inelastic atom scattering by surfaces. Journal of Statistical Physics, 52(5–6):1197– 1201, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011641.

Koonin:1986:AFM

[Koo86] Steven E. Koonin. Auxiliary-field Monte Carlo methods. Journal of Statistical Physics, 43(5–6):985–990, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02628325.

Kopelman:1986:RPF

[Kop86] Raoul Kopelman. Rate processes on fractals: Theory, simulations, and experiments. Journal of Statistical Physics, 42(1– 2):185–200, January 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010846.

Korshunov:1986:PTTb

[Kor86] S. E. Korshunov. Phase transitions in two-dimensional uniformly frustrated XY models. II. General scheme. Journal of Statistical Physics, 43(1–2):17–32, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010570.

Kostin:1986:VPD

[Kos86] M. D. Kostin. Velocity of propagation in diffusional quantum theory. Journal of Statistical Physics, 45(3–4):765–767, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01021095.

Klauder:1985:SCN

[KP85] John R. Klauder and Wesley P. Petersen. Spectrum of certain non-self-adjoint operators and solutions of Langevin equations with complex drift. Journal of Statistical Physics, 39(1–2):53–72, April 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01007974.

Kolb:1986:PTE

[KP86] M. Kolb and K. A. Penson. Phase transitions in electron systems with short-range pairing interactions: Ground state. Journal of Statistical Physics, 44(1–2):129–152, July 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010908.

Kaiser:1989:SCM

[KP89a] Christian Kaiser and Ingo Peschel. Surface and corner magnetizations in the two-dimensional Ising model. Journal of Statistical Physics, 54(3–4):567–579, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019769.

Koch:1989:SRR

[KP89b] Hans Koch and Jacques Piasko. Some rigorous results on the Hopfield neural network model. Journal of Statistical Physics, 55 (5–6):903–928, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01041071.

Khoruzhenko:1989:LLH

[KPS89] B. A. Khoruzhenko, L. A. Pastur, and M. V. Shcherbina. Largen limit of the Heisenberg model: The decorated lattice and the disordered chain. Journal of Statistical Physics, 57(1–2):41–52, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023633.

Kusmartsev:1985:SSB

[KR85] F. V. Kusmartsev and E. I. Rashba. Spontaneous symmetry breaking of optimum fluctuations in semiconductors. Journal of Statistical Physics, 38(1–2):313–327, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01017864.

Kratky:1981:IDS

[Kra81] Karl W. Kratky. Intersecting disks (and spheres) and statistical mechanics. I. Mathematical basis. Journal of Statistical Physics, 25(4):619–634, August 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01022357.

Kratky:1982:NGE

[Kra82a] Karl W. Kratky. A new graph expansion of virial coefficients. Journal of Statistical Physics, 27(3):533–551, March 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011091.

Kratky:1982:OGR

[Kra82b] Karl W. Kratky. Overlap graph representation of B_{6} and B_{7}. Journal of Statistical Physics, 29(1):129–138, September 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008253.

Kratky:1985:EIF

[Kra85] Karl W. Kratky. Explicit and implicit finiteness corrections of virial coefficients. Journal of Statistical Physics, 39(3–4):379– 403, May 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01018669.

Kraichnan:1988:RDH

[Kra88a] Robert H. Kraichnan. Reduced descriptions of hydrodynamic turbulence. Journal of Statistical Physics, 51(5–6):949–963, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014893.

Kratky:1988:PTH

[Kra88b] Karl W. Kratky. Is the percolation transition of hard spheres a thermodynamic phase transition? Journal of Statistical Physics, 52(5–6):1413–1421, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011656.

Kremen:1980:CGF

[Kre80] Alexandr Kremen. Coding of Gibbs function flows from nucleotide pools. Journal of Statistical Physics, 23(4):483–494, October 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011577.

Kirillov:1989:VDT

[KRS89] A. B. Kirillov, D. C. Radulescu, and D. F. Styer. Vasserstein distances in two-state systems. Journal of Statistical Physics, 56(5– 6):931–937, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016786.

Kotulski:1981:LSN

[KS81] Z. Kotulski and K. Sobczyk. Linear systems and normality. Journal of Statistical Physics, 24(2):359–373, February 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01013306.

Khokhlov:1985:TLC

[KS85] A. R. Khokhlov and A. N. Semenov. On the theory of liquidcrystalline ordering of polymer chains with limited flexibility. Journal of Statistical Physics, 38(1–2):161–182, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017855.

Kirsch:1986:LTP

[KS86] W. Kirsch and B. Simon. Lifshitz tails for periodic plus random potentials. Journal of Statistical Physics, 42(5–6):799–808, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010445.

Kremp:1985:NRT

[KSB85] D. Kremp, M. Schlanges, and Th. Bornath. Nonequilibrium real time Green’s functions and the condition of weakening of initial correlation. Journal of Statistical Physics, 41(3–4):661–670, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009026.

Kincaid:1989:NMC

[KSF89] J. M. Kincaid, K. B. Shon, and G. Fescos. New methods for calculating the dew/bubble curves of classical model fluids. Journal of Statistical Physics, 57(3–4):937–963, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01022842.

Kramli:1987:HCC

[KSS87] András Krámli, Nándor Simányi, and Domokos Szász. Heat conduction in caricature models of the Lorentz gas. Journal of Statistical Physics, 46(1–2):303–318, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010348.

Krieg:1989:PCP

[KST89] R. Krieg, K.-J. Schmitt, and C. Toepffer. m-Particle correlations in the n-particle McKean model. Journal of Statistical Physics, 57 (1–2):267–288, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023643.

Kawasaki:1985:DGF

[KT85] Kyozi Kawasaki and Paolo Tombesi. Domain growth of first-order phase transitions: An exactly solvable model. Journal of Statistical Physics, 38(1–2):207–215, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01017858.

Kawabata:1986:MCM

[KTB86] C. Kawabata, M. Takeuchi, and A. R. Bishop. Monte Carlo molecular dynamics simulations for two-dimensional magnets. Journal of Statistical Physics, 43(5–6):869–871, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02628313.

Keller:1988:CGA

[KTP88] U. Keller, B. Thomas, and H.-J. Pohley. Comment on a genetic application of square-lattice Kauffman models. Journal of Statistical Physics, 52(3–4):1129–1132, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019744.

Korshunov:1986:PTTa

[KU86] S. E. Korshunov and G. V. Uimin. Phase transitions in twodimensional uniformly frustrated XY models. I. Antiferromagnetic model on a triangular lattice. Journal of Statistical Physics, 43(1–2):1–16, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010569.

Kung:1986:QMC

[Kun86] David S. Kung. Quantum Monte Carlo of coupled fermion chains. Journal of Statistical Physics, 43(5–6):873–874, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02628314.

Kuroda:1983:PSI

[Kur83a] Koji Kuroda. Phase separations in Ising model with free boundary condition. Journal of Statistical Physics, 30(1):1–13, January 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010865.

Kurtze:1983:YLE

[Kur83b] Douglas A. Kurtze. The Yang–Lee edge singularity in onedimensional Ising and n-vector models. Journal of Statistical Physics, 30(1):15–35, January 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010866.

Kurten:1988:ISM

[Kür88] K. E. Kürten. An immune system model in discrete time based on the analogy with the central nervous system. Journal of Statistical Physics, 52(1–2):489–495, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01016430.

Kiefer:1982:TSD

[KV82] J. Kiefer and P. B. Visscher. Tail shortening by discrete hydrodynamics. Journal of Statistical Physics, 27(2):389–405, February 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008945.

Kutner:1985:IUD

[KvB85] R. Kutner and H. van Beijeren. Influence of a uniform driving force on tracer diffusion in a one-dimensional hard-core lattice gas. Journal of Statistical Physics, 39(3–4):317–325, May 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01018665.

Kuz:1987:TST

[KVGZ87] V. A. Kuz, M. A. Vila, A. N. Garazo, and G. J. Zarragoicoechea. Theory of surface tension and its application to simple fluids. Journal of Statistical Physics, 49(5–6):1209–1219, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017567.

Klein:1981:CEG

[KW81] Douglas J. Klein and Terry L. Welsher. Cluster expansion and generalized transfer matrices for the statistical mechanics of linear chains. Journal of Statistical Physics, 24(4):555–586, April 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011375.

Kalle:1982:MCT

[KW82] C. Kalle and V. Winkelmann. Monte Carlo technique for very large Ising models. Journal of Statistical Physics, 28(4):639–648, August 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011873.

Knobloch:1983:BFS

[KW83] E. Knobloch and K. A. Wiesenfeld. Bifurcations in fluctuating systems: The center-manifold approach. Journal of Statistical Physics, 33(3):611–637, December 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01018837.

Khandekar:1988:SMP

[KW88] D. C. Khandekar and F. W. Wiegel. Statistical mechanical properties of polymer configurations which enclose a constant area. Journal of Statistical Physics, 53(5–6):1073–1080, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023858.

Kottalam:1987:ABL

[KWL87a] J. Kottalam, Bruce J. West, and Katja Lindenberg. Analogy between the Lorenz strange attractor and a bistable stochastic oscillator. Journal of Statistical Physics, 46(1–2):119–133, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010335.

Kottalam:1987:ESR

[KWL87b] J. Kottalam, Bruce J. West, and Katja Lindenberg. Extension of statistical replacement to systems with time-correlated fluctuations. Journal of Statistical Physics, 49(1–2):347–367, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009968.

Kostelich:1985:BRB

[KY85] Eric Kostelich and James A. Yorke. Book review: Nonlinear oscillations, dynamical systems, and bifurcations of vector fields. Journal of Statistical Physics, 38(3–4):805–806, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010491.

Klein:1988:PTO

[KY88a] David Klein and Wei-Shih Yang. Phase transition for a onedimensional lattice gas with hard core. Journal of Statistical Physics, 52(3–4):735–745, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01019727.

Kuniba:1988:LSP

[KY88b] Atsuo Kuniba and Tetsu Yajima. Local state probabilities for solvable restricted solid-on-solid models: A_{n}, D_{n}, D_{n}^{(1)}, and A_{n}^{(1)}. Journal of Statistical Physics, 52(3–4):829–883, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01019732.

Klein:1989:SCI

[KY89] David Klein and Wei-Shih Yang. Some correlation inequalities for Ising antiferromagnets. Journal of Statistical Physics, 57(5–6): 1049–1058, December 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020047.

Kesten:1987:SIS

[KZ87] Harry Kesten and Yu Zhang. Strict inequalities for some critical exponents in two-dimensional percolation. Journal of Statistical Physics, 46(5–6):1031–1055, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011155.

Ladd:1989:DAC

[LA89] Anthony J. C. Ladd and Berni J. Alder. Decay of angular correlations in hard-sphere fluids. Journal of Statistical Physics, 57(3– 4):473–482, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01022817.

Ladd:1987:SRD

[LAA87] Anthony J. C. Ladd, W. Edward Alley, and Berni J. Alder. Structural relaxation in dense hard-sphere fluids. Journal of Statistical Physics, 48(5–6):1147–1156, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009538.

Lam:1988:PAK

[Lam88] P. M. Lam. A percolation approach to the Kauffman model. Journal of Statistical Physics, 50(5–6):1263–1269, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01019165.

Lam:1989:CPC

[Lam89] P. M. Lam. Collapse of percolation clusters — a transfer matrix study. Journal of Statistical Physics, 54(3–4):1081–1089, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019788.

Landsberg:1984:EPT

[Lan84a] P. T. Landsberg. The ellipsoid paradox of thermodynamics. Journal of Statistical Physics, 34(1–2):357, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01770368.

Landsberg:1984:EAE

[Lan84b] P. T. Landsberg. Is equilibrium always an entropy maximum? Journal of Statistical Physics, 35(1–2):159–169, April 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017372.

Lanford:1984:FEC

[Lan84c] Oscar E. Lanford III. Functional equations for circle homeomorphisms with golden ratio rotation number. Journal of Statistical Physics, 34(1–2):57–73, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01770349.

Landauer:1988:MNS

[Lan88] Rolf Landauer. Motion out of noisy states. Journal of Statistical Physics, 53(1–2):233–248, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011555.

Landauer:1989:DNI

[Lan89] Rolf Landauer. Dissipation and noise immunity in computation, measurement, and communication. Journal of Statistical Physics, 54(5–6):1509–1517, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01044731.

Larsson:1986:MKR

[Lar86] Thomas A. Larsson. Migdal–Kadanoff renormalization group for the O(n) model. Journal of Statistical Physics, 43(3–4):671–686, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020659.

Lavaud:1982:EGMb

[Lav82a] Michel Lavaud. Estimates of general Mayer graphs. I: Construction of upper bounds for a given graph by means of sets of subgraphs. Journal of Statistical Physics, 27(3):593–627, March 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011094.

Lavaud:1982:EGMc

[Lav82b] Michel Lavaud. Estimates of general Mayer graphs III: Upper bounds obtained by means of spanning n-trees. Journal of Statistical Physics, 27(4):745–766, April 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013446.

Lavaud:1982:EGMa

[Lav82c] Michel Lavaud. Estimates of general Mayer graphs. IV. On the computation of Gaussian integrals by the star-mesh transformation. Journal of Statistical Physics, 27(1):57–64, January 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011739.

Lavaud:1987:BSD

[Lav87] Michel Lavaud. Behavior at small distances and low temperatures of the ion-ion distribution function of the two-dimensional Coulomb gas. Journal of Statistical Physics, 49(5–6):1191–1208, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017566.

Lawler:1988:LES

[Law88] Gregory F. Lawler. Loop-erased self-avoiding random walk in two and three dimensions. Journal of Statistical Physics, 50(1– 2):91–108, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01022989.

Lax:1986:MC

[Lax86] Peter D. Lax. Mathematics and computing. Journal of Statistical Physics, 43(5–6):749–756, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02628302.

Lu:1987:PSQ

[LB87] Jian Ping Lu and Joseph L. Birman. Percolation and scaling on a quasilattice. Journal of Statistical Physics, 46(5–6):1057–1066, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011156.

Lutton:1983:BPM

[LBF83] Jean-Luc Lutton, Ernesto Bonomi, and Marc Roy Feix. Blocking probability for a multistage clos connecting network. Journal of Statistical Physics, 31(1):155–167, April 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010928.

Lei:1985:GDP

[LCG85] Lin Lei, Shu Changqing, and Xu Gang. Generation and detection of propagating solitons in shearing liquid crystals. Journal of Statistical Physics, 39(5–6):633–652, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008357. See erra-

tum [LCG86].

Lei:1986:EGD

[LCG86] Lin Lei, Shu Changqing, and Xu Gang. Erratum: Generation and detection of propagating solitons in shearing liquid crystals. Journal of Statistical Physics, 43(1–2):391, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01010589. See [LCG85].

Leutheusser:1983:VHS

[LCY83] E. Leutheusser, D. P. Chou, and Sidney Yip. Van Hove selfcorrelation function of a hard-disk fluid: Enskog theory and computer simulation. Journal of Statistical Physics, 32(3):523– 538, September 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008953.

Lam:1987:CTC

[LD87] P. M. Lam and J. A. M. S. Duarte. Collapse transition and cyclomatic number distribution of directed lattice animals. Journal of Statistical Physics, 49(1–2):245–260, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009961.

LeMehaute:1984:TPF

[Le 84] Alain Le Mehaute. Transfer processes in fractal media. Journal of Statistical Physics, 36(5–6):665–676, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012930.

Lebowitz:1981:IDM

[Leb81] Joel Lebowitz. This issue is dedicated to the memory of Pierre Résibois. Journal of Statistical Physics, 24(1):3, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01007629.

Lebowitz:1984:F

[Leb84] Joel L. Lebowitz. Foreword. Journal of Statistical Physics, 34(5– 6):659, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01009432;   http://link.
springer.com/article/10.1007/BF01009432.

Lebowitz:1985:EPa

[Leb85a] Joel L. Lebowitz. Editor’s preface. Journal of Statistical Physics, 38(1–2):i–1, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01017843.

Lebowitz:1985:EPb

[Leb85b] Joel L. Lebowitz. Editor’s preface. Journal of Statistical Physics, 39(5–6):ii, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01008343;
http://link.springer.com/article/10.1007/BF01008343.

Lebowitz:1986:BRS

[Leb86a] Joel L. Lebowitz. Book review: Statistical physics II: Nonequilibrium statistical mechanics. Journal of Statistical Physics, 44 (3–4):697–699, August 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011313.

Lebowitz:1986:PSMa

[Leb86b] Joel L. Lebowitz. Program of the 54th Statistical Mechanics Meeting. Journal of Statistical Physics, 43(3–4):717–725, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020664.

Lebowitz:1986:PSMb

[Leb86c] Joel L. Lebowitz. Program of the 55th Statistical Mechanics Meeting. Journal of Statistical Physics, 45(5–6):1115–1119, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020593.

Lebowitz:1987:Pa

[Leb87] Joel L. Lebowitz. Preface. Journal of Statistical Physics, 46 (5–6):801–802, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011141.

Lebowitz:1988:ENa

[Leb88a] Joel L. Lebowitz. Editor’s note. Journal of Statistical Physics, 50(5–6):1287, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019171.

Lebowitz:1988:ENb

[Leb88b] Joel L. Lebowitz. Editor’s note. Journal of Statistical Physics, 53 (3–4):1019, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014243.

Lebowitz:1988:P

[Leb88c] Joel L. Lebowitz. Preface. Journal of Statistical Physics, 51 (5–6):741–742, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014881.

Lebowitz:1988:PSM

[Leb88d] Joel L. Lebowitz. Program of the 58th Statistical Mechanics Meeting. Journal of Statistical Physics, 51(3–4):729–737, May 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01028481.

Lebowitz:1989:BRS

[Leb89a] Joel L. Lebowitz. Book review: Statistical mechanics: Rigorous results. Journal of Statistical Physics, 56(3–4):559, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01044455.

Lebowitz:1989:PSMa

[Leb89b] Joel L. Lebowitz. Program of the 59th Statistical Mechanics Meeting. Journal of Statistical Physics, 54(3–4):1103–1108, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019792.

Lebowitz:1989:PSMb

[Leb89c] Joel L. Lebowitz. Program of the 60th Statistical Mechanics Meeting. Journal of Statistical Physics, 55(5–6):1361–1370, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01041097.

Lebowitz:1989:PSS

[Leb89d] Joel L. Lebowitz. Program of the 61st Semiannual Statistical Mechanics Meeting. Journal of Statistical Physics, 56(5–6): 977–983, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016792.

Lecar:1985:BRR

[Lec85] Harold Lecar. Book review: Random walks in biology. Journal of Statistical Physics, 40(3–4):609–610, August 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01017189.

Lee:1983:BRE

[Lee83] B. K. Lee. Book review: Eighth symposium on thermophysical properties. Journal of Statistical Physics, 32(1):203–204, July 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009429.

Lee:1987:DEC

[Lee87] T. D. Lee. Difference equations and conservation laws. Journal of Statistical Physics, 46(5–6):843–860, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011145.

Lendi:1988:MIE

[Len88] K. Lendi. Measure of irreversibility and entropy production in open quantum systems. Journal of Statistical Physics, 50(5–6): 1103–1109, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019155.

Lesche:1982:IRE

[Les82] Bernhard Lesche. Instabilities of Rényi entropies. Journal of Statistical Physics, 27(2):419–422, February 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008947.

Leuthausser:1987:SME

[Leu87] Ira Leuthäusser. Statistical mechanics of Eigen’s evolution model. Journal of Statistical Physics, 48(1–2):343–360, July 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010413.

Levy:1986:NSB

[Lév86] Yves-Emmanuel Lévy. A note on Sinaï and Bunimovich’s Markov partition for billiards. Journal of Statistical Physics, 45(1– 2):63–68, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01033077.

Levine:1988:SWF

[Lev88] R. D. Levine. The statistical wave function. Journal of Statistical Physics, 52(5–6):1203–1220, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011642.

Lipscombe:1988:SII

[LFtH88] T. Lipscombe, A. L. Frenkel, and D. ter Haar. Stability of isotropic incompressible turbulence against weak mean flow perturbations. Journal of Statistical Physics, 53(1–2):95–108, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011547.

Lai:1987:ASM

[LG87] Pik-Yin Lai and Yadin Y. Goldschmidt. Application of statistical mechanics to combinatorial optimization problems: The chromatic number problem and q-partitioning of a graph. Journal of Statistical Physics, 48(3–4):513–529, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019685.

Lifshitz:1985:SCI

[LGP85] I. M. Lifshitz, S. A. Gredeskul, and L. A. Pastur. Spectral characteristics of the impurity band in the structural disorder model. Journal of Statistical Physics, 38(1–2):37–63, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017847.

Ladd:1985:LGS

[LH85] Anthony J. C. Ladd and William G. Hoover. Lorentz gas shear viscosity via nonequilibrium molecular dynamics and Boltzmann’s equation. Journal of Statistical Physics, 38(5–6):973–988, March 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010425.

Lam:1988:RET

[LH88a] P. M. Lam and Alex Hansen. Resistivity exponent of twodimensional lattice animals. Journal of Statistical Physics, 52 (1–2):447–451, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016425.

Landsberg:1988:CVG

[LH88b] P. T. Landsberg and P. Harshman. Canonical versus grand canonical occupation numbers for simple systems. Journal of Statistical Physics, 53(1–2):475–482, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011567.

Laidlaw:1988:FPF

[LHFW88] W. G. Laidlaw, G. R. Hamilton, R. B. Flewwelling, and W. G. Wilson. Fractal patterns of fluid domains for displacement processes in porous media. Journal of Statistical Physics, 53(3–4): 713–731, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014222.

Hao:1982:HCB

[lHyZ82] Bai lin Hao and Shu yu Zhang. Hierarchy of chaotic bands. Journal of Statistical Physics, 28(4):769–792, August 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011880.

Libchaber:1984:EOC

[Lib84] A. Libchaber. Experiments on the onset of chaotic behavior. Journal of Statistical Physics, 34(5–6):1047, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009457.

Lindenberg:1984:BRI

[Lin84] Katja Lindenberg. Book review: Introduction to statistical mechanics and thermodynamics. Journal of Statistical Physics, 37(5–6):711–712, December 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010504.

Lindman:1985:NSI

[Lin85] E. L. Lindman. Numerical simulation of ion acoustic turbulence. Journal of Statistical Physics, 39(5–6):769–782, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008365.

Lin:1987:SMI

[Lin87] K. Y. Lin. Spontaneous magnetization of the Ising model on a 4–8 lattice. Journal of Statistical Physics, 49(1–2):269–279, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009963.

Lindenberg:1988:BRI

[Lin88] Katja Lindenberg. Book review: Introduction to modern statistical mechanics. Journal of Statistical Physics, 52(3–4):1133–1134, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019745.

Lin:1989:TSC

[Lin89a] K. Y. Lin. Three-spin correlation of the Ising model on the generalized checkerboard lattice. Journal of Statistical Physics, 56(5– 6):631–643, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016771.

Lindenberg:1989:BRN

[Lin89b] Katja Lindenberg. Book review: Nonlinear stochastic systems in physics and mechanics. Journal of Statistical Physics, 54(3–4): 1099–1100, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019790.

Lawrence:1985:MDD

[LKMJ85] Albert Lawrence, Nung Soo Kim, James McDaniel, and Michael Jack. Multifluxon dynamics in driven Josephson junctions. Journal of Statistical Physics, 39(5–6):563–581, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008352.

Lindenberg:1986:BR

[LL86] Katja Lindenberg and Katja Lindenberg. Book review. Journal of Statistical Physics, 45(5–6):1091–1094, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020592.

Lanford:1983:AWS

[LLE^{+}83] Oscar E. Lanford, Thomas Lundgren, J. P. Eckmann, Pierre Collet, D. G. Aronson, Giovanni Gallavotti, C. Conley, Richard McGehee, Charles Tresser, Oscar E. Lanford, L. Gross, C. M. Newman, William G. Faris, John Palmer, Craig Tracy, and et al. Abstracts of the workshop on statistical mechanics, dynamical systems, and turbulence. Journal of Statistical Physics, 32(1):169–202, July 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009428.

Lemarchand:1988:IHB

[LLS88] A. Lemarchand, H. Lemarchand, and E. Sulpice. Interaction of a Hopf bifurcation and a symmetry-breaking bifurcation: Stochastic potential and spatial correlations. Journal of Statistical Physics, 53(3–4):613–654, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01014217.

Lebowitz:1984:PFF

[LM84] J. L. Lebowitz and Ph. A. Martin. On potential and field fluctuations in classical charged systems. Journal of Statistical Physics, 34(1–2):287–311, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01770360.

Luciani:1986:RMC

[LM86] J. F. Luciani and P. Mora. Resummation methods of the Chapman–Enskog expansion for a strongly inhomogeneous plasma. Journal of Statistical Physics, 43(1–2):281–302, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010582.

Lebowitz:1987:EEF

[LM87] Joel L. Lebowitz and Christian Maes. The effect of an external field on an interface, entropic repulsion. Journal of Statistical Physics, 46(1–2):39–49, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010329.

Lin:1988:SMI

[LM88] K. Y. Lin and S. K. Ma. Spontaneous magnetization of the Ising model on a layered square lattice. Journal of Statistical Physics, 51(3–4):491–499, May 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01028468.

Laidlaw:1987:SFP

[LMJ87] Don Laidlaw, Gary MacKay, and Naeem Jan. Some fractal properties of the percolating backbone in two dimensions. Journal of Statistical Physics, 46(3–4):507–515, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01013371.

Lehr:1984:CNL

[LMN84] Walter Lehr, Jonathan Machta, and Mark Nelkin. Current noise and long time tails in biased disordered random walks. Journal of Statistical Physics, 36(1–2):15–29, July 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01015723.

Livi:1988:SOD

[LMR88] Roberto Livi, Amos Maritan, and Stefano Ruffo. The spectrum of a one-dimensional hierarchical model. Journal of Statistical Physics, 52(3–4):595–608, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01019719.

Lemarchand:1984:SAS

[LN84] H. Lemarchand and G. Nicolis. Stochastic analysis of symmetrybreaking bifurcations: Master equation approach. Journal of Statistical Physics, 37(5–6):609–629, December 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010498.

Lima:1985:STS

[LN85] R. Lima and J. Naudts. Spontaneous translation symmetry breakdown for quantum lattice systems. Journal of Statistical Physics, 39(1–2):157–165, April 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01007977.

Lebowitz:1987:Pb

[LN87] Joel L. Lebowitz and Gregoire Nicolis. Preface. Journal of Statistical Physics, 48(5–6):957, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009525.

Luck:1988:LTL

[LN88] J. M. Luck and Th. M. Nieuwenhuizen. Lifshitz tails and longtime decay in random systems with arbitrary disorder. Journal of Statistical Physics, 52(1–2):1–22, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01016401.

Larkin:1985:CCQ

[LO85] A. I. Larkin and Yu. N. Ovchinnikov. The crossover from classical to quantum regime in the problem of the decay of the metastable state. Journal of Statistical Physics, 41(3–4):425–443, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009016.

Loh:1986:MCM

[Loh86] E. Loh, Jr. A Monte Carlo method for quantum spins using boson world lines. Journal of Statistical Physics, 43(5–6):875–876, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02628315.

Lomdahl:1985:SJJ

[Lom85] P. S. Lomdahl. Solitons in Josephson junctions: An overview. Journal of Statistical Physics, 39(5–6):551–561, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008351.

Lopes:1988:CPG

[Lop88] Artur Oscar Lopes. The complex potential generated by the maximal measure for a family of rational maps. Journal of Statistical Physics, 52(3–4):571–575, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01019717.

Lorenz:1984:VNS

[Lor84] Edward N. Lorenz. A very narrow spectral band. Journal of Statistical Physics, 36(1–2):1–14, July 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01015722.

Luck:1988:RTL

[LOS88] J. M. Luck, H. Orland, and U. Smilansky. On the response of a two-level quantum system to a class of time-dependent quasiperiodic perturbations. Journal of Statistical Physics, 53(3– 4):551–564, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014213.

Luck:1986:PSO

[LP86] J. M. Luck and D. Petritis. Phonon spectra in one-dimensional quasicrystals. Journal of Statistical Physics, 42(3–4):289–310, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01127714.

Lebowitz:1989:EP

[LP89] Joel L. Lebowitz and Errico Presutti. Editors’ preface. Journal of Statistical Physics, 55(3–4):475, May 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01041594.

Lewis:1984:SPP

[LPdS84] J. T. Lewis, J. V. Pulè, and P. de Smedt. The superstability of pair-potentials of positive type. Journal of Statistical Physics, 35 (3–4):381–385, May 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014391.

Livi:1987:CBN

[LPRV87a] Roberto Livi, Marco Pettini, Stefano Ruffo, and Angelo Vulpiani. Chaotic behavior in nonlinear Hamiltonian systems and equilibrium statistical mechanics. Journal of Statistical Physics, 48(3– 4):539–559, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019687.

Livi:1987:LEH

[LPRV87b] Roberto Livi, Antonio Politi, Stefano Ruffo, and Angelo Vulpiani. Liapunov exponents in high-dimensional symplectic dynamics. Journal of Statistical Physics, 46(1–2):147–160, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010337.

Lebowitz:1985:PDC

[LPS85] Joel L. Lebowitz, Mohan K. Phani, and Daniel F. Styer. Phase diagram of Cu–Au-type alloys. Journal of Statistical Physics, 38 (1–2):413–431, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01017871.

Lebowitz:1988:MMH

[LPS88] Joel L. Lebowitz, Errico Presutti, and Herbert Spohn. Microscopic models of hydrodynamic behavior. Journal of Statistical Physics, 51(5–6):841–862, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014887.

Lebowitz:1987:OPS

[LPT87] J. L. Lebowitz, J. K. Percus, and J. Talbot. On the orientational properties of some one-dimensional model systems. Journal of Statistical Physics, 49(5–6):1221–1234, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01017568.

Landau:1981:EGC

[LPW81] L. Landau, J. Fernando Perez, and W. F. Wreszinski. Energy gap, clustering, and the Goldstone theorem in statistical mechanics. Journal of Statistical Physics, 26(4):755–766, December 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010937.

Lewis:1988:ETS

[LR88a] J. T. Lewis and G. A. Raggio. The equilibrium thermodynamics of a spin-boson model. Journal of Statistical Physics, 50(5– 6):1201–1220, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019161.

Lima:1988:SLA

[LR88b] Ricardo Lima and Stefano Ruffo. Scaling laws for all Liapunov exponents: Models and measurements. Journal of Statistical Physics, 52(1–2):259–265, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01016413.

Lebowitz:1988:SMN

[LRS88] Joel L. Lebowitz, Harvey A. Rose, and Eugene R. Speer. Statistical mechanics of the nonlinear Schrödinger equation. Journal of Statistical Physics, 50(3–4):657–687, February 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01026495.

Lebowitz:1989:SMN

[LRS89] Joel L. Lebowitz, Harvey A. Rose, and Eugene R. Speer. Statistical mechanics of the nonlinear Schrödinger equation. II. Mean field approximation. Journal of Statistical Physics, 54(1–2):17–56, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023472.

Lavenda:1982:TCG

[LS82a] B. H. Lavenda and E. Santamato. Thermodynamic criteria governing irreversible processes under the influence of small thermal fluctuations. Journal of Statistical Physics, 29(2):345–361, October 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020791.

Lebowitz:1982:MBF

[LS82b] J. L. Lebowitz and H. Spohn. Microscopic basis for Fick’s law for self-diffusion. Journal of Statistical Physics, 28(3):539–556, July 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01008323.

Lebowitz:1982:SSS

[LS82c] J. L. Lebowitz and H. Spohn. Steady state self-diffusion at low density. Journal of Statistical Physics, 29(1):39–55, September 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008247.

Lavis:1984:RGS

[LS84] D. A. Lavis and B. W. Southern. Renormalization group study of a three-dimensional lattice model with directional bonding. Journal of Statistical Physics, 35(5–6):489–506, June 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010822.

Lebowitz:1987:AOT

[LS87] J. L. Lebowitz and R. H. Schonmann. On the asymptotics of occurrence times of rare events for stochastic spin systems. Journal of Statistical Physics, 48(3–4):727–751, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019694.

Law:1989:FFT

[LS89a] B. M. Law and J. V. Sengers. Fluctuations in fluids out of thermal equilibrium. Journal of Statistical Physics, 57(3–4):531–547, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01022821.

Loss:1989:NQS

[LS89b] D. Loss and H. Schoeller. A new quantum statistical evaluation method for time correlation functions. Journal of Statistical Physics, 54(3–4):765–795, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019775.

Loss:1989:QSK

[LS89c] D. Loss and H. Schoeller. Quantum-statistical kinetic equations. Journal of Statistical Physics, 56(1–2):175–201, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01044240.

Lima-Santos:1989:NRS

[LSM89] A. Lima-Santos and E. C. Marino. Nonperturbative renormalization of the sine–Gordon/ Coulomb gas system for β^{2} < 8π: A functional integral approach. Journal of Statistical Physics, 55 (1–2):157–170, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01042595.

Lindenberg:1980:LRW

[LSSW80a] Katja Lindenberg, V. Seshadri, K. E. Shuler, and George H. Weiss. Lattice random walks for sets of random walkers. First passage times. Journal of Statistical Physics, 23(1):11–25, July 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014427.

Lindenberg:1980:ETS

[LSSW80b] Katja Lindenberg, V. Seshadri, K. E. Shuler, and Bruce J. West. Equal-time second-order moments of a harmonic oscillator with stochastic frequency and driving force. Journal of Statistical Physics, 23(6):755–765, December 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008517.

Liverani:1986:IKE

[LT86] C. Liverani and G. Turchetti. Improved KAM estimates for the Siegel radius. Journal of Statistical Physics, 45(5–6):1071–1086, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020589.

Lefever:1987:IFN

[LTL87] R. Lefever, J. Wm. Turner, and L. A. Lugiato. Influence of frequency noise on nascent hysteresis in optical bistability. Journal of Statistical Physics, 48(5–6):1045–1058, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009532.

Lu:1987:ECS

[Lu87] Binglin Lu. The evolution of the cluster size distribution in a coagulation system. Journal of Statistical Physics, 49(3–4):669–684, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009352.

Luczka:1986:EPD

[Luc86a] J. Luczka. Exact probability distribution for soluble model with quadratic noise. Journal of Statistical Physics, 42(5–6): 1009–1018, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010459.

Luczka:1986:RPQ

[Luc86b] J. Luczka. Relaxation problem with quadratic noise. Journal of Statistical Physics, 45(1–2):309–317, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01033093.

Luczka:1987:RPQ

[Luc87] J. Luczka. Relaxation problem with a quadratic noise: Analysis. Journal of Statistical Physics, 47(3–4):505–526, May 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01007523.

Luzzi:1980:RFT

[LV80] Roberto Luzzi and Aurea R. Vasconcellos. Response function theory for far-from-equilibrium statistical systems. Journal of Statistical Physics, 23(5):539–559, November 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011729.

Luciani:1988:BDC

[LV88] J. F. Luciani and A. D. Verga. Bistability driven by correlated noise: Functional integral treatment. Journal of Statistical Physics, 50(3–4):567–597, February 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01026491.

Leegwater:1989:HCF

[LvB89a] Jan A. Leegwater and Henk van Beijeren. Hydrodynamic correlation functions of hard-sphere fluids at short times. Journal of Statistical Physics, 57(3–4):595–629, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01022824.

Leegwater:1989:OPC

[LvB89b] Jan A. Leegwater and Henk van Beijeren. The orientational pair correlation functions in a dense hard sphere fluid at long times. Journal of Statistical Physics, 57(1–2):383–391, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023650.

Levesque:1983:SPT

[LW83] D. Levesque and J. J. Weis. Surface properties of the threedimensional one-component plasma. Journal of Statistical Physics, 33(3):549–557, December 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01018833.

Levesque:1985:OSD

[LW85] D. Levesque and J.-J. Weis. Orientational structure of dipolar hard spheres near a hard neutral wall. Journal of Statistical Physics, 40(1–2):29–38, July 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010525.

Lindenberg:1986:FBO

[LW86] Katja Lindenberg and Bruce J. West. The first, the biggest, and other such considerations. Journal of Statistical Physics, 42(1– 2):201–243, January 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010847.

Lin:1988:MIM

[LW88] K. Y. Lin and F. Y. Wu. Magnetization of the Ising model on the generalized checkerboard lattice. Journal of Statistical Physics, 52 (3–4):669–678, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019723.

LuValle:1988:ATS

[LWS88] M. J. LuValle, T. L. Welsher, and K. Svoboda. Acceleration transforms and statistical kinetic models. Journal of Statistical Physics, 52(1–2):311–330, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01016417.

Leutheusser:1983:DCH

[LYAA83] E. Leutheusser, Sidney Yip, B. J. Alder, and W. E. Alley. Dynamical correlations in a hard-disk fluid: Generalized Enskog theory. Journal of Statistical Physics, 32(3):503–521, September 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008952.

Maassen:1984:RTE

[Maa84] Hans Maassen. Return to thermal equilibrium by the solution of a quantum Langevin equation. Journal of Statistical Physics, 34 (1–2):239–261, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01770357.

MacGowan:1983:SCM

[Mac83a] David MacGowan. The self-consistent mean spherical approximation for the one-component plasma. Journal of Statistical Physics, 32(1):123–139, July 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009425.

Machta:1983:PLD

[Mac83b] Jonathan Machta. Power law decay of correlations in a billiard problem. Journal of Statistical Physics, 32(3):555–564, September 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01008956.

Machta:1983:RGA

[Mac83c] Jonathan Machta. Renormalization group approach to random walks on disordered lattices. Journal of Statistical Physics, 30 (2):305–314, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012305.

Machta:1986:ALT

[Mac86] Jonathan Machta. Anomalous long time tails due to trapping. Journal of Statistical Physics, 42(5–6):941–948, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010455.

Machacek:1987:DSR

[Mac87] Martin Machácek. Dynamical systems with relaxation time. Journal of Statistical Physics, 47(5–6):949–952, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01206169.

Madras:1988:EPS

[Mad88] Neal Madras. End patterns of self-avoiding walks. Journal of Statistical Physics, 53(3–4):689–701, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014220.

Magnen:1987:PSA

[Mag87] Jacques Magnen. Phase space analysis and particle structure in field theory. Journal of Statistical Physics, 47(5–6):851–865, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01206161.

Maier:1987:BDS

[Mai87] Robert S. Maier. Bounds on the density of states of random Schrödinger operators. Journal of Statistical Physics, 48(3–4):

425–447, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019681.

Majewski:1987:AND

[Maj87] W. A. Majewski. Asymptotically normal dynamical semigroups. Journal of Statistical Physics, 47(5–6):953–957, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01206170.

Majewski:1989:RES

[Maj89] Władysław A. Majewski. Return to equilibrium and stability of dynamics (semigroup dynamics case). Journal of Statistical Physics, 55(1–2):417–433, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01042609.

Malakis:1982:TTS

[Mal82] A. Malakis. Two- and three-spin triangular Ising model: Variational approximations. Journal of Statistical Physics, 27(1):1–17, January 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011736.

Mandelbrot:1984:CEB

[Man84a] Benoît B. Mandelbrot. Comment on the equivalence between fracton/spectral dimensionality, and the dimensionality of recurrence. Journal of Statistical Physics, 36(5– 6):541–543, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012919;   http:
//www.springerlink.com/content/k046263428n04707/.

Mandelbrot:1984:FPS

[Man84b] Benoît B. Mandelbrot. Fractals in physics: Squig clusters, diffusions, fractal measures, and the unicity of fractal dimensionality. Journal of Statistical Physics, 34(5–6):895–930, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009448;   http://www.springerlink.com/content/xq15g34856h83h81/

.

Mandelbrot:1984:SSS

[Man84c] Benoit B. Mandelbrot. Squig sheets and some other squig fractal constructions. Journal of Statistical Physics, 36(5–6): 519–539, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012918.

Mareschal:1981:TIO

[Mar81] Michel Mareschal. H-theorem for an infinite, one-dimensional, hard-point system. Journal of Statistical Physics, 24(1):139–157, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01007640.

Martin:1985:LES

[Mar85] Olivier Martin. Lyapunov exponents of stochastic dynamical systems. Journal of Statistical Physics, 41(1–2):249–261, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020611.

Martinez:1987:PPA

[Mar87] Servet Martínez. The PMC processes associated to the (1/2, 1/2)- Bernoulli shift are also Bernoulli. Journal of Statistical Physics, 47(3–4):527–541, May 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01007524.

Mascagni:1989:BRB

[Mas89] Michael Mascagni. Book review: Simulated annealing and Boltzmann machines: A stochastic approach to combinatorial optimization and neural computing. Journal of Statistical Physics, 56 (3–4):553–555, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01044453.

Matsumoto:1984:NIO

[Mat84] K. Matsumoto. Noise-induced order II. Journal of Statistical Physics, 34(1–2):111–127, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01770351.

Mayer:1985:SPA

[May85] Dieter H. Mayer. Scaling properties of Z^{k−1} actions on the circle. Journal of Statistical Physics, 38(3–4):785–803, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010490.

Mazo:1981:BMF

[Maz81] Robert M. Mazo. On the Brownian motion of a frequencymodulated oscillator. Journal of Statistical Physics, 24(1):39–44, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01007633.

Mazur:1981:EIC

[MB81] P. Mazur and D. Bedeaux. An electrothermal instability in a conducting wire: Homogeneous and inhomogeneous stationary states for an exactly solvable model. Journal of Statistical Physics, 24 (1):215–233, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01007645.

Montroll:1984:LSD

[MB84] Elliott W. Montroll and John T. Bendler. On Lévy (or stable) distributions and the Williams–Watts model of dielectric relaxation. Journal of Statistical Physics, 34(1–2):129–162, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01770352.

Munoz:1989:AGF

[MB89] Gerardo Muñoz and William S. Burgett. Auxiliary ghost fields in statistical dynamics. Journal of Statistical Physics, 56(1–2):59– 68, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01044231.

Machta:1987:BR

[MBAWP87] Jonathan Machta, Daniel Ben-Avraham, Frederik W. Wiegel, and Adrian Parsegian. Book reviews. Journal of Statistical Physics, 46 (1–2):425–432, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010352.

Mohanty:1982:BLR

[MBG82] Udayan Mohanty, Biman Bagchi, and Julian H. Gibbs. Bimodality and long-range order in ideal Bose systems. Journal of Statistical Physics, 28(4):685–710, August 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011876.

Muthukumar:1981:CDD

[MC81] M. Muthukumar and R. I. Cukier. Concentration dependence of diffusion-controlled processes among stationary reactive sinks. Journal of Statistical Physics, 26(3):453–469, November 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011428.

Marchetti:1985:ISD

[MCDK85] M. C. Marchetti, E. G. D. Cohen, J. R. Dorfman, and T. R. Kirkpatrick. Initial state dependence of nonlinear kinetic equations: The classical electron gas. Journal of Statistical Physics, 41 (1–2):75–93, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020604.

McKean:1987:GKT

[McK87] H. P. McKean. Geometry of KDV (2): Three examples. Journal of Statistical Physics, 46(5–6):1115–1143, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011159.

McLaughlin:1981:PDB

[McL81] John B. McLaughlin. Period-doubling bifurcations and chaotic motion for a parametrically forced pendulum. Journal of Statistical Physics, 24(2):375–388, February 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013307.

McLennan:1982:KEQ

[McL82a] James A. McLennan. Kinetic equations for a quantum gas with bound states. Journal of Statistical Physics, 28(3):521–537, July 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008322.

McLennan:1982:TCO

[McL82b] James A. McLennan. Triple-collision operator for a quantum gas with bound states. Journal of Statistical Physics, 28(2):257–290, June 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012606.

McLennan:1989:BED

[McL89] James A. McLennan. Boltzmann equation for a dissociating gas. Journal of Statistical Physics, 57(3–4):887–905, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01022839.

Marchetti:1983:TPF

[MD83] M. Cristina Marchetti and James W. Dufty. Tagged particle fluctuations in uniform shear flow. Journal of Statistical Physics, 32 (2):255–277, August 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012710.

Maasjost:1982:SPA

[ME82] Winfried Maasjost and Klaus ElsÄsser. Stochastic particle acceleration. Journal of Statistical Physics, 28(4):793–814, August 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011881.

Mandel:1987:SPT

[ME87] Paul Mandel and Thomas Erneux. The slow passage through a steady bifurcation: Delay and memory effects. Journal of Statistical Physics, 48(5–6):1059–1070, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009533.

Meirovitch:1983:MCS

[Mei83] Hagai Meirovitch. A Monte Carlo study of the entropy, the pressure, and the critical behavior of the hard-square lattice gas. Journal of Statistical Physics, 30(3):681–698, March 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009683.

Meijer:1988:ICL

[Mei88a] Paul H. E. Meijer. Influence of the chain length of long molecules on the equation of state in binary gas-liquid mixtures. Journal of Statistical Physics, 53(1–2):543–548, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011572.

Meijer:1988:P

[Mei88b] Paul H. E. Meijer. Preface. Journal of Statistical Physics, 53 (1–2):3–4, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011540.

Meunier:1984:CTI

[Meu84] C. Meunier. Continuity of type-I intermittency from a measuretheoretical point of view. Journal of Statistical Physics, 36(3– 4):321–365, August 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010987.

Meunier:1985:NCW

[Meu85] C. Meunier. Nonlinear coupling of waves in a plasma in a strong dissipation limit. Journal of Statistical Physics, 40(5–6): 759–782, September 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009899.

Machta:1984:LTT

[MEvBD84] J. Machta, M. H. Ernst, H. van Beijeren, and J. R. Dorfman. Long time tails in stationary random media II: Applications. Journal of Statistical Physics, 35(3–4):413–442, May 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014394.

Mezincescu:1987:LSP

[Mez87] G. A. Mezincescu. Lifschitz singularities for periodic operators plus random potentials. Journal of Statistical Physics, 49(5–6): 1181–1190, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01017565.

Milton:1983:CFD

[MF83] G. W. Milton and Michael E. Fisher. Continuum fluids with a discontinuity in the pressure. Journal of Statistical Physics, 32 (2):413–438, August 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012719.

Martin:1983:NPS

[MG83] Ph. A. Martin and Ch. Gruber. A new proof of the Stillinger– Lovett complete shielding condition. Journal of Statistical Physics, 31(3):691–710, June 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01019506.

Madden:1988:DFF

[MG88] William G. Madden and Eduardo D. Glandt. Distribution functions for fluids in random media. Journal of Statistical Physics, 51(3–4):537–558, May 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01028471.

Moseley:1989:DMC

[MGJ89] L. L. Moseley, P. W. Gibbs, and Naeem Jan. Dynamic Monte Carlo renormalization group. II. Journal of Statistical Physics, 57 (1–2):421–427, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023656.

Masoliver:1987:GDI

[MGL87] J. Masoliver, L. Garrido, and J. Llosa. Geometrical derivation of the intrinsic Fokker–Planck equation and its stationary distribution. Journal of Statistical Physics, 46(1–2):233–248, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010343.

Movaghar:1983:THM

[MGP^{+}83] B. Movaghar, M. Grünewald, B. Pohlmann, D. Würtz, and W. Schirmacher. Theory of hopping and multiple-trapping transport in disordered systems. Journal of Statistical Physics, 30 (2):315–334, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012306.

Mansour:1988:SFS

[MGTM88] M. Malek Mansour, A. L. Garcia, J. W. Turner, and M. Mareschal. On the scattering function of simple fluids in finite systems. Journal of Statistical Physics, 52(1–2):295–309, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01016416.

Monta:1981:HOS

[MH81] T. Monta and T. Horiguchi. Higher-order susceptibilities of the regular and the random Ising model on the Cayley tree. I. Journal of Statistical Physics, 26(4):665–681, December 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010932.

Milchev:1986:FSS

[MHB86] A. Milchev, D. W. Heermann, and K. Binder. Finite-size scaling analysis of the Φ^{4} field theory on the square lattice. Journal of Statistical Physics, 44(5–6):749–784, September 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011906.

Moran:1987:DPL

[MHB87] Bill Moran, William G. Hoover, and Stronzo Bestiale. Diffusion in a periodic Lorentz gas. Journal of Statistical Physics, 48(3– 4):709–726, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019693.

Meunier:1989:IPT

[MHV89] C. Meunier, D. Hansel, and A. Verga. Information processing in three-state neural networks. Journal of Statistical Physics, 55 (5–6):859–901, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01041070.

Miekisz:1989:FCI

[Mie89] Jacek Miekisz. Frustration without competing interactions. Journal of Statistical Physics, 55(1–2):351–355, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01042605.

Morriss:1986:SFT

[MIH86] Gary P. Morriss, Dennis J. Isbister, and Barry D. Hughes. Shear flow in the two-body Boltzmann gas. II. Small and large γ expansion of the shear viscosity. Journal of Statistical Physics, 44 (1–2):107–128, July 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010907.

Missarov:1985:EWR

[Mis85] M. D. Missarov. The equations of Wilson’s renormalization group in dimension 4 and analytic renormalization. Journal of Statistical Physics, 38(5–6):851–860, March 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010418.

Misra:1987:FKF

[Mis87] B. Misra. Fields as Kolmogorov flows. Journal of Statistical Physics, 48(5–6):1295–1320, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009547.

Miyazawa:1983:EFP

[MJ83] Sanzo Miyazawa and Robert L. Jernigan. Equilibrium folding pathways for model proteins. Journal of Statistical Physics, 30 (2):549–559, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012328.

Mon:1985:STU

[MJ85] K. K. Mon and David Jasnow. Surface tension and universality in the three-dimensional Ising model. Journal of Statistical Physics, 41(1–2):273–280, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020613.

Masters:1983:ERR

[MK83a] A. J. Masters and T. Keyes. An Enskog repeated-ring kinetic equation; long-time tails and the Brownian limit. Journal of Statistical Physics, 33(1):149–180, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009754.

Mercer:1983:EKH

[MK83b] John Mercer and T. Keyes. Expansion of the kinetic hierarchy for a massive particle: The repeated ring and Fokker–Planck equations. Journal of Statistical Physics, 32(1):35–51, July 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009418.

Masters:1984:LTB

[MK84] A. Masters and T. Keyes. The long time and Brownian limits for tagged particle motion in liquids. Journal of Statistical Physics, 36(3–4):401–433, August 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010989.

Marchetti:1985:KTL

[MK85a] M. C. Marchetti and T. R. Kirkpatrick. Kinetic theory of long time tails in velocity correlation functions in a moderately dense electron gas. Journal of Statistical Physics, 41(3–4):621–660, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009025.

Masters:1985:TRD

[MK85b] A. J. Masters and T. Keyes. Translational and rotational diffusion of an anisotropic particle in a molecular liquid: Long-time tails and Brownian limit. Journal of Statistical Physics, 39(1–2):215– 239, April 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01007980.

Mareschal:1987:OFN

[MK87] M. Mareschal and E. Kestemont. Order and fluctuations in nonequilibrium molecular dynamics simulations of twodimensional fluids. Journal of Statistical Physics, 48(5–6):1187– 1201, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009540.

Mazur:1989:ALS

[MK89] Peter Mazur and Joel Keizer. Anomalous light scattering at the surface of growing ice crystals and thermodynamics of irreversible processes. Journal of Statistical Physics, 57(3–4):

497–509, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01022819.

Meijer:1986:SMD

[MKB86] Paul H. E. Meijer, Mustafa Keskin, and Erik Bodegom. A simple model for the dynamics towards metastable states. Journal of Statistical Physics, 45(1–2):215–232, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01033088.

Marchetti:1987:HTEa

[MKD87a] M. C. Marchetti, T. R. Kirkpatrick, and J. R. Dorfman. Hydrodynamic theory of electron transport in a strong magnetic field. Journal of Statistical Physics, 46(3–4):679–708, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013380.

Marchetti:1987:HTEb

[MKD87b] M. C. Marchetti, T. R. Kirkpatrick, and J. R. Dorfman. Hydrodynamic theory of electron transport in a strong magnetic field. Journal of Statistical Physics, 49(3–4):871–872, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009366.

Marchetti:1985:KEW

[MKDC85] M. C. Marchetti, T. R. Kirkpatrick, J. R. Dorfman, and E. G. D. Cohen. Kinetic equation for a weakly interacting classical electron gas. Journal of Statistical Physics, 41(1–2):37–74, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020603.

Mayer-Kress:1981:INL

[MKH81] G. Mayer-Kress and H. Haken. The influence of noise on the logistic model. Journal of Statistical Physics, 26(1):149–171, September 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01106791.

Mayer-Kress:1989:SCN

[MKK89] Gottfried Mayer-Kress and Kunihiko Kaneko. Spatiotemporal chaos and noise. Journal of Statistical Physics, 54(5–6): 1489–1508, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01044730.

Marro:1985:NPT

[MLSK85] J. Marro, J. L. Lebowitz, H. Spohn, and M. H. Kalos. Nonequilibrium phase transition in stochastic lattice gases: Simulation of a three-dimensional system. Journal of Statistical Physics, 38(3– 4):725–733, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010487.

Masoliver:1983:KFO

[MM83] J. Masoliver and J. Marro. Kinetics of a finite one-dimensional mixture of hard rods with different masses. Journal of Statistical Physics, 31(3):565–575, June 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01019498.

Martinelli:1984:RMA

[MM84] F. Martinelli and D. Merlini. A refined Mermin argument for the two-dimensional jellium. Journal of Statistical Physics, 34 (1–2):313–318, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01770361.

Marchand:1986:EPD

[MM86] J.-P. Marchand and Ph. A. Martin. Exclusion process and droplet shape. Journal of Statistical Physics, 44(3–4):491–504, August 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011307.

Martinelli:1987:LCC

[MM87] F. Martinelli and L. Micheli. On the large-coupling-constant behavior of the Liapunov exponent in a binary alloy. Journal of Statistical Physics, 48(1–2):1–18, July 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010397.

Marchand:1988:EPD

[MM88] J.-P. Marchand and Ph. A. Martin. Exclusion process and droplet shape. Journal of Statistical Physics, 50(1–2):469–471, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023007.

Mao:1987:CSP

[mMH87] Jian min Mao and Bambi Hu. Corrections to scaling for period doubling. Journal of Statistical Physics, 46(1–2):111–117, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010334.

Messager:1984:CFC

[MMSP84] A. Messager, S. Miracle-Sole, and C. E. Pfister. On classical ferromagnets with a complex external field. Journal of Statistical Physics, 34(1–2):279–286, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01770359.

Malyshev:1984:UGF

[MN84] V. A. Malyshev and I. V. Nickolaev. Uniqueness of Gibbs fields via cluster expansions. Journal of Statistical Physics, 35 (3–4):375–379, May 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014390.

Malyshev:1985:TDL

[MNT85] V. A. Malyshev, I. V. Nickolaev, and Yu. A. Terlecky. Temperature dynamics of the locally perturbed classical ideal gas. Journal of Statistical Physics, 40(1–2):133–146, July 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010529.

Mazenko:1982:ARS

[MO82] Gene F. Mazenko and Enis Oḡuz. Application of the real space dynamic renormalization group method to the one-dimensional spin-exchange model. Journal of Statistical Physics, 28(3):497– 520, July 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01008321.

Marchioro:1983:TEI

[MO83] C. Marchioro and E. Omerti. Time evolution of an infinite number of vortices in a strip. Journal of Statistical Physics, 33(1):133–148, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009753.

Monroe:1980:NMI

[Mon80] James L. Monroe. Nonconcavity of the magnetization in Ising ferromagnets. Journal of Statistical Physics, 23(3):315–319, September 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011371.

Monroe:1983:ZPF

[Mon83] James L. Monroe. Zeros of the partition function using theorems of Ruelle. Journal of Statistical Physics, 33(1):77–89, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009749.

Monroe:1985:UBC

[Mon85] James L. Monroe. Upper bounds on the critical temperature for various Ising models. Journal of Statistical Physics, 40(1–2):249– 257, July 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010536.

Monroe:1988:RHS

[Mon88] James L. Monroe. Results from the Holsztynski–Slawny reduction method for ferromagnetic Ising models. Journal of Statistical Physics, 51(1–2):195–203, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01015326.

Morita:1984:CRM

[Mor84] T. Morita. Consistent relations in the method of reducibility in the cluster variation method. Journal of Statistical Physics, 34 (1–2):319–328, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01770362.

Martinelli:1989:SRP

[MOS89] Fabio Martinelli, Enzo Olivieri, and Elisabetta Scoppola. Small random perturbations of finite- and infinite-dimensional dynamical systems: Unpredictability of exit times. Journal of Statistical Physics, 55(3–4):477–504, May 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01041595.

Moussa:1982:SCS

[Mou82] P. Moussa. Subdominant critical singularities in the bcc Ising model. Journal of Statistical Physics, 27(4):711–719, April 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013443.

Marchioro:1982:TEI

[MP82] Carlo Marchioro and Mario Pulvirenti. Time evolution of infinite one-dimensional Coulomb system. Journal of Statistical Physics, 27(4):809–822, April 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013449.

Marchetti:1989:KTP

[MP89] D. H. U. Marchetti and J. Fernando Perez. The Kosterlitz– Thouless phase transition in two-dimensional hierarchical Coulomb gases. Journal of Statistical Physics, 55(1–2):141–156, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01042594.

Muller:1987:DSC

[MPH87] Stefan C. Müller, Theo Plesser, and Benno Hess. Distinctive sites in chemical waves: The spiral core and the collision area of two annuli. Journal of Statistical Physics, 48(5–6):991–1004, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009528.

Mello:1988:SRM

[MPK88] Pier A. Mello, Pedro Pereyra, and Narendra Kumar. A soluble random-matrix model for relaxation in quantum systems. Journal of Statistical Physics, 51(1–2):77–94, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01015321.

Marchioro:1978:VPI

[MPPT78] C. Marchioro, A. Pellegrinotti, M. Pulvirenti, and L. Triolo. Velocity of a perturbation in infinite lattice systems. Journal of Statistical Physics, 19(5):499–510, November 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01011695. See erratum [MPPT80].

Marchioro:1980:EVP

[MPPT80] C. Marchioro, A. Pellegrinotti, M. Pulvirenti, and L. Triolo. Erratum: Velocity of a perturbation in infinite lattice systems. Journal of Statistical Physics, 22(6):743, June 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01013939. See [MPPT78].

Mayer:1983:SAA

[MR83] D. Mayer and G. Roepstorff. Strange attractors and asymptotic measures of discrete-time dissipative systems. Journal of Statistical Physics, 31(2):309–326, May 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011584.

Machta:1986:DCR

[MR86] Jonathan Machta and Bruce Reinhold. Decay of correlations in the regular Lorentz gas. Journal of Statistical Physics, 42(5– 6):949–959, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010456.

Mayer:1987:RTG

[MR87] D. Mayer and G. Roepstorff. On the relaxation time of Gauss’s continued-fraction map I. The Hilbert space approach (Koopmanism). Journal of Statistical Physics, 47(1–2):149–171, April 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009039.

Mayer:1988:RTG

[MR88a] D. Mayer and G. Roepstorff. On the relaxation time of Gauss’ continued-fraction map. II. The Banach space approach (transfer operator method). Journal of Statistical Physics, 50(1–2): 331–344, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01022997.

Miller:1988:SMB

[MR88b] B. N. Miller and K. Ravishankar. Stochastic modeling of a billiard in a gravitational field: Power law behavior of Lyapunov exponents. Journal of Statistical Physics, 53(5–6):1299–1314, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023870.

Miller:1981:DMR

[MS81a] Bruce N. Miller and William E. Stein. Diffusion modeling of the Rayleigh piston. Journal of Statistical Physics, 26(3):539–553, November 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011434.

Morriss:1981:CFD

[MS81b] G. P. Morriss and E. R. Smith. Correlation functions for diatomic symmetric molecules from the RISM equation. Journal of Statistical Physics, 24(4):607–616, April 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011378.

Martina:1982:SMP

[MS82a] Esteban Martina and G. Stell. System of magnetic particles in an external magnetic field: Surface and bulk structure. Journal of Statistical Physics, 27(2):407–417, February 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008946.

Messer:1982:SMI

[MS82b] Joachim Messer and Herbert Spohn. Statistical mechanics of the isothermal Lane–Emden equation. Journal of Statistical Physics, 29(3):561–578, November 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01342187.

Meyer:1983:CGP

[MS83a] J. Meyer and J. Schröter. Comments on the grad procedure for the Fokker–Planck equation. Journal of Statistical Physics, 32(1):53– 69, July 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009419.

Montroll:1983:MEF

[MS83b] Elliott W. Montroll and Michael F. Shlesinger. Maximum entropy formalism, fractals, scaling phenomena, and 1/f noise: A tale of tails. Journal of Statistical Physics, 32(2):209–230, August 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012708.

Mittag:1984:YLZ

[MS84] Laurence Mittag and Michael J. Stephen. Yang–Lee zeros of the Potts model. Journal of Statistical Physics, 35(3–4):303–320, May 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014386.

Muller:1986:IAF

[MS86a] V. F. Müller and J. Schiemann. Infrared asymptotic freedom of a hierarchical ϕ_{3}^{6} lattice theory. Journal of Statistical Physics, 43 (1–2):123–142, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010575.

Murthy:1986:TMC

[MS86b] Ganpathy Murthy and R. Shankar. Tests for Monte Carlo renormalization studies on Ising models. Journal of Statistical Physics, 42(3–4):275–288, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01127713.

Madras:1987:NLL

[MS87] Neal Madras and Alan D. Sokal. Nonergodicity of local, length-conserving Monte Carlo algorithms for the self-avoiding walk. Journal of Statistical Physics, 47(3–4):573–595, May 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01007527.

Madras:1988:PAH

[MS88a] Neal Madras and Alan D. Sokal. The pivot algorithm: A highly efficient Monte Carlo method for the self-avoiding walk. Journal of Statistical Physics, 50(1–2):109–186, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01022990.

Martinelli:1988:RCT

[MS88b] F. Martinelli and E. Scoppola. Random composition of two rational maps: Singularity of the invariant measure. Journal of Statistical Physics, 50(5–6):1021–1042, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019151.

Matkowsky:1984:UET

[MST84] B. J. Matkowsky, Z. Schuss, and C. Tier. Uniform expansion of the transition rate in Kramers’ problem. Journal of Statistical Physics, 35(3–4):443–456, May 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014395.

Movaghar:1982:TDE

[MSW82] B. Movaghar, G. W. Sauer, and D. Würtz. Time decay of excitations in the one-dimensional trapping problem. Journal of Statistical Physics, 27(3):473–485, March 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011087.

McGuire:1982:APS

[MT82] James B. McGuire and Colin J. Thompson. Asymptotic properties of sequences of iterates of nonlinear transformations. Journal of Statistical Physics, 27(1):183–200, January 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011746.

Matsumoto:1983:NIOa

[MT83a] K. Matsumoto and I. Tsuda. Noise-induced order. Journal of Statistical Physics, 31(1):87–106, April 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010923.

Matsumoto:1983:NIOb

[MT83b] K. Matsumoto and I. Tsuda. Noise-induced order. Journal of Statistical Physics, 33(3):757, December 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01018842.

Martinez:1984:DCM

[MT84] S. Martinez and E. Tirapegui. Description of a class of Markov processes ‘equivalent’ to K-shifts. Journal of Statistical Physics, 37(1–2):173–186, October 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012910.

Martinez:1987:RSM

[MT87] S. Martinez and E. Tirapegui. The role of stable manifolds and information in the Prigogine theory of irreversibility. Journal of Statistical Physics, 48(5–6):1283–1294, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009546.

Mariz:1985:PDI

[MTA85] Ananias M. Mariz, Constantino Tsallis, and E. L. Albuquerque. Phase diagram of the Ising model on a Cayley tree in the presence of competing interactions and magnetic field. Journal of Statistical Physics, 40(3–4):577–592, August 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01017186.

Mansour:1987:CFS

[MTG87] M. Malek Mansour, John W. Turner, and Alejandro L. Garcia. Correlation functions for simple fluids in a finite system under nonequilibrium constraints. Journal of Statistical Physics, 48(5–6):1157–1186, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009539.

Mukamel:1982:ELR

[Muk82] Shaul Mukamel. On the expansion of linear response functionsapplication to electrical conductivity of disordered metals. Journal of Statistical Physics, 27(2):317–338, February 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008941.

Mukamel:1983:DED

[Muk83] Shaul Mukamel. On the dynamics of excitations in disordered systems. Journal of Statistical Physics, 30(1):179–184, January 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010873.

Murmann:1987:HLO

[Mür87] Michael G. Mürmann. The hydrodynamic limit of a onedimensional nearest neighbor gradient system. Journal of Statistical Physics, 48(3–4):769–788, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019696.

Muthukumar:1983:MCR

[Mut83] M. Muthukumar. Monte Carlo renormalization group calculations for polymers. Journal of Statistical Physics, 30(2):457–465, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012319.

Murthy:1982:ARO

[MV82] K. P. N. Murthy and M. C. Valsakumar. Asymptotic results on occupancy times of random walks. Journal of Statistical Physics, 29(4):773–779, December 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011790.

Mauricio:1986:SEI

[MV86] F. Mauricio and S. Velasco. Stochastic effects on the isothermal explosion arising in the oxidation of iron (II) by nitric acid. Journal of Statistical Physics, 43(3–4):521–536, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020651.

Marro:1987:NDP

[MV87] J. Marro and J. L. Vallés. Nonequilibriurn discontinuous phase transitions in a fast ionic conductor model: Coexistence and spinodal lines. Journal of Statistical Physics, 49(1–2):121–137, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009957.

Maes:1988:DVC

[MV88a] D. Maes and C. Van den Broeck. Discrete versus continuous-time random walks. Journal of Statistical Physics, 50(5–6):1089–1102, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019154.

Marro:1988:NDP

[MV88b] J. Marro and J. L. Vallés. Nonequilibrium discontinuous phase transitions in a fast ionic conductor model: Coexistence and spinodal lines. Journal of Statistical Physics, 51(1–2):323, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01015336.

Meunier:1988:NB

[MV88c] C. Meunier and A. D. Verga. Noise and bifurcations. Journal of Statistical Physics, 50(1–2):345–375, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01022998.

Masoliver:1988:SFP

[MW88] Jaume Masoliver and George H. Weiss. Some first passage time problems for shot noise processes. Journal of Statistical Physics, 50(1–2):377–382, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01022999.

Martin:1980:CFC

[MY80] Ph. A. Martin and T. Yalcin. The charge fluctuations in classical Coulomb systems. Journal of Statistical Physics, 22(4):435–463, April 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012866.

Mikhailov:1985:QFM

[MY85] A. S. Mikhailov and V. V. Yashin. Quantum field methods in the theory of diffusion-controlled reactions. Journal of Statistical Physics, 38(1–2):347–359, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01017866.

Minlos:1989:ADC

[MZ89] R. A. Minlos and E. A. Zhizhina. Asymptotics of decay of correlations in the ANNNI model at high temperatures. Journal of Statistical Physics, 56(5–6):957–963, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01016789.

Nagle:1985:PGP

[Nag85] John F. Nagle. Proof of gridlock in a polymer model. Journal of Statistical Physics, 38(3–4):531–540, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010476.

Nakagomi:1981:MTN

[Nak81] Teruaki Nakagomi. Mesoscopic thermodynamics of nonequilibrium open systems. I. Negentropy consumption and residual entropy. Journal of Statistical Physics, 26(3):567–611, November 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011436.

Nakazawa:1986:SDE

[Nak86] Hiroshi Nakazawa. Stochastic differential equations in quantum statistical mechanics. Observables and multiple Wiener integrals. Journal of Statistical Physics, 45(5–6):1049–1069, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020588.

Narnhofer:1982:LCD

[Nar82] Heide Narnhofer. Local coexistence of different phases. Journal of Statistical Physics, 29(4):761–771, December 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011789.

Nauenberg:1985:RWR

[Nau85] Michael Nauenberg. Random walk in a random medium in one dimension. Journal of Statistical Physics, 41(5–6):803–810, December 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010004.

Nauenberg:1987:FBD

[Nau87] Michael Nauenberg. Fractal boundary of domain of analyticity of the Feigenbaum function and relation to the Mandelbrot set. Journal of Statistical Physics, 47(3–4):459–475, May 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01007520.

Nicolis:1987:IRS

[NB87] G. Nicolis and F. Baras. Intrinsic randomness and spontaneous symmetry-breaking in explosive systems. Journal of Statistical Physics, 48(5–6):1071–1090, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009534.

Novick-Cohen:1985:NCH

[NC85] Amy Novick-Cohen. The nonlinear Cahn–Hilliard equation: Transition from spinodal decomposition to nucleation behavior. Journal of Statistical Physics, 38(3–4):707–723, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010486.

Nemirovsky:1988:DDS

[NCF88] Adolfo M. Nemirovsky and Maurício D. Coutinho-Filho. From dilute to dense self-avoiding walks on hypercubic lattices. Journal of Statistical Physics, 53(5–6):1139–1153, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01023861.

Nieuwenhuizen:1985:ENH

[NE85] Th. M. Nieuwenhuizen and M. H. Ernst. Excess noise in a hopping model for a resistor with quenched disorder. Journal of Statistical Physics, 41(5–6):773–801, December 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010003.

Neaderhouser:1980:CBS

[Nea80] Carla C. Neaderhouser. Convergence of block spins defined by a random field. Journal of Statistical Physics, 22(6):673–684, June 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013936.

Negele:1986:MCS

[Neg86] J. W. Negele. Monte Carlo studies of nuclear many-particle systems. Journal of Statistical Physics, 43(5–6):991–1015, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02628326.

Nelkin:1989:WDW

[Nel89a] Mark Nelkin. What do we know about self-similarity in fluid turbulence? Journal of Statistical Physics, 54(1–2):1–15, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023471.

Nelson:1989:SMF

[Nel89b] David R. Nelson. Statistical mechanics of flux lines in high-T_{c} superconductors. Journal of Statistical Physics, 57(3–4):511–530, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01022820.

Newman:1986:SCE

[New86] C. M. Newman. Some critical exponent inequalities for percolation. Journal of Statistical Physics, 45(3–4):359–368, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01021076.

Newman:1987:ACE

[New87] Charles M. Newman. Another critical exponent inequality for percolation: β ≥ 2/δ. Journal of Statistical Physics, 47(5–6):695– 699, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01206153.

Noskowicz:1987:DFR

[NG87] S. H. Noskowicz and I. Goldhirsch. Distribution functions for random walk processes on networks: An analytic method. Journal of Statistical Physics, 48(1–2):255–290, July 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010409.

Nordholm:1982:GVW

[NGH82] Sture Nordholm, James Gibson, and Malcolm A. Hooper. Generalized van der Waals theory VIII. An improved analysis of the liquid/gas interface. Journal of Statistical Physics, 28(2):391–406, June 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012613.

Nguyen:1987:CLC

[Ngu87a] Bao Gia Nguyen. Correlation length and its critical exponent for percolation processes. Journal of Statistical Physics, 46(3– 4):517–523, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013372.

Nguyen:1987:GEP

[Ngu87b] Bao Gia Nguyen. Gap exponents for percolation processes with triangle condition. Journal of Statistical Physics, 49(1– 2):235–243, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009960.

Nguyen:1988:TCS

[Ngu88] Bao Gia Nguyen. Typical cluster size for two-dimensional percolation processes. Journal of Statistical Physics, 50(3–4):715–726, February 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026497.

Neumann:1988:DMS

[NH88] Avidan U. Neumann and Shlomo Havlin. Distributions and moments of structural properties for percolation clusters. Journal of Statistical Physics, 52(1–2):203–236, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01016410.

Normand:1988:PCD

[NHH88] J.-M. Normand, H. J. Herrmann, and M. Hajjar. Precise calculation of the dynamical exponent of two-dimensional percolation. Journal of Statistical Physics, 52(1–2):441–446, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016424.

Nienhuis:1984:CBT

[Nie84] Bernard Nienhuis. Critical behavior of two-dimensional spin models and charge asymmetry in the Coulomb gas. Journal of Statistical Physics, 34(5–6):731–761, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009437.

Nienhuis:1988:PER

[Nie88] G. Nienhuis. Photon emission as a random event. Journal of Statistical Physics, 53(1–2):417–438, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011564.

Nishimori:1981:SQN

[Nis81] Hidetoshi Nishimori. Spin quantum number in the ground state of the Mattis–Heisenberg model. Journal of Statistical Physics, 26 (4):839–845, December 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010945.

Nieuwoudt:1984:LRB

[NKD84] J. C. Nieuwoudt, T. R. Kirkpatrick, and J. R. Dorfman. Longrange boundary effects in simple fluids. Journal of Statistical Physics, 34(1–2):203–223, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01770355.

Nishimori:1989:LRO

[NKO^{+}89] Hidetoshi Nishimori, Kenn Kubo, Yukiyasu Ozeki, Yasuhiro Tomita, and Tatsuya Kishi. Long-range order in the XXZ model. Journal of Statistical Physics, 55(1–2):259–277, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01042599.

Nieuwenhuizen:1985:SBD

[NL85] Th. M. Nieuwenhuizen and J. M. Luck. Singular behavior of the density of states and the Lyapunov coefficient in binary random harmonic chains. Journal of Statistical Physics, 41(5–6):745–771, December 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010002.

Nieuwenhuizen:1987:LSR

[NL87] Th. M. Nieuwenhuizen and J. M. Luck. Lifshitz singularities in random harmonic chains: Periodic amplitudes near the band edge and near special frequencies. Journal of Statistical Physics, 48(3– 4):393–424, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019680.

Nieuwenhuizen:1986:SFL

[NLC^{+}86] Th. M. Nieuwenhuizen, J. M. Luck, J. Canisius, J. L. van Hemmen, and W. J. Ventevogel. Special frequencies and Lifshitz singularities in binary random harmonic chains. Journal of Statistical Physics, 45(3–4):395–417, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01021078.

Nielaba:1989:BQP

[NLSV89] P. Nielaba, J. L. Lebowitz, H. Spohn, and J. L. Vallés. Behavior of a quantum particle in contact with a classical heat bath. Journal of Statistical Physics, 55(3–4):745–767, May 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01041606.

Nieuwoudt:1984:TPD

[NM84] Johan Nieuwoudt and Shaul Mukamel. Transport and percolation in disordered systems — a self-consistent time-local approach. Journal of Statistical Physics, 36(5–6):677–686, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012931.

Nicolis:1980:SAM

[NMM80] G. Nicolis and M. Malek-Mansour. Systematic analysis of the multivariate master equation for a reaction-diffusion system. Journal of Statistical Physics, 22(4):495–512, April 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012869.

Nakashima:1989:SBT

[NNK89] Katsuya Nakashima, Tatsuzo Nagai, and Kyozi Kawasaki. Scaling behavior of two-dimensional domain growth: Computer simulation of vertex models. Journal of Statistical Physics, 57(3–4): 759–787, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01022831.

Nicolis:1989:CDM

[NNN89] G. Nicolis, C. Nicolis, and John S. Nicolis. Chaotic dynamics, Markov partitions, and Zipf’s law. Journal of Statistical Physics, 54(3–4):915–924, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019781.

Nappo:1989:RDM

[NOR89] G. Nappo, E. Orlandi, and H. Rost. A reaction-diffusion model for moderately interacting particles. Journal of Statistical Physics, 55(3–4):579–600, May 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01041598.

Nossal:1983:SAB

[Nos83] Ralph Nossal. Stochastic aspects of biological locomotion. Journal of Statistical Physics, 30(2):391–400, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012313.

Nossal:1984:BRM

[Nos84] R. Nossal. Book review: Measurement of suspended particles by quasi-elastic light scattering. Journal of Statistical Physics, 35 (5–6):765–766, June 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010833.

Nossal:1986:BRBb

[Nos86a] Ralph Nossal. Book review: 1. Chaos. 2. Deterministic Chaos. Journal of Statistical Physics, 44(3–4):707–708, August 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011316.

Nossal:1986:BRBa

[Nos86b] Ralph Nossal. Book review: Dynamic light scattering: Applications of photon correlation spectoscopy. Journal of Statistical Physics, 44(1–2):273–274, July 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010917.

Nicolo:1982:FOP

[NR82] Francesco Nicoló and Charles Radin. A first-order phase transition between crystal phases in the shift model. Journal of Statistical Physics, 28(3):473–478, July 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008318.

Nijboer:1988:EPP

[NR88] B. R. A. Nijboer and Th. W. Ruijgrok. On the energy per particle in three- and two-dimensional Wigner lattices. Journal of Statistical Physics, 53(1–2):361–382, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011562.

Newman:1980:CFE

[NS80] C. M. Newman and L. S. Schulman. Complex free energies and metastable lifetimes. Journal of Statistical Physics, 23(2): 131–148, August 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012588.

Newman:1981:ICP

[NS81] C. M. Newman and L. S. Schulman. Infinite clusters in percolation models. Journal of Statistical Physics, 26(3):613–628, November 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011437.

Nelkin:1981:DVF

[NT81] Mark Nelkin and A.-M. S. Tremblay. Deviation of 1/f voltage fluctuations from scale-similar Gaussian behavior. Journal of Statistical Physics, 25(2):253–268, June 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01022186.

Narnhofer:1989:MPQ

[NT89] H. Narnhofer and W. Thirring. Mixing properties of quantum systems. Journal of Statistical Physics, 57(3–4):811–825, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01022834.

Narnhofer:1988:TEQ

[NTW88] H. Narnhofer, W. Thirring, and H. Wiklicky. Transitivity and ergodicity of quantum systems. Journal of Statistical Physics, 52(3– 4):1097–1112, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019741.

Napiorkowski:1986:ATF

[NZ86] M. Napiórkowski and U. Zaus. Average trajectories and fluctuations from noisy, nonlinear maps. Journal of Statistical Physics, 43(1–2):349–368, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010587.

OCarroll:1984:APCb

[OB84] Michael O’Carroll and Wilson Dantas Barbosa. Analyticity properties and a convergent expansion for the inverse correlation length of the low-temperature d-dimensional Ising model. Journal of Statistical Physics, 34(3–4):609–614, February 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01018561.

Owczarek:1987:CIR

[OB87] A. L. Owczarek and R. J. Baxter. A class of interaction-round-a-face models and its equivalence with an ice-type model. Journal of Statistical Physics, 49(5–6):1093–1115, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01017562.

OCarroll:1984:APCa

[O’C84a] Michael O’Carroll. Analyticity properties and a convergent expansion for the inverse correlation length of the high-temperature d-dimensional Ising model. Journal of Statistical Physics, 34(3– 4):597–608, February 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01018560.

OCarroll:1984:CEA

[O’C84b] Michael O’Carroll. Convergent expansions for asymptotically degenerate inverse correlation lengths of classical lattice spin systems. Journal of Statistical Physics, 37(3–4):439–449, November 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011843.

Ojima:1988:EPP

[OHI88] I. Ojima, H. Hasegawa, and M. Ichiyanagi. Entropy production and its positivity in nonlinear response theory of quantum dynamical systems. Journal of Statistical Physics, 50(3–4):633–655, February 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026494.

Oppenheim:1982:BR

[OI82] I. Oppenheim and W. Israel. Book review. Journal of Statistical Physics, 28(1):227–229, May 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011634.

Ojima:1989:EPN

[Oji89] Izumi Ojima. Entropy production and nonequilibrium stationarity in quantum dynamical systems. Physical meaning of van Hove limit. Journal of Statistical Physics, 56(1–2):203–226, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01044241.

Okabe:1986:TBM

[Oka86] Yasunori Okabe. On the theory of Brownian motion with the Alder–Wainwright effect. Journal of Statistical Physics, 45(5–6): 953–981, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020584.

Okninski:1988:CDM

[Okn88] A. Okniński. Chaos in discrete maps, deterministic scattering, and nondifferentiable functions. Journal of Statistical Physics, 52 (3–4):577–594, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019718.

Olivieri:1988:CEL

[Oli88] Enzo Olivieri. On a cluster expansion for lattice spin systems: A finite-size condition for the convergence. Journal of Statistical Physics, 50(5–6):1179–1200, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019160.

Omnes:1988:LRQa

[Omn88a] Roland Omnès. Logical reformulation of quantum mechanics. I. Foundations. Journal of Statistical Physics, 53(3–4):893–932, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014230.

Omnes:1988:LRQb

[Omn88b] Roland Omnès. Logical reformulation of quantum mechanics. II. Interferences and the Einstein–Podolsky–Rosen experiment. Journal of Statistical Physics, 53(3–4):933–955, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014231.

Omnes:1988:LRQc

[Omn88c] Roland Omnès. Logical reformulation of quantum mechanics. III. Classical limit and irreversibility. Journal of Statistical Physics, 53(3–4):957–975, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014232.

Omnes:1989:LRQ

[Omn89] Roland Omnès. Logical reformulation of quantum mechanics. IV. Projectors in semiclassical physics. Journal of Statistical Physics, 57(1–2):357–382, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01023649.

Onogi:1986:CPA

[OMS86] Toshiyuki Onogi, Seiji Miyashita, and Masuo Suzuki. Critical properties of an approximant of two-dimensional quantum XYZ models. Journal of Statistical Physics, 45(1–2):183–200, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01033086.

Oppenheim:1981:BR

[OP81] Irwin Oppenheim and Adrian Parsegian. Book review. Journal of Statistical Physics, 25(4):771–773, August 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01022366.

Orfanopoulos:1987:EEF

[OP87] B. A. Orfanopoulos and J. K. Percus. Effect of an external field on a modification of Luttinger’s many-fermion model. Journal of Statistical Physics, 46(3–4):525–533, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01013373.

Ord:1989:GOD

[OP89] G. N. Ord and J. K. Percus. A generalization of the onedimensional δ-function Bose gas. Journal of Statistical Physics, 56(5–6):681–693, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01016774.

Oppenheim:1980:BRM

[Opp80] Irwin Oppenheim. Book review: The maximum entropy formalism. Journal of Statistical Physics, 23(1):127, July 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014432.

Oppenheim:1982:BRN

[Opp82] Irwin Oppenheim. Book review: Nonequilibrium thermodynamics and its statistical foundation. Journal of Statistical Physics, 27(4):825–826, April 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013451.

Oppenheim:1986:BRM

[Opp86a] Irwin Oppenheim. Book review: Multicritical phenomena. Journal of Statistical Physics, 44(5–6):1005, September 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011920.

Oppenheim:1986:BRW

[Opp86b] Irwin Oppenheim. Book review: The wonderful world of stochastics: A tribute to Elliott w. Montroll. Journal of Statistical Physics, 43(3–4):715–716, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020663.

Oppenheim:1988:BRBa

[Opp88a] I. Oppenheim. Book review: Simple models of equilibrium and nonequilibrium phenomena. Journal of Statistical Physics, 50 (1–2):467, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023006.

Oppenheim:1988:BRBb

[Opp88b] Irwin Oppenheim. Book review: Statistical thermodynamics of non-equilibrium processes. Journal of Statistical Physics, 50 (5–6):1281, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019168.

Oppenheim:1989:BRC

[Opp89] I. Oppenheim. Book review: A commentary on thermodynamics. Journal of Statistical Physics, 54(3–4):1101, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01019791.

Orbach:1984:DFS

[Orb84] R. Orbach. Dynamics of fractal structures. Journal of Statistical Physics, 36(5–6):735–748, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012935.

Omerti:1986:NEM

[ORD86] E. Omerti, M. Ronchetti, and D. Dürr. Numerical evidence for mass dependence in the diffusive behavior of the ‘Heavy Particle’ on the line. Journal of Statistical Physics, 44(3–4):339–346, August 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011302.

Oevel:1981:BEM

[OS81] W. Oevel and J. Schröter. Balance equations for micromorphic materials. Journal of Statistical Physics, 25(4):645–662, August 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01022359.

Oji:1984:MSM

[OS84] Herbert Oji and Henrik Smith. Magnetotransport in simple metals. An exactly soluble model. Journal of Statistical Physics, 34(3–4):381–398, February 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01018551.

Ott:1984:DCA

[OWY84] E. Ott, W. D. Withers, and J. A. Yorke. Is the dimension of chaotic attractors invariant under coordinate changes? Journal of Statistical Physics, 36(5–6):687–697, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012932.

Oono:1987:CDS

[OY87] Y. Oono and C. Yeung. A cell dynamical system model of chemical turbulence. Journal of Statistical Physics, 48(3–4):593–644, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019690.

Plischke:1988:MCS

[PA88] Michael Plischke and Farid F. Abraham. Monte Carlo simulation using the Fourier transform of the interatomic potential. Journal of Statistical Physics, 52(5–6):1353–1358, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011652.

Pfeifer:1984:SBS

[PAF84] Peter Pfeifer, David Avnir, and Dina Farin. Scaling behavior of surface irregularity in the molecular domain: From adsorption studies to fractal catalysts. Journal of Statistical Physics, 36(5– 6):699–716, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012933.

Pfeifer:1985:ESB

[PAF85] Peter Pfeifer, David Avnir, and Dina Farin. Erratum: Scaling behavior of surface irregularity in the molecular domain: From adsorption studies to fractal catalysts. Journal of Statistical Physics, 39(1–2):263, April 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01007983.

Pandey:1984:CSI

[Pan84] R. B. Pandey. Computer simulation of irreversible kinetic gelation: Crossover between random and kinetic growth, radical conservation, poisons, and chain-preferred reactions. Journal of Statistical Physics, 34(1–2):163–190, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01770353.

Pandey:1989:CSC

[Pan89] R. B. Pandey. Computer simulation of a cellular automata model for the immune response in a retrovirus system. Journal of Statistical Physics, 54(3–4):997–1010, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019785.

Papanicolaou:1983:BRR

[Pap83] George Papanicolaou. Book review: Random fields. Journal of Statistical Physics, 33(1):231, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009758.

Parisi:1980:FTA

[Par80] Giorgio Parisi. Field-theoretic approach to second-order phase transitions in two- and three-dimensional systems. Journal of Statistical Physics, 23(1):49–82, July 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01014429.

Park:1982:CEC

[Par82] Yong Moon Park. The cluster expansion for classical and quantum lattice systems. Journal of Statistical Physics, 27(3):553–576, March 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011092.

Park:1985:QSM

[Par85] Yong Moon Park. Quantum statistical mechanics for superstable interactions: Bose–Einstein statistics. Journal of Statistical Physics, 40(1–2):259–302, July 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010537.

Park:1987:EUG

[Par87a] Yong Moon Park. Existence and uniqueness of Gibbs states for a statistical mechanical polyacetylene model. Journal of Statistical Physics, 46(3–4):567–598, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013376.

Parsegian:1987:BRF

[Par87b] V. Adrian Parsegian. Book review: Fluid interfacial phenomena. Journal of Statistical Physics, 49(1–2):393–394, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009971.

Park:1989:DEI

[Par89] Yong Moon Park. Direct estimates on intersection probabilities of random walks. Journal of Statistical Physics, 57(1–2):319–331, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023646.

Pastur:1982:DSM

[Pas82] L. A. Pastur. Disordered spherical model. Journal of Statistical Physics, 27(1):119–151, January 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011743. See erra-

tum [Pas83].

Pastur:1983:EDS

[Pas83] L. A. Pastur. Erratum: Disordered spherical model. Journal of Statistical Physics, 33(2):485, November 1983. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01009807. See [Pas82].

Penrose:1983:KNL

[PB83] O. Penrose and A. Buhagiar. Kinetics of nucleation in a lattice gas model: Microscopic theory and simulation compared. Journal of Statistical Physics, 30(1):219–241, January 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010876.

Phythian:1986:VAS

[PC86] R. Phythian and W. D. Curtis. A variational approach to stochastic nonlinear problems. Journal of Statistical Physics, 42(5–6): 1019–1046, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010460.

Phythian:1987:DSB

[PC87] R. Phythian and W. D. Curtis. Diffusion in a symmetric bistable potential: A variational approach. Journal of Statistical Physics, 46(1–2):217–231, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010342.

Prigogine:1981:PR

[PD81] I. Prigogine and G. Dewel. Pierre Résibois: 1936–1979. Journal of Statistical Physics, 24(1):7–19, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01007631.

Pearce:1981:MFB

[Pea81] Paul A. Pearce. Mean-field bounds on the magnetization for ferromagnetic spin models. Journal of Statistical Physics, 25(2):309– 320, June 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01022189.

Pearle:1985:TIT

[Pea85] Philip Pearle. On the time it takes a state vector to reduce. Journal of Statistical Physics, 41(3–4):719–727, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009030.

Pellegrinotti:1988:EPD

[Pel88] A. Pellegrinotti. Evidence for the Poisson distribution for quasienergies in the quantum kicked-rotator model. Journal of Statistical Physics, 53(5–6):1327–1336, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01023872.

Pellegrinotti:1989:EPD

[Pel89] A. Pellegrinotti. Evidence for the Poisson distribution for quasienergies in the quantum Kicked-Rotator Model. Journal of Statistical Physics, 55(1–2):471–472, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01042614.

Penrose:1986:PTF

[Pen86] O. Penrose. Phase transitions on fractal lattices with long-range interactions. Journal of Statistical Physics, 45(1–2):69–88, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/

10.1007/BF01033078. See corrigendum [Pen87].

Penrose:1987:CPT

[Pen87] O. Penrose. Corrigendum: Phase transitions on fractal lattices with long-range interactions. Journal of Statistical Physics, 48 (3–4):947–950, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019706. See [Pen86].

Percus:1980:SDA

[Per80] J. K. Percus. Shielded distribution approximation for a wallbounded classical fluid. Journal of Statistical Physics, 23(6): 657–699, December 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008515.

Percus:1982:ODC

[Per82] J. K. Percus. One-dimensional classical fluid with nearestneighbor interaction in arbitrary external field. Journal of Statistical Physics, 28(1):67–81, May 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011623.

Percus:1984:BRM

[Per84] J. K. Percus. Book review: Molecular theory of capillarity. Journal of Statistical Physics, 34(3–4):647–649, February 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01018563.

Percus:1986:SSM

[Per86] J. K. Percus. Some solvable models of nonuniform classical fluids. Journal of Statistical Physics, 42(5–6):921–930, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010452.

Percus:1987:SAW

[Per87] J. K. Percus. Some aspects of the wetting transition1. Journal of Statistical Physics, 47(5–6):801–825, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01206159.

Percus:1988:FEM

[Per88] J. K. Percus. Free energy models for nonuniform classical fluids. Journal of Statistical Physics, 52(5–6):1157–1178, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011639.

Percus:1989:CMI

[Per89] J. K. Percus. Collective modes in Ising lattices. Journal of Statistical Physics, 55(5–6):1263–1277, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01041086.

Peschanskii:1985:GSE

[Pes85] V. G. Peschanskii. On galvanomagnetic size effects in metals. Journal of Statistical Physics, 38(1–2):253–266, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017861.

Petrosky:1987:CCS

[Pet87] T. Y. Petrosky. The cometary cloud in the solar system and the Résibois–Prigogine singular invariants of motion. Journal of Statistical Physics, 48(5–6):1363–1372, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009550.

Privman:1983:FSE

[PF83] Vladimir Privman and Michael E. Fisher. Finite-size effects at first-order transitions. Journal of Statistical Physics, 33(2): 385–417, November 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009803.

Palmer:1985:LHD

[PF85] R. G. Palmer and H. L. Frisch. Low-and high-dimension limits of a phase separation model. Journal of Statistical Physics, 38 (5–6):867–872, March 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010420.

Pfister:1982:PTA

[Pfi82] C.-E. Pfister. Phase transitions in the Ashkin–Teller model. Journal of Statistical Physics, 29(1):113–116, September 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008251.

Pfister:1987:PDH

[Pfi87] Charles-Ed. Pfister. Phase diagram of the half-infinite Ising model. Journal of Statistical Physics, 47(5–6):761–771, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01206157.

Paveri-Fontana:1989:NGM

[PFvdMZ89] Stefano L. Paveri-Fontana, C. V. M. van der Mee, and P. F. Zweifel. A neutral gas model for electron swarms. Journal of Statistical Physics, 57(1–2):247–265, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01023642.

Polewczak:1983:SRA

[PG83] Jacek Polewczak and William Greenberg. Some remarks about continuity properties of local Maxwellians and an existence theorem for the BGK model of the Boltzmann equation. Journal of Statistical Physics, 33(2):307–316, November 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009799.

Personnaz:1986:BCL

[PGDT86] L. Personnaz, I. Guyon, G. Dreyfus, and G. Toulouse. A biologically constrained learning mechanism in networks of formal neurons. Journal of Statistical Physics, 43(3–4):411–422, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020645.

Pierre:1987:NRN

[PGS87] L. Pierre, T. Giamarchi, and H. J. Schulz. A new random-number generator for multispin Monte Carlo algorithms. Journal of Statistical Physics, 48(1–2):135–149, July 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010404.

Perez-Hernandez:1981:ESM

[PHB81] W. Pérez-Hernández and L. Blum. Explicit solution of the mean spherical model for ions and dipoles. Journal of Statistical Physics, 24(3):451–454, March 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012816.

Phillies:1982:UUS

[Phi82] George D. J. Phillies. Upon the utility of spherical harmonic expansions in the evaluation of cluster integrals. Journal of Statistical Physics, 28(4):673–683, August 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011875.

Piasecki:1981:REL

[Pia81a] J. Piasecki. The runaway effect in a Lorentz gas. Journal of Statistical Physics, 24(1):45–58, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01007634.

Piasecki:1981:SDF

[Pia81b] J. Piasecki. Self-diffusion in fluids with weak long-range forces. Journal of Statistical Physics, 26(2):375–396, October 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013178.

Piasecki:1983:AFI

[Pia83] J. Piasecki. Approach to field-induced stationary state in a gas of hard rods. Journal of Statistical Physics, 30(1):185–193, January 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010874.

Piasecki:1984:DCT

[Pia84] J. Piasecki. Dynamics of a charged test particle in a hard rod fluid. Journal of Statistical Physics, 35(5–6):635–643, June 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010827.

Piasecki:1987:LTR

[Pia87] J. Piasecki. Local H-theorem for the revised Enskog equation. Journal of Statistical Physics, 48(5–6):1203–1211, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009541.

Picco:1983:ABS

[Pic83] P. Picco. On the absence of breakdown of symmetry for the plane rotator model with long-range random interaction. Journal of Statistical Physics, 32(3):627–648, September 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008960.

Picco:1984:UBD

[Pic84] P. Picco. Upper bound on the decay of correlations in the plane rotator model with long-range random interaction. Journal of Statistical Physics, 36(3–4):489–516, August 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010993.

Papatriantafillou:1989:IDI

[PIP89] C. Papatriantafillou, M. E. Issigoni, and C. E. Paraskevaidis. Influence of disorder on the interface sharpness of an Ising ferromagnetic system. Journal of Statistical Physics, 54(3–4):755–763, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019774.

Piasecki:1987:MOD

[PL87] J. Piasecki and D. Levesque. From monotonic to oscillatory decay of correlations: Analytical approximation for the two-dimensional, one-component plasma. Journal of Statistical Physics, 47(3–4):489–503, May 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01007522.

Platkowski:1986:GET

[Pla86] Tadeusz Platkowski. On a global existence theorem for the Enskog equation. Journal of Statistical Physics, 45(1–2):349–352, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01033095.

Penrose:1981:GCF

[PLM^{+}81] O. Penrose, Joel L. Lebowitz, J. Marro, M. H. Kalos, and A. Sur. Growth of clusters in a first-order phase transition. Journal of Statistical Physics, 24(4):703, April 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011380.

Penrose:1984:KFO

[PLM^{+}84] O. Penrose, J. Lebowitz, J. Marro, M. Kalos, and J. Tobochnik. Kinetics of a first-order phase transition: computer simulations and theory. Journal of Statistical Physics, 34(3–4):399–426, February 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01018552.

Percus:1982:BR

[PM82] Jerome Percus and Robert M. Miura. Book reviews. Journal of Statistical Physics, 28(3):615–618, July 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008330.

Palffy-Muhoray:1984:TRO

[PMBB^{+}84] P. Palffy-Muhoray, R. Barrie, B. Bergersen, I. Carvalho, and M. Freeman. Tunneling resistivity of a one-dimensional random lattice and the Petersburg problem. Journal of Statistical Physics, 35(1–2):119–130, April 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01017369.

Pacault:1987:BOC

[POD87] A. Pacault, Q. Ouyang, and P. De Kepper. Bistable and oscillating chemical reactions. Journal of Statistical Physics, 48(5–6): 1005–1016, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009529.

Poland:1984:UNT

[Pol84] Douglas Poland. On the universality of the nonphase transition singularity in hard-particle systems. Journal of Statistical Physics, 35(3–4):341–353, May 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014388.

Polewczak:1988:CSN

[Pol88a] Jacek Polewczak. Classical solution of the nonlinear Boltzmann equation in all R^{3}: Asymptotic behavior of solutions. Journal of Statistical Physics, 50(3–4):611–632, February 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01026493.

Pollett:1988:PEQ

[Pol88b] P. K. Pollett. On the problem of evaluating quasistationary distributions for open reaction schemes. Journal of Statistical Physics, 53(5–6):1207–1215, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01023864.

Poland:1989:KAO

[Pol89a] Douglas Poland. The kinetics of adsorption to a one-dimensional lattice with nearest-neighbor exclusion: A series expansion study. Journal of Statistical Physics, 55(1–2):389–415, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01042608.

Polewczak:1989:GEM

[Pol89b] Jacek Polewczak. Global existence in L^{1} for the modified nonlinear Enskog equation in R^{3}. Journal of Statistical Physics, 56 (1–2):159–173, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01044239.

Pomeau:1981:SPP

[Pom81] Y. Pomeau. A solvable problem in Poisson statistics. Journal of Statistical Physics, 24(1):189–197, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01007643.

Pomeau:1985:NBE

[Pom85] Y. Pomeau. A note on the Boltzmann equation for hard spheres. Journal of Statistical Physics, 40(1–2):361–364, July 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010540.

Piasecki:1982:LEB

[PP82] J. Piasecki and Y. Pomeau. Large energy behavior of the velocity distribution for the hard-sphere gas. Journal of Statistical Physics, 28(2):235–255, June 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012605.

Percus:1983:PTF

[PP83] O. E. Percus and J. K. Percus. Phase transition in a fourdimensional random walk with application to medical statistics. Journal of Statistical Physics, 30(3):755–783, March 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009686.

Parsons:1987:QDA

[PP87] R. W. Parsons and P. K. Pollett. Quasistationary distributions for autocatalytic reactions. Journal of Statistical Physics, 46(1– 2):249–254, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010344.

Parris:1984:MCP

[PPK84] P. E. Parris, S. M. Phatak, and V. M. Kenkre. Motion and capture in the presence of cooperative trap interactions II: Exact calculations for perfect absorbers in one dimension. Journal of Statistical Physics, 35(5–6):749–760, June 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010831.

Papatriantafillou:1989:ETB

[PPKP89] C. Papatriantafillou, C. E. Paraskevaidis, C. Krikos, and P. Panagopoulos. Exact thermodynamic behavior of a generalized spin-isospin Ising system on the Bethe lattice. Journal of Statistical Physics, 54(1–2):191–200, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01023478.

Pomeau:1984:ISM

[PPP84] Y. Pomeau, A. Pumir, and P. Pelce. Intrinsic stochasticity with many degrees of freedom. Journal of Statistical Physics, 37(1– 2):39–49, October 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012904.

Prakash:1983:HTD

[Pra83] Chetan Prakash. High-temperature differentiability of lattice Gibbs states by Dobrushin uniqueness techniques. Journal of Statistical Physics, 31(1):169–228, April 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010929.

Priezzhev:1981:SDTa

[Pri81a] V. B. Priezzhev. The statistics of dimers on a three-dimensional lattice. I. An exactly solvable model. Journal of Statistical Physics, 26(4):817–828, December 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010943.

Priezzhev:1981:SDTb

[Pri81b] V. B. Priezzhev. The statistics of dimers on a three-dimensional lattice. II. An improved lower bound. Journal of Statistical Physics, 26(4):829–837, December 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010944.

Priezzhev:1986:IWC

[Pri86] V. B. Priezzhev. Interacting walkers on the Cayley tree, and polymer statistics. Journal of Statistical Physics, 44(5–6):921– 932, September 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011914.

Procaccia:1984:FST

[Pro84] Itamar Procaccia. Fractal structures in turbulence. Journal of Statistical Physics, 36(5–6):649–663, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012929.

Perez:1982:GCS

[PS82a] J. Fernando Perez and R. H. Schonmann. On the global character of some restricted equilibrium conditions — a remark on metastability. Journal of Statistical Physics, 28(3):479–485, July 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008319.

Petrosky:1982:DQK

[PS82b] T. Y. Petrosky and W. C. Schieve. The derivation of the quantum kinetic equation and the two-time resolvent method. Journal of Statistical Physics, 28(4):711–744, August 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011877.

Privman:1982:ACF

[PS82c] V. Privman and L. S. Schulman. Analytic continuation at firstorder phase transitions. Journal of Statistical Physics, 29(2): 205–229, October 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020783.

Presutti:1985:TEO

[PS85] E. Presutti and E. Scacciatelli. Time evolution of a onedimensional point system: A note on Fritz’s paper. Journal of Statistical Physics, 38(3–4):647–653, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010482.

Perram:1987:MDF

[PS87] J. W. Perram and E. R. Smith. Microscopic derivation of fluctuation formulas for calculating dielectric constants by simulation. Journal of Statistical Physics, 46(1–2):179–190, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010339.

Pearce:1988:CTH

[PS88a] Paul A. Pearce and Katherine A. Seaton. A classical theory of hard squares. Journal of Statistical Physics, 53(5–6):1061–1072, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023857.

Privman:1988:DESa

[PS88b] V. Privman and N. M. Svrakić. Difference equations in statistical mechanics. I. Cluster statistics models. Journal of Statistical Physics, 51(5–6):1091–1110, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014901.

Privman:1988:DESb

[PS88c] V. Privman and N. M. Svrakić. Difference equations in statistical mechanics. II. Solid-on-solid models in two dimensions. Journal of Statistical Physics, 51(5–6):1111–1126, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014902.

Privman:1988:RUR

[PS88d] V. Privman and N. M. Svrakić. Restoration of universality for the rod-to-coil transition scaling in the infinite-dimensionality limit: Exact results for directed walks. Journal of Statistical Physics, 50(1–2):81–89, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01022988.

Privman:1989:ADT

[PS89] V. Privman and N. M. Svrakić. Asymptotic degeneracy of the transfer matrix spectrum for systems with interfaces: Relation to surface stiffness and step free energy. Journal of Statistical Physics, 54(3–4):735–754, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019773.

Pandey:1984:DRS

[PSMZ84] R. B. Pandey, D. Stauffer, A. Margolina, and J. G. Zabolitzky. Diffusion on random systems above, below, and at their percolation threshold in two and three dimensions. Journal of Statistical Physics, 34(3–4):427–450, February 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01018553.

Pandey:1987:MCE

[PSZ87] R. B. Pandey, D. Stauffer, and J. G. Zabolitzky. Monte Carlo evidence for the deviation from the Alexander–Orbach rule in threedimensional percolation. Journal of Statistical Physics, 49(3–4): 849–854, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009361.

Peschel:1986:KPC

[PT86] I. Peschel and T. T. Truong. The kinetic Potts chain and related Potts models with competing interactions. Journal of Statistical Physics, 45(1–2):233–244, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01033089.

Pisani:1987:GCT

[PT87] Carmelo Pisani and Colin J. Thompson. Generalized classical theory of magnetism. Journal of Statistical Physics, 46(5–6):971–982, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011152.

Pitsianis:1989:DOD

[PTB^{+}89] N. Pitsianis, Ph. Tsalides, G. L. Bleris, A. Thanailakis, and H. C. Card. Deterministic one-dimensional cellular automata. Journal of Statistical Physics, 56(1–2):99–112, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01044234.

Pulvirenti:1982:TES

[Pul82] Mario Pulvirenti. On the time evolution of the states of infinitely extended particles systems. Journal of Statistical Physics, 27 (4):693–709, April 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013442.

Paladin:1989:HDT

[PV89] G. Paladin and S. Vaienti. Hausdorff dimensions in twodimensional maps and thermodynamic formalism. Journal of Statistical Physics, 57(1–2):289–299, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01023644.

Perez:1981:PTR

[PW81] J. Fernando Perez and W. F. Wreszinski. Phase transitions and reflection positivity for a class of quantum lattice systems. Journal of Statistical Physics, 26(2):401–420, October 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01013180.

Pietronero:1984:SMD

[PW84] L. Pietronero and H. J. Wiesmann. Stochastic model for dielectric breakdown. Journal of Statistical Physics, 36(5–6): 909–916, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012949.

Packard:1985:TDC

[PW85] Norman H. Packard and Stephen Wolfram. Two-dimensional cellular automata. Journal of Statistical Physics, 38(5–6):901–946, March 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010423.

Perk:1986:NSM

[PW86] J. H. H. Perk and F. Y. Wu. Nonintersecting string model and graphical approach: Equivalence with a Potts model. Journal of Statistical Physics, 42(5–6):727–742, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010443.

Presutti:1988:MSFa

[PW88a] Errico Presutti and W. David Wick. Macroscopic stochastic fluctuations in a one-dimensional mechanical system. Journal of Statistical Physics, 51(5–6):871–876, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01014889.

Presutti:1988:MSFb

[PW88b] Errico Presutti and W. David Wick. Macroscopic stochastic fluctuations in a one-dimensional mechanical system. Journal of Statistical Physics, 52(1–2):497–502, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01016431.

Perez:1984:MSM

[PWvH84] J. F. Perez, W. F. Wreszinski, and J. L. van Hemmen. The mean spherical model in a random external field and the replica method. Journal of Statistical Physics, 35(1–2):89–98, April 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017366.

Parisi:1985:FTG

[PYC85] Giorgio Parisi and Zhang Yi-Cheng. Field theories and growth models. Journal of Statistical Physics, 41(1–2):1–16, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020601.

Chen:1981:HECa

[qC81] Tian quan Chen. Hilbert-Enskog-Chapman expansion in the turbulent kinetic theory of gases. I. Journal of Statistical Physics, 25(3):491–514, July 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010802.

Chen:1981:HECb

[qCeY81] Tian quan Chen and Miao en Yuan. Hilbert-Enskog-Chapman expansion in the turbulent kinetic theory of gases. II. Journal of Statistical Physics, 25(3):515–567, July 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010803.

Radin:1981:GSS

[Rad81] Charles Radin. The ground state for soft disks. Journal of Statistical Physics, 26(2):365–373, October 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013177. See note

[Wag83].

Radin:1984:CGS

[Rad84] Charles Radin. Classical ground states in one dimension. Journal of Statistical Physics, 35(1–2):109–117, April 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01017368.

Radin:1986:CCG

[Rad86] Charles Radin. Correlations in classical ground states. Journal of Statistical Physics, 43(3–4):707–712, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020661.

Raggio:1988:FES

[Rag88] G. A. Raggio. The free energy of the spin-boson model. Journal of Statistical Physics, 53(3–4):565–581, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014214.

Rainwater:1981:DFE

[Rai81] James C. Rainwater. Delta-function expansion of Mayer function with application to virial coefficients. Journal of Statistical Physics, 26(2):423–424, October 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01013182.

Rammal:1984:RWS

[Ram84a] R. Rammal. Random walk statistics on fractal structures. Journal of Statistical Physics, 36(5–6):547–560, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012921.

Ramshaw:1984:DPR

[Ram84b] John D. Ramshaw. Dielectric polarization in random media. Journal of Statistical Physics, 35(1–2):49–75, April 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017364.

Ramshaw:1985:ALA

[Ram85] John D. Ramshaw. Augmented Langevin approach to fluctuations in nonlinear irreversible processes. Journal of Statistical Physics, 38(3–4):669–680, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010484.

Ramshaw:1986:EDN

[Ram86] John D. Ramshaw. Elementary derivation of nonlinear transport equations from statistical mechanics. Journal of Statistical Physics, 45(5–6):983–999, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020585.

Ramshaw:1989:PSD

[Ram89] John D. Ramshaw. Phase space density representations in fluid dynamics. Journal of Statistical Physics, 56(1–2):149–158, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01044238.

Rapaport:1985:FNM

[Rap85] D. C. Rapaport. The fractal nature of molecular trajectories in fluids. Journal of Statistical Physics, 40(5–6):751–758, September 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009898.

Rapaport:1987:BRB

[Rap87] D. C. Rapaport. Book review: Monte Carlo methods. Volume I: Basics. Journal of Statistical Physics, 49(3–4):867–868, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009364.

Rapaport:1988:BRI

[Rap88] Dennis Rapaport. Book review: An introduction to computer simulation methods. Journal of Statistical Physics, 53(5–6): 1341–1343, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023874.

Rapaport:1989:BRM

[Rap89] Dennis C. Rapaport. Book review: Monte Carlo simulation in statistical physics: An introduction. Journal of Statistical Physics, 57(5–6):1139–1140, December 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020054.

Ravishankar:1988:APS

[Rav88] K. Ravishankar. Asymptotic properties of a simple random motion. Journal of Statistical Physics, 53(3–4):977–983, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014233.

Ravishankar:1989:PLS

[Rav89] K. Ravishankar. Power law scaling of the top Lyapunov exponent of a product of random matrices. Journal of Statistical Physics, 54 (1–2):531–537, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023493.

Rubi:1988:BMF

[RB88] J. M. Rubi and D. Bedeaux. Brownian motion in a fluid in elongational flow. Journal of Statistical Physics, 53(1–2):125–136, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011549.

Reynolds:1986:MPC

[RBHL86] P. J. Reynolds, R. N. Barnett, B. L. Hammond, and W. A. Lester, Jr. Molecular physics and chemistry applications of quantum Monte Carlo. Journal of Statistical Physics, 43(5–6):1017–1026, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02628327.

Richetti:1987:CBZ

[RDRS87] P. Richetti, P. De Kepper, J. C. Roux, and Harry L. Swinney. A crisis in the Belousov–Zhabotinskii reaction: Experiment and simulation. Journal of Statistical Physics, 48(5–6): 977–990, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009527.

Rebbi:1986:LSC

[Reb86] Claudio Rebbi. Large-scale calculations for hadron spectroscopy. Journal of Statistical Physics, 43(5–6):1117–1127, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02628336.

Redner:1981:ODI

[Red81] S. Redner. One-dimensional Ising chain with competing interactions: Exact results and connection with other statistical models. Journal of Statistical Physics, 25(1):15–23, May 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008476.

Redner:1982:FPC

[Red82] S. Redner. A FORTRAN program for cluster enumeration. Journal of Statistical Physics, 29(2):309–315, October 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020788.

Redner:1989:BRC

[Red89] Sidney Redner. Book review: Chance and matter. Journal of Statistical Physics, 55(1–2):469–470, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01042613.

Reichl:1980:MMFa

[Rei80a] L. E. Reichl. Microscopic modes in a Fermi superfluid. I. Linearized kinetic equations. Journal of Statistical Physics, 23(1):83– 110, July 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014430.

Reichl:1980:MMFb

[Rei80b] L. E. Reichl. Microscopic modes in a Fermi superfluid. II. Dispersion relations. Journal of Statistical Physics, 23(1):111–125, July 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014431.

Reichl:1982:BR

[Rei82] L. E. Reichl. Book review. Journal of Statistical Physics, 29(3): 643–644, November 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01342194.

Reichl:1984:BRS

[Rei84] L. E. Reichl. Book review: Statistical mechanics: Methods and applications. Journal of Statistical Physics, 34(1–2):351–352, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01770366.

Reichl:1988:BRS

[Rei88a] L. E. Reichl. Book review: Studies in network thermodynamics. Journal of Statistical Physics, 50(1–2):465, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023005.

Reichl:1988:TFR

[Rei88b] L. E. Reichl. Transition in the Floquet rates of a driven stochastic system. Journal of Statistical Physics, 53(1–2):41–48, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011543.

Requardt:1982:APD

[Req82] Manfred Requardt. About the poor decay of certain crosscorrelation functions in the statistical mechanics of phase transitions in the static and dynamical regime. Journal of Statistical Physics, 29(1):117–127, September 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008252.

Requardt:1983:DCT

[Req83] Manfred Requardt. The decay of correlation of the two-particle distribution function in a phase-separating layer and the possibility of spatial phase coexistence. Journal of Statistical Physics, 31(3):679–689, June 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019505.

Requardt:1988:WGE

[Req88] Manfred Requardt. Wall and gravitational effects on the fine structure of interface layers at two-phase coexistence: Some rigorous results. Journal of Statistical Physics, 50(3–4):737–754, February 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026499.

Reznikova:1981:CLT

[Rez81] A. Ja. Reznikova. The central limit theorem for the spectrum of the random one-dimensional Schrödinger operator. Journal of Statistical Physics, 25(2):291–308, June 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01022188.

Romanelli:1988:DCN

[RFH88] L. Romanelli, M. A. Figliola, and F. A. Hirsch. Deterministic chaos and natural phenomena. Journal of Statistical Physics, 53(3–4):991–994, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014235.

Randelman:1986:CTS

[RG86a] R. E. Randelman and Gary S. Grest. N-city traveling salesman problem: Optimization by simulated annealings. Journal of Statistical Physics, 45(5–6):885–890, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020579.

Rudnick:1986:BPF

[RG86b] Joseph Rudnick and George Gaspari. Bond percolation on a finite lattice: The one-state Potts model reconsidered. Journal of Statistical Physics, 42(5–6):833–860, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010448.

Rudnick:1985:FSS

[RGJ85] Joseph Rudnick, Hong Guo, and David Jasnow. Finite-size scaling and the renormalization group. Journal of Statistical Physics, 41(3–4):353–373, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009013.

Roux:1989:POD

[RHG89] Stéphane Roux, Alex Hansen, and Etienne Guyon. Propagation of order in the dilute antiferromagnetic three-state Potts model. Journal of Statistical Physics, 55(1–2):341–350, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01042604.

Roux:1988:RHM

[RHHG88] Stéphane Roux, Alex Hansen, Hans Herrmann, and Etienne Guyon. Rupture of heterogeneous media in the limit of infinite disorder. Journal of Statistical Physics, 52(1–2):237–244, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01016411.

Reiss:1986:TTN

[RHM86] H. Reiss, Audrey Dell Hammerich, and E. W. Montroll. Thermodynamic treatment of nonphysical systems: Formalism and an example (single-lane traffic). Journal of Statistical Physics, 42(3– 4):647–687, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01127733.

Richards:1983:RWI

[Ric83] Peter M. Richards. Random walk to and interaction with an impurity. Journal of Statistical Physics, 30(2):497–507, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012323.

Riela:1985:TTT

[Rie85] Giovanni Riela. Transitions torichaos through collisions with hyperbolic orbits. Journal of Statistical Physics, 41(1–2):201–224, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020609.

Ray:1988:CBH

[RK88a] T. S. Ray and W. Klein. Crossover and the breakdown of hyperscaling in long-range bond percolation. Journal of Statistical Physics, 53(3–4):773–794, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01014225.

Rothman:1988:ICA

[RK88b] Daniel H. Rothman and Jeffrey M. Keller. Immiscible cellularautomaton fluids. Journal of Statistical Physics, 52(3–4):1119– 1127, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019743.

Runnels:1980:CTH

[RL80] L. K. Runnels and J. L. Lebowitz. Circle theorem for hard-core binary lattice gases. Journal of Statistical Physics, 23(1):1–10, July 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014426.

Ramshaw:1986:ALD

[RL86] John D. Ramshaw and Katja Lindenberg. Augmented Langevin description of multiplicative noise and nonlinear dissipation in Hamiltonian systems. Journal of Statistical Physics, 45(1–2): 295–307, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01033092.

Rosinberg:1986:SML

[RLB86] M. L. Rosinberg, J. L. Lebowitz, and L. Blum. A solvable model for localized adsorption in a Coulomb system. Journal of Statistical Physics, 44(1–2):153–182, July 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010909.

Rajagopal:1983:NDR

[RNRT83] A. K. Rajagopal, K. L. Ngai, R. W. Rendell, and S. Teitler. Nonexponential decay in relaxation phenomena. Journal of Statistical Physics, 30(2):285–292, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012303.

Rodenhausen:1989:ERB

[Rod89] Hermann Rodenhausen. Einstein’s relation between diffusion constant and mobility for a diffusion model. Journal of Statistical Physics, 55(5–6):1065–1088, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01041079.

Roepstorff:1981:SMC

[Roe81] Gert Roepstorff. Solvable models of classical lattice gases. Journal of Statistical Physics, 26(4):783–794, December 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010939.

Rosenfeld:1984:CMS

[Ros84] Yaakov Rosenfeld. Comments on the mean spherical approximation for hard-core and soft potentials and an application to the one-component plasma. Journal of Statistical Physics, 37(1– 2):215–236, October 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012912.

Rosenfeld:1986:CVM

[Ros86] Yaakov Rosenfeld. Comments on the variational modifiedhypernetted-chain theory for simple fluids. Journal of Statistical Physics, 42(3–4):437–457, February 1986. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01127720.

Ross:1987:CMS

[Ros87] Charles R. Ross II. CVM modeling of the square Ising lattice with one next-nearest-neighbor interaction. Journal of Statistical Physics, 49(1–2):261–268, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009962.

Rothman:1989:NVL

[Rot89] Daniel H. Rothman. Negative-viscosity lattice gases. Journal of Statistical Physics, 56(3–4):517–524, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01044449.

Roux:1987:GBM

[Rou87] Stéphane Roux. Generalized Brownian motion and elasticity. Journal of Statistical Physics, 48(1–2):201–213, July 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010406.

Rowlands:1985:QMC

[Row85] S. Rowlands. Quantum mechanical coherence in human red blood cells. Journal of Statistical Physics, 39(5–6):543–549, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008350.

Rowlinson:1986:PCG

[Row86] J. S. Rowlinson. The properties of connected graphs and some corrections to the list of Uhlenbeck and Ford. Journal of Statistical Physics, 42(5–6):931–933, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010453.

Radii:1985:SDC

[RP85] Remo Radii and Antonio Politi. Statistical description of chaotic attractors: The dimension function. Journal of Statistical Physics, 40(5–6):725–750, September 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009897.

Rodriguez:1985:MED

[RPSS85] M. A. Rodriguez, L. Pesquera, M. San Miguel, and J. M. Sancho. Master equation description of external Poisson white noise in finite systems. Journal of Statistical Physics, 40(5–6): 669–724, September 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009896.

Reiss:1982:EBC

[RR82] Howard Reiss and Gloria R. Reiss. Examination of bounds on the chemical potential, using a one-dimensional system. Journal of Statistical Physics, 29(4):795–802, December 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011792.

Romero-Rochin:1988:RPW

[RRO88] Victor Romero-Rochin and Irwin Oppenheim. Relaxation properties of weakly coupled classical systems. Journal of Statistical Physics, 53(1–2):307–322, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011559.

Roepstorff:1984:MAD

[RS84] G. Roepstorff and L. S. Schulman. Metastability and analyticity in a dropletlike model. Journal of Statistical Physics, 34 (1–2):35–56, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01770348.

Roerdink:1985:APMa

[RS85a] J. B. T. M. Roerdink and K. E. Shuler. Asymptotic properties of multistate random walks. I. Theory. Journal of Statistical Physics, 40(1–2):205–240, July 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010534.

Roerdink:1985:APMb

[RS85b] J. B. T. M. Roerdink and K. E. Shuler. Asymptotic properties of multistate random walks. II. Applications to inhomogeneous periodic and random lattices. Journal of Statistical Physics, 41(3– 4):581–606, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009023.

Radulescu:1987:DSP

[RS87a] Dan C. Radulescu and Daniel F. Styer. The Dobrushin–Shlosman phase uniqueness criterion and applications to hard squares. Journal of Statistical Physics, 49(1–2):281–295, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009964.

Roskies:1987:HTE

[RS87b] Ralph Z. Roskies and Penny D. Sackett. High-temperature expansions to fifteenth order. Journal of Statistical Physics, 49(3–4): 447–490, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009345.

Rossky:1986:QSA

[RSK86] Peter J. Rossky, Jürgen Schnitker, and Robert A. Kuharski. Quantum simulations of aqueous systems. Journal of Statistical Physics, 43(5–6):949–965, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF02628322.

Rodriguez:1983:FPL

[RSRD83] Rosalío F. Rodríguez, E. Salinas-Rodríguez, and James W. Dufty. Fokker–Planck and Langevin descriptions of fluctuations in uniform shear flow. Journal of Statistical Physics, 32(2):279–298, August 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012711.

Rogers:1981:ALR

[RT81] Jeffrey B. Rogers and Colin J. Thompson. Absence of long-range order in one-dimensional spin systems. Journal of Statistical Physics, 25(4):669–678, August 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01022361.

Richey:1986:BMS

[RT86] Matthew P. Richey and Craig A. Tracy. Z_{n} Baxter model: Symmetries and the Belavin parametrization. Journal of Statistical Physics, 42(3–4):311–348, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01127715.

Rubin:1982:CQS

[Rub82] Morton H. Rubin. On the control of quantum statistical systems. Journal of Statistical Physics, 28(1):177–188, May 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011631.

Rubin:1984:TDO

[Rub84] Robert J. Rubin. Transport in a disordered one-dimensional system: A fractal view. Journal of Statistical Physics, 36(5–6): 615–624, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012926.

Rubinstein:1986:MDS

[Rub86] Jacob Rubinstein. On the macroscopic description of slow viscous flow past a random array of spheres. Journal of Statistical Physics, 44(5–6):849–863, September 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011910.

Ruelle:1981:MTF

[Rue81] David Ruelle. Must thermodynamic functions be piecewise analytic? Journal of Statistical Physics, 26(2):397–399, October 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013179.

Ruelle:1986:LRA

[Rue86a] David Ruelle. Locating resonances for Axiom A dynamical systems. Journal of Statistical Physics, 44(3–4):281–292, August 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011300.

Ruelle:1986:BLT

[Rue86b] Philippe Ruelle. Brownian local time and quantum mechanics. Journal of Statistical Physics, 43(3–4):609–620, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020655.

Ruiz:1989:ISP

[Rui89] Jean Ruiz. Interface sharpness in the Potts model. Journal of Statistical Physics, 56(5–6):621–629, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01016770.

Rujan:1982:ODLa

[Ruj82a] P. Ruján. Order and disorder lines in systems with competing interactions: I. Quantum spins at T = 0. Journal of Statistical Physics, 29(2):231–245, October 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020784.

Rujan:1982:ODLb

[Ruj82b] P. Ruján. order and disorder lines in systems with competing interactions: II. The IRF model. Journal of Statistical Physics, 29(2):247–262, October 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020785.

Rujan:1984:ODL

[Ruj84] P. Ruján. Order and disorder lines in systems with competing interactions. III. Exact results from stochastic crystal growth. Journal of Statistical Physics, 34(3–4):615–646, February 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01018562.

Rujan:1987:CAS

[Ruj87] Pàl Rujàn. Cellular automata and statistical mechanical models. Journal of Statistical Physics, 49(1–2):139–222, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009958.

Robledo:1981:RBD

[RV81] A. Robledo and C. Varea. On the relationship between the density functional formalism and the potential distribution theory for nonuniform fluids. Journal of Statistical Physics, 26(3):513–525, November 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011432.

Risken:1985:EEF

[RV85] H. Risken and K. Voigtlaender. Eigenvalues and eigenfunctions of the Fokker–Planck equation for the extremely underdamped Brownian motion in a double-well potential. Journal of Statistical Physics, 41(5–6):825–863, December 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010006.

Robert:1984:FIP

[RW84] M. Robert and B. Widom. Field-induced phase separation in one dimension. Journal of Statistical Physics, 37(3–4):419–437, November 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011842.

Requardt:1986:PDC

[RW86] M. Requardt and H. J. Wagner. Poor decay of correlations in inhomogeneous fluids and solids and their relevance for the physics of phase boundaries. Journal of Statistical Physics, 45(5–6): 815–841, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020576.

Ryter:1987:NIT

[Ryt87] D. Ryter. Noise-induced transitions in a double-well potential at low friction. Journal of Statistical Physics, 49(3–4):751–765, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009355.

Rucklidge:1988:MMI

[RZ88] Alastair Rucklidge and Stéphane Zaleski. A microcanonical model for interface formation. Journal of Statistical Physics, 51(1– 2):299–307, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01015333.

Saleur:1986:CTD

[Sal86] H. Saleur. Collapse of two-dimensional linear polymers. Journal of Statistical Physics, 45(3–4):419–438, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01021079.

Saleur:1988:CFC

[Sal88] H. Saleur. Correlation functions of the critical Ashkin–Teller model on a torus. Journal of Statistical Physics, 50(3–4):475–508, February 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026488.

Sarkar:1984:SOT

[Sar84] Subir Kumar Sarkar. Supercritical ordered trajectories with quadratically irrational winding number. Journal of Statistical Physics, 37(3–4):385–405, November 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011840.

Sawada:1984:TVP

[Saw84] Yasuji Sawada. A thermodynamic variational principle in nonlinear systems far from equilibrium. Journal of Statistical Physics, 34(5–6):1039–1045, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009456.

Sadiq:1984:DFT

[SB84] A. Sadiq and K. Binder. Dynamics of the formation of twodimensional ordered structures. Journal of Statistical Physics, 35(5–6):517–585, June 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010824.

Schwartz:1988:LBM

[SB88] Moshe Schwartz and Ram Brustein. From Lagrangian to Brownian motion. Journal of Statistical Physics, 51(3–4):585–613, May 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01028474.

Silverberg:1988:DTC

[SBS88] M. Silverberg and A. Ben-Shaul. Dynamic and thermodynamic consequences of adsorbate lateral interactions in surface reaction kinetics. Journal of Statistical Physics, 52(5–6):1179– 1195, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011640.

Schmitz:1985:FFUa

[SC85a] R. Schmitz and E. G. D. Cohen. Fluctuations in a fluid under a stationary heat flux. I. General theory. Journal of Statistical Physics, 39(3–4):285–316, May 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01018664.

Schmitz:1985:FFUb

[SC85b] R. Schmitz and E. G. D. Cohen. Fluctuations in a fluid under a stationary heat flux II. Slow part of the correlation matrix. Journal of Statistical Physics, 40(3–4):431–482, August 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01017182.

Schmitz:1987:FFU

[SC87] R. Schmitz and E. G. D. Cohen. Fluctuations in a fluid under a stationary heat flux. III. Brillouin lines. Journal of Statistical Physics, 46(1–2):319–348, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010349.

Shlesinger:1988:LTS

[SC88] Michael F. Shlesinger and Michael A. Coplan. On the long-tail solar wind electron velocity distribution. Journal of Statistical Physics, 52(5–6):1423–1428, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011657.

Scalapino:1986:STS

[Sca86] D. J. Scalapino. Simulations: A tool for studying quantum condensed matter systems. Journal of Statistical Physics, 43 (5–6):757–770, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02628303.

Schattke:1983:SSM

[Sch83] W. Schattke. Special solution of the mode-coupling equations for a two-component magnetic system. Journal of Statistical Physics, 30(3):607–615, March 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009679.

Schlijper:1984:EVM

[Sch84] A. G. Schlijper. Exact variational methods and cluster-variation approximations. Journal of Statistical Physics, 35(3–4):285–301, May 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014385.

Schlijper:1985:SVA

[Sch85a] A. G. Schlijper. On some variational approximations in twodimensional classical lattice systems. Journal of Statistical Physics, 40(1–2):1–27, July 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010524.

Schonmann:1985:MCP

[Sch85b] Roberto H. Schonmann. Metastability for the contact process. Journal of Statistical Physics, 41(3–4):445–464, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009017.

Schrader:1985:ECR

[Sch85c] Robert Schrader. On the Euler characteristics of random surfaces. Journal of Statistical Physics, 40(3–4):533–561, August 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017184.

Schwabl:1985:PSC

[Sch85d] F. Schwabl. Propagation of sound at continuous structural phase transitions. Journal of Statistical Physics, 39(5–6):719–737, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008362.

Schneider:1986:MSD

[Sch86a] Toni Schneider. Mapping of stochastic dynamics onto associated quantum models and (d+1)-dimensional classical and static systems. Journal of Statistical Physics, 43(5–6):877–878, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02628316.

Schonmann:1986:ACP

[Sch86b] Roberte H. Schonmann. The asymmetric contact process. Journal of Statistical Physics, 44(3–4):505–534, August 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011308.

Schulman:1986:DQE

[Sch86c] L. S. Schulman. Deterministic quantum evolution through modification of the hypotheses of statistical mechanics. Journal of Statistical Physics, 42(3–4):689–719, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01127734.

Schlogl:1987:VILa

[Sch87a] F. Schlögl. The variance of information loss as a characteristic quantity of dynamical chaos. Journal of Statistical Physics, 46 (1–2):135–146, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010336.

Schlogl:1987:VILb

[Sch87b] F. Schlögl. The variance of information loss as a characteristic quantity of dynamical chaos. Journal of Statistical Physics, 49 (3–4):873, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009367.

Schmitt:1987:PCM

[Sch87c] K.-J. Schmitt. n-particle correlations in the McKean model. Journal of Statistical Physics, 46(1–2):283–302, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010347.

Schilling:1988:SQO

[Sch88a] R. Schilling. Slow quenching for a one-dimensional kinetic Ising model: Residual energy and domain growth. Journal of Statistical Physics, 53(5–6):1227–1235, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01023866.

Schlijper:1988:TPU

[Sch88b] A. G. Schlijper. Tiling problems and undecidability in the cluster variation method. Journal of Statistical Physics, 50(3–4):689–714, February 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026496.

Schonmann:1988:TCI

[Sch88c] Roberto H. Schonmann. On two correlation inequalities for Potts models. Journal of Statistical Physics, 52(1–2):61–67, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016404.

Schenzle:1989:CQN

[Sch89a] Axel Schenzle. Classical and quantum noise in nonlinear optical systems. Journal of Statistical Physics, 54(5–6):1243–1288, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01044715.

Schmitz:1989:SMN

[Sch89b] R. Schmitz. Sound modes in a nonequilibrium fluid with acoustic boundary conditions. Journal of Statistical Physics, 57(3–4): 549–579, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01022822.

Stella:1981:DRK

[SD81] Attilio Stella and Raf Dekeyser. Dynamical renormalization of kinetic van der Waals spin models. Journal of Statistical Physics, 25(3):443–458, July 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010799.

Spohn:1985:QTD

[SD85] Herbert Spohn and Rolf Dümcke. Quantum tunneling with dissipation and the Ising model over R. Journal of Statistical Physics, 41(3–4):389–423, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009015.

Saleur:1986:TMC

[SD86] H. Saleur and B. Derrida. Transfer matrix calculation of the exponent γ for two-dimensional self-avoiding walks. Journal of Statistical Physics, 44(1–2):225–235, July 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010914.

Seke:1983:SET

[Sek83] J. Seke. Spontaneous emission of a two-level system and the influence of the rotating-wave approximation on the final state. I. Journal of Statistical Physics, 33(1):223–229, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009757.

Selke:1989:DTD

[Sel89] Walter Selke. Droplets in two-dimensional Ising and Potts models. Journal of Statistical Physics, 56(5–6):609–620, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016769.

Sengers:1982:BR

[Sen82] J. V. Sengers. Book review. Journal of Statistical Physics, 29 (3):641, November 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01342193.

Sengers:1986:BRM

[Sen86] J. V. Sengers. Book review: Multicritical phenomena. Journal of Statistical Physics, 42(3–4):721, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01127735.

Sengupta:1988:DSA

[Sen88] A. Sengupta. A discretized spectral approximation in neutron transport theory. Some numerical considerations. Journal of Statistical Physics, 51(3–4):657–676, May 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01028477.

Serra:1986:SRQ

[Ser86] Maria Serra. Scattering of radiation by a quasiperiodic twodimensional medium. Journal of Statistical Physics, 42(5–6): 961–977, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010457.

Schimmele:1988:NER

[SF88] L. Schimmele and M. Fähnle. Note on eigenvectors of a renormalization transformation. Journal of Statistical Physics, 50(1– 2):459–460, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023003.

Staaden:1989:CCT

[SF89] U. Staaden and M. Fähnle. Comment on the calculation of thermal averages by long-time Monte Carlo simulations. Journal of Statistical Physics, 57(1–2):405–410, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01023653.

Soulet:1981:CGF

[SG81] Y. Soulet and A. Gomes. Causal Green’s function formalism in statistical mechanics of composite particles. Journal of Statistical Physics, 25(4):695–716, August 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01022363.

Schramm:1987:LTL

[SG87] Peter Schramm and Hermann Grabert. Low-temperature and long-time anomalies of a damped quantum particle. Journal of Statistical Physics, 49(3–4):767–810, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009356.

Sasaki:1988:ECC

[SG88] Kazuo Sasaki and Robert B. Griffiths. Equivalence of certain convex and nonconvex models of spatially modulated structures. Journal of Statistical Physics, 53(5–6):1031–1040, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023855.

Suto:1989:PDL

[SGL89] András Sütö, Christian Gruber, and Pirmin Lemberger. Phase diagrams of lattice systems with residual entropy. II. Low temperature expansion. Journal of Statistical Physics, 56(3–4):261–290, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01044437.

Satija:1982:IOM

[SH82] Indubala I. Satija and P. C. Hohenberg. Irrelevant operators and momentum-shell recursion relations in d = 2 + ϵ dimensions. Journal of Statistical Physics, 28(1):83–97, May 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011624.

Schumaker:1989:MPB

[SH89a] Mark Schumaker and Werner Horsthemke. A modified pitchfork bifurcation as a model of the TI–TII transition in liquid helium counterflow with external noise. Journal of Statistical Physics, 54 (5–6):1189–1206, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01044711.

Sigeti:1989:PRO

[SH89b] David Sigeti and Werner Horsthemke. Pseudo-regular oscillations induced by external noise. Journal of Statistical Physics, 54(5– 6):1217–1222, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01044713.

Stovneng:1989:BLM

[SH89c] J. A. Støvneng and E. H. Hauge. The Büttiker–Landauer model generalized. Journal of Statistical Physics, 57(3–4):841–863, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/

10.1007/BF01022836. See response [BL90].

Shalitin:1982:RBS

[Sha82a] D. Shalitin. Relations between site percolation thresholds. Journal of Statistical Physics, 28(1):99–110, May 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011625.

Shankar:1982:SSV

[Sha82b] R. Shankar. On the solution of some vertex models using factorizable S matrices. Journal of Statistical Physics, 29(4):649–682, December 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011784.

Sharpe:1986:LQG

[Sha86] Stephen R. Sharpe. Lattice QCD: Going beyond the mass spectrum. Journal of Statistical Physics, 43(5–6):1129–1145, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02628337.

Shastry:1988:DST

[Sha88] B. Sriram Shastry. Decorated star-triangle relations and exact integrability of the one-dimensional Hubbard model. Journal of Statistical Physics, 50(1–2):57–79, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01022987.

Shiner:1981:FEF

[Shi81] J. S. Shiner. Fluxes and the elimination of fast-relaxing variables. Journal of Statistical Physics, 26(3):555–565, November 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011435.

Shinomoto:1983:ETH

[Shi83] Shigeru Shinomoto. Equilibrium theory for the hard-core systems. Journal of Statistical Physics, 32(1):105–113, July 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009423.

Schreiber:1986:PAR

[SHKM86] Igor Schreiber, Martin Holodniok, Milan Kubícek, and Milos Marek. Periodic and aperiodic regimes in coupled dissipative chemical oscillators. Journal of Statistical Physics, 43(3–4):489– 519, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020650.

Shlosman:1980:CIA

[Shl80] S. B. Shlosman. Correlation inequalities for antiferromagnets. Journal of Statistical Physics, 22(1):59–64, January 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01007987.

Shlesinger:1984:WWD

[Shl84] Michael F. Shlesinger. Williams—Watts dielectric relaxation: A fractal time stochastic process. Journal of Statistical Physics, 36(5–6):639–648, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012928.

Shlesinger:1985:BRS

[Shl85] Michael F. Shlesinger. Book review: Statistical physics and the atomic theory of matter, from Boyle and Newton to Landau and Onsager. Journal of Statistical Physics, 40(5–6):783–785, September 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009900.

Shlesinger:1988:BRC

[Shl88] Michael F. Shlesinger. Book review: Chaos: Making a new science. Journal of Statistical Physics, 50(5–6):1285–1286, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01019170.

Saleur:1987:TDF

[SI87] H. Saleur and C. Itzykson. Two-dimensional field theories close to criticality. Journal of Statistical Physics, 48(3–4):449–475, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019682.

Sprik:1986:SES

[SIK86] M. Sprik, R. W. Impey, and M. L. Klein. Study of electron solvation in polar solvents using path integral calculations. Journal of Statistical Physics, 43(5–6):967–972, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF02628323.

Silbey:1981:BR

[Sil81] Robert Silbey. Book review. Journal of Statistical Physics, 26 (2):421–422, October 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013181.

Silbey:1982:BR

[Sil82] Robert Silbey. Book review. Journal of Statistical Physics, 28(1): 231, May 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011635.

Simon:1980:MFU

[Sim80] Barry Simon. Mean field upper bound on the transition temperature in multicomponent ferromagnets. Journal of Statistical Physics, 22(4):491–493, April 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012868.

Simon:1981:ACS

[Sim81a] Barry Simon. Absence of continuous symmetry breaking in a one-dimensional n^{−2} model. Journal of Statistical Physics, 26 (2):307–311, October 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013173.

Simon:1981:RFI

[Sim81b] Barry Simon. The rate of falloff of Ising model correlations at large temperature. Journal of Statistical Physics, 26(1): 53–58, September 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01106785.

Simon:1985:LTA

[Sim85] Barry Simon. Lifschitz tails for the Anderson model. Journal of Statistical Physics, 38(1–2):65–76, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01017848.

Simon:1987:ILT

[Sim87] Barry Simon. Internal Lifschitz tails. Journal of Statistical Physics, 46(5–6):911–918, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011147.

Sinai:1982:CIP

[Sin82] Ya. G. Sinai. Commensurate-incommensurate phase transitions in one-dimensional chains. Journal of Statistical Physics, 29(3): 401–425, November 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01342181.

Singh:1983:STC

[Sin83] H. B. Singh. Structure and thermodynamics of the classical onecomponent plasma. Journal of Statistical Physics, 33(2):371–383, November 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009802.

Sinai:1987:ALO

[Sin87] Ya. G. Sinai. Anderson localization for one-dimensional difference Schrödinger operator with quasiperiodic potential. Journal of Statistical Physics, 46(5–6):861–909, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011146.

Siu:1985:OUB

[Siu85] Byron Siu. The Oguchi upper bound on the magnetization for ferromagnetic Ising models. Journal of Statistical Physics, 38(3– 4):519–530, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010475.

Sikkenk:1988:SWD

[SIvL^{+}88] J. H. Sikkenk, J. O. Indekeu, J. M. J. van Leeuwen, E. O. Vossnack, and A. F. Bakker. Simulation of wetting and drying at solid-fluid interfaces on the Delft Molecular Dynamics Processor. Journal of Statistical Physics, 52(1–2):23–44, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01016402.

Shenker:1982:CBK

[SK82] Scott J. Shenker and Leo P. Kadanoff. Critical behavior of a KAM surface: I. Empirical results. Journal of Statistical Physics, 27(4):631–656, April 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013439.

Schobinger:1986:LDS

[SKA86] M. Schöbinger, S. W. Koch, and Farid F. Abraham. Langevin dynamics simulations of phase separation in a one-component system. Journal of Statistical Physics, 42(5–6):1071–1089, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010462.

Sekimoto:1987:AND

[SKO87] Ken Sekimoto, Kyozi Kawasaki, and Hideaki Ohta. Aspects of nucleation and drift processes in a one-dimensional model. Journal of Statistical Physics, 48(5–6):1213–1241, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009542.

Skrypnik:1984:CFI

[Skr84] W. I. Skrypnik. Correlation functions of infinite system of interacting Brownian particles; local in time evolution close to equilibrium. Journal of Statistical Physics, 35(5–6):587–602, June 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010825.

Sung:1982:MTI

[SKS82] Wokyung Sung, John Karkheck, and G. Stell. Microscopic theory of intrinsic shear and bulk viscosities. Journal of Statistical Physics, 27(3):513–532, March 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011090.

Shlesinger:1982:RWI

[SKW82] Michael F. Shlesinger, Joseph Klafter, and Y. M. Wong. Random walks with infinite spatial and temporal moments. Journal of Statistical Physics, 27(3):499–512, March 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011089.

Seshadri:1980:ESW

[SL80] V. Seshadri and Katja Lindenberg. Extrema statistics of Wiener– Einstein processes in one, two, and three dimensions. Journal of Statistical Physics, 22(1):69–79, January 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01007989.

Slawny:1983:MFT

[Sla83] J. Slawny. On the mean field theory bound on the magnetization. Journal of Statistical Physics, 32(2):375–388, August 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012716.

Stocks:1989:ASQ

[SLM89] N. G. Stocks, C. J. Lambert, and P. V. E. McClintock. Analogue simulation of quantum mechanical systems. Journal of Statistical Physics, 54(5–6):1397–1410, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01044723.

Szabo:1984:LPT

[SLW84] Attila Szabo, Gene Lamm, and George H. Weiss. Localized partial traps in diffusion processes and random walks. Journal of Statistical Physics, 34(1–2):225–238, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01770356.

Schmidt:1986:MCC

[SM86] K. E. Schmidt and J. W. Moskowitz. Monte Carlo calculations of atoms and molecules. Journal of Statistical Physics, 43(5– 6):1027–1041, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02628328.

Sinha:1989:ASS

[SM89] K. Sinha and Frank Moss. Analog simulation of a simple system with state-dependent diffusion. Journal of Statistical Physics, 54 (5–6):1411–1423, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01044724.

Smith:1988:MSM

[Smi88] E. R. Smith. A mean spherical model with Coulomb interactions. Journal of Statistical Physics, 50(3–4):813–838, February 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01026501.

Smith:1989:MSM

[Smi89] E. R. Smith. A mean spherical model with Coulomb interactions. II. Correlations at a free surface. Journal of Statistical Physics, 55(1–2):127–139, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01042593.

Sheu:1988:ANS

[SMK88] Jin Sheng Sheu, Jer Ru Maa, and Joseph L. Katz. Adsorption and nucleation on smooth surfaces. Journal of Statistical Physics, 52(5–6):1143–1155, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011638.

Stocks:1989:EEF

[SMM89] N. G. Stocks, R. Mannella, and P. V. E. McClintock. Effect of external fluctuations on the Fréedericksz transition in an analogue simulator. Journal of Statistical Physics, 54(5–6):1383–1396, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01044722.

Shigematsu:1983:ASP

[SMYO83] H. Shigematsu, H. Mori, T. Yoshida, and H. Okamoto. Analytic study of power spectra of the tent maps near band-splitting transitions. Journal of Statistical Physics, 30(3):649–679, March 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009682.

Stanton:1984:RWM

[SN84] C. J. Stanton and Mark Nelkin. Random-walk model for equilibrium resistance fluctuations. Journal of Statistical Physics, 37 (1–2):1–16, October 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012901.

Sokal:1981:MIC

[Sok81a] Alan D. Sokal. More inequalities for critical exponents. Journal of Statistical Physics, 25(1):25–50, May 1981. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008477.

Sokal:1981:RPH

[Sok81b] Alan D. Sokal. Rigorous proof of the high-temperature Josephson inequality for critical exponents. Journal of Statistical Physics, 25(1):51–56, May 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008478.

Sokal:1982:MFB

[Sok82] Alan D. Sokal. Mean-field bounds and correlation inequalities. Journal of Statistical Physics, 28(3):431–439, July 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008316.

Soria:1988:CIT

[Sor88] José L. Soria. Correlation inequalities for two-component hypercubic φ^{4} models. Journal of Statistical Physics, 52(3–4):711–726, August 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019725.

Sornette:1989:ALW

[Sor89] Didier Sornette. Anderson localization and wave absorption. Journal of Statistical Physics, 56(5–6):669–680, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016773.

Svrakic:1988:FSC

[SPA88] N. M. Svrakić, V. Privman, and D. B. Abraham. Finite-size corrections for inclined interfaces in two dimensions: Exact results for Ising and solid-on-solid models. Journal of Statistical Physics, 53(5–6):1041–1059, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01023856.

Speziale:1984:ICE

[Spe84] Charles G. Speziale. On the invariance of constitutive equations according to the kinetic theory of gases. Journal of Statistical Physics, 35(3–4):457–470, May 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014396.

Speer:1986:MCT

[Spe86] Eugene R. Speer. Mayer coefficients in two-dimensional Coulomb systems. Journal of Statistical Physics, 42(5–6):895–920, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010451.

Spencer:1988:LRQ

[Spe88] Thomas Spencer. Localization for random and quasiperiodic potentials. Journal of Statistical Physics, 51(5–6):1009–1019, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014897.

Spigler:1985:NPO

[Spi85] Renato Spigler. Nonlinear parametric oscillations in certain stochastic systems: A random van der Pol oscillator. Journal of Statistical Physics, 41(1–2):175–200, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020608.

Spohn:1981:FAB

[Spo81] Herbert Spohn. Fluctuations around the Boltzmann equation. Journal of Statistical Physics, 26(2):285–305, October 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013172.

Spouge:1983:SDM

[Spo83] John L. Spouge. The size distribution for the A_{g}RB_{f−g} model of polymerization. Journal of Statistical Physics, 31(2):363–378, May 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011587.

Spohn:1985:BRN

[Spo85a] Herbert Spohn. Book review: Non-equilibrium entropy and irreversibility. Journal of Statistical Physics, 39(1–2):259–261, April 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01007982.

Spouge:1985:PRG

[Spo85b] John L. Spouge. Polymers and random graphs: Asymptotic equivalence to branching processes. Journal of Statistical Physics, 38(3–4):573–587, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010478.

Spohn:1986:BRF

[Spo86a] Herbert Spohn. Book review: Fundamental problems in statistical mechanics VI. Journal of Statistical Physics, 42(5–6): 1135–1136, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010467.

Spouge:1986:ERF

[Spo86b] John L. Spouge. Equilibrium ring formation in polymer solutions. Journal of Statistical Physics, 43(1–2):143–196, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010576.

Spohn:1987:TDD

[Spo87] Herbert Spohn. Tracer dynamics in Dyson’s model of interacting Brownian particles. Journal of Statistical Physics, 47(5–6):669– 679, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01206151.

Shapiro:1986:FFS

[SR86] Jonathan Shapiro and Joseph Rudnick. The fully finite spherical model. Journal of Statistical Physics, 43(1–2):51–83, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010572.

SanMiguel:1980:CNA

[SS80] M. San Miguel and J. M. Sancho. A colored-noise approach to Brownian motion in position space. Corrections to the Smoluchowski equation. Journal of Statistical Physics, 22(5):605–624, May 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011341.

Simon:1981:REE

[SS81] Barry Simon and Alan D. Sokal. Rigorous entropy-energy arguments. Journal of Statistical Physics, 25(4):679–694, August 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01022362.

Schulman:1982:PAS

[SS82a] Lawrence S. Schulman and Philip E. Seiden. Percolation analysis of stochastic models of galactic evolution. Journal of Statistical Physics, 27(1):83–118, January 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011742.

Simon:1982:REE

[SS82b] Barry Simon and Alan D. Sokal. Rigorous entropy-energy arguments. Journal of Statistical Physics, 29(1):155, September 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008255.

Sutherland:1983:ESL

[SS83] Bill Sutherland and B. Sriram Shastry. Exact solution of a large class of interacting quantum systems exhibiting ground state singularities. Journal of Statistical Physics, 33(2):477–484, November 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009806.

Sancho:1984:TET

[SS84a] J. M. Sancho and M. San Miguel. Theory of external two-state Markov noise in the presence of internal fluctuations. Journal of Statistical Physics, 37(1–2):151–172, October 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012909.

Simon:1984:FAM

[SS84b] B. Simon and B. Souillard. Franco–American meeting on the mathematics of random and almost periodic potentials. Journal of Statistical Physics, 36(1–2):273–288, July 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01015735.

Shklovskii:1985:EIS

[SS85] B. I. Shklovskii and B. Z. Spivak. Effect of impurity scattering of a tunneling electron on variable-range hopping conduction. Journal of Statistical Physics, 38(1–2):267–272, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01017862.

Shuhov:1986:ODX

[SS86] A. G. Shuhov and Yu. M. Suhov. One-dimensional XY model: Ergodic properties and hydrodynamic limit. Journal of Statistical Physics, 45(3–4):669–694, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01021090.

Scheucher:1988:SKM

[SS88a] Manfred Scheucher and Herbert Spohn. A soluble kinetic model for spinodal decomposition. Journal of Statistical Physics, 53(1– 2):279–294, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011557.

Schoolderman:1988:KTT

[SS88b] A. J. Schoolderman and L. G. Suttorp. Kinetic theory of time correlation functions for a dense one-component plasma in a magnetic field. Journal of Statistical Physics, 53(5–6):1237–1260, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023867.

Schlijper:1989:TSB

[SS89] A. G. Schlijper and B. Smit. Two-sided bounds on the free energy from local states in Monte Carlo simulations. Journal of Statistical Physics, 56(3–4):247–260, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01044436.

Sancho:1982:AES

[SSD82] J. M. Sancho, M. San Miguel, and D. Dürr. Adiabatic elimination for systems of Brownian particles with nonconstant damping coefficients. Journal of Statistical Physics, 28(2):291–305, June 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012607.

Slusher:1985:LSS

[SSM85] R. E. Slusher, C. M. Surko, and C. A. Murray. Light scattering studies of the ion acoustic instability in a positive column plasma. Journal of Statistical Physics, 39(5–6):755–767, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008364.

Sauermann:1981:VPD

[ST81a] G. Sauermann and H. Turschner. Variational principle for derivation of macroscopic equations. Journal of Statistical Physics, 25(3):459–473, July 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010800.

Schrader:1981:ANE

[ST81b] R. Schrader and E. Tränkle. Analytic and numerical evidence from quantum field theory for the hyperscaling relation dv = 2Δ − γ in the d = 3 Ising model. Journal of Statistical Physics, 25(2):269–290, June 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01022187.

Szasz:1981:RWI

[ST81c] D. Szász and A. Telcs. Random walk in an inhomogeneous medium with local impurities. Journal of Statistical Physics, 26 (3):527–537, November 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011433.

Selinger:1984:KBL

[ST84a] Jonathan V. Selinger and U. M. Titulaer. The kinetic boundary layer for the Klein–Kramers equation; a new numerical approach. Journal of Statistical Physics, 36(3–4):293–319, August 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010986.

Szasz:1984:PRW

[ST84b] D. Szász and B. Tóth. Persistent random walks in a onedimensional random environment. Journal of Statistical Physics, 37(1–2):27–38, October 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012903.

Szasz:1987:DTB

[ST87] Domokos Szász and Bálint Tóth. A dynamical theory of Brownian motion for the Rayleigh gas. Journal of Statistical Physics, 47 (5–6):681–693, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01206152.

Sokal:1988:AMG

[ST88] Alan D. Sokal and Lawrence E. Thomas. Absence of mass gap for a class of stochastic contour models. Journal of Statistical Physics, 51(5–6):907–947, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014892.

Seibold:1989:BVP

[ST89a] H. Seibold and C. Toepffer. The boundary value problem in fermion systems. Journal of Statistical Physics, 55(5–6): 1129–1155, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01041082.

Sokal:1989:ECE

[ST89b] Alan D. Sokal and Lawrence E. Thomas. Exponential convergence to equilibrium for a class of random-walk models. Journal of Statistical Physics, 54(3–4):797–828, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019776.

Szabo:1989:SBB

[ST89c] K. G. Szabó and T. Tél. On the symmetry-breaking bifurcation of chaotic attractors. Journal of Statistical Physics, 54(3– 4):925–948, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019782.

Stanley:1984:AFC

[Sta84] H. Eugene Stanley. Application of fractal concepts to polymer statistics and to anomalous transport in randomly porous media. Journal of Statistical Physics, 36(5–6):843–860, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012944.

Stadje:1987:EPD

[Sta87a] W. Stadje. The exact probability distribution of a twodimensional random walk. Journal of Statistical Physics, 46(1– 2):207–216, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010341.

Stauffer:1987:FFK

[Sta87b] D. Stauffer. On forcing functions in Kauffman’s random Boolean networks. Journal of Statistical Physics, 46(3–4):789–794, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013386.

Stadje:1988:ANT

[Sta88a] Wolfgang Stadje. Asymptotic normality in a two-dimensional random walk model for cell motility. Journal of Statistical Physics, 51(3–4):615–635, May 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01028475.

Sugiyama:1988:QMC

[STA88b] G. Sugiyama, L. Terray, and B. J. Alder. Quantum Monte Carlo study of a proton in an electron gas. Journal of Statistical Physics, 52(5–6):1221–1232, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011643.

Stadje:1989:EPD

[Sta89] Wolfgang Stadje. Exact probability distributions for noncorrelated random walk models. Journal of Statistical Physics, 56 (3–4):415–435, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01044444.

Stephen:1981:WFD

[Ste81] Michael J. Stephen. Wave functions in disordered systems. Journal of Statistical Physics, 25(4):663–668, August 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01022360.

Stein:1987:BRE

[Ste87] D. L. Stein. Book review: Evolution, games, and learning. Journal of Statistical Physics, 49(5–6):1255–1256, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017570.

Stein:1989:OCC

[Ste89] N. D. Stein. Oscillatory convection and chaos in a Lorenz-type model of a rotating fluid. Journal of Statistical Physics, 56(5–6): 841–878, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016782.

Stillinger:1980:SCM

[Sti80] F. H. Stillinger. Simple cell model with collapse instability. Journal of Statistical Physics, 23(2):219–230, August 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012592.

Stoller:1986:NAQ

[Sto86] Lincoln Stoller. A new approach to quantum spin-systems: The critical properties of a Heisenberg ferromagnet. Journal of Statistical Physics, 43(5–6):879–881, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF02628317.

Stolzle:1988:UTD

[Stö88] Susanne Stölzle. Universality two-dimensional Kauffman model for parallel and sequential updating. Journal of Statistical Physics, 53(3–4):995–1004, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01014236.

Stroscio:1982:ETS

[Str82] Michael A. Stroscio. Enhancement of turbulence in a stratified fluid by the presence of a shear field. Journal of Statistical Physics, 28(3):607–612, July 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008328.

Streater:1987:BMQ

[Str87] R. F. Streater. A Boltzmann map for quantum oscillators. Journal of Statistical Physics, 48(3–4):753–767, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019695.

Stratonovich:1988:DLN

[Str88] R. L. Stratonovich. Derivation of the linear and nonlinear non-Markov fluctuation-dissipation relations of the first kind for a dynamical model. Journal of Statistical Physics, 53(1–2):5–18, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011541.

Sugar:1986:MCS

[Sug86] R. L. Sugar. Monte Carlo studies with the ST-100 array processor. Journal of Statistical Physics, 43(5–6):1147–1170, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02628338.

Suto:1980:PTI

[Süt80] A. Sütó. Phase transition for Ising frustration potentials. Journal of Statistical Physics, 23(2):203–217, August 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012591.

Suto:1989:SCS

[Süt89] András Sütö. Singular continuous spectrum on a Cantor set of zero Lebesgue measure for the Fibonacci Hamiltonian. Journal of Statistical Physics, 56(3–4):525–531, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01044450.

Suzuki:1986:QSM

[Suz86a] Masuo Suzuki. Quantum statistical Monte Carlo methods and applications to spin systems. Journal of Statistical Physics, 43 (5–6):883–909, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF02628318.

Suzuki:1986:TFD

[Suz86b] Masuo Suzuki. Thermo field dynamics of quantum spin systems. Journal of Statistical Physics, 42(5–6):1047–1070, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010461.

Suzuki:1987:NFS

[Suz87] Masuo Suzuki. Nonlinear fluctuations, separation of procedures, and linearization of processes. Journal of Statistical Physics, 49(5–6):977–992, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01017555.

Suzuki:1988:SEF

[Suz88] Masuo Suzuki. Super-effective-field theory and coherent-anomaly method in cooperative phenomena. Journal of Statistical Physics, 53(1–2):483–497, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011568.

Sinai:1980:DCO

[SV80] Ja. G. Sinai and E. B. Vul. Discovery of closed orbits of dynamical systems with the use of computers. Journal of Statistical Physics, 23(1):27–47, July 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014428.

Schulz:1988:TFS

[SVBO88] U. Schulz, J. Villain, E. Brézin, and H. Orland. Thermal fluctuations in some random field models. Journal of Statistical Physics, 51(1–2):1–27, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01015318.

Salinas:1985:MFI

[SW85] S. R. Salinas and W. F. Wreszinski. On the mean-field Ising model in a random external field. Journal of Statistical Physics, 41(1– 2):299–313, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020615.

Stilck:1987:EPS

[SW87] Jürgen F. Stilck and John C. Wheeler. Equilibrium polymerization in a solvent: Solution on the Bethe lattice. Journal of Statistical Physics, 46(1–2):1–34, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010327.

Stillinger:1988:NOS

[SW88] F. H. Stillinger and T. A. Weber. Nonlinear optimization simplified by hypersurface deformation. Journal of Statistical Physics, 52(5–6):1429–1445, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011658.

Swendsen:1984:MCR

[Swe84] Robert H. Swendsen. Monte Carlo renormalization group. Journal of Statistical Physics, 34(5–6):963–973, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009451.

Szulga:1987:PTO

[SWYM87] Jerzy Szulga, Wojbor A. Woyczynski, Bernard Ycart, and J. Adin Mann, Jr. The phase transition in a one-dimensional lattice of axisymmetric bodies. Journal of Statistical Physics, 46(1–2):67–85, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010331.

Suto:1983:PAM

[SYG83] András Sütö, Tarik Yalcin, and Christian Gruber. A probabilistic approach to the models of spin glasses. Journal of Statistical Physics, 31(3):639–659, June 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01019502.

Sylvester:1983:MCF

[Syl83] Garrett S. Sylvester. Magnetization concavity in ferromagnets. Journal of Statistical Physics, 33(1):91–98, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009750.

So:1984:CSI

[SYOM84] Byon Chol So, Nobuyuki Yoshitake, Hisao Okamoto, and Hazime Mori. Correlations and spectra of an intermittent chaos near its onset point. Journal of Statistical Physics, 36(3–4):367–400, August 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010988.

Shnidman:1988:LGH

[SZ88] Yitzhak Shnidman and R. K. P. Zia. A lattice-gas Hamiltonian for micellar binary solutions. Journal of Statistical Physics, 50(3– 4):839–848, February 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01026502.

Szasz:1980:JDL

[Szá80] Domokos Szász. Joint diffusion on the line. Journal of Statistical Physics, 23(2):231–240, August 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012593.

Szamel:1989:CSP

[Sza89] Grzegorz Szamel. Comment on symmetry properties of the linear Enskog kinetic operators. Journal of Statistical Physics, 55 (1–2):381–387, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01042607.

tHooft:1988:ERB

[’t 88] Gerard ’t Hooft. Equivalence relations between deterministic and quantum mechanical systems. Journal of Statistical Physics, 53 (1–2):323–344, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011560.

Takesue:1989:EPT

[Tak89] Shinji Takesue. Ergodic properties and thermodynamic behavior of elementary reversible cellular automata. I. Basic properties. Journal of Statistical Physics, 56(3–4):371–402, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01044442.

Tarnawski:1987:OOC

[Tar87] Maciek Tarnawski. An order-by-order construction of lowtemperature phase diagrams for classical lattice systems. Journal of Statistical Physics, 47(5–6):757–759, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01206156.

Tasaki:1985:GBC

[Tas85] Hal Tasaki. Gaussian bounds for correlations in lattice spin systems. Journal of Statistical Physics, 38(5–6):873–890, March 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010421.

Tasaki:1987:GCE

[Tas87] Hal Tasaki. Geometric critical exponent inequalities for general random cluster models. Journal of Statistical Physics, 49(3–4): 841–847, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009360.

Tasaki:1989:UCD

[Tas89] Hal Tasaki. On the upper critical dimensions of random spin systems. Journal of Statistical Physics, 54(1–2):163–170, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023476.

Tang:1988:MFT

[TB88] Chao Tang and Per Bak. Mean field theory of self-organized critical phenomena. Journal of Statistical Physics, 51(5–6):797– 802, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014884.

Tartaglia:1984:ASC

[TC84] Piero Tartaglia and Sow-Hsin Chen. Application of stochastic computer simulation to a solution of a nonlinear Fokker–Planck equation governing bacteria chemotaxis. Journal of Statistical Physics, 35(1–2):171–186, April 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01017373.

Tejero:1989:LRI

[TCB89] C. F. Tejero, J. A. Cuesta, and R. Brito. Long-range inverse twospin correlations in one-dimensional Potts lattices. Journal of Statistical Physics, 56(1–2):33–42, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01044229.

Tejero:1987:ODI

[Tej87] C. F. Tejero. One-dimensional inhomogeneous Ising model: A new approach. Journal of Statistical Physics, 48(3–4):531–538, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019686.

Tejero:1989:ZFE

[Tej89] C. F. Tejero. Zero-frequency elastic moduli of uniform fluids. Journal of Statistical Physics, 57(1–2):393–398, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023651.

Tel:1983:ICA

[Tél83] Tamás Tél. Invariant curves, attractors, and phase diagram of a piecewise linear map with chaos. Journal of Statistical Physics, 33 (1):195–221, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009756.

Tel:1988:SDS

[Tél88] Tamás Tél. On the stationary distribution of self-sustained oscillators around bifurcation points. Journal of Statistical Physics, 50 (5–6):897–912, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019146.

Theumann:1986:SML

[TF86] W. K. Theumann and José F. Fontanari. The spherical-model limit in a random field. Journal of Statistical Physics, 45(1– 2):99–112, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01033080.

Thomae:1981:CSP

[TG81] Stefan Thomae and Siegfried Grossmann. Correlations and spectra of periodic chaos generated by the logistic parabola. Journal of Statistical Physics, 26(3):485–504, November 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011430.

Torney:1987:RDL

[TG87] David C. Torney and Byron Goldstein. Rates of diffusion-limited reaction in periodic systems. Journal of Statistical Physics, 49(3– 4):725–750, November 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009354.

Tang:1988:LDC

[TG88] Lei-Han Tang and Robert B. Griffiths. Localized defects in classical one-dimensional models. Journal of Statistical Physics, 53(3– 4):853–892, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014229.

Tracy:1987:MPH

[TGN87] Craig A. Tracy, Larry Grove, and M. F. Newman. Modular properties of the hard hexagon model. Journal of Statistical Physics, 48(3–4):477–502, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01019683.

Tasaki:1984:MFB

[TH84] Hal Tasaki and Takashi Hara. Mean field bound and GHS inequality. Journal of Statistical Physics, 35(1–2):99–107, April 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017367.

Talkner:1987:DDM

[THFT87] P. Talkner, P. Hänggi, E. Freidkin, and D. Trautmann. Discrete dynamics and metastability: Mean first passage times and escape rates. Journal of Statistical Physics, 48(1–2):231–254, July 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010408.

Thompson:1982:LPI

[Tho82a] Colin J. Thompson. Local properties of an Ising model on a Cayley tree. Journal of Statistical Physics, 27(3):441–456, March 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011085.

Thompson:1982:MFE

[Tho82b] Colin J. Thompson. On mean field equations for spin glasses. Journal of Statistical Physics, 27(3):457–472, March 1982. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011086.

Titulaer:1984:DPK

[Tit84] U. M. Titulaer. The density profile for the Klein–Kramers equation near an absorbing wall. Journal of Statistical Physics, 37(5– 6):589–607, December 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010497.

Titulaer:1988:MBO

[Tit88] Urbaan M. Titulaer. On the microscopic background of the Onsager–Casimir reciprocity relations. Journal of Statistical Physics, 53(1–2):61–75, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011545.

Tedeschini-Lalli:1982:TNS

[TL82] Laura Tedeschini-Lalli. Truncated Navier–Stokes equations: continuous transition from a five-mode to a seven-mode model. Journal of Statistical Physics, 27(2):365–388, February 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008944.

Leung:1988:IPD

[tL88] Kwan tai Leung. Interfacial properties of a driven diffusive system. Journal of Statistical Physics, 50(1–2):405–437, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023001.

Leung:1986:FTCa

[tLC86a] Kwan tai Leung and John L. Cardy. Field theory of critical behavior in a driven diffusive system. Journal of Statistical Physics, 44 (3–4):567–588, August 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011310.

Leung:1986:FTCb

[tLC86b] Kwan tai Leung and John L. Cardy. Field theory of critical behavior in a driven diffusive system. Journal of Statistical Physics, 45 (5–6):1087, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020590.

Thompson:1988:AES

[TM88] Colin J. Thompson and J. B. McGuire. Asymptotic and essentially singular solutions of the Feigenbaum equation. Journal of Statistical Physics, 51(5–6):991–1007, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014896.

Thirumalai:1989:PEN

[TM89] D. Thirumalai and Raymond D. Mountain. Probes of equipartition in nonlinear Hamiltonian systems. Journal of Statistical Physics, 57(3–4):789–801, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01022832.

Torney:1986:VRR

[Tor86a] David C. Torney. Variance of the range of a random walk. Journal of Statistical Physics, 44(1–2):49–66, July 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010904.

Torquato:1986:MCB

[Tor86b] S. Torquato. Microstructure characterization and bulk properties of disordered two-phase media. Journal of Statistical Physics, 45(5–6):843–873, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020577.

Tourenne:1983:EBS

[Tou83] Christian J. Tourenne. The entropy of the BKW solution. Journal of Statistical Physics, 32(1):71–80, July 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009420.

Tourenne:1985:MCB

[Tou85] Christian J. Tourenne. A method to calculate Boltzmann entropy. Journal of Statistical Physics, 41(1–2):335–348, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020618.

Tough:1989:ECQ

[Tou89] J. T. Tough. The effects of colored quadratic noise on a turbulent transition in liquid^{4} He. Journal of Statistical Physics, 54(5–6):1173, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01044709.

Tracy:1986:BMC

[Tra86] Craig A. Tracy. Z_{n} Baxter model: Critical behavior. Journal of Statistical Physics, 44(1–2):183–191, July 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010910.

Tracy:1988:UCF

[Tra88] Craig A. Tracy. Universality class of a Fibonacci Ising model. Journal of Statistical Physics, 51(3–4):481–490, May 1988. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01028467.

Truong:1986:SPS

[Tru86] T. T. Truong. Structural properties of a Z(N^{2})-spin model and its equivalent Z(N)-vertex model. Journal of Statistical Physics, 42(3–4):349–379, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01127716.

Takano:1981:MKR

[TS81] Hiroshi Takano and Masuo Suzuki. Migdal–Kadanoff renormalization group approach to the spin-1/2 anisotropic Heisenberg model. Journal of Statistical Physics, 26(4):635–663, December 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010931.

Tsallis:1988:PGB

[Tsa88] Constantino Tsallis. Possible generalization of Boltzmann–Gibbs statistics. Journal of Statistical Physics, 52(1–2):479–487, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016429.

Tirapegui:1988:TII

[TvdB88] E. Tirapegui and C. van den Broeck. A thermochemical instability. II. Inhomogeneous fluctuations. Journal of Statistical Physics, 50(1–2):439–458, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01023002.

Thiran:1989:AD

[TVW89] E. Thiran, D. Verstegen, and J. Weyers. p-adic dynamics. Journal of Statistical Physics, 54(3–4):893–913, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019780.

Tsvelick:1985:ESM

[TW85] A. M. Tsvelick and P. B. Wiegmann. Exact solution of the multichannel Kondo problem, scaling, and integrability. Journal of Statistical Physics, 38(1–2):125–147, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01017853.

Thirumalai:1986:PIM

[TWB86] D. Thirumalai, A. Wallqvist, and B. J. Berne. Path-integral Monte Carlo simulations of electron localization in water clusters. Journal of Statistical Physics, 43(5–6):973–984, June 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02628324.

Uchiyama:1988:BGL

[Uch88] Kôhei Uchiyama. On the Boltzmann–Grad limit for the Broadwell model of the Boltzmann equation. Journal of Statistical Physics, 52(1–2):331–355, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016418.

Ueyama:1980:FHE

[Uey80a] Hiroshi Ueyama. Fluctuating hydrodynamic equations of mixed and of chemically reacting gases. Journal of Statistical Physics, 23 (4):463–482, October 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011576.

Ueyama:1980:SBE

[Uey80b] Hiroshi Ueyama. The stochastic Boltzmann equation and hydrodynamic fluctuations. Journal of Statistical Physics, 22(1):1–26, January 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01007985.

Uimin:1984:LTA

[Uim84] G. Uimin. Low-temperature analysis of the ANNNI model in an external magnetic field: Cascades of phase transitions, complete devil’s staircases. Journal of Statistical Physics, 34(1–2):1–34, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01770347.

Ullah:1981:ESE

[Ull81] Nazakat Ullah. On the expansion of the single eigenvalue probability density function. Journal of Statistical Physics, 24(3): 413–418, March 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012813.

Unnerstall:1989:DCD

[Unn89] Thomas Unnerstall. Dynamics of the current-driven Josephson junction. Journal of Statistical Physics, 54(1–2):379–403, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023485.

Uimin:1984:TDC

[UP84] G. V. Uimin and V. L. Pokrovsky. Two-dimensional commensurate soliton structures. Journal of Statistical Physics, 34(5– 6):775–780, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009439.

Velarde:1981:PPT

[VA81] Manuel G. Velarde and J. Carlos Antoranz. Phenomenological phase transition approach to the two-component Rayleigh– Bénard problem: Tricritical behavior. Journal of Statistical Physics, 24(1):235–241, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01007646.

Vaienti:1989:CPA

[Vai89] S. Vaienti. Computing the pressure for Axiom-A attractors by time series and large deviations for the Lyapunov exponent. Journal of Statistical Physics, 56(3–4):403–413, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01044443.

Valsakumar:1983:SFP

[Val83] M. C. Valsakumar. Solution of Fokker–Planck equation using Trotter’s formula. Journal of Statistical Physics, 32(3): 545–553, September 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008955.

Valsakumar:1985:USD

[Val85] M. C. Valsakumar. Unstable state dynamics: Treatment of colored noise. Journal of Statistical Physics, 39(3–4):347–365, May 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01018667.

VandenBroeck:1983:CFT

[Van83a] C. Van den Broeck. Critical fluctuations in a thermochemical instability. I. Mean field description. Journal of Statistical Physics, 32(1):153–168, July 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009427.

VanDenBroeck:1983:RBW

[Van83b] C. Van Den Broeck. On the relation between white shot noise, Gaussian white noise, and the dichotomic Markov process. Journal of Statistical Physics, 31(3):467–483, June 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019494.

VanVliet:1988:VKO

[Van88] Carolyn M. Van Vliet. On van Kampen’s objections against linear response theory. Journal of Statistical Physics, 53(1–2):49–60, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011544.

Varley:1981:NAP

[Var81] R. L. Varley. A nonequilibrium analog of the Percus–Yevick equation. Journal of Statistical Physics, 24(2):301–324, February 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013303.

Vericat:1980:MSM

[VB80] Fernando Vericat and Lesser Blum. Mean spherical model for hard ions and dipoles: Thermodynamics and correlation functions. Journal of Statistical Physics, 22(5):593–604, May 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011340.

vanBeijeren:1984:MCT

[vB84] Henk van Beijeren. Mode-coupling theory for purely diffusive systems. Journal of Statistical Physics, 35(3–4):399–412, May 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014393.

vanBeijeren:1985:MCT

[vB85] Henk van Beijeren. Mode coupling theory for purely diffusive system. Journal of Statistical Physics, 39(3–4):449–450, May 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01018674.

VandenBroeck:1986:DTO

[VB86] C. Van den Broeck and M. Bouten. Decay times in onedimensional chains. Journal of Statistical Physics, 45(5–6):1031– 1047, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020587.

Vieillefosse:1989:SRO

[VB89] P. Vieillefosse and M. Brajon. Sum rules for the one-component plasma with additional short-range forces. Journal of Statistical Physics, 55(5–6):1169–1183, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01041084.

vanBeijeren:1988:ETN

[vBC88] H. van Beijeren and E. G. D. Cohen. The effects of thermal noise in a Rayleigh–Benard cell near its first convective instability. Journal of Statistical Physics, 53(1–2):77–94, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011546.

vanBeijeren:1980:KTHa

[vBD80a] Henk van Beijeren and J. R. Dorfman. Kinetic theory of hydrodynamic flows. I. The generalized normal solution method and its application to the drag on a sphere. Journal of Statistical Physics, 23(3):335–402, September 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011373.

vanBeijeren:1980:KTHb

[vBD80b] Henk van Beijeren and J. R. Dorfman. Kinetic theory of hydrodynamic flows. II. The drag on a sphere and on a cylinder. Journal of Statistical Physics, 23(4):443–461, October 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011575.

vanBeijeren:1982:KTH

[vBD82] Henk van Beijeren and J. R. Dorfman. Kinetic theory of hydrodynamic flows, III. The torque on a rotating sphere or cylinder. Journal of Statistical Physics, 29(1):139–153, September 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008254.

vanBeijeren:1988:EP

[vBEL88] H. van Beijeren, M. H. Ernst, and J. L. Lebowitz. Editors’ preface. Journal of Statistical Physics, 53(1–2):1, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011539.

vanBeijeren:1987:TDC

[vBK87] Henk van Beijeren and Ryszard Kutner. Tracer diffusion in concentrated lattice gas models. Rectangular lattices with anisotropic jump rates. Journal of Statistical Physics, 49(5–6):1043–1051, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017559.

vanBeijeren:1980:ETC

[vBLLS80] H. van Beijeren, O. E. Lanford III, J. L. Lebowitz, and H. Spohn. Equilibrium time correlation functions in the low-density limit. Journal of Statistical Physics, 22(2):237–257, February 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008050.

vanBeijeren:1983:TPO

[vBS83] Henk van Beijeren and Herbert Spohn. Transport properties of the one-dimensional stochastic Lorentz model: I. Velocity autocorrelation function. Journal of Statistical Physics, 31(2):231– 254, May 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011581.

vanDongen:1987:FCSb

[vD87] P. G. J. van Dongen. Fluctuations in coagulating systems. II. Journal of Statistical Physics, 49(5–6):927–975, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017554.

vanDongen:1988:FRD

[vD88] P. G. J. van Dongen. Fluctuations in reaction-diffusion systems: A new exactly soluble growth model. Journal of Statistical Physics, 53(1–2):221–232, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011554.

vanDongen:1989:SFR

[vD89a] P. G. J. van Dongen. Spatial fluctuations in reaction-limited aggregation. Journal of Statistical Physics, 54(1–2):221–271, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023480.

Vanderhaegen:1989:LRT

[VD89b] D. Vanderhaegen and C. Deutsch. Linear radiation transport in randomly distributed binary mixtures: A one-dimensional and exact treatment for the scattering case. Journal of Statistical Physics, 54(1–2):331–360, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01023483.

vandenBerg:1983:CFB

[vdB83] M. van den Berg. On condensation in the free-boson gas and the spectrum of the Laplacian. Journal of Statistical Physics, 31(3):623–637, June 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019501.

vandenBerg:1984:CIB

[vdBLdS84] M. van den Berg, J. T. Lewis, and P. de Smedt. Condensation in the imperfect boson gas. Journal of Statistical Physics, 37(5– 6):697–707, December 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010502.

vanDongen:1984:KRP

[vDE84] P. G. J. van Dongen and M. H. Ernst. Kinetics of reversible polymerization. Journal of Statistical Physics, 37(3–4):301–324, November 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011836.

vanDongen:1986:OGT

[vDE86] P. G. J. van Dongen and M. H. Ernst. On the occurrence of a gelation transition in Smoluchowski’s coagulation equation. Journal of Statistical Physics, 44(5–6):785–792, September 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011907.

vanDongen:1987:FCSa

[vDE87] P. G. J. van Dongen and M. H. Ernst. Fluctuations in coagulating systems. Journal of Statistical Physics, 49(5–6):879–926, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017553.

vanDongen:1988:SSS

[vDE88] P. G. J. van Dongen and M. H. Ernst. Scaling solutions of Smoluchowski’s coagulation equation. Journal of Statistical Physics, 50 (1–2):295–329, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01022996.

vanDamme:1989:CBC

[vDG89] Ruud van Damme and Bernard J. Geurts. Complexes of block copolymers in solution: A graph-theoretical approach. Journal of Statistical Physics, 57(5–6):1099–1122, December 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020050.

vanEnter:1985:BCD

[vE85] Aernout C. D. van Enter. Bounds on correlation decay for longrange vector spin glasses. Journal of Statistical Physics, 41(1– 2):315–321, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020616.

vanEnter:1987:PSA

[vE87a] Aernout C. D. van Enter. Proof of Straley’s argument for bootstrap percolation. Journal of Statistical Physics, 48(3–4):943–945, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019705.

vanEnter:1987:UBC

[vE87b] Aernout C. D. van Enter. Upper bounds on correlation decay for one-dimensional long-range spin-glass models. Journal of Statistical Physics, 47(5–6):905–910, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01206164.

vanEnter:1989:RDN

[vEF89] A. C. D. van Enter and R. Fernández. A remark on different norms and analyticity for many-particle interactions. Journal of Statistical Physics, 56(5–6):965–972, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01016790.

vanEnter:1983:TLL

[vEvH83] A. C. D. van Enter and J. L. van Hemmen. The thermodynamic limit for long-range random systems. Journal of Statistical Physics, 32(1):141–152, July 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009426.

vanEnter:1985:APT

[vEvH85] A. C. D. van Enter and J. L. van Hemmen. Absence of phase transitions in certain one-dimensional long-range random systems. Journal of Statistical Physics, 39(1–2):1–13, April 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01007972.

Vogelsang:1987:DFC

[VH87] R. Vogelsang and C. Hoheisel. Determination of the friction coefficient via the force autocorrelation function. A molecular dynamics investigation for a dense Lennard-Jones fluid. Journal of Statistical Physics, 47(1–2):193–207, April 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009042.

Vogelsang:1989:FCL

[VH89] R. Vogelsang and C. Hoheisel. The friction coefficient of a Lennard-Jones fluid from the random force autocorrelation function determined as a memory function by molecular dynamics calculations. Journal of Statistical Physics, 54(1–2):315–329, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023482.

vanHemmen:1984:SWQ

[vHBW84] J. L. van Hemmen, A. A. S. Brito, and W. F. Wreszinski. Spin waves in quantum ferromagnets. Journal of Statistical Physics, 37 (1–2):187–213, October 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012911.

vanHemmen:1988:NNNa

[vHGHK88a] J. L. van Hemmen, D. Grensing, A. Huber, and R. Kühn. Nonlinear neural networks. I. General theory. Journal of Statistical Physics, 50(1–2):231–257, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01022994.

vanHemmen:1988:NNNb

[vHGHK88b] J. L. van Hemmen, D. Grensing, A. Huber, and R. Kühn. Nonlinear neural networks. II. Information processing. Journal of Statistical Physics, 50(1–2):259–293, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01022995.

vanHemmen:1988:MSP

[vHZ88] J. L. van Hemmen and V. A. Zagrebnov. Models of a structural phase transition with general anharmonicity and disorder. Journal of Statistical Physics, 53(3–4):835–852, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014228.

Vidal:1987:EST

[Vid87] C. Vidal. Experimental study of traveling waves and target patterns in oscillatory reacting media. Journal of Statistical Physics, 48(5–6):1017–1030, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009530.

Vieillefosse:1985:SRP

[Vie85] P. Vieillefosse. Sum rules and perfect screening conditions for the one-component plasma. Journal of Statistical Physics, 41(5–6): 1015–1035, December 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010013.

Vigfusson:1982:TRS

[Vig82] J. O. Vigfusson. Time relaxation of the solutions of master equations for large systems. Journal of Statistical Physics, 27(2): 339–353, February 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008942.

Vilgis:1987:DSH

[Vil87] T. A. Vilgis. On the dynamics of simple hierarchical systems. Journal of Statistical Physics, 47(1–2):133–148, April 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009038.

Visscher:1981:EFP

[Vis81] P. B. Visscher. Exact fixed points in discrete hydrodynamics. Journal of Statistical Physics, 25(2):211–227, June 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01022183.

Visscher:1985:RGD

[Vis85] P. B. Visscher. Renormalization-group derivation of Navier– Stokes equation. Journal of Statistical Physics, 38(5–6):989–1013, March 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010426.

vanKampen:1981:BR

[vK81a] N. G. van Kampen. Book review. Journal of Statistical Physics, 26(4):847–848, December 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010946.

vanKampen:1981:IVS

[vK81b] N. G. van Kampen. Itô versus Stratonovich. Journal of Statistical Physics, 24(1):175–187, January 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01007642.

vanKampen:1981:VNL

[vK81c] N. G. van Kampen. The validity of nonlinear Langevin equations. Journal of Statistical Physics, 25(3):431–442, July 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010798.

vanKampen:1986:BMM

[vK86] N. G. van Kampen. Brownian motion on a manifold. Journal of Statistical Physics, 44(1–2):1–24, July 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010902.

vanKampen:1987:CEA

[vK87a] N. G. van Kampen. Chapman–Enskog as an application of the method for eliminating fast variables. Journal of Statistical Physics, 46(3–4):709–727, February 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013381.

vanKampen:1987:IFE

[vK87b] N. G. van Kampen. Intrinsic fluctuations in explosive reactions. Journal of Statistical Physics, 46(5–6):933–948, March 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011150.

vanKampen:1989:LLE

[vK89] N. G. van Kampen. Langevin-like equation with colored noise. Journal of Statistical Physics, 54(5–6):1289–1308, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01044716.

Vinokurov:1985:RBV

[VKK85] L. I. Vinokurov, A. V. Kats, and V. M. Kontorovich. The relation between the velocity and mass distributions. The role of collisionless relaxation processes. Journal of Statistical Physics, 38 (1–2):217–229, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01017859.

vanLeeuwen:1989:DIE

[vL89] J. M. J. van Leeuwen. Does the interface experience the van der Waals loop? Journal of Statistical Physics, 57(3–4):433–453, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01022815.

Valles:1986:NPT

[VM86] J. L. Vallés and J. Marro. Nonequilibrium phase transitions in stochastic lattice systems: Influence of the hopping rates. Journal of Statistical Physics, 43(3–4):441–461, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020647.

Valles:1987:NSO

[VM87] J. L. Vallés and J. Marro. Nonequilibrium second-order phase transitions in stochastic lattice systems: A finite-size scaling analysis in two dimensions. Journal of Statistical Physics, 49(1– 2):89–119, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009956.

Valles:1988:NSO

[VM88] J. L. Vallés and J. Marro. Nonequilibrium second-order phase transitions in stochastic lattice systems: A finite-size scaling analysis in two dimensions. Journal of Statistical Physics, 51(1–2): 321, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01015335.

Valsakumar:1983:LNL

[VMA83] M. C. Valsakumar, K. P. N. Murthy, and G. Ananthakrishna. On the linearization of nonlinear Langevin-type stochastic differential equations. Journal of Statistical Physics, 30(3):617–631, March 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009680.

VandenBroeck:1982:APC

[VMB82] C. Van den Broeck, M. Malek Mansour, and F. Baras. Asymptotic properties of coupled nonlinear Langevin equations in the limit of weak noise. I: Cusp bifurcation. Journal of Statistical Physics, 28(3):557–575, July 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008324.

Voorhees:1985:TOR

[Voo85] P. W. Voorhees. The theory of Ostwald ripening. Journal of Statistical Physics, 38(1–2):231–252, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01017860.

Voss:1984:MFA

[Vos84] Richard F. Voss. Multiparticle fractal aggregation. Journal of Statistical Physics, 36(5–6):861–872, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012945.

Voigtlaender:1985:SFP

[VR85] K. Voigtlaender and H. Risken. Solutions of the Fokker–Planck equation for a double-well potential in terms of matrix continued fractions. Journal of Statistical Physics, 40(3–4):397–429, August 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01017181.

Veerman:1989:RAM

[VT89] J. J. P. Veerman and F. M. Tangerman. Renormalization of Aubry–Mather Cantor sets. Journal of Statistical Physics, 56 (1–2):83–98, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01044233.

VanAssche:1986:NLY

[VTB86] W. Van Assche, G. Turchetti, and D. Bessis. On the nature of the Lee–Yang measure for Ising ferromagnets. Journal of Statistical Physics, 43(1–2):85–108, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010573.

Vichniac:1986:AQI

[VTH86] G. Y. Vichniac, P. Tamayo, and H. Hartman. Annealed and quenched inhomogeneous cellular automata (INCA). Journal of Statistical Physics, 45(5–6):875–883, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020578.

Valles:1989:DNL

[VtLZ89] J. L. Vallés, K. t. Leung, and R. K. P. Zia. Driven nonequilibrium lattice systems with shifted periodic boundary conditions. Journal of Statistical Physics, 56(1–2):43–58, July 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01044230.

Vannimenus:1989:CTA

[VV89] J. Vannimenus and X. G. Viennot. Combinatorial tools for the analysis of ramified patterns. Journal of Statistical Physics, 54(5– 6):1529–1538, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01044733.

vanVelzen:1987:BPT

[vVE87] G. A. van Velzen and M. H. Ernst. Bond percolation in two and three dimensions: Numerical evaluation of time-dependent transport properties. Journal of Statistical Physics, 48(3–4):677–708, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019692.

Wagner:1983:CTD

[Wag83] Heinz-Jürgen Wagner. Crystallinity in two dimensions: A note on a paper of C. Radin. Journal of Statistical Physics, 33(3): 523–526, December 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01018831. See [Rad81].

Wagner:1989:MST

[Wag89] Heinz-Jürgen Wagner. The microscopic stress tensor field in particle systems with many-body interactions. Journal of Statistical Physics, 55(5–6):1293–1302, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01041089.

Wallace:1984:P

[Wal84] D. J. Wallace. Preface. Journal of Statistical Physics, 34(5–6): 661, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01009433;   http://link.
springer.com/article/10.1007/BF01009433.

Watanabe:1980:KEK

[Wat80] Masaaki Watanabe. Kinetic equation in the kinetic region of the dilute and nonuniform electron plasma. Journal of Statistical Physics, 22(5):553–562, May 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011338.

Watanabe:1989:BSA

[Wat89] Hiroshi Watanabe. Block spin approach to ϕ_{3}^{4} field theory. Journal of Statistical Physics, 54(1–2):171–190, January 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01023477.

Wagner:1987:ADC

[WB87] N. Wagner and I. Balberg. Anomalous diffusion and continuum percolation. Journal of Statistical Physics, 49(1–2):369–382, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009969.

Wright:1999:EME

[WB99] J. D. Wright and David Brydges. Erratum: Mayer Expansions and the Hamiltonian–Jacobi Equation. II. Fermions, Dimensional Reduction Formulas. Journal of Statistical Physics, 97(5–6):1027, December 1999. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1023/A%3A1004666114935. See [BW88b].

Widom:1983:SOC

[WBKS83] Michael Widom, David Bensimon, Leo P. Kadanoff, and Scott J. Shenker. Strange objects in the complex plane. Journal of Statistical Physics, 32(3):443–454, September 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008949.

Wang:1989:CSD

[WBL89] Jian-Sheng Wang, Kurt Binder, and Joel L. Lebowitz. Computer simulation of driven diffusive systems with exchanges. Journal of Statistical Physics, 56(5–6):783–819, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01016780.

Wright:1981:SOZ

[WC81] C. C. Wright and P. T. Cummings. Solution of the Ornstein– Zernike equation for a softcore Yukawa fluid. II. Numerical results. Journal of Statistical Physics, 24(2):389–401, February 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013308.

Weitz:1989:BRR

[WC89] David Weitz and Antonio Coniglio. Book review: Random fluctuations and pattern growth: Experiments and models. Journal of Statistical Physics, 56(5–6):973–976, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01016791.

Wenzel:1985:SES

[WCD85] R. G. Wenzel, J. L. Carlsten, and K. J. Drühl. Soliton experiments in stimulated Raman scattering. Journal of Statistical Physics, 39(5–6):621–632, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008356.

Weiss:1986:FPT

[WDG86] George H. Weiss, Edmund A. DiMarzio, and Richard J. Gaylord. First passage time densities for random walk spans. Journal of Statistical Physics, 42(3–4):567–572, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01127728.

Wood:1982:LSD

[WE82] William W. Wood and Jerome J. Erpenbeck. On the linearity of the self-diffusion process. Journal of Statistical Physics, 27 (1):37–56, January 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011738.

Webman:1984:PTE

[Web84] Itzhak Webman. Propagation and trapping of excitations on percolation clusters. Journal of Statistical Physics, 36(5–6): 603–614, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012925.

Weiss:1981:FPT

[Wei81] George H. Weiss. First passage time problems for one-dimensional random walks. Journal of Statistical Physics, 24(4):587–594, April 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011376.

Weiss:1982:BRA

[Wei82] George H. Weiss. Book review: Approximation of population processes. Journal of Statistical Physics, 27(4):823, April 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013450.

Weiss:1983:BRS

[Wei83] George H. Weiss. Book review: Stochastic processes in physics and chemistry. Journal of Statistical Physics, 31(2):423–424, May 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011591.

Weiss:1984:BRH

[Wei84a] George H. Weiss. Book review: Handbook of stochastic methods for physics, chemistry, and the natural sciences. Journal of Statistical Physics, 36(1–2):289–290, July 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01015736.

Weiss:1984:FPT

[Wei84b] George H. Weiss. First passage times for correlated random walks and some generalizations. Journal of Statistical Physics, 37(3–4): 325–330, November 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011837.

Weiss:1985:BR

[Wei85a] George H. Weiss. Book review. Journal of Statistical Physics, 39(3–4):443–444, May 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01018672.

Weiss:1985:CBS

[Wei85b] N. O. Weiss. Chaotic behavior in stellar dynamos. Journal of Statistical Physics, 39(5–6):477–491, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008346.

Weiss:1986:BRB

[Wei86a] George H. Weiss. Book review: Brownian motion and stochastic flow systems. Journal of Statistical Physics, 44(1–2):277, July 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010919.

Weiss:1986:BRE

[Wei86b] George H. Weiss. Book review: Enigmas of chance, an autobiography. Journal of Statistical Physics, 44(3–4):703–705, August 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011315.

Weiss:1986:BRH

[Wei86c] George H. Weiss. Book review: Handbook of stochastic methods for physics, chemistry, and the natural sciences. Journal of Statistical Physics, 45(5–6):1089, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020591.

Weiss:1986:CTE

[Wei86d] George H. Weiss. Competitive trapping effects in a set of partially absorbing traps. Journal of Statistical Physics, 44(5–6): 933–939, September 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011915.

Weiss:1986:OTM

[Wei86e] George H. Weiss. Overview of theoretical models for reaction rates. Journal of Statistical Physics, 42(1–2):3–36, January 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010838.

Weinberger:1987:MNS

[Wei87a] Ed Weinberger. A model of natural selection that exhibits a dynamic phase transition. Journal of Statistical Physics, 49(5–6): 1011–1028, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01017557.

Weiss:1987:BRK

[Wei87b] George H. Weiss. Book review: The kind of motion we call heat, and fluctuation phenomena. Journal of Statistical Physics, 49 (5–6):1259, December 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01017572.

Weiss:1989:BRP

[Wei89] George H. Weiss. Book review: Principles of radiophysics. Journal of Statistical Physics, 55(5–6):1325–1326, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01041093.

Wertheim:1984:FHDa

[Wer84a] M. S. Wertheim. Fluids with highly directional attractive forces. I. Statistical thermodynamics. Journal of Statistical Physics, 35 (1–2):19–34, April 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01017362.

Wertheim:1984:FHDb

[Wer84b] M. S. Wertheim. Fluids with highly directional attractive forces. II. Thermodynamic perturbation theory and integral equations. Journal of Statistical Physics, 35(1–2):35–47, April 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017363.

Wertheim:1986:FHDa

[Wer86a] M. S. Wertheim. Fluids with highly directional attractive forces. III. Multiple attraction sites. Journal of Statistical Physics, 42(3– 4):459–476, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01127721.

Wertheim:1986:FHDb

[Wer86b] M. S. Wertheim. Fluids with highly directional attractive forces. IV. Equilibrium polymerization. Journal of Statistical Physics, 42(3–4):477–492, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01127722.

Wertheim:1988:CTD

[Wer88] M. S. Wertheim. Coincidence theorem for the direct correlation function of hard-particle fluids. Journal of Statistical Physics, 52(5–6):1367–1387, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011654.

Westcott:1982:RWL

[Wes82] Mark Westcott. Random walks on a lattice. Journal of Statistical Physics, 27(1):75–82, January 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011741.

West:1983:BRN

[Wes83a] Bruce J. West. Book review: Nonlinear phenomena in physics and biology. Journal of Statistical Physics, 31(2):425–428, May 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011592.

West:1983:BRS

[Wes83b] Bruce J. West. Book review: Stability of thermodynamic systems. Journal of Statistical Physics, 33(1):233–236, October 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009759.

West:1984:BRB

[Wes84a] Bruce J. West. Book review: Nonequilibrium problems in the physical sciences and biology. Volume II: Long-time prediction in dynamics. Journal of Statistical Physics, 34(1–2):347–349, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01770365.

West:1984:BRR

[Wes84b] Bruce J. West. Book review: Regular and stochastic motion. Journal of Statistical Physics, 35(5–6):761–763, June 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010832.

West:1985:BRN

[Wes85a] Bruce J. West. Book review: Noise induced transitions, theory, and applications in physics, chemistry, and biology. Journal of Statistical Physics, 38(5–6):1099–1101, March 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010431.

West:1985:BRS

[Wes85b] Bruce J. West. Book review: Statistical physics and chaos in fusion plasma. Journal of Statistical Physics, 41(1–2):349–350, October 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020619.

West:1986:BRBa

[Wes86a] Bruce J. West. Book review: Chaos and statistical methods. Journal of Statistical Physics, 42(3–4):723, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01127736.

West:1986:BRBb

[Wes86b] Bruce J. West. Book review: Chemical oscillations, waves and turbulence. Journal of Statistical Physics, 42(5–6):1137–1138, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010468.

West:1987:BRS

[Wes87] Bruce J. West. Book review: Statistical optics. Journal of Statistical Physics, 48(1–2):361–362, July 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010414.

West:1988:BRF

[Wes88] Bruce J. West. Book review: Frontiers of nonequilibrium statistical physics. Journal of Statistical Physics, 53(3–4):1007–1009, November 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014238.

West:1989:BRC

[Wes89] Bruce J. West. Book review: Convection and chaos in fluids. Journal of Statistical Physics, 55(5–6):1323–1324, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01041092.

Williams:1988:ASS

[WF88] G. O. Williams and H. L. Frisch. Anomalous scaling in systems partially controlled by diffusion. Journal of Statistical Physics, 52(5–6):1233–1245, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011644.

Weiss:1984:TRW

[WH84] George H. Weiss and Shlomo Havlin. Trapping of random walks on the line. Journal of Statistical Physics, 37(1–2):17–25, October 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012902.

Weiss:1985:SPR

[WHB85] George H. Weiss, Shlomo Havlin, and Armin Bunde. On the survival probability of a random walk in a finite lattice with a single trap. Journal of Statistical Physics, 40(1–2):191–199, July 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01010532. See comment [dH85a].

Whittington:1983:SAW

[Whi83] S. G. Whittington. Self-avoiding walks with geometrical constraints. Journal of Statistical Physics, 30(2):449–456, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012318.

Whittle:1989:SRD

[Whi89] P. Whittle. The statistics of random directed graphs. Journal of Statistical Physics, 56(3–4):499–516, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01044448.

Weiss:1989:PNM

[WHM89] George H. Weiss, Shlomo Havlin, and Ofer Matan. Properties of noninteger moments in a first passage time problem. Journal of Statistical Physics, 55(1–2):435–439, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01042610.

Wick:1985:DPT

[Wic85] W. David Wick. A dynamical phase transition in an infinite particle system. Journal of Statistical Physics, 38(5–6):1015–1025, March 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010427.

Wick:1989:HLN

[Wic89] W. David Wick. Hydrodynamic limit of a nongradient interacting particle process. Journal of Statistical Physics, 54(3–4):873–892, February 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019779.

Widom:1988:NIT

[Wid88a] B. Widom. Note on the interfacial tension of phase-separated polymer solutions. Journal of Statistical Physics, 52(5–6):1343– 1351, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011651.

Widom:1988:SST

[Wid88b] B. Widom. Scaling of the surface tension of phase-separated polymer solutions. Journal of Statistical Physics, 53(1–2):523–529, October 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011570.

Wiesenfeld:1985:NPN

[Wie85] Kurt Wiesenfeld. Noisy precursors of nonlinear instabilities. Journal of Statistical Physics, 38(5–6):1071–1097, March 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010430.

Weidenmuller:1984:SDM

[WJS84] H. A. Weidenmüller and Zhang Jing-Shang. Stationary diffusion over a multidimensional potential barrier: A generalization of Kramers’ formula. Journal of Statistical Physics, 34(1– 2):191–201, January 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01770354.

Williams:1984:NMC

[WK84] G. O. Williams and M. H. Kalos. A new multispin coding algorithm for Monte Carlo simulation of the Ising model. Journal of Statistical Physics, 37(3–4):283–299, November 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011835.

Whitlock:1983:SVM

[WKC83] P. A. Whitlock, M. H. Kalos, and G. V. Chester. A Schwinger variational method for the Bloch equation. Journal of Statistical Physics, 32(2):389–405, August 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012717.

West:1989:PFD

[WKL89] Bruce J. West, Raoul Kopelman, and Katja Lindenberg. Pattern formation in diffusion-limited reactions. Journal of Statistical Physics, 54(5–6):1429–1439, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01044727.

Wada:1988:RPP

[WKUK88] Koh Wada, Makoto Kaburagi, Takashi Uchida, and Ryoichi Kikuchi. Reformulation of the path probability method and its application to crystal growth models. Journal of Statistical Physics, 53(5–6):1081–1101, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01023859.

Wang:1988:PTU

[WL88] Jian-Sheng Wang and Joel L. Lebowitz. Phase transitions and universality in nonequilibrium steady states of stochastic Ising models. Journal of Statistical Physics, 51(5–6):893–906, June 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014891.

Wu:1983:KAS

[WM83] C. H. Wu and E. W. Montroll. Kinetics of adsorption on stepped surfaces and the determination of surface diffusion constants. Journal of Statistical Physics, 30(2):537–547, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012327.

Wilbertz:1988:SDP

[WMvBL88] Harald Wilbertz, Jan Michels, Henk van Beijeren, and Jan Adriaan Leegwater. Self-diffusion of particles interacting through a square-well or square-shoulder potential. Journal of Statistical Physics, 53(5–6):1155–1177, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01023862.

West:1986:BRBc

[WO86] Bruce J. West and Irwin Oppenheim. Book review: Chaotic behavior in quantum systems. Journal of Statistical Physics, 45(1– 2):353–355, October 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01033096.

Wolynes:1982:BRF

[Wol82] Peter Guy Wolynes. Book review: Functional integration: Theory and applications. Journal of Statistical Physics, 28(2):409–410, June 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012615.

Wolfram:1986:CAF

[Wol86] Stephen Wolfram. Cellular automaton fluids 1: Basic theory. Journal of Statistical Physics, 45(3–4):471–526, November 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01021083.

Wood:1989:LTT

[Woo89] W. W. Wood. Long-time tails of the Green–Kubo integrands for a binary mixture. Journal of Statistical Physics, 57(3–4):675–727, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01022828.

Wiegel:1982:SMR

[WP82] Frederik W. Wiegel and Alan S. Perelson. Statistical mechanics of red blood cell aggregation: The distribution of rouleaux in thermal equilibrium. Journal of Statistical Physics, 29(4): 813–848, December 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011794.

Wiegel:1987:BR

[WP87] Frederik W. Wiegel and Adrian Parsegian. Book reviews. Journal of Statistical Physics, 46(1–2):429–432, January 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010353.

Weiss:1980:ICS

[WR80] George H. Weiss and Robert J. Rubin. Internal configurations of span-constrained random walks. Journal of Statistical Physics, 22(1):97–109, January 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01007992.

Wreszinski:1983:VSM

[Wre83] Walter F. Wreszinski. Vanishing spontaneous magnetization for quantum mechanical models of a spin glass. Journal of Statistical Physics, 32(2):407–412, August 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012718.

West:1983:AGR

[WRL83] Bruce J. West, Galina Rovner, and Katja Lindenberg. Approximate Gaussian representation of evolution equations I. Single degree of freedom nonlinear equations. Journal of Statistical Physics, 30(3):633–648, March 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009681.

Weiss:1982:END

[WS82] George H. Weiss and Michael F. Shlesinger. On the expected number of distinct points in a subset visited by an n-step random walk. Journal of Statistical Physics, 27(2):355–363, February 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008943.

West:1983:RWD

[WS83] Bruce J. West and Michael F. Shlesinger. Random walk of dislocations following a high-velocity impact. Journal of Statistical Physics, 30(2):527–535, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012326.

West:1984:FIW

[WS84] Bruce J. West and Michael F. Shlesinger. The fractal interpretation of the weak scattering of elastic waves. Journal of Statistical Physics, 36(5–6):779–786, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012937.

Wei:1988:SGS

[WS88] I-Chien Wei and Robert L. Scott. In search of the Griffiths shield region. Journal of Statistical Physics, 52(5–6):1315– 1324, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011649.

Weiss:1983:OSF

[WSL83] George H. Weiss, Kurt E. Shuler, and Katja Lindenberg. Order statistics for first passage times in diffusion processes. Journal of Statistical Physics, 31(2):255–278, May 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011582.

Widder:1989:KBL

[WT89a] M. E. Widder and U. M. Titulaer. The kinetic boundary layer around an absorbing sphere and the growth of small droplets. Journal of Statistical Physics, 55(5–6):1109–1128, June 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01041081.

Widder:1989:MMB

[WT89b] M. E. Widder and U. M. Titulaer. The moment method for boundary layer problems in Brownian motion theory. Journal of Statistical Physics, 56(3–4):471–498, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01044447.

Wiesenfeld:1989:PS

[WTB89] Kurt Wiesenfeld, Chao Tang, and Per Bak. A physicist’s sandbox. Journal of Statistical Physics, 54(5–6):1441–1458, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01044728.

Wu:1980:ERD

[Wu80] F. Y. Wu. Exact results for a dilute Potts model. Journal of Statistical Physics, 23(6):773–782, December 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008519.

Wu:1985:EST

[Wu85] F. Y. Wu. Exact solution of a triangular Ising model in a nonzero magnetic field. Journal of Statistical Physics, 40(5–6): 613–620, September 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009892.

Wu:1986:TDI

[Wu86] F. Y. Wu. Two-dimensional Ising model with crossing and fourspin interactions and a magnetic field i(π/2)kT. Journal of Statistical Physics, 44(3–4):455–463, August 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011305.

Wu:1988:PMG

[Wu88] F. Y. Wu. Potts model and graph theory. Journal of Statistical Physics, 52(1–2):99–112, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01016406.

Welford:1982:EPT

[WW82] W. T. Welford and R. Winston. The ellipsoid paradox in thermodynamics. Journal of Statistical Physics, 28(3):603–606, July 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008327.

Wehner:1986:NIE

[WW86a] M. F. Wehner and W. G. Wolfer. A new integral equation for the radial distribution function of a hard sphere fluid. Journal of Statistical Physics, 42(3–4):493–508, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01127723.

Wehner:1986:PHS

[WW86b] M. F. Wehner and W. G. Wolfer. The pressure of a hard sphere fluid on a curved surface. Journal of Statistical Physics, 42(3– 4):509–521, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01127724.

Weiss:1987:BR

[WW87] George H. Weiss and F. Y. Wu. Book review. Journal of Statistical Physics, 48(3–4):951–953, August 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019707.

Weiss:1988:MSR

[WW88a] George H. Weiss and Haim Weissman. A measure of the symmetry of random walks. Journal of Statistical Physics, 52(1–2):287– 293, July 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01016415.

Wu:1988:BEG

[WW88b] X. N. Wu and F. Y. Wu. Blume–Emery–Griffiths model on the honeycomb lattice. Journal of Statistical Physics, 50(1–2):41–55, January 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01022986.

Weissman:1989:TPC

[WWH89] Haim Weissman, George H. Weiss, and Shlomo Havlin. Transport properties of the continuous-time random walk with a longtailed waiting-time density. Journal of Statistical Physics, 57(1– 2):301–317, October 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023645.

Wimp:1989:HLP

[WZ89] Jet Wimp and Doron Zeilberger. How likely is Pólya’s drunkard to stay in x ≥ y ≥ z ? Journal of Statistical Physics, 57(5–6): 1129–1135, December 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020052.

Wansleben:1984:MCS

[WZK84] Stephan Wansleben, John G. Zabolitzky, and Claus Kalle. Monte Carlo simulation of Ising models by multispin coding on a vector computer. Journal of Statistical Physics, 37(3–4):271–282, November 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011834.

Xu:1989:BSV

[XS89] Jiasai Xu and George Stell. Bulk and shear viscosities of a polydisperse hard-sphere fluid. Journal of Statistical Physics, 57(3– 4):921–936, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01022841.

Yip:1982:ETC

[YAA82] Sidney Yip, W. E. Alley, and B. J. Alder. Evaluation of time correlation functions from a generalized Enskog equation. Journal of Statistical Physics, 27(1):201–217, January 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011747.

Yang:1987:DST

[Yan87] Wei-Shih Yang. Debye screening for two-dimensional Coulomb systems at high temperatures. Journal of Statistical Physics, 49 (1–2):1–32, October 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009952.

Yu:1989:PSS

[YAR89] Anthony W. Yu, Govind P. Agrawal, and Rajarshi Roy. Power spectra and spatial pattern dynamics of a ring laser. Journal of Statistical Physics, 54(5–6):1223–1241, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01044714.

Yeung:1989:ECS

[Yeu89] C. Yeung. Explicit construction of steady state of a model of chemical turbulence. Journal of Statistical Physics, 55(1–2):357– 380, April 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01042606.

Yamazaki:1980:DRS

[YHM80] Yoshitake Yamazaki, Henk J. Hilhorst, and Günther Meissner. Differential real-space renormalization of the d-dimensional Gaussian model. Journal of Statistical Physics, 23(5):609–625, November 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011732.

Yamazaki:1985:SDP

[YHOF85] Yoshitake Yamazaki, Arno Holz, Moyuru Ochiai, and Yoshiichi Fukuda. Static and dynamic properties of XY systems with extended defects in cubic anisotropic crystallines. Journal of Statistical Physics, 41(3–4):497–530, November 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009020.

Yip:1989:CSC

[Yip89] Sidney Yip. Commentary on the self-consistent mode-coupling approximation. Journal of Statistical Physics, 57(3–4):665–673, November 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01022827.

Yamaguchi:1981:SSN

[YKS81] Y. Yamaguchi, K. Kometani, and H. Shimizu. Self-synchronization of nonlinear oscillations in the presence of fluctuations. Journal of Statistical Physics, 26(4):719–743, December 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010935.

Yoshida:1983:ASC

[YMS83] T. Yoshida, H. Mori, and H. Shigematsu. Analytic study of chaos of the tent map: Band structures, power spectra, and critical behaviors. Journal of Statistical Physics, 31(2):279–308, May 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011583.

Yokoi:1986:DPC

[YNS86] Carlos S. O. Yokoi, John F. Nagle, and Silvio R. Salinas. Dimer pair correlations on the brick lattice. Journal of Statistical Physics, 44(5–6):729–747, September 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011905.

Young:1984:SG

[You84] A. P. Young. Spin glasses. Journal of Statistical Physics, 34 (5–6):871–881, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009446.

Yukhnovsky:1985:TSP

[YS85] I. R. Yukhnovsky and Yu. V. Shulepov. Theory of the spontaneous polarization of the adsorbed monolayer of polar molecules. The collective variables method. Journal of Statistical Physics, 38(3–4):541–572, February 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01010477.

Zabolitzky:1988:CPR

[Zab88] John G. Zabolitzky. Critical properties of rule 22 elementary cellular automata. Journal of Statistical Physics, 50(5–6):1255–1262, March 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019164.

Zagrebnov:1982:SPK

[Zag82] V. A. Zagrebnov. Spectral properties of Kirkwood–Salsburg and Kirkwood–Ruelle operators. Journal of Statistical Physics, 27 (3):577–591, March 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011093.

Zahradnik:1987:ALT

[Zah87] Milos Zahradník. Analyticity of low-temperature phase diagrams of lattice spin models. Journal of Statistical Physics, 47(5–6):725– 755, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01206155.

Zia:1988:SCC

[ZAT88] R. K. P. Zia, J. E. Avron, and J. E. Taylor. The Summertop construction: Crystals in a corner. Journal of Statistical Physics, 50(3–4):727–736, February 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01026498.

Zumofen:1989:LWT

[ZBKS89] G. Zumofen, A. Blumen, J. Klafter, and M. F. Shlesinger. Lévy walks for turbulence: A numerical study. Journal of Statistical Physics, 54(5–6):1519–1528, March 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01044732.

Zabolitzky:1986:PCE

[ZBS86] J. G. Zabolitzky, D. J. Bergman, and D. Stauffer. Precision calculation of elasticity for percolation. Journal of Statistical Physics, 44(1–2):211–223, July 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010913.

Zanon:1988:WDE

[ZD88] N. Zanon and B. Derrida. Weak disorder expansion of Liapunov exponents in a degenerate case. Journal of Statistical Physics, 50(3–4):509–528, February 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01026489.

Zegarlinski:1986:EGM

[Zeg86] Bogusław Zegarliński. Extremality and the global Markov property II: The global Markov property for non-FKG maximal Gibbs measures. Journal of Statistical Physics, 43(3–4):687–705, May 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020660.

Zegarlinski:1987:SGL

[Zeg87] B. Zegarliński. Spin glasses with long-range interaction at high temperatures. Journal of Statistical Physics, 47(5–6):911–930, June 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01206165.

Zhang:1988:PCF

[Zha88] M. Q. Zhang. Pair correlation function for Ising spins with competing dynamics. Journal of Statistical Physics, 53(5–6): 1217–1225, December 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023865.

Zhang:1989:FVM

[Zha89a] M. Q. Zhang. A fast vectorized multispin coding algorithm for 3D Monte Carlo simulations using Kawasaki spin-exchange dynamics. Journal of Statistical Physics, 56(5–6):939–950, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016787.

Zhang:1989:DPH

[Zha89b] Yi-Cheng Zhang. Directed polymers in Hartree–Fock approximation. Journal of Statistical Physics, 57(5–6):1123–1128, December 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020051.

Zia:1986:EES

[Zia86] R. K. P. Zia. Exact equilibrium shapes of Ising crystals on triangular/honeycomb lattices. Journal of Statistical Physics, 45(5– 6):801–813, December 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020575.

Ziff:1980:KP

[Zif80] Robert M. Ziff. Kinetics of polymerization. Journal of Statistical Physics, 23(2):241–263, August 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012594.

Zippelius:1981:EEN

[ZL81] Annette Zippelius and Manfred Lücke. The effect of external noise in the Lorenz model of the Bénard problem. Journal of Statistical Physics, 24(2):345–358, February 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01013305.

Zolafdek:1986:LST

[Zol86] H. Zolafdek. Localization of the spectrum of transfer matrix of Ising model. Journal of Statistical Physics, 43(1–2):109–122, April 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010574.

Zoladek:1987:ODR

[Zol87] H. Zoladek. One-dimensional random walk with self-interaction. Journal of Statistical Physics, 47(3–4):543–550, May 1987. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01007525.

Zhang:1989:IIM

[ZP89] M. Q. Zhang and J. K. Percus. Inhomogeneous Ising model on a multiconnected network. Journal of Statistical Physics, 56(5–6): 695–708, September 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016775.

Zhabotinsky:1987:MND

[ZR87] A. M. Zhabotinsky and A. B. Rovinsky. Mechanism and nonlinear dynamics of an oscillating chemical reaction. Journal of Statistical Physics, 48(5–6):959–975, September 1987. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009526.

Zurek:1980:DPB

[ZS80] W. H. Zurek and W. C. Schieve. Deviations from the Poisson behavior of equilibrium fluctuations in a closed system. Journal of Statistical Physics, 22(3):289–296, March 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014645.

Zheng:1982:TPS

[ZS82] W. M. Zheng and W. C. Schieve. Transition probabilities and stochastic equations for the mean field Ising model. Journal of Statistical Physics, 29(2):375–385, October 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020793.

Zhou:1988:HSF

[ZS88] Yaoqi Zhou and George Stell. The hard-sphere fluid: New exact results with applications. Journal of Statistical Physics, 52(5–6): 1389–1412, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011655.

Zvyagin:1985:QGS

[ZT85] A. A. Zvyagin and V. M. Tsukernik. The quantum ground state of a Heisenberg ferromagnet with an ‘easy-plane’ type anisotropy. Journal of Statistical Physics, 38(1–2):405–411, January 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01017870.

Zwanzig:1981:BRS

[Zwa81] Robert Zwanzig. Book review: Studies in statistical mechanics. Journal of Statistical Physics, 26(3):629–630, November 1981. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011438.

Zwanzig:1982:NMD

[Zwa82] Robert Zwanzig. Non-Markoffian diffusion in a one-dimensional disordered lattice. Journal of Statistical Physics, 28(1):127–133, May 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011627.

Zwanzig:1983:CDC

[Zwa83] Robert Zwanzig. From classical dynamics to continuous time random walks. Journal of Statistical Physics, 30(2):255–262, February 1983. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012300.

Zhang:1988:PLD

[ZWLV88] M. Q. Zhang, J.-S. Wang, J. L. Lebowitz, and J. L. Vallés. Power law decay of correlations in stationary nonequilibrium lattice gases with conservative dynamics. Journal of Statistical Physics, 52(5–6):1461–1478, September 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011660.